Module Catalog International AVS Center

Date: November 9, 1995

2

1. International AVS Center Module Catalog

This catalog summarizes the public-domain AVS modules available from the International AVS Center as of May 1, 1993. The catalog separates the AVS modules into four different sections – data input, filter, mapper, and data output – and alphabetically organizes modules within each section.

A postscript version of this module catalog is available on the International AVS Center ftp site. It is updated continuously and is suitable for printing on any postscript printer. For help on acquiring this information, please read the following Obtaining Modules section.

2. International AVS Center

“The North Carolina Supercomputing Center was chosen [as the International AVS Center] primarily because of its unparalleled capability to gather and distribute modules and act as a user/vendor communication channel via its CONCERT network, which is ideal to meet the information and module distribution needs of the AVS users and vendors,” says Paul Esdale of Advanced Visual Systems Inc.

The International AVS Center serves as a catalyst for expanding the AVS user base and for increasing the functionality of AVS by fostering discipline-specific module development and development of new AVS applications. As a world clearinghouse for AVS information, the International AVS Center collects, ports, and distributes user-contributed, public-domain AVS modules and acts as liaison between AVS users and vendors.

In addition to fostering collaborations among AVS users with related interests, the International AVS Center coordinates an international AVS users group, publishes the quarterly AVS Network News magazine, operates an AVS showroom, and serves as a porting facility for distributing AVS applications and modules across heterogeneous platforms.

3. AVS Consortium

The AVS Consortium provides growth, direction, and promotion strategies for AVS to become the de facto standard visualization environment. It also manages the International AVS Center’s activities and the allocation of its resources.

In operation since October 1, 1991, the AVS Consortium currently comprises representatives from various AVS vendors, including Advanced Visual Systems Inc., CONVEX Computer Corporation, Digital Equipment Corporation, Hewlett-Packard, IBM, SUN Microsystems, and Wavetracer. These vendors’ cooperative effort to provide a common interface for their users’ benefit exemplifies true vision and marks a tremendous milestone in visualization. Additional AVS vendors may join the consortium in the near future.

4. Obtaining Modules

AVS modules may be obtained by standard ftp protocol, email, or standard mail. The International AVS Center ftp site is located on the Internet at avs.ncsc.org (128.109.178.23). When connected to the avs server, log in as anonymous and enter your email address as the password. Instructions on navigating through the ftp site’s multiple directories are in your local man pages on ftp.

Several introductory files are at the ftp site’s top level, including an FAQ, this module catalog, and an AVS README file that includes detailed information on the International AVS Center ftp site. The top level also includes four directories that correspond to this catalog’s data, filters, mappers, and renderers sections.

Each of these four directories has subdirectories for individual modules that include module source code, module documentation, sample networks, and Makefiles for each platform to which the International AVS Center staff port the module (e.g., make.CONVEX, make.DEC, make.IBM, make.HP, make.Sun, and make.Wave).

For users who do not have ftp protocol on their system, an email system is available for obtaining modules. The three Internet email address are avs@ncsc.org for general questions and module submissions; avsemail@ncsc.org for current catalog and README retrieval; and avsorder@ncsc.org for automated module retrieval that includes all files associated with the retrieved module.

Until completion of current construction of porting facilities for modules on tape, tape requests continue to be serviced by email and standard mail. For instructions on obtaining modules on tape, send email to avs@ncsc.org or standard mail to the International AVS Center, Post Office Box 12889, 3021 Cornwallis Road, Research Triangle Park, North Carolina 27709-2889.

5. Submitting Modules

Donating a module to the International AVS Center benefits the entire AVS user community by facilitating AVS use to visualize complex scientific phenomena. A module that is not donated may be duplicated by someone else, wasting that person’s time and possibly hindering further development of other module capabilities for all users’ benefit. All modules listed in this catalog are free submissions by module authors.

Modules can be submitted by ftp, email, or tape methods that parallel the preceding module retrieval methods. Modules can be placed on the ftp site within the SUBMIT directory, sent by email to avs@ncsc.org, or put on tape and mailed to the International AVS Center.

All submissions must include the entire source code for the module, a Makefile for the appropriate platforms upon which the module was developed, and documentation on the module’s usage. A sample AVS network demonstrating module usage also is requested.

Submitting Modules vi

INTERNATIONAL AVS CENTER MODULE CATALOG

secno = ount0

DATA INPUT MODULES

International AVS Center

DATA INPUT MODULES

3D axis

Author :  Anonymous  This module displays the coordinate
Kubota Pacific Computer  axis of the 3D graphic display.

Inc

Number :  1152
Language :  C
Submitted :  03/12/92
Version :  1.000
Last Updated :  03/12/92
Ported To :  Kubota DEC IBM Sun

ACE to UCD

Author :  Ian Curington  This is intended as a source code
Advanced Visual Systems,  example to show how to allocate and
Inc.  manage a simple UCD output port
Number :  1637  structure, from within a free-running
Language :  C  AVS coroutine. This simulates time
Submitted :  09/24/93  varying data on a 2D quadrilateral
Version :  1.000  mesh. It has no parameters or input
Last Updated :  09/24/93  ports, for simplicity all values are
Ported To :  HP Kubota Sun  hand-coded. The UCD output structure

contains a 2D mesh of quadrilateral cells, with connected shared nodes. The geometric and connectivity info is only built the first time, then only the values change from each time step.

INTERNATIONAL AVS CENTER MODULE CATALOG

AFS extract

Author :  Jeff Wang  This module allows users to extract
MCNC, Environmetal  data from AIRS AFS (AIRS Facility
Programs Group  System) files. The AIRS AFS format are
Number :  1732  refered to User’s Guide For the Urban
Language :  C  Airshed Model, EPA-450/4-90-007D. Users
Submitted :  12/15/93  may read through the very begining of
Version :  1.000  the c codes provided to get an idea how
Last Updated :  12/15/93  the data is constructed. This is one of
Ported To :  Sun  a suite of modules we used to plot the

emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

AMS extract

Author :  Jeff Wang  This module allows users to extract
MCNC, Environmetal  data from AIRS AMS (AIRS Area and
Programs Group  Mobile Subsystem) files. The AIRS AMS
Number :  1733  format are refered to User’s Guide For
Language :  C  the Urban Airshed Model,
Submitted :  12/15/93  EPA-450/4-90-007D. Users may read
Version :  1.000  through the very begining of the C
Last Updated :  12/15/93  codes provided to get an idea how the
Ported To :  Sun  data is constructed. This is one of a

suite of modules we used to plot the emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

ANIMATE DUAL

Author :  Kathy Pearson  Unlike the AVS module, “animated
MCNC Environmental  integer“, this animator co-routine
Programs  keeps up with the range of the values
Number :  1793  available. It also allows the user to
Language :  C  set a delta value for going forward
Submitted :  04/21/94  (positive) or backward (negative).
Version :  1.000  Additionally, the user can advance by
Last Updated :  04/21/94  single steps through data or
Ported To :  DEC HP Kubota Sun  continuously. When the end of the range

is reached, the animator will wrap around the end of the range to the opposite end by delta. A second set of variables allows the synchronous control of two viewing windows. Output from the animator will cease for “Pause Seconds“ after each previous output. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

ARC MAP

Author :  Kathy Pearson  The ARCMAP module creates an AVS
MCNC Environmental  geometry by reading input from an
Programs  ARCinfo gen file. The gen file
Number :  1792  structure is that of ARC segment
Language :  C  number, a list of x-y coordinates of
Submitted :  04/21/94  the segment polygon, and an ’END’
Version :  1.000  statement. The MCNC Environmental
Last Updated :  04/21/94  Programs Group has created this module
Ported To :  DEC HP IBM Kubota Sun  as part of a suite of AVS demo modules,

networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Application Menu

Author :  Joseph Fall  The ApplMenu module reads a menu
Essential Planning  description file and creates an AVS
Systems  menu that conforms to this description.
Number :  1829  It also controls this menu and issues
Language :  C  the appropriate CLI commands when
Submitted :  07/16/94  various menu choices are made by a
Version :  1.000  user.
Last Updated :  07/16/94
Ported To :  DEC HP Kubota Sun

INTERNATIONAL AVS CENTER MODULE CATALOG

BGF Protein Disp

Author :  Advanced Visual Systems, This module is both a reader and a
Inc.  mapper. It reads ASCII Biograph protein
Number :  1679  files, and generates a geometric object
Language :  C  for the geometry viewer. It has
Submitted :  09/29/93  advanced modes, such as creating a
Version :  1.000  spline ribbon representation of the
Last Updated :  09/29/93  backbone chain.
Ported To :  HP IBM Kubota Sun

BP Coroutine

Author :  Ian Curington  This module is a source example of how
Advanced Visual Systems  to write a free-running asynchronous
Number :  1325  coroutine that provides data to AVS on
Language :  C  the output port when data is ready. In
Submitted :  08/31/92  its present form, it free-runs, that is
Version :  1.000  continually sends data to AVS as fast
Last Updated :  08/31/92  as it can be processed. The project
Ported To :  DEC Sun HP  specific version of this had external

calls to obtain data from a data acquisition instrument.

ConstImage

Author :  Stefano Brozzi, Andrea  This module was built due to the need
Menozzi, Giulio E.  to be able to generate a constant image
Destri  to input as bias or as initial state in
Universita’ degli Studi  the module aCNN. The module accept as
di Parma, 43100 Parma,  input a field, tipically the image, and
Italy  assumig the dimensions of this field
Number :  1873  gives as output a field 2-D scalar byte
Language :  C  of a constant value.
Submitted :  09/06/95
Version :  1.000
Last Updated :  09/06/95
Ported To :  DEC ultrix DEC alpha

DATA INPUT MODULES

DMOL to field

Author :  Mark Reed  This module reads the graphical output
North Carolina  files generated by DMol from Biosym
Supercomputing Center  (see REFERENCES included in
Number :  1857  documentation), and generates an AVS
Language :  C  field. This field can then be
Submitted :  10/17/94  manipulated to display an isosurface of
Version :  1.000  the electronic charge or use slicers to
Last Updated :  10/17/94  view the data.
Ported To :  DEC alpha HP IBM Sun

Dials Contro

Author :  Anonymous  This simple module provides rotation,
Stardent Computer  scale, and translation to any object in
Number :  1027  the geometry viewer scene, with
Language :  C  separate controls in the form of dial
Submitted :  10/10/91  parameters. The module connects in
Version :  1.000  PARALLEL with other geometric objects
Last Updated :  10/10/91  to the geometry viewer. It operates on
Ported To :  Sun Kubota IBM DEC HP  the named object in the scene.

Dials Control

Author :  James Peters (Convex),  This simple module provides rotation,
Ian Curington (AVS  scale, and translation to any object in
Inc.), and Stephen Pass  the geometry viewer scene, with
(???)  separate controls in the form of dial
Number :  1745  parameters. The module connects in
Language :  C  PARALLEL with other geometric inputs to
Submitted :  12/15/93  the geometry viewer. It operates on the
Version :  1.000  named object in the scene.
Last Updated :  12/15/93
Ported To :  Kubota Sun

Dipole

Author :  John Sheehan  This is an example module to show how
Advanced Visual Systems  one can go from cylindrical or
Number :  1701  spherical to cartesian coordinates.
Language :  C  This has been written for specific
Submitted :  09/30/93  datasets, but it should serve the
Version :  1.000  purpose. Anyways, the trick is to
Last Updated :  09/30/93  calculate the points array properly in
Ported To :  HP IBM Kubota Sun  order to feed field2mesh, in most

cases. See also - Read.

Edit File

Author :  Jeff Wang  This module allows users to run a vi
MCNC, Ennvironmetal  editor on a remote machine inside AVS.
Programs Group  The user selects the file to edit and
Number :  1720  inputs the IP number for the local
Language :  C  machine.
Submitted :  09/30/93
Version :  1.000
Last Updated :  09/30/93
Ported To :  HP Kubota Sun

Edit String

Author :  Ian Curington  The Edit String module puts up a LONG
Advanced Visual Systems  dialog box, allows typing of any
Number :  1328  string, then with either “OK” or
Language :  C  “CANCEL” controls, will output this
Submitted :  08/31/92  string to the output port of the
Version :  1.000  module. The next time the “Edit” button
Last Updated :  08/31/92  is pushed, the dialog box reappears,
Ported To :  with the old string, which can be

edited. This module is for long expressions, lists of ints, etc, where the normal string type in widget is too small. The dialog box comes up with nicer colours than the default grey.

Extrusion

Author :  Feizal Mirza  This module reads an ASCII file
University of Texas CHPC  consisting of lists of x and y points
Number :  1313  and constructs a 3-D geometry object
Language :  C  using the points in the file as
Submitted :  08/30/92  vertices. It uses the value specified
Version :  1.000  by the user in the depth dial for the
Last Updated :  08/30/92  depth of the object.
Ported To :  DEC Sun HP

FIND DATA AND MA

Author :  Kathy Pearson  FIND DATA AND MAP – locates data and
MCNC Environmental  map based on user choice from a DATA
Programs  LIST. The MCNC Environmental Programs
Number :  1796  Group has created this module as part
Language :  C  of a suite of AVS demo modules,
Submitted :  04/21/94  networks, and sample data sets to
Version :  1.000  demonstrate the functionality of its
Last Updated :  04/21/94  visualization prototype for an
Ported To :  DEC HP IBM Sun  Environmental Decision Support System

(EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

FIND MAP

Author :  Kathy Pearson  FIND MAP – locates a map based on user
MCNC Environmental  choice from a MAP LIST. The MCNC
Programs  Environmental Programs Group has
Number :  1797  created this module as part of a suite
Language :  C  of AVS demo modules, networks, and
Submitted :  04/21/94  sample data sets to demonstrate the
Version :  1.000  functionality of its visualization
Last Updated :  04/21/94  prototype for an Environmental Decision
Ported To :  DEC HP IBM Kubota Sun  Support System (EDSS). You can download

these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

FITS read

Author :  D. LE CORRE  This DATA INPUT module reads a FITS
CONVEX S.A  (Flexible Image Transport System) file
Number :  1306  into a field, and optionally to a text
Language :  CandFORTRAN  window. For primary data or image
Submitted :  08/30/92  extensions, the number of dimensions is
Version :  2.000  derived from the NAXIS keyword. For
Last Updated :  11/04/92  Ascii and binary tables, the number of
Ported To :  IBM  dimensions may be 1 or 2, depending on

a parameter. Random Groups do not seem to be supported (I did’nt see any example of them). Version 2 has code to port to the IBM platform.

GET MODEL MAP

Author :  Kathy Pearson  GET MODEL MAP – generates map and crop
MCNC Environmental  information based on selected map. The
Programs  MCNC Environmental Programs Group has
Number :  1799  created this module as part of a suite
Language :  C  of AVS demo modules, networks, and
Submitted :  04/21/94  sample data sets to demonstrate the
Version :  1.000  functionality of its visualization
Last Updated :  04/21/94  prototype for an Environmental Decision
Ported To :  DEC HP IBM Sun  Support System (EDSS). You can download

these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Generate label

Author :  Penny Rheingans  Generate label creates a geom object
US EPA Scientific  describing a label object. This label
Visualization Center  is a title, that is it has a fixed
Number :  1300  position on the screen, regardless of
Language :  C  viewpoint transformations. This module
Submitted :  08/26/92  provides a portion of the functionality
Version :  1.000  of labels in the geometry viewer, but
Last Updated :  08/26/92  adds the ability to generate labels
Ported To :  DEC Sun HP  automatically. This is useful for

putting up frame or time-step numbers.

Get Terrain Data

Author :  Jeff Wang  This module allows users to read in
MCNC, Environmetal  sorted ASCII terrain data and generate
Programs Group  an 2-D irregular field. Together with
Number :  1780  the “Draw Map” module on the ftp site,
Language :  C  users can create a 2-D terrain plot
Submitted :  02/09/94  over the political boundary map. Sample
Version :  1.000  data for this module can be found via
Last Updated :  02/09/94  anonymous ftp to avs.ncsc.org, in
Ported To :  HP Sun  /sample data/avs data/Get Terrain D

Data/dma.dat.

Graph slider

Author :  Hideo Miyachi  Graph slider can shift the range in x
Kubota Corporation  axis in graph viewer. By default, you
Number :  1614  can specify the max and min value
Language :  C  respectively. You can set the range and
Submitted :  05/01/93  min value in this module so that if you
Version :  1.000  change only min value the graph slides.
Last Updated :  05/01/93  This module has no input and output
Ported To :  Kubota HP  ports. That means this module accesses

CLI. This module is very simple and short code. You can use as a refernce for CLI access.

Image Sequen

Author :  Brian Kaplan  The Image Sequencer module is used to
Center for Innovative  read a series of images that have been
Computer Applications  previously saved to disk and output
Number :  1134  them as AVS image data. This can be
Language :  C  very useful for displaying a series of
Submitted :  02/23/92  images, recording them to video, or
Version :  1.000  compositing or filtering a series of
Last Updated :  02/23/92  images. The images can be in any of the
Ported To :  Kubota HP  formats supported by the Image

Sequencer. Image formats currently supported are AVS image format, MIFF, RLE, and RLA/B.

LBF extract

Author :  Jeff Wang  This module allows users to extract
MCNC, Environmetal  data from link base mobile emission
Programs Group  files. The link base emission format
Number :  1734  are refered to User’s Guide For the
Language :  C  Urban Airshed Model, EPA-450/4-90-007D.
Submitted :  12/15/93  Users may read through the very
Version :  1.000  begining of the C codes provided to get
Last Updated :  12/15/93  an idea how the data is constructed.
Ported To :  Sun  This is one of a suite of modules we

used to plot the emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Label Axis

Author :  Roy Wirthlin  The Label Axis module generates axes
Advanced Visual Systems  with tic marks and labels in 3D space.
Inc  The axis can be generated with tic
Number :  1289  marks and tic labels along each axis.
Language :  C  This module very simular to the
Submitted :  07/03/92  supported module generate axis but
Version :  1.000  porvides more flexibility and
Last Updated :  07/03/92  funcionality to the user.
Ported To :  Sun HP

Life

Author :  Steve Thorpe  Life is an implementation of the “Life”
International AVS Center  problem. This is an algorithm where
Number :  1277  “neighbors”, or cells, propagate or
Language :  C  die, based on the proximity of other
Submitted :  05/27/92  living neighbors. In this case, a
Version :  1.000  living neighbor is represented by a
Last Updated :  05/27/92  simulation of a processor in the
Ported To :  DEC IBM Sun Kubota HP  massively parrallel Wavetracer being

“turned on”. The thoughts behind the algorithm are - If there are too many neighbors, a cell will die due to overcrowding. If there are too few neighbors, a cell will die due to lack of resources. It is meant to be used with the Life WT AVS module, which is a parrallel implementation of Life. As the number of processors increases, the performance of Life WT greatly supercedes that of Life, due to the benefits of the Wavetracer’s multiple processors operating in parrallel. The output from this coroutine module is an AVS image, which can be sent to the display image module for viewing as the computation progresses. The user sees a visualization of the processors that are “alive” at that moment in the computation.

Life WT

Author :  Steve Thorpe  Life WT is an implementation of the
International AVS Center  “Life” problem. This is an algorithm
Number :  1280  where “neighbors”, or cells, propagate
Language :  multiC  or die, based on the proximity of other
Submitted :  06/02/92  living neighbors. In this case, a
Version :  1.000  living neighbor is represented by a
Last Updated :  06/02/92  processor in the massively parrallel
Ported To :  IBM Wavetracer  Wavetracer being “turned on”. The

thoughts behind the algorithm are If there are too many neighbors, a cell will die due to overcrowding. If there are too few neighbors, a cell will die due to lack of resources. The user is able to control the number of processors in the X direction, the number of processors in the Y direction, the number of generations to compute, and the number of generations to compute between outputs of an image. A oneshot parameter, “Go for it..” is supplied to begin the execution of the module. Life WT is written in multiC, a fully compatible extension of the ANSI C programming language. It was developed using a Data Transport Computer from Wavetracer, Inc.

Load Simulation

Author :  Prashant Gupta and  The Loadsim module will create a ball
Steven Smith  and stick geometry representation of
North Carolina State  molecular dynamics or Monte Carlo data.
University  The ability to interactively select
Number :  1729  positions in the data file as well as
Language :  C  individual chains is advantageous for
Submitted :  12/14/93  visualizing chain motions. Although
Version :  1.000  written for visualizing simple
Last Updated :  12/14/93  chain-like molecules it can be extended
Ported To :  Sun  to other systems.

MOSS LINES

Author :  Scott Goodyear  MOSS Polylines - Reader for MOSS
Number :  1843  format polyline vector lists. The
Language :  C  Arc/Info MOSS export files are an ASCII
Submitted :  07/16/94  format file readable by the US
Version :  1.000  Department of the Interior’s MOSS
Last Updated :  07/16/94  public domain GIS. The file contains
Ported To :  DEC HP IBM Sun  points, lines, or closed polygon loops,

and a 30-character attribute field, called the “subject” value. The Arc/Info coverages are converted into the MOSS files using the “ARCMOSS” utility in the data conversion utility subsystem. This module has a file browser, for selecting the ASCII input MOSS file, which is scanned for closed polygon areas. The polylines are converted into AVS geometry polylines for use by the AVS geometry viewer. The polyline does not have a color defined in this module, such that it can be interactively altered in the viewer, and saved with a network file.

MOSS Markers

Author :  Scott Goodyear  MOSS Markers - Reader for MOSS format
Number :  1841  markers. The Arc/Info MOSS export files
Language :  C  are an ASCII format file readable by
Submitted :  07/16/94  the US Department of the Interior’s
Version :  1.000  MOSS public domain GIS. The file
Last Updated :  07/16/94  contains points, lines, or closed
Ported To :  DEC HP Kubota IBM Sun  polygon loops, and a 30-character

attribute field, called the “subject” value. The Arc/Info coverages are converted into the MOSS files using the “ARCMOSS” utility in the data conversion utility subsystem. This module has a file browser, for selecting the ASCII input MOSS file, which is scanned for markers. The markers are converted into AVS geometry for use by the AVS geometry viewer. Each marker is shown as a sphere, with radius controlled by the dial. The marker does not have a color defined in this module, such that it can be interactively altered in the viewer, and saved with a network file.

MOSS POLYGONS

Author :  Scott Goodyear  MOSS Polygons - Reader for MOSS format
Number :  1842  polygon coverages. The Arc/Info MOSS
Language :  C  export files are an ASCII format file
Submitted :  07/16/94  readable by the US Department of the
Version :  1.000  Interior’s MOSS public domain GIS. The
Last Updated :  07/16/94  file contains points, lines, or closed
Ported To :  DEC HP IBM Sun  polygon loops, and a 30-character

attribute field, called the “subject” value. The Arc/Info coverages are converted into the MOSS files using the “ARCMOSS” utility in the data conversion utility subsystem. This module has a file browser, for selecting the ASCII input MOSS file, which is scanned for closed polygon areas. The polygons are converted into AVS geometry for use by the AVS geometry viewer. The polygon does not have a color defined in this module, such that it can be interactively altered in the viewer, and saved with a network file.

Menger sponge

Author :  Larry Gelberg  Menger sponge produces a Menger Sponge
Advanced Visual Systems  (Sierpinski Sponge) as a byte volume. A
Number :  1710  Menger Sponge (sometimes known as a
Language :  C  Sierpinski Sponge) is a recursively
Submitted :  09/30/93  defined volume. It is the 3D analog to
Version :  1.000  a Sierpinski Carpet which is can be
Last Updated :  09/30/93  poorly defined as a recursive 3x3 grid
Ported To :  HP IBM Kubota Sun  where the center square of each level

of the recursion is missing. A more formal definition can be found in Benoit Mandelbrot’s book “The Fractal Geometry of Nature“ on page 144. You can see a Sierpinski Carpet if you orthoslice the sponge and look at slice 0 (any direction). CURDLE (at the IAC) is another example of a module which produces recursively defined “fractal” data sets.

Mouse Trap

Author :  Ian Curington  The “Mouse Trap” module accesses the
Advanced Visual Systems  X-Window of the Geometry Viewer
Number :  1670  display, and changes the mouse button
Language :  C  definitions. It is intended as an AVS
Submitted :  09/29/93  application developers example, to
Version :  1.000  create custom interface controls. It
Last Updated :  09/29/93  intercepts all mouse events, and then
Ported To :  HP  selectively sends commands on to AVS

for action. At present, it ignores all except the right mouse button, which behaves as a “pick”, like the normal AVS left mouse button, while ignoring rotation and translation requests. The implementation is to make an invisible input only window, exactly covering the geometry viewer window. Events from this invisible window are detected, masking the geometry viewer window native events. Commands are then sent via the CLI mechanism. As a coroutine, this module can also take events from user interface parameters, although none are present in this version.

Mouse Trap IV

Author :  Ian Curington  The “Mouse Trap IV” module accesses the
Advanced Visual Systems  X-Window of the Image Viewer display,
Number :  1828  and changes the mouse button
Language :  C  definitions. It intercepts all mouse
Submitted :  07/15/94  events, and then selectively sends
Version :  1.000  commands on to AVS for action. This
Last Updated :  07/15/94  version ignores all normal image viewer
Ported To :  DEC HP IBM Kubota Sun  mouse controls, and causes special

behavior instead. The left button will “zoom in” the image, while the right button will “zoom out” again, each by a preset amount. It is intended as an AVS application developers example, to create custom interface controls. This module also shows how to change the mouse cursor shape while over viewer window. This is intended as a source example for further development work.

QUAL field

Author :  Marc Curry  The QUAL field module is used to
North Carolina  selectively grab data elements from a
Supercomputing Center  large output plot file and map them to
Number :  1321  the geometry of an example river. The
Language :  C  geometry can be changed (with some
Submitted :  08/30/92  effort) to match that of any river. The
Version :  1.000  codes that generate the output plot
Last Updated :  08/30/92  files used by this module are QUAL2E
Ported To :  DEC Sun HP  and QUAL2E-UNCAS which were developed

by Linfield Brown (Dept. of Civil Eng, Tufts Univ.) and Thomas Barnwell (EPA, Athens GA.).

RADM Credits

Author :  Terry Myerson  This module outputs a series of “title”
International AVS Center  labels, creating a credit page for an
(NCSC)  AVS application. This module provides
Number :  1111  an example for creating a single
Language :  C  geometry object with multiple labels
Submitted :  12/06/91  within it.
Version :  1.000
Last Updated :  12/06/91
Ported To :  Kubota DEC IBM Sun HP

READ ANY IMAGE

Author :  Terry Myerson, Steve  This module reads an image into an AVS
Thorpe, and SDSC  Network in a variety of formats. All
International AVS Center  formats which the San Diego
(NCSC) and San Diego  Supercomputing Center’s Image Tools
Supercomputer Center  support, this module supports. With
Number :  1211  this module you can read in a
Language :  C  postscript file, and edit it with the
Submitted :  05/21/92  standard tools of AVS . Any of the
Version :  3.000  following image file formats can be
Last Updated :  07/05/93  read in by this module eps Encapsulated
Ported To :  HP IBM DEC Sun  PostScript file, gif Compuserve

Graphics image file, hdf Hierarchical Data File, icon Sun Icon and Cursor file, iff Sun TAAC Image File Format, mpnt Apple Macintosh MacPaint file, pbm Portable Bit Map file, pcx ZSoft IBM PC Paintbrush file, pgm Portable Gray Map file, pic PIXAR picture file, pict Apple Macintosh QuickDraw/PICT file, pix Alias image file, pnm Portable aNy Map file, ppm Portable Pixel Map file, ps PostScript file, ras Sun Rasterfile, rgb SGI RGB image file, rla Wavefront raster image file, rle Utah Run length encoded image file, rpbm Raw Portable Bit Map file, rpgm Raw Portable Gray Map file, rpnm Raw Portable aNy Map file, rppm Raw Portable Pixel Map file, synu Synu image file, tiff Tagged image file, x Stardent AVS X image file, xbm X11 bitmap file, xwd X Window System window dump image file, and any others supported by SDSC’s Image Tools. Version 3.0 is much easier to compile than before, as it doesn’t link with the Image Tools - it calls the imconv utility directly using a system call.

READ TIMESTAMP

Author :  Kathy Pearson  READ TIMESTAMP – read the input field
MCNC Environmental  and timestamp file and output a subset
Programs  field and a subset timestamp string
Number :  1813  (suitable as recursive inputs to this
Language :  C  module once the timestamp string is
Submitted :  04/21/94  written into a file). The MCNC
Version :  1.000  Environmental Programs Group has
Last Updated :  04/21/94  created this module as part of a suite
Ported To :  DEC HP IBM Sun  of AVS demo modules, networks, and

sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

READ TIN

Author :  Scott Goodyear  NAME READ TIN - Reader for Arc/Info
Number :  1845  Triangulated Region Mesh The READ TIN
Language :  C  module accesses external ASCII files
Submitted :  07/16/94  defining a triangulated irregular
Version :  1.000  network (TIN) structure from Arc/Info.
Last Updated :  07/16/94  This is generally used for a surface
Ported To :  DEC HP Sun  representation derived from irregulary

space sample points and breakline features. The TIN data set includes topological relationships between points and their proximal triangles. Each sample point has an x,y coordinate and a surface, or z-value. These points are connected by edges to form a set of nonoverlapping triangles userd to represent the surface. TIN files are also called “irregular triangular mesh” or “irregular triangular surface model” in some contexts. “TIN” is the name of the ESRI Arc/Info application module that models, creates, displays, and exports TIN files. Within the TIN application the UNGENERATETIN command with the NET option should generate files compatable with this reader module. An optional image input port will apply a full colour texture map onto the TIN geometry, such that the geometry viewer can show the texture overlay on the terrain. The second file browser (WorldImageFile) allows a small coordinate range correction file to specify the mapping between the optional input raster texture image and the TIN coordinate system. This module requires AVS 5.01, with texture per object support.

READ WIND TIMEST

Author :  Kathy Pearson  READ WINDTIMESTAMP – read the input
MCNC Environmental  wind field and timestamp file and
Programs  output a subset field and a subset
Number :  1814  timestamp string (suitable as recursive
Language :  C  inputs to this module once the
Submitted :  04/21/94  timestamp string is written into a
Version :  1.000  file). The MCNC Environmental Programs
Last Updated :  04/21/94  Group has created this module as part
Ported To :  DEC HP IBM Sun  of a suite of AVS demo modules,

networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

RdUcdAnim

Author :  D. LE CORRE  This DATA input module animates a
CONVEX S.A  sequence of data applied to a UCD
Number :  1316  structure. The base format is exactly
Language :  C  the same as the standard UCD, except
Submitted :  08/30/92  that you can specify a 6-th value on
Version :  1.200  the first line (if you use ASCII
Last Updated :  12/30/92  format) , giving the maximum number of
Ported To :  Sun HP  time steps to consider. By default,

allocation is provided for at most 256 time steps.

ReadLat

Author :  Dominique Le Corre  This module reads an Explorer Lattice
Tethys S.A.  file and generates a field. It supports
Number :  1596  both ASCII and binary files. The file
Language :  C  name parameter only shows file names
Submitted :  03/04/93  with extensions .lat and .cmap, but you
Version :  1.000  can enter any file name. Components are
Last Updated :  03/04/93  labeled Channel 0, Channel 1, etc. and
Ported To :  DEC IBM Kubota HP  units also.

ReadRadiance

Author :  Stephen H. Price for  This data input module reads a *.pic
Advanced Visual Systems, file created with Radiance 2.1 (or
Inc.  greater) radiosity package. It outputs
Number :  1727  a Field 2D 4-vector b yte field for
Language :  C  viewing with the image viewer module.
Submitted :  12/14/93  An sample image file is included.
Version :  1.000
Last Updated :  12/14/93
Ported To :  Sun

ReadSeqAnim

Author :  Ion Barosan  The read seq input modules sends a
Eindhoven University of  sequence of files’ names, to a string
Technology, The  parameter ports. Its purpose is to
Netherlands  generate a sequence of files names, a
Number :  1584  sequence which will be used as input
Language :  C  for different data input modules. It
Submitted :  02/11/93  has four parameters, called “File
Version :  2.000  name“, ”Animation On/Off“, ”Current
Last Updated :  03/04/93  step“, and ”Nr.frames“. In a typical
Ported To :  IBM Kubota Sun  network with read seq, before you can

connect the character string to the receiving module,you must make that receiving module’s parameter port visible.

Read 2

Author :  John Sheehan  This is an example module to show how
Advanced Visual Systems  one can go from cylindrical or
Number :  1702  spherical to cartesian coordinates.
Language :  C  This has been written for specific
Submitted :  09/30/93  datasets, but it should serve the
Version :  1.000  purpose. Anyways, the trick is to
Last Updated :  09/30/93  calculate the points array properly in
Ported To :  HP IBM Kubota Sun  order to feed field2mesh, in most

cases. See also - Dipole.

Read CFD

Author :  KGT (Kubota Graphics  You can use this module for example in
Technologies) Inc.  the field of CFD. The “Read CFD” module
Number :  1882  reads a sample data. The format of a
Language :  CandFORTRAN  sample data(CFD) is discussed below in
Submitted :  09/07/95  the “Read CFD.txt” file. It is the
Version :  1.000  most populer format of CFD.
Last Updated :  09/07/95

Read DXF

Author :  John Sheehan  Read DXF is a utility module for
Advanced Visual Systems  converting 2D and 3D DXF files into an
Number :  1281  AVS Geometry Data Type. If level
Language :  C  information is available in the DXF
Submitted :  06/02/92  file, Read DXF will use this to create
Version :  1.000  separate objects. DXF is the file type
Last Updated :  06/02/92  for autoCAD output. Supported Entities
Ported To :  DEC Kubota HP  LINE POLYLINE 3DFACE BLOCK LEVEL Read

DXF can be easily modified to handle more primitives, ie. points, arcs, and circles. All of the DXF header information is overlooked. If Read DXF encounters a primitive it doesn’t understand it dumps out a geometry of what it had generated to that point, plus a message with the code number , the command to interpret, and the current layer name.

Read EJ65

Author :  Ian Curington  The EJ65 file format interface is for
Advanced Visual Systems  3D body fitted coordinate flowfield and
Number :  1680  mesh data. This module is a Reader for
Language :  FORTRAN  the EJ65 geometry and field data, in
Submitted :  09/29/93  FORTRAN. It accesses the files via
Version :  1.000  interactive browsers, and reads the
Last Updated :  09/29/93  binary files directly into the AVS
Ported To :  HP  session. The two browsers are for the

mesh and flow field files. They are combined on one output port. The module is automatic, and follow the input array dimensions, but must be compiled with an interal array size to match the largest expected field. Two types of input are supported, native floating point format of the host machine (double precision), and if the host is bigendian, then native CONVEX non-IEEE binary input is supported, using the optional mode. By default, the flow solution data is moved 1/2 a cell to the node positions. If requested, it is left at the cell centers, and nodal values are calculated by nearest cell interpolation. The interpolation mode is untested. The Freeze button us used to set new flow or mesh files, or other parameters on the module, before attempting the read. Both files are accessed, regardless of the parameter change. The EJ65 file format is in use at a number of sites, including British Aerospace, PLC, UK

Read Hologram

Author :  Ian Curington  A well-commented Reader for 2D field
Advanced Visual Systems  data, in FORTRAN, that combines ASCII
Number :  1322  file inport in fixed FORTRAN FORMAT,
Language :  FORTRAN  and does numeric computation on the
Submitted :  08/31/92  data before sending it to the output
Version :  1.000  port. It allows multiplication of a
Last Updated :  08/31/92  coefficient on one particular row,
Ported To :  HP  defined by a channel dial. This is

intended as a working example of a FORTRAN read module, to be modified by others.

Read SDRC

Author :  Steve Larkin, Ian  Read SDRC reads a file which is in the
Curington, and Geoff  SDRC I-DEAS universal format converts
Cook  it into an AVS UCD format. At present
University of Manchester  only two types of node data record are
and AVS Inc  implemented, static and transient. A
Number :  1590  static record has one set of values at
Language :  C  each node. A transient record has a set
Submitted :  02/21/93  of values at each node for different
Version :  1.000  times. Please see the International AVS
Last Updated :  02/21/93  Center’s sample dataset subdirectory
Ported To :  DEC IBM Kubota HP  for two datafiles you might want to try

with this module - crankshaft.unv and trans-therm.unv. PARAMETERS

Read SEG Y

Author :  John Tee  READ SEG Y is an data input module
AVS Inc. Consultant  which reads seismic data stored in the
Number :  1753  SEG Y format. SEG “Y” is a format
Language :  CandFORTRAN  devised by the Society of Exploration
Submitted :  12/15/93  Geophysicists (SEG), for transfer of
Version :  1.000  demultiplexed seismic data on 9-track
Last Updated :  12/15/93  tape. SEG Y was published in 1974, and
Ported To :  Kubota  superceded the SEG “Ex” format. SEG Y

defines, for each tape reel, a 3200

byte EBCDIC text header, followed by a 400 byte binary header. The data follows as a variable number of traces. Each trace has a 240 byte header, followed by a variable number of samples, each represented as a 4-byte real value.

Read Unversal FI

Author :  Hiro Kawanishi  Read Unversal File is a utility module
Digital Equipment  for conversion of I-DEAS Unversal Files
Corporation JAPAN  into AVS UCD data. This module only
Number :  1562  extracts Stress data and node-element
Language :  C  information.
Submitted :  12/30/92
Version :  1.000
Last Updated :  12/30/92
Ported To :  IBM Kubota

Sin Wav

Author :  Steve Thorpe  This little module was written while
International AVS Center  playing with the graph viewer. A simple
Number :  1600  module to spit out a 1-D field was
Language :  C  desired, and this did the trick. It
Submitted :  11/30/92  computes a sin wave over the range
Version :  1.000  1..npoints degrees, and spits out a
Last Updated :  11/30/92  “field 1-D 1-space 1-vector uniform
Ported To :  DEC Kubota HP IBM  float“ suitable for display with the

graph viewer module. This may be a useful starter module if you want to learn how to get your 1-D data into AVS.

Sonify

Author :  Brian Kaplan  The Sonify module will synthesize
Center for Innovative  sounds and can be controlled by other
Computer Applications  modules via command string input. It is
Number :  1635  most useful for data sonification where
Language :  C  a data-producing module will send
Submitted :  07/05/93  commands to Sonify based on the data
Version :  1.000  represented.
Last Updated :  07/05/93

Splash

Author :  Marc Curry  This module is a simple C subroutine
North Carolina  that will output a variety of
Supercomputing Center  surfaces/colors depending upon what 3-D
Number :  1585  mathematical function you choose it to
Language :  C  have. The basic framework for the
Submitted :  02/11/93  existing module example simply computes
Version :  1.000  the distance from the center of a
Last Updated :  02/11/93  100x100 grid (R = sqrt(r1*r1 + r2*r2))
Ported To :  DEC IBM Kubota HP  and multiplies it by a sinusoidal

factor (cos(scale * R)) where scale is a scaling parameter, used to give it the waviness (z-dimension) we want. You can make it very interesting by attaching the “animated float” module to the widget input port instead of the float dial. The example network provided does this. This is a great example to learn how to manipulate AVS fields.

Tick Marks

Author :  Jeffrey A Thingvold  The tick mark module creates sheets
National Center for  with major and minor line grids along
Supercomputing  the three primary axes. It makes
Applications  polytriangle objects, not lines, so
Number :  1338  they can have variable thickness. It
Language :  C  also has label annotation on the 3D
Submitted :  08/31/92  object.
Version :  1.000
Last Updated :  08/31/92
Ported To :  DEC Sun HP

Vector Label

Author :  Charles J Williams III  This module produces labels that are
Naval Research Lab  geometries. They can be scaled,
Number :  1563  rotated, and translated. Other
Language :  C  geometric operators work as well, such
Submitted :  12/30/92  as color, perspective, and material
Version :  1.000  properties. The source for the vector
Last Updated :  12/30/92  information for the fonts is derived
Ported To :  IBM Kubota HP  from the Hershey Fonts, stored in

comp.sources.unix, volume 4, on your nearest archive site. The Hershey Fonts were originally created by Dr. A. V. Hershey while working at the U.S. National Bureau of Standards. The format of the font data in the Internet distribution was originally created by James Hurt at Cognition, Inc.

Vernier Dial

Author :  Ian Curington  This module is a C coroutine that
Advanced Visual Systems  presents a special X-window with a
Number :  1331  large dial. The dial is similar to the
Language :  C  standard AVS dial widget, but has
Submitted :  08/31/92  higher resolution and a secondary
Version :  1.000  needle that moves much slower than the
Last Updated :  08/31/92  primary needle. The dial can work in
Ported To :  accumulator mode (around and around) or

with fixed low and high end stops. In immediate mode it sends an output float to the output port continuously while the needle is moving, otherwise only when the mouse button lifts up.

WriteSeqAnim

Author :  Ion Barosan  The WriteSeqAnim module sends a
Eindhoven University of  sequence of file names to a string
Technology, The  parameter port. Its purpose is to
Netherlands  generate a sequence of file names; the
Number :  1598  sequence will be used as input for
Language :  C  different DATA OUTPUT modules. Before
Submitted :  03/04/93  you can connect the character string to
Version :  1.000  the receiving module, you must make
Last Updated :  03/04/93  that receiving module’s parameter port
Ported To :  DEC IBM Kubota  visible. To make a parameter port

visible, call up the module’s Editor Window panel by pressing the middle or right mouse button on the module icon dimple. Next, look under the “Parameters” list to find the parameter you want to plug into. Position the mouse cursor over that parameter’s button and press any mouse button. When the Parameter’s Editor Window appears, click any mouse button over its “Port Visible“ switch. A blue-green (teal) parameter port should appear on the module icon. Connect this parameter port to the character string module icon in the usual way one connects modules.

Zoomer

Author :  Jeff Wang  This module allows users to set the
MCNC, Environmetal  coordinate of the “zoom box” and UTM
Programs Group  zone of the view area. This is one of a
Number :  1743  suite of modules we used to plot the
Language :  C  emission data for the state of North
Submitted :  12/15/93  Carolina. This group of modules allows
Version :  1.000  displaying points, lines and polygons
Last Updated :  12/15/93  over a map of the United States. The
Ported To :  Kubota Sun  module suite is stored as a group in a

compressed tar file at the

International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

animate file

Author :  Terry Myerson  anim fname is used to output a series
International AVS Center  of filenames for input into a reader
(NCSC)  module. The module inputs an integer
Number :  1135  and a filename base, and output a
Language :  C  filename in the form “$base.%3d”.
Submitted :  02/24/92  This module is very useful for a series
Version :  2.000  of files containing a time series of
Last Updated :  11/04/92  data. Bug fixes and extensions for
Ported To :  Kubota DEC IBM Sun HP  version 2.0 added by Wes Bethel, LBL.

animate floa

Author :  Wes Bethel  The animate float module automatically
Lawrence Berkeley  modifies teh floating point parameters.
Laboratory  It is used to create simple animations
Number :  1009  or to drive user simulation code. This
Language :  C  module is similar to standard animate
Submitted :  10/10/91  float module except for the following
Version :  1.000  modifications (1) The maximum number of
Last Updated :  10/10/91  steps has been modified to be minimum
Ported To :  Kubota DEC IBM Sun HP  distance detector 10000, rather than

the standard 100 (2) The “mode” choice has been removed. This module operates in a manner analogous to “animate float“’s ”one time“ mode. That is, the module executes the number of times indicated by steps. (3) A restart switch has been added. This is useful for restarting the module from step 0 ( without having to kill then reinistate the module)..

animate inte

Author :  Wes Bethel  The animate integer module
Lawrence Berkeley  automatically modifies the integer
Laboratory  parameters. It is used to create simple
Number :  1010  animations or to drive user simulation
Language :  C  code. This module is similar to the
Submitted :  10/10/91  standard animate integer module except
Version :  1.000  for the following modifications (1) The
Last Updated :  10/10/91  maximum number of steps has been
Ported To :  Kubota DEC IBM Sun HP  modified to be 10000, rather than the

standard 100. (2) The “mode” choice has been removed. This module operates in a manner analgous to “animated float”’s “one time” mode. That is, the module executes the number of times indicated by “steps”. (3) A restart switch has been added. This is useful for restarting the module from step 0 (without having to kill then reinstantiate the module).

animated boolean

Author :  William Hunter  The animated boolean module
GEC Ferranti DSL,  automatically modifies toggle
Edinborough, Scotland  parameters. This can be used to create
Number :  1183  simple animations or to drive user
Language :  C  simulation code. You plug animated
Submitted :  05/04/92  boolean into another module’s toggle
Version :  1.000  parameter port, type in a number of
Last Updated :  05/04/92  steps (default 10). When you turn off
Ported To :  DEC IBM Sun Kubota HP  sleep, it starts the animation.

animated boolean can be connected to multiple modules.

animated str

Author :  Wes Bethel  The animated string module
Lawrence Berkeley  automatically sends string parameters.
Laboratory  Unlike the “animated integer” and
Number :  1011  “animated float” modules, which compute
Language :  C  their respective output value, the
Submitted :  10/10/91  animated string module requires the
Version :  1.000  user to supply a file containing the
Last Updated :  10/10/91  strings which are sent to downstream
Ported To :  Kubota DEC IBM Sun HP  modules.

animated track

Author :  Ian Curington  Animated Track uses the “track” widget
Advanced Visual Systems  to control transformations by producing
Number :  1329  a 4x4 transformation matrix, or using
Language :  C  direct geometry object transformation
Submitted :  08/31/92  control. The track control is in
Version :  1.000  “immediate” mode, so that
Last Updated :  08/31/92  transformations are send dynamically in
Ported To :  DEC Sun HP  a stream to the rest of the network.

The module has both Field and Geom output ports, so that either one or both may be used. The module is a coroutine, so that animation sequences could be added by inserting matrix modification code to the inner loop.

application mana

Author :  John Stephen  The user selects a file which is then
Tessella Support  played as an AVS CLI script. The effect
Services  is identical to the AVS CLI command
Number :  1638  “script -play myfile.scr”.
Language :  C
Submitted :  09/24/93
Version :  1.000
Last Updated :  09/24/93
Ported To :  HP Sun

awais

Author :  Steve Thorpe  There is now a WAIS (Wide Area
International AVS Center  Information Servers) server running at
Number :  1208  the International AVS Center. WAIS
Language :  C  allows a user to ask a question to a
Submitted :  05/15/92  server, which provides a ranked list of
Version :  1.000  documents that may help answer that
Last Updated :  05/15/92  question. The user can then peruse
Ported To :  DEC Sun IBM HP  through the documents that seem useful.

It provides a convienient interface to large amounts of data. AWAIS is a WAIS client, of which there are several freely available. For example, I might be interested in JPEG files, so I might ask the server the question “Can I read or write JPEG files using AVS?“ A typical response from a WAIS client that you could type your question into, would be a document list ranked from “most hits” on down. Each list item shows that document’s rank, its score (1000 is for the “most hits”), its name (this may show only partially), and the number of lines in that document. I could then select any of these documents for viewing.

cabinet maker

Author :  Larry Gelberg  The cabinet maker module is one of a
Advanced Visual Systems  suite of four AVS modules which can be
Number :  1631  used to design simple woodworking
Language :  C  projects. For more information on the
Submitted :  07/05/93  purpose of the project, please read the
Version :  1.000  file called “Poster Session” and the
Last Updated :  07/05/93  man pages for each module. The four
Ported To :  HP IBM Kubota DEC Sun  modules are “cabinet maker” - generates

structures made of boxes, “taper” - taperer - generates structures made of tapered boxes, “extruder” - creates capped, extruded tubes from lists of X/Y pairs, and “grid gen” - creates measured grids so that accurate models can be constructed. Due to the current structure of the IAC’s directories, this suite has been split into 4 separate directories on the ftp site. However, a compressed tar archive of all 4 mods is included with each of the directories as wood mod suite.tar.Z. Also note that there are several images contributed with this module suite. These have been stored on the ftp site in the directory at avs.ncsc.org called sample data/avs data/woodworking ima ages.

ccyl pipe

Author :  Ken Flurchick  This module defines the cylindrical
North Carolina  geometry coordinates which can used by
Supercomputing Center  “volume bounds” (a supported module) to
Number :  1581  draw a hull. This module also passes
Language :  C  the geometry field to “newfilla”
Submitted :  02/11/93  (available via anonymous ftp) which in
Version :  1.000  turn fills the cylindrical geometry
Last Updated :  02/11/93  with data.
Ported To :  DEC IBM Kubota HP

choice 2

Author :  Jonathan Cox  The choice module allows the selection
Department of Computer  and output of choice parameter values.
Science, University of  The output can be linked to choice
Manchester  parameters of other modules. This
Number :  1711  allows a single choice widget to
Language :  C  control more than one choice
Submitted :  09/30/93  parameters. It can also be used to
Version :  1.000  restrict the range of choices
Last Updated :  09/30/93  available. The available choices are
Ported To :  HP  set by the choice str parameter. This

should be set to the list of the choices delimited by colons.

choice alias

Author :  John Stephen  This module navigates a menu structure
Tessella Support  which can be designed and configured by
Services  the user. Each menu level is described
Number :  1642  by a single ASCII text file. A menu
Language :  C  level contains several buttons with
Submitted :  09/24/93  user-configurable names. Each button
Version :  1.000  results in one of three basic kinds of
Last Updated :  09/24/93  action in the module - (1) load a new
Ported To :  HP Kubota Sun  menu level (2) pass a command to

downstream modules (3) go to a file to find a series of commands to be passed to down stream modules. Option 3 allows multiple commands to be played in background as the result of a single button press. The file containing the commands to be played in background is the second type of file associated with this system. Thus there are two types of file - (1) the MENU FILE to define a menu level (2) the AUTOMATIC FILE which contains a series of commands to be executed

choice selec

Author :  Ian Curington  This module takes a string input port,
Advanced Visual Systems  reconfigures a choice list parameter
Number :  1744  based on the strings and separators in
Language :  C  the input port, then outputs the
Submitted :  12/15/93  sub-string on a string output port,
Version :  1.000  along with a zero-based integer from
Last Updated :  12/15/93  the index down the choice list.
Ported To :  Kubota Sun

color cube

Author :  Larry Gelberg  COLOR CUBE generates a color cube in
Stardent Computer  either RGB (red-green-blue) or HSV
Number :  1018  (hue-saturation-value) space. The
Language :  C  RESOLUTION knob controls the number of
Submitted :  10/10/91  voxels in each direction (setting to 16
Version :  1.000  creates a 16x16x16 or 4096 voxel data
Last Updated :  10/10/91  set). The ALPHA knob sets a constant
Ported To :  Kubota DEC IBM Sun HP  alpha for all the voxels - values of 16

are pretty transparent - anything above about 32 becomes pretty opaque.

color palette

Author :  John Tee  COLOR PALETTE is used to create a
AVS Inc. Consultant  colormap suitable for the display of
Number :  1643  “label” images, ie scalar fields
Language :  C  containing pixel values in the range 1
Submitted :  09/24/93  to n where n is the number of
Version :  1.000  distinguishable categories of pixels.
Last Updated :  09/24/93  Such images arise from multispectral
Ported To :  HP IBM Sun  classification operations. COLOR

PALETTE creates a colormap in which the hue values are equally spaced, and an amount of “jitter” may be introduced to the saturation and value in order to improve the differentiation of similar hues. The saturation alternates between 1. and 1. - satn jitter. The value is successively 1., 1. - val jitter/2, 1.

- val jitter.

DATA INPUT MODULES

component menu

Author :  John Stephen  This module controls selection of part
Tessella Support  id’s based on integer part lists, in
Services  and out. The selection is by browser,
Number :  1644  and can be added, inverted, or cleared.
Language :  C  This can be used with a UCD application
Submitted :  09/24/93  to select various material groupings
Version :  1.000  for display or processing.
Last Updated :  09/24/93
Ported To :  HP IBM Sun

compress file

Author :  Lee J. Silverman  The compress/uncompress module pair is
Brown University  designed to allow datasets or images or
Number :  1588  other files to be stored and kept in
Language :  C  compressed format, but still be read
Submitted :  02/21/93  into an AVS network.
Version :  1.000
Last Updated :  02/21/93
Ported To :  DEC IBM HP

coord2geom

Author :  Mark Reed  This module reads molecular geometry
North Carolina  information from a variety of file
Supercomputing Center  formats, namely, xyz, DMol, MOPAC,
Number :  1856  MOL2, Chem36, and pdb. The user can
Language :  C  specify the units of the coordinates,
Submitted :  10/17/94  and the display type via widgits.
Version :  1.000
Last Updated :  10/17/94
Ported To :  DEC alpha HP Sun

crystal

Author :  Keith Refson  This is not *quite* a new module,
Earth Sciences  rather it is a successor to
Department, Oxford  “read shak” which now becomes
University  obsolete. CRYSTAL is a crystal
Number :  1833  structure viewer for AVS. It accepts
Language :  C  files in one of several formats -
Submitted :  07/16/94  currently SCHAKAL, SHALXL and PDB,
Version :  1.000  although I hope that some enterprising
Last Updated :  07/16/94  person may add more readers. The
Ported To :  Sun  SCHAKAL and SHELXL formats fully

support generation of complex structures from an assymetric unit with space-group symmetry operations which can be typed in almost verbatim from the “International Tables”. It has facilities for replication of unit cells, slicing the displayed structure using a miller plane, calculation of interatomic distances and variations on the ball-and-stick/space-filling display. The distribution includes a small number of example datasets, though I can supply more on request. This actually comes in 2 pieces. crystal - an AVS module for displaying crystal structure or molecular models, and cryst to geom - an offline filter for conversion of various crystal structure files to geometry.

curdle

Author :  Larry Gelberg  CURDLE creates a fractal image of a
Stardent Computer  ’curdle’ or Cantor dust. This kind of
Number :  1025  fractal is based on a square ’kernel’
Language :  C  whose size is controlled by the
Submitted :  10/10/91  KERNEL SIZE parameter. Every element
Version :  1.000  in the kernel gets recursively turned
Last Updated :  10/10/91  into a new copy of the kernel. The
Ported To :  Kubota DEC IBM Sun HP  number of levels of recursion are

controlled by the RECURSION parameter. Every element in the kernel which is ACTIVE gets evaluated further. Inactive kernel elements do not get evaluated. You activate kernel elements by dialing the CELL NUMBER to a particular cell within the kernel and hitting the ACTIVE button.

cylinders

Author :  Eric Sills  Cylinders reads a file containing
North Carolina  cylinder center point coordinates
Supercomputing Center  (x,y,z), cylinder radius, and cylinder
Number :  1298  height, one per line of the input file.
Language :  C  Cylinders builds a polyhedron object
Submitted :  08/26/92  for each cylinder (surface normals are
Version :  1.000  computed for shading).
Last Updated :  08/26/92
Ported To :  DEC Sun HP

draw arrow

Author :  anonymous  This modules creates a 3 dimensional
Number :  1888  arrow for use in the geometry viewer.
Language :  C  The arrow can then be moved and rotated
Submitted :  09/07/95  so that it points at a particular
Version :  1.000  object or interesting data value. The
Last Updated :  09/07/95  arrow can be straight, or it can have a
Ported To :  DEC ultrix DEC alpha  bend in the shaft. The arrow is created

each time that the ’Make Arrow’ button is pressed, removing any previous arrow in the process. Therefore, to create two arrows, you must use two modules.

draw polygon

Author :  anonymous  This module reads an ascii input file
Number :  1889  and creates a set of polygons in a
Language :  C  single geometry object. The module
Submitted :  09/07/95  expects the file to have an extention
Version :  1.000  of “.ply” or “.txt”. Polygons can have
Last Updated :  09/07/95  different numbers of vertices. The
Ported To :  DEC ultrix  information that must be supplied in

the ascii file is the number of polygons, the number of vertices in each polygon, and the coordinate information for the vertices. Additionally, normal information may be supplied for each polygon. In this module, it is assumed that the polygons are flat, and therefore only require 1 normal per polygon, instead of one per vertex. If the normal information is not present, the vertices are assumed to be arranged counterclockwise from the “outside” viewpoint.

endif

Author :  John Tee  Two modules, “if” and “endif” can be
AVS Inc. Consultant  used to make a logical branch in a flow
Number :  1336  network. The “if” module copies the
Language :  FORTRAN  input to one of two output ports, based
Submitted :  08/31/92  on a boolean flag, while “endif” copies
Version :  1.000  one of two input ports data to the
Last Updated :  08/31/92  output port, closing the condition.
Ported To :  DEC HP  Different processing modules can then

be placed along the two logical branches of the network. Both modules are in Fortran, and dynamically allocate new space for the output fields.

extruder

Author :  Larry Gelberg  The extruder module is one of a suite
Advanced Visual Systems  of four AVS modules which can be used
Number :  1632  to design simple woodworking projects.
Language :  C  For more information on the purpose of
Submitted :  07/05/93  the project, please read the file
Version :  1.000  called “Poster Session” and the man
Last Updated :  07/05/93  pages for each module. The four modules
Ported To :  HP IBM Kubota DEC Sun  are “cabinet maker” - generates

structures made of boxes, “taperer” - generates structures made of tapered boxes, “extruder” - creates capped, extruded tubes from lists of X/Y pairs, and “grid gen” - creates measured grids so that accurate models can be constructed. Due to the current structure of the IAC’s directories, this suite has been split into 4 separate directories on the ftp site. However, a compressed tar archive of all 4 mods is included with each of the directories as wood mod suite.tar.Z. Also note that there are several images contributed with this module suite. These have been stored on the ftp site in the directory at avs.ncsc.org called sample data/avs data/woodworking ima ages.

fast animate

Author :  Keith Refson  The fast animate module implements an
Oxford University  animation of a sequence of AVS geometry
Number :  1181  files by reading them into memory and
Language :  C  sending the geometries in rapid
Submitted :  04/30/92  succession to its output port. By
Version :  1.000  storing the geometries in memory it can
Last Updated :  04/30/92  achieve substantially better animation
Ported To :  DEC IBM Sun Kubota HP  rates than the use of a cycle object in

a script would. Furthermore the animation rate is independent of the length of the sequence until physical memory is exhausted. The module was developed for visuali- zation of molecular dynamics simulations where the geometries are “ball and stick” representations of molecular assemblies, but is not restricted to such.

field2 from Math

Author :  Oleg Perelet, Wolfram  The “field2 from Math” module allows a
Research  two-dimensional scalar field to be
Number :  1542  supplied to AVS. When the “field2 from
Language :  C  Math“ module is invoked an xterm is
Submitted :  10/14/92  started running Mathematica with the
Version :  1.000  input and output of the xterm directed
Last Updated :  10/14/92  to and from the Mathematica session
Ported To :  Sun DEC Kubota  main loop. Entering Mathematica

commands into the window will result in the appropriate Mathematica results. The session is initialised with a package of Mathematica commands which have started a MathLink communication channel to AVS. Other initialisations are made so that the Mathematica command AVSWriteField[ data] will write a three-dimensional scalar field to AVS. Please note that at the present time the IAC does NOT have Mathematica installed on a , Sun, or Kubota platforms. As a result this module was placed on the ftp site using the Makefiles provided by the author, without testing.

field3 from Math

Author :  Oleg Perelet  The “field3 from Math” module allows a
Wolfram Research  three-dimensional scalar field to be
Number :  1540  supplied to AVS. When the “field3 from
Language :  C  Math“ module is invoked an xterm is
Submitted :  10/14/92  started running Mathematica with the
Version :  1.000  input and output of the xterm directed
Last Updated :  10/14/92  to and from the Mathematica session
Ported To :  Sun DEC Kubota  main loop. Entering Mathematica

commands into the window will result in the appropriate Mathematica results. The session is initialised with a package of Mathematica commands which have started a MathLink communication channel to AVS. Other initialisations are made so that the Mathematica command AVSWriteField[ data] will write a three-dimensional scalar field to AVS. Please note that at the present time the IAC does not have Mathematica installed on a , Sun, or Kubota platform. As a result this module was placed on the ftp site using the Makefiles provided by the author, without testing.

file path and na

Author :  John Stephen  This module takes an interactive file
Tessella Support  browser widget, and outputs the actual
Services  filename and the full absolute
Number :  1655  directory path as separate string
Language :  C  output ports.
Submitted :  09/29/93
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Sun

file to field

Author :  Phil McDonald  The file to field module creates a
NOAA/ERL/Forecast  uniform field from an ASCII or binary
Systems Laboratory  file. The field parameters that are
Number :  1267  normally contained in the header of a
Language :  C  field file are specified interactively
Submitted :  05/22/92  through the file to field parameters.
Version :  1.100
Last Updated :  04/21/93
Ported To :  HP Sun Kubota

flight path

Author :  Ian Curington  The Flight Path Module takes a scatter
Advanced Visual Systems  field position list and animates the
Number :  1656  camera path along the trajectory.
Language :  C  Derived from camera.c in
Submitted :  09/29/93  /usr/avs/examples. The scatter path
Version :  1.000  field is in the form of “field 1D
Last Updated :  09/29/93  3-space irregular float“ where the
Ported To :  HP IBM Kubota Sun  coordinate information in the field

specifies the path. The data values are ignored. It moves %top, not the camera, so multiple camera views can be set up to watch the path. The file descriptor or read field modules can be used to get the path from an external file. Animated Integer can be used to control time, for a moving sequence.

fractal 3d f

Author :  Anonymous  Produce mid-point displacement fractal
Stardent Computer  in two or three dimensions. Controls
Number :  1036  are max width, power, and reseed value.
Language :  C
Submitted :  10/10/91
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

fractal field

Author :  David E. Harris  This module generates a 2D scalar byte
Georgia Tech Research  field whose elements are statistically
Institute  self-similar (fractal) in distribution
Number :  1205  without regard to scale. (I.e. size.
Language :  C  Not to be confused with the intensity
Submitted :  05/11/92  scale value parameter.) A gaussian
Version :  1.000  distributed random field is first
Last Updated :  05/11/92  created and then transformed to the
Ported To :  Kubota  frequency domain by FFT. The result is

then filtered to agree with the fractal power law and inverse transformed. The second ouput port provides access to the gaussian random field used to generate the fractal field.

gaussian field

Author :  David E. Harris  This module generates a 1-, 2-, or 3-D
Georgia Tech Research  fields of arbitrary size whose elements
Institute  are gaussian-distributed random
Number :  1204  floating point values.
Language :  C
Submitted :  05/11/92
Version :  1.000
Last Updated :  05/11/92
Ported To :  Kubota

generate contine

Author :  Ed Bender  This module is useful for decorating
Convex Computer  globes that represent the Earth. Given
Corporation  an input radius (Earth radius) a
Number :  1594  geometry is generated consisting of
Language :  C  multiple polylines that show the
Submitted :  03/04/93  outlines of the continents and islands
Version :  3.000  of the earth. Optionally, a single
Last Updated :  03/09/93  sphere is generated also at a slightly
Ported To :  Kubota IBM DEC  smaller radius than the continental

outlines. Version 2.0 has fixes by NCSC’s Will Ivey for DEC version. Version 3.0 has fixes by ORNL’s Ray Flanery for Kubota version.

generate gri

Author :  Larry Gelberg  Generate grid creates a colored image
Stardent Computer  of a grid of one color superimposed
Number :  1037  over a background color. You can
Language :  C  control the colors of the grid with the
Submitted :  10/10/91  FOREGROUND HUE, SAT(uration), and VALUE
Version :  1.000  dials, the color of the background with
Last Updated :  10/10/91  BACKGROUND HUE, SAT(uration), and VALUE
Ported To :  Kubota DEC Sun HP  dials, the thickness of the grid lines

with the GRID WIDTH dial, the size of the squares between the grid liines with the GRID SPACING dial, and the resolution of the output image with the RESOLUTION typeins.

genloadmd

Author :  Ken Flurchick and Mark  This module is designed to read in
Reed  position and velocity information from
North Carolina  the file specified as “Input File”, for
Supercomputing Center  a group of point particles in three
Number :  1603  dimensions. The program represents each
Language :  C  point with a sphere at the position
Submitted :  04/09/93  specifed and color coded by the
Version :  1.000  magnitude of the velocity. The colors
Last Updated :  04/09/93  0-255 are specified in the colors.h
Ported To :  DEC IBM Kubota  file and range from blue (slowest) to

red (fastest). Initially the module will compute the minimum and maximum speeds from the data set read in and echo these back to the parameter widgets. The user can change the size of the spheres and the min and max values used to determine the colors of the sphere. Note that if the minimum speed is greater than or equal to the maximum speed then a message is echoed back to standard output (the screen) and no action is taken. Sphere color is determined by a linear interpolation between the minimum and maximum speeds of the particle velocity magnitude.If the velocity magnitude of a particular sphere exceeds the maximum speed then it is displayed as white while those less than the minimum are displayed as black (note - change the background color if this also is black).

geom from Math

Author :  Oleg Perelet  The “geom from Math” module allows
Wolfram Research  three-dimensional graphics commands in
Number :  1541  Mathematica to supply AVS with an input
Language :  C  geometry. When the “geom from Math”
Submitted :  10/14/92  module is invoked an xterm is started
Version :  1.000  running Mathematica with the input and
Last Updated :  10/14/92  output of the xterm directed to and
Ported To :  Sun DEC Kubota  from the Mathematica session main loop.

Entering Mathematica commands into the window will result in the appropriate Mathematica results. The session is initialised with a package of Mathematica code which starts a MathLink communication channel to AVS. Other initialisations are made so the typical Mathematica three dimensional graphics commands such as Plot3D or ParametricPlot3D send their results into AVS rather than rendering them in the more typical Mathematica ways. Please note that at the present time the IAC does not have Mathematica installed on a , Sun, or Kubota platform. As a result this module was placed on the ftp site using the Makefiles provided by the author, without testing.

geometry sphere

Author :  Marc Curry  This geometry sphere module
North Carolina  demonstrates how to convert a set of
Supercomputing Center  vertices into a geomtry format for
Number :  1320  viewing within AVS. Also, some simple
Language :  C  widgets are provided to take care of
Submitted :  08/30/92  (what were) program parameters. The
Version :  2.000  sphere is actually a triangle mesh
Last Updated :  10/02/92  approximating a sphere by recursive
Ported To :  Sun  subdivision into more and more

triangles. The first approximation is an octahedron, with each level of refinement increasing the number of triangles by a factor of 4.

graduate32

Author :  David Haith  Generates shaded 2D images for creating
Power-Media  synthetic backgrounds. Parameters
Number :  1041  control degree of ramp, starting and
Language :  C  ending values. Gradation is top to
Submitted :  10/10/91  bottom, with horizontal variation
Version :  1.000  constant. RGB color specs are given for
Last Updated :  10/10/91  both top and bottom. A nice noise
Ported To :  Kubota DEC IBM Sun HP  source is available for mixing with the

ramp. Output can be seen with “display image“.

grid generator

Author :  Larry Gelberg  The grid generator module is one of a
Advanced Visual Systems  suite of four AVS modules which can be
Number :  1633  used to design simple woodworking
Language :  C  projects. For more information on the
Submitted :  07/05/93  purpose of the project, please read the
Version :  1.000  file called “Poster Session” and the
Last Updated :  07/05/93  man pages for each module. The four
Ported To :  HP IBM Kubota DEC Sun  modules are “cabinet maker” - generates

structures made of boxes, “taperer” - generates structures made of tapered boxes, “extruder” - creates capped, extruded tubes from lists of X/Y pairs, and “grid gen” - creates measured grids so that accurate models can be constructed. Due to the current structure of the IAC’s directories, this suite has been split into 4 separate directories on the ftp site. However, a compressed tar archive of all 4 mods is included with each of the directories as wood mod suite.tar.Z. Also note that there are several images contributed with this module suite. These have been stored on the ftp site in the directory at avs.ncsc.org called sample data/avs data/woodworking ima ages.

harwell colo

Author :  Anonymous  Harwell is a special color map
Stardent Computer  generator. It provides a standard
Number :  1042  fringe color set, based on a transfer
Language :  C  equation. It is gamma corrected, and
Submitted :  10/10/91  has essentially three regions, with
Version :  2.000  linear interpolation between the
Last Updated :  09/30/93  regions. The regions are blue –¿ green
Ported To :  Kubota DEC IBM Sun HP  –¿ yellow –¿ red The cutoff points

are controlled by the three float dial parameters. The gamma is also adjustable. The default values are designed for the “standard fringe”. Version 2.0 update created by Ian Curington of Advanced Visual Systems, Incorporated.

if 2

Author :  John Tee  Two modules, “if” and “endif” can be
AVS Inc. Consultant  used to make a logical branch in a flow
Number :  1337  network. The “if” module copies the
Language :  FORTRAN  input to one of two output ports, based
Submitted :  08/31/92  on a boolean flag, while “endif” copies
Version :  1.000  one of two input ports data to the
Last Updated :  08/31/92  output port, closing the condition.
Ported To :  HP  Different processing modules can then

be placed along the two logical

branches of the network. Both modules are in Fortran, and dynamically allocate new space for the output fields.

image snap

Author :  Ian Curington  This module uses the CLI to cause image
Advanced Visual Systems  zoom and pan operations on one image to
Number :  1661  be echoed to a second image. A “snap”
Language :  C  button is used to cause the
Submitted :  09/29/93  transformations. Two string type-in
Version :  1.000  controls specifiy the original and
Last Updated :  09/29/93  target image names in the image viewer.
Ported To :  HP IBM Kubota Sun

int positiv

Author :  Richard Albrecht  This is a short module to adjust the
Max-Planck-Institute for  ’tracer’ to ’downsize’. It works in
Biochemistry  connection with ’downsize points’. The
Number :  1725  first parameter is a replacement of
Language :  C  ’integer’. Only positive values are
Submitted :  12/02/93  allowed. The second parameter is
Version :  1.000  multiplied with the first to give an
Last Updated :  12/02/93  adjustment of the alpha value in
Ported To :  Kubota Sun  ’tracer’. Now it is possible to use

’downsize’ with the ’tracer’. The size and the opacity of the image is independent from the downsize value.

integer list

Author :  John Stephen  Converts a list of integers within a
Tessella Support  string (blank-delimited) to a field of
Services  integers (a 1D array).
Number :  1663
Language :  C
Submitted :  09/29/93
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Kubota Sun

integer step

Author :  Young S. Ham  The integer step module supplies an
Naval Research  integer value to one or more receiving
Laboratory  modules from an intial integer. The
Number :  1771  current integer is incremented by
Language :  C  step size by means of a mouse click on
Submitted :  02/09/94  a button. This module is very useful
Version :  1.000  when viewing images or to fire an input
Last Updated :  02/09/94  value in order is desired. One does not
Ported To :  HP Kubota Sun  have to keep typing an integer or play

with the dial widget for a new input value.

jigsaw

Author :  Ian Curington  Jigsaw creates a simple geometric
Advanced Visual Systems  object of extruded shapes, with input
Number :  1665  file control over contour of shapes. An
Language :  C  AVS field controls the height, color,
Submitted :  09/29/93  or visibility of each block. This
Version :  1.000  module is similar to “Read Blokjes” at
Last Updated :  09/29/93  the IAC. Some degree of concave outline
Ported To :  HP IBM Kubota Sun  structure is allowed. Each piece has a

unique name, and can be manipulated independenty in the geometry viewer, if lock mode is not selected. The color range is defined externally in the colormap data structure.

lathe

Author :  Ian Currington  This module is a C co-routine, that
Stardent Computer  includes an X-window application. It
Number :  1048  has an interactive pallet window, for
Language :  C  defining the profile of a “lathe”
Submitted :  10/10/91  object. This profile is turned into a
Version :  1.000  3D geometric object by surface of
Last Updated :  10/10/91  revolution.
Ported To :  Sun HP

lines to geom

Author :  Tim O. Robinson  This module reads an ascii input file
UC Berkeley  and creates a set of 3D line segments
Number :  1272  in a single geometry object. The module
Language :  C  expects the file to have an extention
Submitted :  05/27/92  of “.xyz”. Each line segment is assumed
Version :  1.000  to have the same number of vertexes.
Last Updated :  05/27/92  Duplicate vertexes can be used to
Ported To :  DEC IBM Sun Kubota HP  overcome this but they are not ignored

in this version.

load md

Author :  David Bock  AVS subroutine module source to read
North Carolina  into memory a series of time steps from
Supercomputing Center  a specified input file and provide the
Number :  1109  user with the ability to select an
Language :  C  individual time step or cycle through a
Submitted :  12/05/91  series of selected time steps. Module
Version :  1.000  has been used to visualized the
Last Updated :  12/05/91  dynamics of a given set of molecules
Ported To :  Kubota DEC IBM Sun HP  over a specified period of time. Module

simply loads the user-specified time steps of molecular coordinates and generates a sphere for every molecule at the stored coordinates.

load mdn

Author :  Ken Flurchick  This is an enhancement of the load md
North Carolina  module already available. This module
Supercomputing Center  is designed to read in one time step of
Number :  1583  data from a series of input files and
Language :  C  provide the user with the ability to
Submitted :  02/11/93  select an individual time step or cycle
Version :  1.000  through a series of selected time
Last Updated :  02/11/93  steps. Module has been used to
Ported To :  DEC IBM Kubota HP  visualized the dynamics of a given set

of molecules over a specified period of time. Module simply loads the user-specified time steps of molecular coordinates and generates a sphere for every molecule at the stored coordinates. This module has the option of specifying the data file through animated filename, for multiple files, or with a file browser widget for a single file. The default is to use the animated filename input port, but if this value is set to $NULL then the module will use the file browser.

macro menu

Author :  John Stephen, Tessella  This module navigates a menu structure
Support Services  which can be designed and configured by
and Ian Curington,  the user. Each menu level is described
Advanced Visual Systems,  by a single ASCII text file. A menu
Inc.  level contains several buttons with
Number :  1719  user-configurable names. Each button
Language :  C  results in one of three basic kinds of
Submitted :  09/30/93  action in the module - load a new menu
Version :  1.000  level, pass a command to downstream
Last Updated :  09/30/93  modules, or go to a file to find a
Ported To :  HP Kubota Sun  series of commands to be passed to down

stream modules. Option 3 allows multiple commands to be played in background as the result of a single button press. The file containing the commands to be played in background is the second type of file associated with this system. Thus there are two types of file - the MENU FILE to define a menu level, and the AUTOMATIC FILE which contains a series of commands to be executed.

mandelbrot

Author :  Larry Gelberg  Generates the mandelbrot set, at a
Stardent Computer  specific location, for a certain number
Number :  1108  of iterations, at a certain size, all
Language :  C  determined by paramter widgets.
Submitted :  11/27/91
Version :  1.000
Last Updated :  11/27/91
Ported To :  Kubota DEC IBM Sun HP

mask generator

Author :  David E. Harris  This module generates a field that is
Georgia Tech Research  intended to be interpreted as a binary
Institute  mask for filtering FFTs, a gray scale
Number :  1201  image, or a structuring element for
Language :  C  various morphological operations. The
Submitted :  05/08/92  field can be any size from 2x2 pixels
Version :  1.000  on up and an ellipse of any size major
Last Updated :  05/08/92  and minor axes from one pixel on up.
Ported To :  DEC Kubota  The background can be either black or

white and the intensity value of the ellipse can be any gray level. The position of the ellipse within the field is arbitrary and can intersect the edges of the field or be outside the field entirely.

menu example

Author :  Larry Gelberg  This is an example of how you can use
Stardent Computer  the layout feature of AVS to implement
Number :  1053  multiple level menus. When the
Language :  C  selection “level 1” is madea menu
Submitted :  10/10/91  showing the choices level2a1, level2a2,
Version :  1.000  level2a3 is displayed. When the
Last Updated :  10/10/91  selection “level 2” is made, the
Ported To :  Kubota DEC IBM Sun HP  choices level2b1, level2b2, level2b3

are displayed. There are several ways in which this is accomplished in the source code (1) the call AVSadd parameter prop() is used to set up the layout (2) the call AVScommand is used to issue CLIO calls from within the module body.

microscope

Author :  Hennekens Didier  This avs subroutine transforms a
University of  picture of atomic force microscope in
Mons-Hainaut (Belgium)  the “scalar uniform field” structure as
Number :  1283  defined by AVS. This picture consists
Language :  C  of a set of heights on a square uniform
Submitted :  06/07/92  grid. The informations on this picture
Version :  1.000  are contained in a ascii file. The
Last Updated :  06/07/92  expected ascii format is very simple.
Ported To :  DEC IBM HP  It consist of a header of three lines

(the file name and two experimental parameters of the atomic force microscope - the z scale and the z dsclale) followed by data.

mk vel2

Author :  Ian Curington  This module generates a synthetic 3D
Advanced Visual Systems  field, with random perturbations in
Number :  1716  both coordinates and data values. The
Language :  C  3D scalar field is the square of the
Submitted :  09/30/93  distance from a point on the center of
Version :  1.000  the bottom plane. An isosurface of this
Last Updated :  09/30/93  is approximately a parabola. This was
Ported To :  HP IBM Kubota Sun  used in a demonstration of seismic

velocity modelling.

mk vel

Author :  Ian Curington  This module generates a synthetic 3D
Advanced Visual Systems  field, with random perturbations in
Number :  1715  both coordinates and data values. The
Language :  C  3D scalar field is the square of the
Submitted :  09/30/93  distance from a point on the center of
Version :  1.000  the bottom plane. An isosurface of this
Last Updated :  09/30/93  is approximately a parabola. This was
Ported To :  HP IBM Kubota Sun  used in a demonstration of seismic

velocity modelling.

multiformat read

Author :  A. Jacobson  This module reads an image from disk
UCLA Dept. Pharmacology into AVS in any of a variety of
/ Crump Institute  formats. Some of the formats which this
Number :  1849  module supports are from San Diego
Language :  C  Supercomputing Center’s image tools
Submitted :  10/12/94  package, others are supported from John
Version :  1.000  Cristy’s ImageMagick package, still
Last Updated :  10/12/94  others require tools from GNU’s
Ported To :  DEC HP IBM Sun  Ghostscript, and the NetPBM package.

Depending on how the module is configured, it can read and about 30 to 40 image formats, including , AVS X, Alias, Apple Macintosh MacPaint, Apple Macintosh QuickDraw/PICT, CCIR 601, Compuserve Graphics, Flexible Image Transport System (monochrome only), Hierarchical Data File, JFIF Rev1.02, Khoros VIFF, Magick, Mark VanDeWettering’s MTV ray tracer image format, Microsoft Windows bitmap, Microsoft Windows cursor, Microsoft Windows icon, PIXAR picture, Photo CD, PostScript, Radiance, Radiance, SDSC Synu, SGI RGB, Sun Icon and Cursor, Sun Rasterfile, Sun TAAC Image, Tagged, Truevision Targa, Utah Run length encoded, Wavefront raster, X Window System window dump, X11 bitmap, X11 pixmap, ZSoft IBM PC Paintbrush, binary or ASCII PBM Portable Bit Map, binary or ASCII PBM Portable Gray Map, binary or ASCII PBM Portable Pixel Map, binary or ASCII PBM Portable aNy Map. This module contains a superset of the formats in the READ ANY IMAGE module, with several new formats added and improved support for several of the original file types. Several sample images (gif, pict, ppm, ps, rgb, rle, tiff, xwd, and yuv) are included with this module for your testing purposes. Please note that this module calls several executable files, including imconv, convert, and gzip. Platform specific binaries for several of these programs are available from the IAC ftp site with this module. See the .txt file of this module or the file multiio.h for further details on the sources for this module’s tools. The tools are stored at the IAC in one large compressed tar file, which is 11MB and is stored at avs.ncsc.org in /misc tools/multiio tools/multiio to ools.tar.Z. NOTE - THIS MODULE IS AWESOME ... - Steve

59

DATA INPUT MODULES

named toggle

Author :  John Stephen  This creates a “toggle” parameter
Tessella Support  button, whose value is available as a
Services  toggle output port, for remote control
Number :  1672  of other modules with toggle
Language :  C  parameters. The name on the button is
Submitted :  09/29/93  controlled dynamically from a second
Version :  1.000  string input parameter.
Last Updated :  09/29/93
Ported To :  HP IBM Kubota Sun

newfilla

Author :  Ken Flurchick  Inputs a field from “ccyl pipe”
North Carolina  (available via anonymous ftp) and fills
Supercomputing Center  a cylindrical geometry. This module
Number :  1582  will read each time step of data from a
Language :  C  separate file using “animate filename”
Submitted :  02/11/93
Version :  1.000
Last Updated :  02/11/93
Ported To :  DEC IBM Kubota HP

noise source

Author :  David Haith  Generates random 2D fields for creating
Power-Media  synthetic backgrounds, or for mixing
Number :  1105  with other images for various effects.
Language :  C  Parameters control size of image, and
Submitted :  11/26/91  character of random field.
Version :  1.000
Last Updated :  11/26/91
Ported To :  Kubota

pdb to geom

Author :  Advanced Visual Systems, The pdb to geom module reads the
Inc.  description of a molecule from a file
Number :  1551  in the Brookhaven Protein Data Bank
Language :  C  (PDB) data format. Typically, such
Submitted :  11/04/92  files have a filename suffix. The
Version :  1.000  output is an AVS description of the
Last Updated :  11/04/92  molecule.
Ported To :  Kubota HP

phoenics int

Author :  Michael J. Rangitsch  This module converts a phi.dat data
Number :  1175  file produced by PHOENICS into an AVS
Language :  FORTRAN  field. The type of the field is
Submitted :  04/16/92  computed by the header information in
Version :  1.000  the phi.dat file. Cartesian,
Last Updated :  04/16/92  cylindrical and BFC cases are all
Ported To :  Kubota IBM HP  supported. Null velocity fields are

produced for 1 and 2 dimensional cases to fill out three velocity components needed for modules such as the particle advector.

pick center

Author :  Ian Curington  This module allows the user to pick on
Advanced Visual Systems  anything in the GeometryViewer scene
Number :  1752  that is not already registered for
Language :  C  upstream picking, and will set the
Submitted :  12/15/93  current rotate / zoom center to be this
Version :  1.000  picked point. It is most useful if you
Last Updated :  12/15/93  see a large complexe scene, pick on the
Ported To :  Kubota Sun  piece you want to zoom into, then zoom

into it. Without this module, if it was on the edge of the object it would disappear from view as soon as you zoomed in.

place camera

Author :  Jeff Vroom  This module is a prototype of a camera
Advanced Visual Systems  placement system. Two geometry viewers
Number :  1676  are used, where the first has the
Language :  C  object list plus an iconic
Submitted :  09/29/93  representation of the camera, with view
Version :  1.000  frustrum shown in wire frame. The
Last Updated :  09/29/93  second view is the scene that camera
Ported To :  HP IBM Kubota Sun  “sees” as placed in the first scene.

The module shows how to manipulate and control camera and scene transformation matrices.

plate

Author :  Jeff Vroom  This is a simulation of a vibrating
Advanced Visual Systems  tympanic membrane written as a
Number :  1063  co-routine. The coefficients A,B,C,D,
Language :  C  and E relate to the various harmonic
Submitted :  10/10/91  strenths. The “sleep” button turns the
Version :  1.000  simulation on and off.
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

plato super

Author :  Larry Gelberg  This module generates GEOM versions of
Stardent Computer  the five platonic solids. It also
Number :  1080  allows for both primary and secondary
Language :  C  stellations of the faces and random
Submitted :  10/10/91  coloring of the vertecies. The
Version :  1.000  ’Stellate 1’ knob pokes the centroid of
Last Updated :  10/10/91  each face in or out. The ’Stellate 2’
Ported To :  Kubota DEC IBM Sun HP  knob then does the same thing to each

of the resulting triangles. The coloring selections provide either pre-chosen coloring schemes or ither two or three random colors per triangle. Hitting these buttons again and again will give you continuously different colorings.

plot xyz

Author :  Wes Bethel  This module reads an ASCII file
Lawrence Berkeley  consisting of lists of x,y,z points and
Laboratory  constructs a polyline geometry object
Number :  1064  using the points in the file as
Language :  C  vertices.
Submitted :  10/10/91
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

plot xyz col

Author :  Wes Bethel  This module reads an ASCII file
Lawrence Berkeley  consisting of lists of x,y,z points and
Laboratory  constructs a polyline geometry object
Number :  1065  using the points in the file as
Language :  C  vertices. Additionally, color
Submitted :  10/10/91  information is read in on a per-vertex
Version :  1.000  basis, and the line segments are color
Last Updated :  10/10/91  coded as a function of the color index
Ported To :  Sun HP  specified in the user file and the

input colormap.

r tif ra

Author :  Richard Albrecht  r tif ra reads 24 bit, 8 bit grey and
Max-Planck-Institute for  8 bit colormap TIFF-files. This is a
Biochemistry  special reader for uncompressed
Number :  1852  TIFF-files. It converts 24 bit, 8 bit
Language :  C  grey, and 8 bit colormap TIFF-images to
Submitted :  10/17/94  an AVS image. The output is in any case
Version :  1.000  a normal AVS image. Opacity of the
Last Updated :  10/17/94  image is always set to 255, whereas the
Ported To :  DEC alpha HP IBM Sun  red, green and blue channels are copied

directly (24 bit) or set to the values of the colormap. The Colormap itself (if is) is not copied. A short report is copied to ’stdout’ with the toggle.

r tifvol ra

Author :  Richard Albrecht  r tifvol ra reads a serie of
Max-Planck-Institute for  TIFF-files of the form *.nnn This is a
Biochemistry  special reader for uncompressed
Number :  1853  TIFF-files. It reads only 8 bit grey
Language :  C  TIFF-images. A Colormap is ignored.
Submitted :  10/17/94  With a modification of the normal
Version :  1.000  AVS-field descriptor ’*.fld’ a complete
Last Updated :  10/17/94  image serie is read in. The filenames
Ported To :  DEC alpha HP IBM Sun  must be in MS-DOS style ( cccccccc.nnn

). The 3 characters after the dot are the file number ( *.000, *.001 ... ). The output is an AVS-field.

readFLOW3D

Author :  Michael Rangitsch  The read flow3d module ascesses the
Number :  1174  dump files via a browser. Prior to
Language :  C  selecting the filename, use the
Submitted :  04/16/92  parameter toggles to identify which
Version :  1.000  turbulence model, which combustion
Last Updated :  04/16/92  model, the number of scalar species
Ported To :  Kubota HP  were used in the model and whether

density or viscosity are stored, if the convection coefficients are desired or if the problem was in cylindrical coordinates. The reader has no way of knowing from the dump file alone.

read 16 bit

Author :  Larry Gelberg  READ 16 BIT IMAGE reads a 16 bit
Stardent Computer  integer binary image (2D collection of
Number :  1103  bytes) off of disk into an AVS integer
Language :  C  (32 bit) field. It does this by reading
Submitted :  11/22/91  the 16 bit integers into a “short”
Version :  1.000  buffer and then casting them into
Last Updated :  11/22/91  32-bit integers. If you do not know
Ported To :  Kubota DEC IBM Sun HP  what the resolution of the images are,

you can use the WIDTH and HEIGHT controllers like “vertical” and “horizontal” hold to experiment with until the image looks like you expoect.

read ACR NEMA

Author :  Andrea Marchetti  The read ACR-NEMA module reads ACR-NEMA
Istituto di Elaborazione  SPI data files and converts them into
dell’Informazione del  AVS field format. SPI is a standard
CNR  based on the old ACR/NEMA standard used
Number :  1885  by Siemens and Philips. SPI make use of
Language :  C  ACR/NEMA data elements nd groups. SPI
Submitted :  09/07/95  store the integer in little endian
Version :  1.000  format (Least Significant before Most
Last Updated :  09/07/95  Significant).
Ported To :  DEC ultrix DEC alpha

INTERNATIONAL AVS CENTER MODULE CATALOG

read AIFF

Author :  Charles J Williams III  This module is meant to be a companion
Naval Research Lab  to the Write MooV module. It operates
Number :  1543  rather simply, reading in an AIFF file
Language :  C  (this is a rather Macintosh format,
Submitted :  10/14/92  although many programs can translate
Version :  1.000  understand AIFF) and generating a 1D
Last Updated :  10/14/92  field.
Ported To :  Kubota

read Dore im

Author :  Anonymous  Reads a raster Dore image file output.

Kubota Pacific Computer

Inc

Number :  1153
Language :  C
Submitted :  03/12/92
Version :  1.000
Last Updated :  03/12/92
Ported To :  Kubota DEC IBM Sun HP

read F3D

Author :  Michael Rangitsch  This module reads an ascii formatted
Number :  1537  multiblock FLOW3D solution file (from
Language :  C  Harwell FLOW3D version 3) and stores
Submitted :  10/03/92  the cfd data as a ucd data set. The
Version :  5.000  dummy nodes stored at the edges of each
Last Updated :  07/15/94  computational block are removed by the
Ported To :  DEC HP Sun  module. The n-vector data at each

computational cell is stored as cell data, to use most of the ucd modules, this must be converted to nodal data. The cells of the ucd structure are all hexahedra. See the sample data directory on the International AVS Center’s anonymous ftp site for two sample data sets to read with this module. Version 2.000 has several enhancements and debugging features added. Version 3.000 adds geometry representation of 2 dimensional patches. There are 7 types of 2D patches defined by FLOW3D. This code probably doesn’t work for 3d patches. Version 4.0 handles unanticipated data sets, and processes cylindrical data sets and single block data sets, as well as several other enhancements. Version 5.0 handles unformatted as well as formatted data.

read FLOW3D

Author :  Anonymous  The read flow3d module is used for
Stardent Computer  reading/converting CFD data and Mesh
Number :  1035  files from the Harwell “FLOW3d” code
Language :  C  into AVS field data types. The
Submitted :  10/10/91  read flow3d module ascesses the dump
Version :  1.000  files for both turbulence and
Last Updated :  10/10/91  temperature information, separately.
Ported To :  Sun

read GOCAD

Author :  Ian Curington  This is a data import reader module for
Advanced Visual Systems  the GOCAD (ASCII) file format. The
Number :  1844  GOCAD application models, edits, and
Language :  C  describes 3D geologic structures,
Submitted :  07/16/94  containing horizon surfaces, faults,
Version :  1.000  and wells. The GOCAD application is
Last Updated :  07/16/94  authored by a group in Nice, France.
Ported To :  DEC HP Sun  This module reads the contents of one

GOCAD file, and converts this to the AVS geometry data type, which can be sent to the Geometry Viewer for interactive display. Axis specific scaling is allowed by using the floating point parameters, supplied with dial controls by default. The module ignores all other lines from the GOCAD file except for vertices and trianges. It looks at the first character of every line. If the the line starts with “V”, then it reads the id, and the position x, y, z. If the line starts with “T”, it reads three index numbers for which set of verts to tie together as triangles.

read HDF SDS

Author :  Bill Sherman  The “read HDF SDS” module reads
National Center for  Scientific Data Sets (SDS) from a HDF
Supercomputing  files, outputting up to four AVS field
Applications  data sets. HDF is the Hierarchical Data
Number :  1179  Format developed at the National Center
Language :  C  for Supercomputing Applications (NCSA).
Submitted :  04/22/92  The SDS is only one of several types of
Version :  1.000  data that may be contained in an HDF
Last Updated :  04/22/92  file, but it is the only type
Ported To :  DEC Sun IBM  recognized by this module.

read Leica confo

Author :  A. Jacobson  Read Leica Confocal module, designed to
UCLA Dept. Pharmacology read in output from the Leica CLSM in
/ Crump Institute  it’s native file format (not TIFF). The
Number :  1855  module is looking for a
Language :  C  imagename.info file which it parses
Submitted :  10/17/94  for the image plane numbers and plane
Version :  1.000  sizes. It uses these values to open
Last Updated :  10/17/94  each image file in order. The contents
Ported To :  HP Sun  are deposited in a 3D scalar byte field

at the output port. Lots of error

checking is done to make sure the .info file is not corrupted, and the data files read correctly.

read abekas

Author :  Ian Currington  This module provides I/O capability to
Stardent Computer  iamges in the native machine format
Number :  1002  used by the Abekas A60 digital video
Language :  C  disk recording system. The image file
Submitted :  10/10/91  specification is for the PAL ( 50
Version :  1.000  Hz,625 line ) version, but can easily
Last Updated :  10/10/91  be modified to support the NTSC
Ported To :  Kubota DEC IBM Sun HP  version, or both. The Abekas A60 holds

750 frames, or 30 seconds of live broadcast standard video imagery, in the digital storage format specified by the CCIR-601 specification. Although the disk format is Y-uv, the disk has firmware processing to encode/decode to 3-byte Red-Green-Blue images on-the-fly during the tcp/ip file transfers.

read blokjes

Author :  Ian Curington  Read Blokjes is a module that imports
Advanced Visual Systems  blocks, or bricks, and creates
Number :  1332  geometric objects appropriate for the
Language :  C  Geometry Viewer. The format is from a
Submitted :  08/31/92  CFD package called FLOWTHERM from
Version :  1.000  Flowmerics Ltd., Kingston, Surrey,
Last Updated :  08/31/92  England, KT2 5AA. This module is only
Ported To :  DEC Sun HP  one part of a more extensive interface

system. As each block is separate, the property on each one can be edited, changing colour, transparency, rendering mode, etc. Each block has a name, type, and start and end corner points.

read compressed

Author :  Wes Bethel  The “read compressed image” module
Lawrence Berkeley  provides the same functionality as the
Laboratory  AVS-supplied “read image” module, with
Number :  1128  the added twist that the image is in a
Language :  C  “compressed” format. An AVS image may
Submitted :  03/15/92  be compressed using the UNIX(tm)
Version :  2.000  compress utility, or by using the
Last Updated :  05/15/92  sibling module “write compressed
Ported To :  IBM Sun Kubota HP  image.“ Compression ratios of 1 to 10

are easily achieved using this

compression scheme.

read dem

Author :  anonymous  This AVS module reads in an ARC/Info
Number :  1887  Digital Elevation Model (DEM) into AVS.
Language :  C
Submitted :  09/07/95
Version :  1.000
Last Updated :  09/07/95
Ported To :  DEC ultrix DEC alpha

DATA INPUT MODULES

read dyna3d

Author :  David Parker  DYNA3D is a finite element analysis
Stardentand Livermore  program developed by This module
Software Technology  interprets the ASCII input file for
Number :  1100  this program and generates GEOM
Language :  CandFORTRAN  representations of the desired aspects
Submitted :  11/22/91  of the data sets.
Version :  1.000
Last Updated :  11/22/91
Ported To :  IBM

read eps image

Author :  Terry Myerson  This module reads an eps image in the
International AVS Center  eps format into an AVS Network.
(NCSC)  Encapsulated PostScript color,
Number :  1213  grayscale, and black-and- white image
Language :  C  files are used by a variety of
Submitted :  05/21/92  PostScript tools in order to include
Version :  2.000  PostScript diagrams and images within
Last Updated :  05/21/92  other documents. Encapsulated
Ported To :  Sun DEC IBM  PostScript files are recognized by the

follow- ing filename suffixes .eps, .epi, .epsi, and .epsf.

read gif2

Author :  John Langner  read gif2 is yet another module to
Stardent Computer  read a Compuserve GIF image file.
Number :  1303
Language :  C
Submitted :  07/11/92
Version :  1.000
Last Updated :  07/11/92
Ported To :  IBM Sun HP

read gif

Author :  Larry Gelberg  This module reads a GIF format file and
Stardent Computer  creates a 32-bit ARGB image in the AVS
Number :  1101  “field 2D 4-vector byte” format. If the
Language :  C  GIF file has an embedded color-map,
Submitted :  11/22/91  that is used, otherwise, it creates a
Version :  1.000  linear grey-scale ramp.
Last Updated :  11/22/91
Ported To :  Kubota DEC IBM Sun HP

read gif image

Author :  Terry Myerson  This module reads an image in the gif
International AVS Center  format into an AVS Network. GIF
(NCSC)  (Graphics Interchange Format) is
Number :  1214  CompuServe’s standard for generalized
Language :  C  color raster images. This standard is a
Submitted :  05/21/92  mechanism to exchange and display
Version :  2.000  high-quality, high- resolution graphics
Last Updated :  05/21/92  images. CompuServe gif files are
Ported To :  Sun DEC IBM  recognized by the filename suf- fixes

.gif and .giff.

read grd to fiel

Author :  Andy Bingham  The read grd to fld module reads a GMT
Oxford University, Dept.  v1.0 grd file from disk and outputs the
of Earth Sciences  file as “field 2D scalar float 2-coord
Number :  1699  uniform“. The output data is in the AVS
Language :  C  field format. The field extents are set
Submitted :  09/30/93  to the values of east, west, south and
Version :  1.000  north from the input file.
Last Updated :  09/30/93
Ported To :  HP IBM Sun

read hdf image

Author :  Terry Myerson  This module reads an image in the hdf
International AVS Center  format into an AVS Network. hdf is a
(NCSC)  generic, tagged Hierarchical Data File
Number :  1215  format developed by the National Center
Language :  C  for Supercomputing Applica- tions
Submitted :  05/21/92  (NCSA). hdf files may contain images,
Version :  2.000  scientific data sets, and miscellaneous
Last Updated :  05/21/92  data items. Such files may be created
Ported To :  Sun DEC IBM  by several NCSA tools. See the NCSA

documentation for details on how to use these tools. hdf format handling within the SDSC image library is limited to images of certain depths and storage methods. hdf files are recognized by these filename suffixes .hdf, .df, and .ncsa.

read icon image

Author :  Terry Myerson  This module reads an image in the icon
International AVS Center  format into an AVS Network. icon image
(NCSC)  files are used by Sun Microsystem’s
Number :  1216  SunView win- dow system, NeWS window
Language :  C  system, OpenWindows NeWS tool set, and
Submitted :  05/21/92  X11 XView tool set for the storage of
Version :  1.000  icons, cursors, fill patterns, and
Last Updated :  05/21/92  pieces of widgets (like button check-
Ported To :  Sun DEC IBM  marks). Sun icon files can be most

easily generated using Sun’s iconedit(1) icon and cursor editor. The Sun operating sys- tem release includes a directory of standard icons, cursors, background patterns, and widget pieces in icon format in the directory /usr/include/images. See the Sun documentation set for details on how to use the tools dealing with Sun icon files. Sun icon files are recognized by these filename suffixes .icon, .cursor, and .pr.

read iff image

Author :  Terry Myerson  This module reads an image in the iff
International AVS Center  format into an AVS Network. iff image
(NCSC)  files are generated by Sun Microsystems
Number :  1217  TAAC software such as voxvu(1) and
Language :  C  cloudvu(1). See the TAAC-1 Application
Submitted :  05/21/92  Accelerator Software Reference Manual
Version :  1.000  for information on how to use these
Last Updated :  05/21/92  programs. Note that image files
Ported To :  Sun DEC IBM  compressed with the Sun-TAAC utility

make movie(1) cannot be read by the SDSC image library. iff files are recognized by the following filename suffix .iff.

read irregul

Author :  Wes Bethel  This module reads data in ascii format
Lawrence Berkeley  into an “irregular field” data
Laboratory  structure. The disk file is structured
Number :  1074  to contain information about the field
Language :  C  at the beginning of the file, followed
Submitted :  10/10/91  by the data.
Version :  1.000
Last Updated :  10/10/91
Ported To :  DEC Sun HP

read jpeg

Author :  John Langner  Read a file containing an image
Stardent Computer  compressed with the JPEG standard. See
Number :  1099  Also - The best and most readily
Language :  C  available introduction to the JPEG
Submitted :  11/22/91  compression algorithm is Wallace’s
Version :  1.000  article in the April ’91 CACM Wallace,
Last Updated :  11/22/91  Gregory K. “The JPEG Still Picture
Ported To :  IBM Sun DEC HP  Compression Standard“, Communications

of the ACM, April 1991 (v. 34 no. 4), pp. 30-44.

read mpnt image

Author :  Terry Myerson  This module reads an image in the mpnt
International AVS Center  format into an AVS Network. The
(NCSC)  MacPaint mpnt file is the standard
Number :  1218  Apple Macintosh mono- chrome bitmap
Language :  C  image file format. It can be read by
Submitted :  05/21/92  many Macintosh graphics applications,
Version :  2.000  and many Macintosh applica- tions that
Last Updated :  05/21/92  can export bitmap graphics do so in the
Ported To :  Sun DEC IBM  MacPaint mpnt file format. mpnt files

are recognized by these filename suffixes .mpnt, .macp, and .pntg.

read netcdf

Author :  Wright, E.L.  This module reads data from netCDF
U.S Geological Survey  files into uniform or rectilinear AVS
Number :  1564  fields. The data variables must be of
Language :  C  dimension 8 or less. Variables that are
Submitted :  01/22/93  one-dimensional and have the same name
Version :  2.000  as a dimension name are assumed to be
Last Updated :  07/15/94  coordinate variables. Coordinate
Ported To :  DEC Sun  variables, if present, define the

coordinate values along each axis of the data space. A maximum of 8 data variables that are NOT coordinate variables may be present.

read netcdf example

Author :  Wright, E.L.  The read netcdf module is a working
U.S Geological Survey  example of a module that can read
Number :  1210  netCDF files. This version can only
Language :  C  read netCDF files that contain specific
Submitted :  05/20/92  dimension, variable, and attribute
Version :  1.000  names. Also, the coordinate values for
Last Updated :  05/20/92  the irregular field output are computed

in a way that is data specific. The netCDF files to be read by this version of read netcdf must contain the following dimension names xpos, ypos, zpos, and time, where the size of dimension time is declared as UNLIMITED. The following variables must be in the file float variables sigma(zpos), x(ypos,xpos), y(ypos,xpos), and depth(ypos,xpos) and short variables elev(time,ypos,xpos) and salt(time,zpos,ypos,xpos). Variables elev and salt must have the following float attributes scale factor and add offset, where the scale factor is to be multiplied first and then the offset added to the data.

read p 2

Author :  John Langner  Read a PC Paintbrush PCX image file.

Stardent Computer

Number :  1290
Language :  C
Submitted :  07/08/92
Version :  1.000
Last Updated :  07/08/92
Ported To :  IBM Sun HP

read pbm image

Author :  Terry Myerson  This module reads an image in the pbm
International AVS Center  format into an AVS Network. pbm image
(NCSC)  files are used by various tools in Jef
Number :  1219  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pbm files are
Version :  2.000  recognized by the filename suffix .pbm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

read pcx image

Author :  Terry Myerson  This module reads an image in the pcx
International AVS Center  format into an AVS Network. The pcx
(NCSC)  image file format was invented by ZSoft
Number :  1220  for use in its IBM PC PC Paintbrush
Language :  C  tool series. It has become a defacto
Submitted :  05/21/92  standard in the IBM PC world and is
Version :  2.000  regularly used for the storage of
Last Updated :  05/21/92  monochrome and color pixel information
Ported To :  Sun DEC IBM  by paint-type tools. See the

documentation for each of the IBM PC tools for details on how to use them. ZSoft pcx files are recognized by these filename suffixes .pcx and .pcc.

read pgm image

Author :  Terry Myerson  This module reads an image in the pgm
International AVS Center  format into an AVS Network. pgm image
(NCSC)  files are used by various tools in Jef
Number :  1221  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pgm files are
Version :  2.000  recognized by the filename suffix .pgm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

read pic image

Author :  Terry Myerson  This module reads an image in the pic
International AVS Center  format into an AVS Network. pic image
(NCSC)  files are generated by PIXAR
Number :  1222  programming software, such as
Language :  C  ChapVolumes and ChapReyes, the PIXAR
Submitted :  05/21/92  Image Runtime Library called Pirl, and
Version :  2.000  the PIXAR rendering tool RenderMan. See
Last Updated :  05/21/92  the PIXAR documentation set for details
Ported To :  Sun DEC IBM  on how to use these applications and

tools. PIXAR’s pic file format is sometimes referred to as PICIO in PIXAR documentation. pic and PICIO mean the same thing. PIXAR’s xpic is not the same as pic and is Φn Φo Φt supported by the SDSC image library. ΦN Φo Φt Φe PIXAR’s RenderMan always saves its image files with .pic filename suffixes. However, depending upon output defaults, RenderMan can generate .pic files with PICIO (same as pic) data or TIFF data. .pic files with TIFF data will confuse the SDSC image library. For you to avoid generating this type of file, we recommend that you configure RenderMan defaults to generate .pic files with PICIO data. PIXAR pic files are recognized by the following filename SDSC Last change March 4, 1991 1 IMPIC(3IM) SDSC IMAGE LIBRARY IMPIC(3IM) suffixes .pic, .picio, and .pixar.

read pict image

Author :  Terry Myerson  This module reads an image in the pict
International AVS Center  format into an AVS Network. The PICT
(NCSC)  file is the standard Apple Macintosh
Number :  1223  image file format. It can be read by
Language :  C  almost any Macintosh graphics
Submitted :  05/21/92  application. Most Macintosh
Version :  2.000  applications that can export graphics
Last Updated :  05/21/92  do so in the PICT file format. pict
Ported To :  Sun DEC IBM  files are recognized by these filename

suffixes .pict and .pict2.

read pix image

Author :  Terry Myerson  This module reads an image in the pix
International AVS Center  format into an AVS Network. pix image
(NCSC)  files are generated by the rendering
Number :  1224  and painting tools of Alias Research,
Language :  C  Inc., such as renderer, raytracer, and
Submitted :  05/21/92  paint. See the Alias documentation set
Version :  2.000  for details on how to use these tools.
Last Updated :  05/21/92  ΦN Φo Φt Φe Alias quickpaint,
Ported To :  Sun DEC IBM  available on Silicon Graphics, Inc.,

IRIS workstations, uses Silicon Graphic’s rgb image file format rather than the Alias pix image file format. See the imrgb(3IM) man page for details on the rgb format. Alias pix files are recognized by these following filename suffixes .alias and .pix.

read plot3d

Author :  Michael Rangitsch  This module reads either a formatted or
Dow Chemical  unformatted multiblock PLOT3D data set
Number :  1854  consisting of the grid and data vector
Language :  C  (Q file). No corrections are made for
Submitted :  10/17/94  block edge overlap and no co-incident
Version :  1.000  node removal is attempted between
Last Updated :  10/17/94  blocks. The cells of the ucd structure
Ported To :  DEC alpha HP IBM Sun  created are all hexahedra. Version 1.0

of new plot3d reads in only one of the possible plot3d file types, a multiblock (with or without iblanks) stored in the whole field format. The Q vector is labeled with the default names of the primative variables. A second module similar to the unsupported ’extract cfd values’ module is planned to compute derived properties from the primative variables. A crude test is made to determine whether or not a file is formatted. This may fail on some platforms.

DATA INPUT MODULES

read pnm image

Author :  Terry Myerson  This module reads an image in the pnm
International AVS Center  format into an AVS Network. pnm image
(NCSC)  files are used by various tools in Jef
Number :  1225  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pnm files are
Version :  2.000  recognized by the filename suffix .pnm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

read points

Author :  Penny Rheingans  Read observation locations and variable
US EPA Scientific  values from an ASCII file. From these
Visualization Center  observations build a 1D irregular
Number :  1311  (scatter) field with one element for
Language :  C  each observation.
Submitted :  08/30/92
Version :  1.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

read ras image

Author :  Terry Myerson  This module reads an image in the ras
International AVS Center  format into an AVS Network. ras image
(NCSC)  files are used by various Sun
Number :  1227  Microsystems Inc. tools, such as
Language :  C  screendump(1) and screenload(1). See
Submitted :  05/21/92  the Sun documentation set for details
Version :  2.000  on how to use these tools. Sun ras
Last Updated :  05/21/92  files are recognized by any of the
Ported To :  Sun DEC IBM  following filename suffixes .ras, .scr,

.sr, and .sun.

read rectilinear

Author :  Wes Bethel  This module reads data in ascii format
Lawrence Berkeley  into an “rectilinear field” data
Laboratory  structure. The disk file is structured
Number :  1296  to contain information about the field
Language :  C  at the beginning of the file, followed
Submitted :  08/04/92  by the data.
Version :  1.000
Last Updated :  08/04/92
Ported To :  DEC IBM Sun HP

read rgb image

Author :  Terry Myerson  This module reads an image in the rgb
International AVS Center  format into an AVS Network. rgb image
(NCSC)  files are generated by Silicon
Number :  1228  Graphics, Inc., software such as
Language :  C  icut(1) and snapshot(1). See the
Submitted :  05/21/92  Silicon Graphics documentation for
Version :  2.000  information on how to use these and
Last Updated :  05/21/92  other Silicon Graphics programs.
Ported To :  Sun DEC IBM  Silicon Graphics rgb files are

recognized by the following filename suffixes .rgb, .iris, and .sgi.

read rla image

Author :  Terry Myerson  This module reads an image in the rla
International AVS Center  format into an AVS Network. rla image
(NCSC)  files are generated by Wavefront
Number :  1229  Technologies, Inc., software such as
Language :  C  image and tdv. See the Wavefront
Submitted :  05/21/92  documentation for information on how to
Version :  2.000  use these and other Wavefront programs.
Last Updated :  05/21/92  rla is a subset of the newer Wavefront
Ported To :  Sun DEC IBM  rlb specification. Programs that can

read and write rlb files can also read and write rla files. The SDSC image library reader/writer is written to accommodate both types of files. Wavefront rla files are recognized by the following filename suffixes .rla and .rlb.

read rle image

Author :  Terry Myerson  This module reads an image in the rle
International AVS Center  format into an AVS Network. rle image
(NCSC)  files are generated by the tools of
Number :  1230  Utah’s Raster Toolkit. See the Utah
Language :  C  documentation set for details on how to
Submitted :  05/21/92  use these tools. Utah rle files are
Version :  2.000  only recognized by the following
Last Updated :  05/21/92  filename suffix .rle
Ported To :  Sun DEC IBM

DATA INPUT MODULES

read rpbm image

Author :  Terry Myerson  This module reads an image in the rpbm
International AVS Center  format into an AVS Network. rpbm image
(NCSC)  files are used by various tools in Jef
Number :  1231  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpbm files are
Version :  2.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpbm.
Ported To :  Sun DEC IBM

read rpgm image

Author :  Terry Myerson  This module reads an image in the rpgm
International AVS Center  format into an AVS Network. rpgm image
(NCSC)  files are used by various tools in Jef
Number :  1232  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpgm files are
Version :  2.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpgm.
Ported To :  Sun DEC IBM

read rpnm image

Author :  Terry Myerson  This module reads an image in the rpnm
International AVS Center  format into an AVS Network. rpnm image
(NCSC)  files are used by various tools in Jef
Number :  1233  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpnm files are
Version :  2.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpnm.
Ported To :  Sun DEC IBM

read rppm image

Author :  Terry Myerson  This module reads an image in the rppm
International AVS Center  format into an AVS Network. rppm image
(NCSC)  files are used by various tools in Jef
Number :  1234  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rppm files are
Version :  2.000  recognized by the filename suffix
Last Updated :  05/21/92  .rppm.
Ported To :  Sun DEC IBM

read semper

Author :  Michel DROZ  Read Semper accesses an image file
Number :  1339  format from the SEMPER image processing
Language :  C  system. The entire file is read into
Submitted :  08/31/92  memory, and sent to the Image output
Version :  1.000  port for further processing by other
Last Updated :  08/31/92  AVS modules. The Y scan line order is
Ported To :  DEC Sun HP  reversed compared to other AVS images,

so the image will appear upside down unless reversed with “mirror”. The images are generally in “xxx.pic” named file.

read shak

Author :  Keith Refson  The read shak module and the
Department of Earth  shak to geom filter are crys-
Sciences, Oxford  tallography front-ends for AVS. Both
University  read an input file containing a
Number :  1176  description of the atomic co-ordinates
Language :  C  of the structure or molecule to be
Submitted :  04/16/92  visualized and output AVS geometry
Version :  3.000  representing a ball-and-stick or
Last Updated :  09/30/93  space-filling model of it. The read
Ported To :  Sun Kubota  shak module is suited to interactive

viewing whereas the shak to geom filter is useful for off- line creation of a sequence of .geom files for animation by fast animate(6). Please note - there are a number of sample data file for this module available via anonymous ftp from avs.ncsc.org, in the directory sample data/avs data/read shak.

read sunras

Author :  Wes Bethel  Reads an image stored in the SUN
Lawrence Berkeley  Rasterfile format, placing the results
Laboratory  into an AVS image field. Most sun
Number :  1132  rasterfile formats are supported (the
Language :  C  exceptions being RT FORMAT TIFF and
Submitted :  01/29/92  RT FORMAT IFF). Run-length-encoded,
Version :  2.000  BGR and RGB formats are supported. Most
Last Updated :  01/29/92  depths, as well are supported. The
Ported To :  Sun HP  exception is that depth=1 (monochrome)

is NOT supported. This module will also decompress files which have been compressed using the UNIX(tm) “compress” utility (by checking for the existance of a “.Z” suffix.

read synu image

Author :  Terry Myerson  This module reads an image in the synu
International AVS Center  format into an AVS Network. synu is the
(NCSC)  image file format output by SDSC’s synu
Number :  1235  (Syn- thetic Universe) portable
Language :  C  renderer. synu files are recognized
Submitted :  05/21/92  only by the following filename suffix
Version :  2.000
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

read tiff image

Author :  Terry Myerson  This module reads an image in the tiff
International AVS Center  format into an AVS Network. tiff is a
(NCSC)  generic Tagged Image File Format
Number :  1236  developed by Aldus and Microsoft in
Language :  C  conjunction with leading scanner and
Submitted :  05/21/92  printer manufacturers. tiff files may
Version :  2.000  contain images and miscellaneous other
Last Updated :  05/21/92  image-related items. Such files may be
Ported To :  Sun DEC IBM  created and manipulated by a variety of

Tagged Image File Format tools. See the Tagged Image File Format documentation for details on how to use these tools. tiff support within the SDSC image library is limited to images of certain depths and storage methods. tiff files are recognized by the following filename suf- fixes .tiff and .tif.

INTERNATIONAL AVS CENTER MODULE CATALOG

read tiff sectio

Author :  Joe Panico  This module reads single planes
Dept. of Neuroscience,  (slices) from a multi-plane TIFF file
University of  and provides at the output port an AVS
Pennsylvania  image (2-D scalar field). A dial widget
Number :  1616  allows one to interactively dial
Language :  C  through the stack of slices.
Submitted :  05/01/93
Version :  1.000
Last Updated :  05/01/93
Ported To :  IBM

read tiff vol

Author :  Joe Panico  This module reads in a multi-plane TIFF
Dept. of Neuroscience,  file (a volume) and provides at the
University of  output port an AVS volume (3-D scalar
Pennsylvania  field).
Number :  1617
Language :  C
Submitted :  05/01/93
Version :  1.000
Last Updated :  05/01/93
Ported To :  IBM

read ucd points

Author :  Penny Rheingans  Read observation locations and variable
US EPA Scientific  values from an ASCII file. From these
Visualization Center  observations build a ucd structure
Number :  1310  field with one node for each
Language :  C  observation. The cells of the structure
Submitted :  08/30/92  are of type UCD POINT.
Version :  1.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

read uniform

Author :  Wes Bethel  This module reads data in ascii format
Lawrence Berkeley  into a “uniform field” data structure.
Laboratory  The disk file is structured to contain
Number :  1075  information about the field at the
Language :  C  beginning of the file, followed by the
Submitted :  10/10/91  data.
Version :  2.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

read xbm image

Author :  Terry Myerson  This module reads an image in the xbm
International AVS Center  format into an AVS Network. xbm bitmap
(NCSC)  image files are generated by MIT’s X
Number :  1237  Window Sys- tem, version 11 (hereafter
Language :  C  referred to as X11). xbm files are
Submitted :  05/21/92  generated by the X11 bitmap(1) bitmap
Version :  2.000  editor and used by most X11 tools to
Last Updated :  05/21/92  define cursors, icons, and other mono-
Ported To :  Sun DEC IBM  chrome glyphs. See the X11

documentation set for details on how to use tools and subroutines dealing with X11 bitmaps. X11 xbm files are recognized by the following filename suf- fixes .xbm and .bm.

read xwd image

Author :  Terry Myerson  This module reads an image in the xwd
International AVS Center  format into an AVS Network. xwd window
(NCSC)  dump image files are used by the xwd(1)
Number :  1238  and xwud(1) tools of MIT’s X Window
Language :  C  System, version 11 (hereafter referred
Submitted :  05/21/92  to as X11). See the X11 documentation
Version :  2.000  set for details on how to use these
Last Updated :  05/21/92  tools. X11 xwd files are recognized by
Ported To :  Sun DEC IBM  the following filename suf- fixes .xwd

and .x11.

river bed

Author :  Ken Flurchick  This module is used to construct the
North Carolina  outline of a river bed. The length
Supercomputing Center  parameter specifies the length of the
Number :  1579  river in grid cells and coordinates.
Language :  FORTRAN  The river starts at y coordinate 0 and
Submitted :  02/11/93  extends downward to y coordinate length
Version :  1.000  at y index length. The river is
Last Updated :  02/11/93  constructed in the xy plane. Width
Ported To :  IBM Kubota HP  specifies the width of the river in the

x direction in grid cells. Width has nothing to do with the shape of the river. The outline of the river is built from a set of sinusoidal line segments of the form A * Sin (wt) + O, where A - Amplitude w - Omega t - Length O - Offset. The data set that is constructed is filled with 0’s initially. This module subroutine can be conected to the “river fill” coroutine to create a simulation of river contaminants.

shaker

Author :  Janet Jensen  Shaker reads an ascii data file
Number :  1144  defining a molecule of atoms, and
Language :  C  allows the user to modify the
Submitted :  03/03/92  vibrational mode and frequency for that
Version :  2.000  molecule. This module includes a
Last Updated :  05/19/92  command line filter to interface the
Ported To :  DEC IBM Sun Kubota HP  avs module with gaussian 90 output.

sololoadmd

Author :  Ken Flurchick and Mark  sololoadmd - read velocity and
Reed  coordinate information for N point
North Carolina  objects and represent them as spheres
Supercomputing Center  color coded by velocity This module is
Number :  1604  a modified version of the genloadmd
Language :  C  module that is intended for use when
Submitted :  04/09/93  memory restrictions don’t allow the
Version :  1.000  entire data set to be read into an
Last Updated :  04/09/93  array. Therefore this modules only
Ported To :  DEC IBM Kubota  reads in one data set at a time and

then discards it when a new set is read in. It has several enhancements to facilitate use with a single data set at a time. Note also that the data is expected in a slightly different format than in the genloadmd mdoule. This module reads in position and velocity information for a group of point particles in three dimensions, from a file specified as an input parameter. This parameter can be connected to either the file browser or the animated filename module depending on the user requirements. The program represents each point with a sphere at the position specifed and color coded by the magnitude of the velocity. The colors 0-255 are specified in the colors.h file and range from blue (slowest) to red (fastest). Initially the module will compute the minimum and maximum speeds from the data set read in and echo these back to the parameter widgets. The user can change the size of the spheres and the min and max values used to determine the colors of the sphere. Note that if the minimum speed is greater than or equal to the maximum speed then a message is echoed back to standard output (the screen) and no action is taken. Sphere color is determined by a linear interpolation between the minimum and maximum speeds of the particle velocity magnitude.If the velocity magnitude of a particular sphere exceeds the maximum speed then it is displayed as white while those less than the minimum are displayed as black (note - change the background color if this also is black).

spaceball master

Author :  John Stephen  spaceball master - use small geomviewer
Tessella Support  window to control views. This module is
Services  used with spaceball slaves, to create a
Number :  1717  small geometry viewer icon window, that
Language :  C  can be interactively transformed, then
Submitted :  09/30/93  used to cause mirror movements in other
Version :  1.000  geometry viewer scenes. See also -
Last Updated :  09/30/93  spaceball slave.
Ported To :  HP IBM Sun

spaceball slave

Author :  John Stephen  This module is used with spaceball
Tessella Support  master, to create a small geometry
Services  viewer icon window, that can be
Number :  1718  interactively transformed, then used to
Language :  C  cause mirror movements in other
Submitted :  09/30/93  geometry viewer scenes. Individual
Version :  1.000  views can be “snapped” to conform to
Last Updated :  09/30/93  other views, or cause other views to
Ported To :  HP IBM Sun  snap to current one. See also -

spaceball slave.

sphere to geom

Author :  Tim O. Robinson  This module began as a simple wrap
UC Berkeley  around the sphere.c filter but has been
Number :  1273  extended to include some interesting
Language :  C  features. It uses the same ascii file
Submitted :  05/27/92  format as /usr/avs/filter/sphere.c,
Version :  1.000  however, the user may also add a list
Last Updated :  05/27/92  of lines “bonds” to the end of the file
Ported To :  DEC IBM Sun Kubota HP  and these will be picked up and added

to the geometry output. The min and max dials have the effect of 3D cropping on location within the bounds while retaining the boundary box.

string example

Author :  Larry Gelberg  The string example module sends a
Advanced Visual Systems  single user-specified filename string
Number :  1553  to one or more string parameter ports
Language :  C  on one or more receiving modules. Its
Submitted :  11/05/92  purpose is to allow you to
Version :  1.000  simultaneously control filename
Last Updated :  11/05/92  parameter input to more than one module
Ported To :  DEC IBM Kubota HP  using only a single File Browser input

widget. This code was based on the AVS “file browser” module.

string list

Author :  Phil McDonald  String list is a coroutine module that
NOAA/ERL/Forecast  can be used to input a series of
Systems Laboratory  strings from a list, a single string
Number :  1279  from the list, or a string from the
Language :  C  user. If a list of strings is supplied,
Submitted :  05/29/92  the strings may be input
Version :  1.000  “automatically” (read as fast as the
Last Updated :  05/29/92  network will accept them) or stepped
Ported To :  Kubota  through one at a time. The list may be

editted by skipping over individual strings. If the user-supplied string contains an asterisk, the string is used as a template – each list string is substituted for the asterisk wildcard to create the actual string that is output.

string macros

Author :  Ian Curington  This module provides a string macro
Advanced Visual Systems  facility for annotated choice lists
Number :  1685  anywhere strings are used in an AVS
Language :  C  network. The module reads a
Submitted :  09/29/93  configuration macro list file, sets up
Version :  1.000  a choice list of macros, then puts the
Last Updated :  09/29/93  contents of the macro file at the
Ported To :  HP IBM Sun  output port.

strip chart

Author :  Larry Gelberg  STRIP CHART archives (collects) values
Advanced Visual Systems  from some other routine and packages
Number :  1115  them up into 1D fields appropriate for
Language :  C  sending to the graph viewer. There are
Submitted :  01/06/92  two internal arrays, each 1024 element
Version :  1.000  long. One is for integer values and the
Last Updated :  01/06/92  other is for floating point values.
Ported To :  Kubota DEC IBM Sun HP  Each array is static and has its own

counter associated with it. When the counters pass 1024, they wrap back to 0.

taperer

Author :  Larry Gelberg  The taperer module is one of a suite of
Advanced Visual Systems  four AVS modules which can be used to
Number :  1634  design simple woodworking projects. For
Language :  C  more information on the purpose of the
Submitted :  07/05/93  project, please read the file called
Version :  1.000  “Poster Session” and the man pages for
Last Updated :  07/05/93  each module. The four modules are
Ported To :  HP IBM Kubota DEC Sun  “cabinet maker” - generates structures

made of boxes, “taperer” - generates structures made of tapered boxes, “extruder” - creates capped, extruded tubes from lists of X/Y pairs, and “grid gen” - creates measured grids so that accurate models can be constructed. Due to the current structure of the IAC’s directories, this suite has been split into 4 separate directories on the ftp site. However, a compressed tar archive of all 4 mods is included with each of the directories as wood mod suite.tar.Z. Also note that there are several images contributed with this module suite. These have been stored on the ftp site in the directory at avs.ncsc.org called sample data/avs data/woodworking ima ages.

teapot

Author :  Anonymous  This module creates a geom version of
Stardent Computer  the famous Newell teapot as described
Number :  1085  by Jim Blinn. It creates polygons based
Language :  C  on subdiving bi-cubic patch definitions
Submitted :  10/10/91  and then lets you mangle the teapot in
Version :  1.000  all sorts of creative ways. The
Last Updated :  10/10/91  ’Offset’ knob controls the ’bloat’ of
Ported To :  Kubota DEC IBM Sun HP  the teapot (offsetting the polygons in

the direction of their surface

normals). The ’Twist’ knob controls how far around the selected ’Twist Axis’ the teapot will be wrenched by and the ’Mesh Res’ knob controls the resolution of the bi-cubic patch division. While this is generally meant to be a fun demonstration module, it is a good road map for a module which turns bi-cubic patches into renderable polytriangle strips.

track cursor

Author :  Anonymous  Track Cursor uses X Window System calls
Stardent Computer  to follow the cursor around the screen
Number :  1139  and grab N*M pixels off the screen
Language :  C  (with the cursor positioned at the
Submitted :  02/26/92  upper left hand corner of the snagged
Version :  1.000  window). It then outputs the (32-bit,
Last Updated :  02/26/92  true color) window contents as an AVS
Ported To :  Sun HP  field 2D 4-vector byte which may be

displayed using DISPLAY IMAGE, etc.

transform xyz

Author :  Lisa Traub  This module reads an ASCII file
US Geological Survey  consisting of lists of x,y,z points,
Number :  1726  performs a scale and offset on each
Language :  C  dimension and constructs a polyline
Submitted :  12/14/93  geometry object using the points in the
Version :  1.000  file as vertices. This program was
Last Updated :  12/14/93  developed through the combining of Wes
Ported To :  Sun  Bethel’s plot xyz and Convex’s

transform points modules.

tree

Author :  Larry Gelberg  TREE generates a 2D binary tree-like
Stardent Computer  geometry using recursive techniques.
Number :  1104  You can control things like the level
Language :  C  of recursion, the average angle between
Submitted :  11/26/91  the branches, the average branch
Version :  1.000  length, and random Gaussian deviations
Last Updated :  11/26/91  from the average. You can also tell the
Ported To :  Kubota DEC IBM Sun HP  module to generate green “leaves” on

the leaf nodes.

trigger int

Author :  Wes Bethel  The trigger int module waits for a
Lawrence Berkeley  field to be presented on its input
Laboratory  port. When a field (new data) is
Number :  1601  detected, an integer is stuffed into
Language :  C  the output port. The value of the
Submitted :  03/23/93  integer is determined by the “start
Version :  1.000  value“ parmeter for the first
Last Updated :  03/23/93  execution. Subsequent invocations cause
Ported To :  DEC IBM HP  a static variable to be incremented

according to the parameter “increment.” This module is useful in creating networks to facilitate animations based upon a single coroutine.

trigger int 2

Author :  anonymous  The trigger int module waits for a
Number :  1893  field to be presented on its input
Language :  C  port. When a field (new data) is
Submitted :  09/07/95  detected, an integer is stuffed into
Version :  1.000  the output port. The value of the
Last Updated :  09/07/95  integer is determined by the “start
Ported To :  DEC ultrix DEC alpha  value“ parameter for the first

execution. Subsequent invocations cause a static variable to be incremented according to the parameter “increment.” If the Reset Value button is pressed, then the output integer is reset to the start value. This module is useful in creating networks to facilitate animations based upon a single coroutine.

ucd color ra

Author :  Mike French  The range of the UCD scalar node data
Evans and Sutherland UK  is used to set the lo/hi values for the
Number :  1758  colormap. If the UCD min/max value is
Language :  C  not available a warning is printed on
Submitted :  12/15/93  the terminal output for AVS and the
Version :  1.000  range is calculated from the UCD
Last Updated :  12/15/93  dataset. The behaviour of COLOR RANGE
Ported To :  Kubota Sun  UCD is designed to match that of COLOR

RANGE provided with AVS for use on fields. This module is design to be used with the ISOBAND UCD and ISOLINE UCD contouring modules.

ucd comp

Author :  Jonathan Cox  The ucd comp module allows a component
Department of Computer  of model, cell or node data to be
Science, University of  selected from a UCD structure. The
Manchester  selected component is made available on
Number :  1712  an output port. This output can be used
Language :  C  to set the component parameters of
Submitted :  09/30/93  other ucd modules via network
Version :  1.000  connections. This allows a single set
Last Updated :  09/30/93  of radio buttons to control component
Ported To :  HP  selection in a number of UCD modules

within a network. A component of model, cell or nodal data is to be selected. This is controlled through a choice widget. It is likely that in a network this parameter will be not be modified. Therefore, it is envisioned that the network editor will be used to delete this set of radio buttons in most networks. The module also outputs a suitable title string for the selected component.

uncompress file

Author :  Lee J. Silverman  The compress/uncompress module pair is
Brown University  designed to allow datasets or images or
Number :  1589  other files to be stored and kept in
Language :  C  compressed format, but still be read
Submitted :  02/21/93  into an AVS network. There must be
Version :  1.000  enough space for this module to
Last Updated :  02/21/93  uncompress at least one file at a time,
Ported To :  DEC IBM  or it will bomb out. Generally

speaking, a filename will be passed to the uncompress module, either through the browser or via input parameter string. The uncompress module will decompress the file, then pass the filename on to a data-input module such as Read Field.

uv rect

Author :  John Tee  uv rect - Create a rectangle with
AVS Inc. Consultant  longitude and latitude UV coordinates.
Number :  1695  UV RECT generates a geometry which is a
Language :  C  rectangle, with UV coordinates set
Submitted :  09/29/93  according from 0,0 at the top left to
Version :  1.000  “U max” “V max” at the bottom right. UV
Last Updated :  09/29/93  RECT is a demonstration module to show
Ported To :  HP IBM Kubota Sun  how rectangular portions of scanned map

data could be assembled as geometries on to which the map data is texture mapped.

uv sector

Author :  John Tee  uv sector - Create a mesh over a
AVS Inc. Consultant  sector of an annulus. UV SECTOR
Number :  1696  generates a geometry which is a sector
Language :  C  of a circular annulus, with UV
Submitted :  09/29/93  coordinates¿ The geometry of the mesh
Version :  1.000  is determined by the input parameters.
Last Updated :  09/29/93  The Resolution parameter determines the
Ported To :  HP IBM Sun  number of mesh nodes in both

directions. The UV coordinates are set from 0,0 to 1,1 over the extremities of the mesh. The purpose of the module is to demonstrate the ability to warp an image (eg cartographic dataa) by texture mapping it to a geometry whose shape can be varied.

waves demo

Author :  Dr. David Faux  This module is used to teach physics to
Physics Department,  a group of 2nd year students, and
University of Surrey  illustrates the transmission and
Number :  1191  reflection of electromagnetic waves
Language :  FORTRAN  from transparent media. The
Submitted :  05/05/92  transparency of the media represents
Version :  1.000  the magnitude of their refractive
Last Updated :  05/05/92  indices, which can be altered. The
Ported To :  DEC Kubota  incident angle may be changed and all

other angles, reflection and transmission coefficients, amplitudes of waves, phase changes etc are calculated and diplayed. The waves propagate when the animate button is clicked. Interesting physical effects such as the Brewster angle may be demonstrated.

FILTER MODULES

International AVS Center

FILTER MODULES

4D to 3D Sli

Author :  Annette Walsh  Accepts 4D data and extracts a 3D
Number :  1143  “slice”. It is essentially doing the
Language :  FORTRAN  same processing as the “orthogonal
Submitted :  03/03/92  slicer“ but for 4D data rather than 2D
Version :  1.000  or 3D. The slice plane parameter
Last Updated :  03/03/92  determines the value of the array index
Ported To :  IBM HP  to be held constant. This value is

reset to zero each time a new data field is input. The axis parameter selects the dimension (I, J, K or L) in which the array index is to be held constant.

ColorToBW

Author :  Steve Thorpe  This module converts an image from
International AVS Center  color to black and white, keeping the
Number :  1599  alpha channel the same in the resulting
Language :  C  image.
Submitted :  12/15/92
Version :  1.000
Last Updated :  12/15/92
Ported To :  DEC Kubota HP IBM

Color Legend

Stardent Computer  Provides a color map data type for
Number :  1049  other mapping modules to use, and
Language :  C  provides an annotated display of the
Submitted :  10/10/91  current map. It makes a discrete map,
Version :  1.000  with 16 entries, where each entry color
Last Updated :  10/10/91  can be edited individually. The map
Ported To :  Kubota  definition is read from a file (with

file browser) including annotation details.

Cylinder

Author :  D. LE CORRE  This module maps a 2D input field along
CONVEX S.A  a (portion of a) cylinder. If the input
Number :  1199  field is UNIFORM, the y dimension gives
Language :  C  the number of subdivisons of the circle
Submitted :  05/08/92  and the x dimension is the depth of the
Version :  1.000  cylinder. If the input field is
Last Updated :  05/08/92  rectilinear, the x positions are taken
Ported To :  DEC IBM Kubota HP  along the cylinder axis, while the y

values are positions around the circle, given in degrees. direction give the sphere subdivison.

Delta Int

Author :  Stephen H. Price  This module takes in an integer and
Loma Linda University  outputs an integer which is the
Medical Center  differnce based on the value of an
Number :  1629  parameter dial- thus the name
Language :  C  Delta int. The range of the delta is
Submitted :  07/02/93  -200 to +200. This module is really
Version :  1.000  useful if you need to input i.e. N and
Last Updated :  07/02/93  N-1 to two ports down stream. Also
Ported To :  HP IBM Kubota DEC Sun  useful to animate a crop range i.e. 5

to 10, then 6 to 11, then 7 to 12, etc...

Downsize 2

Author :  Ian Curington (AVS  The Downsize module changes the size of
Inc.), and Mike French  the input data set by subsampling the
(E and S)  data. It extracts every Nth element of
Number :  1746  the field along each dimension. This
Language :  C  one works with 1D, 2D, 3D, uniform or
Submitted :  12/15/93  irregular. It has corrections for
Version :  1.000  non-integral downsize coordinates in
Last Updated :  12/15/93  Uniform fields, and will work with 1D
Ported To :  Kubota Sun  scatter data. There is a bug in the

DOWNSIZE modules supplied by AVS Inc. The problem arises when using DOWNSIZE for uniform fields when the downsize value is not an exact factor of all the (dimension-1) values in X, Y and Z. There should be an empty region for the ’remainder’, but the standard version always scales the output to the full extents. The result is that spatial positions within field slices will be slightly wrong. This will be most noticeable when two contour surfaces intersect - the bands will not match exactly. A fixed version of the DOWNSIZE module is supplied with the contouring suite, it’s module name is ’Downsize’, the upper case initial is used to distinguish it from the standard version.

E01SAF

Author :  Steve Larkin  This module provides an interface to
Computer Graphics Unit,  the NAG Fortran routines E01SAF and
University of  E01SBF. The first routine (E01SAF)
Manchester, UK  generates a 2D surface interpolating a
Number :  1568  set of scattered data points, using the
Language :  C  method of Renka and Cline. The second
Submitted :  01/28/93  routine (E01SBF) then evaluates the
Version :  1.000  surface at a given point.
Last Updated :  01/28/93

Extract label

Author :  Penny Rheingans  Extracts the label and units for one
US EPA Scientific  variable in a UCD structure and
Visualization Center  concatenates them into a string.
Number :  1299
Language :  C
Submitted :  08/26/92
Version :  1.000
Last Updated :  08/26/92
Ported To :  DEC Sun HP

Field Concat

Author :  Krzysztof (Chris) S.  Field Concat concatenates Out Veclen
Nowinski  instances of input scalar field into
Institute of Applied  one vector field of dimensions and type
Mathematics and  same as the first scalar field (if a
Mechanics, Warsaw  field of another dimensions and/or type
University  occurs at the input, the module starts
Number :  1704  creation of a new vector field as in
Language :  C  the case of changing of Out Veclen or
Submitted :  09/30/93  resetting). If increment is on, neach
Version :  1.000  new input increments the index
Last Updated :  09/30/93  parameter filling sunsequent components
Ported To :  of the field. The user can control the

index parameter manually usinf set button to fill in new components. If all components of the outfield are set, the module spits it downstream.

Field Formula

Author :  Ronald E. Stogner  The “Field Formula” module accepts a
Computer Sciences  reverse polish notation (RPN) formula,
Corporation  specified by the input string. This
Number :  1545  formula is then applied to all elements
Language :  C  of the field and a field of the same
Submitted :  11/03/92  dimension is output with a vector
Version :  1.000  length of one. This field contains the
Last Updated :  11/03/92  result of the applied formula for all
Ported To :  DEC IBM  locations in the field. “Variables” are

defined by their field labels and are delimited by “[]”’s.

Field Math

Author :  Krzysztof (Chris) S.  The module does some field mathematics
Nowinski  – I am intended to talk about it at
Institute of Applied  the AVS94. The module interprets a
Mathematics and  short (up to 60 lines) program written
Mechanics, Warsaw  in AVSFOL (A Very Simple Field
University  Operation Language). AVSSPL admits two
Number :  1770  types of variables - simple variables
Language :  C  with values being single float numbers;
Submitted :  02/03/94  field variables with values being float
Version :  1.000  arrays of dimensions determined by the
Last Updated :  02/03/94  output field.
Ported To :  HP Sun

GRID MATH

Author :  Kathy Pearson  NAME GRID MATH – create a single
MCNC Environmental  gridded representation of statistical
Programs  functions over multiple time steps. The
Number :  1801  GRID MATH module takes 4-D input and
Language :  C  creates 3-D output based upon the
Submitted :  04/21/94  selected statistical function. The
Version :  1.000  effect is to reduce multiple time
Last Updated :  04/21/94  dimensions to a single one representing
Ported To :  DEC HP IBM Kubota Sun  the chosen math function. The MCNC

Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

GeomXYZtoFld

Author :  Dominique Le Corre  This module takes as input a geometry,
Tethys S.A.  gets all vertices and generates a 1D
Number :  1618  3-space field with the coordinates of
Language :  C  these vertices. It can be used for
Submitted :  05/01/93  instance as the “sampler” input field
Version :  1.000  to hedgehog,streamlines, particle
Last Updated :  05/01/93  advector, ucd streamlines (AVS5),
Ported To :  Kubota HP  scatter dots, etc... Optionally, 2

integer dials (From and To) can be used to limit the output field to a portion of the geometry.

Geom Dup n Mirro

Author :  Janet Haswell  Copy and symtetric mirror geometry
Rutherford Appleton  data. Given axis symetric geometry,
Laboratories  this will duplicate and create mirror
Number :  1657  copies, and produce one geometric
Language :  C  object out of many, performing geometry
Submitted :  09/29/93  concatonation. Used “geom duplicate.c”
Version :  1.000  from Ian Curington (AVS Inc) as a basis
Last Updated :  09/29/93  to generate this module. Take up to 6
Ported To :  HP IBM Kubota Sun  geom input ports, combine into 1

geometry then duplicate and mirror according to flags (can mirror in x and/or y and/or z and all 3). Also allows the user to set transparency level for all objects being mirrored, can also - select wireframe mode, cause AVS to compute the normals, and hide probes (anything with “probe” in the name). Lighting, rendermode, perspective, are controlled by toggles. This was used in to represent a section of a magnetic field, showing the whole from the input of just one quadrant.

Geom Duplicate

Author :  Ian Curington  Merge up to 4 seperate geometries into
Advanced Visual Systems  one, with the multiple geom input
Number :  1335  ports, then replicate the combined
Language :  C  object N times with a positional offset
Submitted :  08/31/92  between them. Each copy is a unique
Version :  1.000  object in the Geometry Viewer, so may
Last Updated :  08/31/92  be transformed or have properties
Ported To :  DEC Sun HP  edited independently.

Geom Unique Name

Author :  Krzysztof (Chris) S.  The module passes through any geometric
Nowinski  input appending to its name a string
Institute of Applied  uniquely describing each instance. When
Mathematics and  no additional input is supplied, the
Mechanics, Warsaw  string is just the current number
University  (increased by one at each call, when
Number :  1706  Increment button is on, reset to 0 when
Language :  C  Reset button is pressed or
Submitted :  09/30/93  alternatively set to the value typed
Version :  1.000  in). Otherwise, the value at input is
Last Updated :  09/30/93  converted to string and appended to the
Ported To :  HP IBM Kubota Sun  name of the input object. This allows

e.g. to obtain a series of isosurfaces named “isosurface 75.0”, “isosurface 75.5”, “isosurface 76.0” etc.

Geos Image

Author :  Terry Myerson  Filters images; originally this was
International AVS Center  intended to be used with satellite
(NCSC)  imagery. The module was written for
Number :  1112  artistic coloration of imagery. If the
Language :  C  red channel is ever greater than the
Submitted :  12/10/91  blue and green channels, individually,
Version :  2.000  then the red and green channels are
Last Updated :  12/10/91  switched in the input image. To insure
Ported To :  Kubota DEC IBM Sun HP  that the difference is not to great,

red and blue can then be added back into the image as a proportion of the red-green difference. The difference after which color is added is the G-R Threshold parameter. The proportion of this difference which is added to the blue and red channels is the Add Factor parameters.

IDL

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1782  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

IMAGE WINDOW

Author :  Terry Myerson  This module allows the user to specify
International AVS  the size of the image they wish to clip
Center(NCSC)  out of the input image, and then lets
Number :  1209  the uses interactively move a window of
Language :  C  that size around the image, in order to
Submitted :  05/15/92  crop out the appropriate section. This
Version :  1.000  is very similar to the crop module, and
Last Updated :  05/15/92  in fact provides the same functionality
Ported To :  Kubota HP  for images, but makes it alot easier if

you know the size of your final image.

Input Formula

Author :  Ronald E. Stogner  The “Input Formula” module accepts a
Computer Sciences  input field and uses labels in the
Corporation  field as variables or “species” in an
Number :  1546  algrebraic expression. The user can
Language :  C  select various parts of the expression
Submitted :  11/03/92  from the available widgets and build a
Version :  1.000  formula interactively. The module
Last Updated :  11/03/92  parses this expression and outputs a
Ported To :  Sun  RPN sting expected by “Field Formula”

when the user selects the “Parse

Formula“ widget.

Int to string

Author :  Penny Rheingans  Converts an integer input to its string
US EPA Scientific  representation. This is really just the
Visualization Center  AVS equivalent of sprintf.
Number :  1301
Language :  C
Submitted :  08/26/92
Version :  1.000
Last Updated :  08/26/92
Ported To :  DEC Sun HP

Intervals Field

Author :  Mike French  A set of data values is taken from the
Evans and Sutherland UK  range of an input data field. The first
Number :  1748  output value is the minimum of the
Language :  FORTRAN  field, the last output value is the
Submitted :  12/15/93  maximum of the field. If the ’Use Zero
Version :  1.000  Level’ flag is OFF, or there is no zero
Last Updated :  12/15/93  point within the range of the data,
Ported To :  Kubota  then all intermediate values are taken

at equal intervals within the data range. This is the default behaviour. If the ’Use Zero Level’ flag is ON and the data range includes zero, then one data level is explicitly set to zero and the remaining levels are partitioned between the positive and negative regions. This module is design to be used with the ISOBAND SLICE, ISOBAND BRICK, ISOLINE SLICE and ISOLINE BRICK modules.

Intervals UCD

Author :  Mike French  A set of N equally spaced data values
Evans and Sutherland UK  are taken from the range of an input
Number :  1759  ucd node data. The first output value
Language :  C  is the minimum of the field, the last
Submitted :  12/15/93  output value is the maximum of the
Version :  1.000  field. Any intermediate values are
Last Updated :  12/15/93  taken at equal intervals within the
Ported To :  Sun  data range. This module is design to be

used with the ISOBAND UCD and ISOLINE UCD modules. Other methods for selecting the data range could be added, e.g. logarithmic, cumulative frequency percentiles, interactive.

Iterate

Author :  Annette Walsh  The first module in a two module pair
Number :  1142  demonstrating looping in an AVS
Language :  FORTRAN  network. The combination of modules
Submitted :  03/03/92  ITERATE and STEPPER performs a very
Version :  1.000  simple iterative convergence
Last Updated :  03/03/92  calculation. Seed data comes into
Ported To :  IBM HP  ITERATE (see Figure 1). If any value in

the input field is less than or equal to the parameter “Constant”, the process has converged and the input field is placed on the “final result” output port. If the process has not converged, ITERATE sends the input data to STEPPER who decrements each value in the field by some amount. STEPPER sends its result back to ITERATE via an AVS upstream connection. ITERATE, having received data on its upstream input port, uses STEPPER’s last result and tests for convergence. Modules connected to the final result output port of ITERATE will only receive data if the convergence criteria is eventually met. NOTE - The upstream connection from STEPPER to ITERATE will be automatically established when the downstream connection is made from ITERATE to STEPPER.

LatLon UTM

Author :  Jeff Wang  This module allows users to do
MCNC, Environmetal  conversion between UTM, or Universal
Programs Group  Transver Mercator, coordinates and the
Number :  1721  regular Latitude/Longitude coordinates.
Language :  C  The Input Data is any of the following
Submitted :  09/30/93  - UTM Easting, UTM Northing UTM Zone,
Version :  1.000  Latitude, Longitude The output data is
Last Updated :  09/30/93  any of the above.

Ported To :

MINMAX 2 3

Author :  Kathy Pearson  The MINMAX module outputs floating
MCNC Environmental  point min and max values for setting
Programs  legend color scale endpoints. The
Number :  1804  values are those selected by the user
Language :  C  or the min-max of the input species
Submitted :  04/21/94  within the input field. Output occurs
Version :  1.000  only on the first invocation and when
Last Updated :  04/21/94  the values change to prevent continuous
Ported To :  DEC HP IBM Kubota Sun  updates of the color table associated

with the color legend. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

MOVING AVG

Author :  Kathy Pearson  The MOVING AVG mod is a filter that
MCNC Environmental  accepts a field 4D 3-coordinate
Programs  1-vector uniform float as input, and
Number :  1806  outputs the same type of field. The
Language :  C  module calculates a moving average of
Submitted :  04/21/94  the input field. The MCNC Environmental
Version :  1.000  Programs Group has created this module
Last Updated :  04/21/94  as part of a suite of AVS demo modules,
Ported To :  DEC HP IBM Kubota Sun  networks, and sample data sets to

demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

MathAVS

Author :  Cleo Borac  MathAVS establishes a data transfer
Number :  1628  communication through MathLink. Having
Language :  C  two modes of operation - ’send’ mode
Submitted :  07/02/93  and ’listen’ mode, Mathematica and AVS
Version :  2.000  may exchange data in turn. Note - the
Last Updated :  12/14/93  International AVS Center does not have
Ported To :  HP IBM Kubota Sun  Mathematica installed on the platforms

listed, so was unable to test out this module completely.

MinMax 2

Author :  Jeff Wang  This Module finds mininum and maxmium
MCNC, Environmetal  of a field. This is one of a suite of
Programs Group  modules we used to plot the emission
Number :  1735  data for the state of North Carolina.
Language :  C  This group of modules allows displaying
Submitted :  12/15/93  points, lines and polygons over a map
Version :  1.000  of the United States. The module suite
Last Updated :  12/15/93  is stored as a group in a compressed
Ported To :  Kubota Sun  tar file at the International AVS

Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Nvect to 2d

Author :  Ian Curington  Converts “field 1D uniform real
Advanced Visual Systems  N-vector“ to ”field 2D uniform real
Number :  1323  scalar“ by making the vector-length
Language :  FORTRAN  into the second dimension
Submitted :  08/31/92
Version :  2.000
Last Updated :  09/30/93
Ported To :  HP

Quantize Image

Author :  John Langner  Quantize Image converts an image with
Stardent Computer  24 bits/pixel to 256 or fewer colors.
Number :  1302  This is done with something resembling
Language :  C  the Heckbert mean cut algorithm.
Submitted :  07/11/92
Version :  1.000
Last Updated :  07/11/92
Ported To :  IBM Sun HP

Reset Legend

Author :  Jeff Wang  This Module recalculates the xoffset,
MCNC, Environmetal  yoffset and scale of the color legend
Programs Group  accoding to the size of the graph. This
Number :  1742  is one of a suite of modules we used to
Language :  C  plot the emission data for the state of
Submitted :  12/15/93  North Carolina. This group of modules
Version :  1.000  allows displaying points, lines and
Last Updated :  12/15/93  polygons over a map of the United
Ported To :  Kubota Sun  States. The module suite is stored as a

group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

SIDE BY SIDE

Author :  Terry Myerson  Side by side pastes input image 1 and
International AVS Center  input image 2 together, side by side,
(NCSC)  to produce the output image. image 1
Number :  1197  and image2 must be the same height. The
Language :  C  dimension of the output image is (width
Submitted :  05/06/92  of image 1 + width of image 2)*(height
Version :  1.000  of image 1). This was used to combine
Last Updated :  05/06/92  two images for comparison.
Ported To :  DEC IBM Sun Kubota HP

Shadow

Author :  Jeffrey A Thingvold  The shadow module generates a set of
National Center for  geometric shadows by taking the polygon
Supercomputing  descriptions of other modules and
Applications  replacing them with 2D versions
Number :  1182  projected onto the major axes as
Language :  C  coloured polytriangle shadows. The
Submitted :  05/04/92  shadow module interprets polytriangles,
Version :  1.000  polyhedra, meshes, but not spheres or
Last Updated :  05/04/92  lines. The orientation is optionally
Ported To :  DEC IBM Sun Kubota HP  controlled by three tri-state buttons,

where the shadow can be placed at the min or max axis position, or inhibited on a particular axis.

SlowDown

Author :  Steve Thorpe  This Filter AVS Module adds a time
International AVS Center  delay in the execution of an AVS
Number :  1723  network. Each time it is executed, it
Language :  C  uses up MilliSeconds worth of time
Submitted :  12/02/93  before passing the input integer to the
Version :  1.000  output integer. This can be useful to
Last Updated :  12/02/93  slow down an animation in order to more
Ported To :  Kubota Sun  closely examine its data.

Sobel operat

Author :  Anonymous  Uses the Sobel Operator for finding
Number :  1079  edges in a 2D byte field. Typical usage
Language :  C  is for finding edges in images prior to
Submitted :  10/10/91  some segmentation operation. The input
Version :  1.000  and output are both field 2d uniform
Last Updated :  10/10/91  scalar byte, and there are no
Ported To :  Kubota DEC IBM Sun HP  parameters. There is a sample network

described in the documentation accompanying the module.

Sphere

Author :  D. LE CORRE  This FILTER module maps a 2D input
CONVEX S.A  field around a sphere. If the input
Number :  1180  field is UNIFORM, the number of
Language :  C  elements in each direction give the
Submitted :  04/28/92  sphere subdivison. If the input field
Version :  3.000  is RECTILINEAR, the coordinates arrays
Last Updated :  12/15/93  are taken as longitude and latitude
Ported To :  Sun Kubota  (expressed in degrees). Thus, you will

only see part of the sphere if the positions do not span 360 degrees. Optionally, you can specify a 2D floating-point scalar field. This scalar value will be added to the sphere radius at each point in the input field and can be used to enter terrain information. Version 2.0 bug fix 6/93 by M.Wheeler of ULCC. Version 3.0 fixes a bug which caused crashes on Sun workstations, and also it now handles byte fields (4 times more compact).

Stepper

Author :  Annette Walsh  Is the second module in a two module
Number :  1141  pair demonstrating looping in an AVS
Language :  FORTRAN  network. The combination of modules
Submitted :  03/03/92  ITERATE and STEPPER performs a very
Version :  1.000  simple iterative convergence
Last Updated :  03/03/92  calculation. Seed data comes into
Ported To :  IBM HP  ITERATE (see Figure 1). A test is made

to see if the process has converged. If the process has not converged, ITERATE sends the input data to STEPPER who decrements each value in the field by a percentage (“percent/step”). STEPPER sends its result back to ITERATE via an AVS upstream connection. ITERATE, having received data on its upstream input port, uses STEPPER’s last result to again test for convergence. NOTE - The upstream connection from STEPPER to ITERATE will be automatically established when the downstream connection is made from ITERATE to STEPPER.

Toggle Black Whi

Author :  Steve Thorpe  Toggle Black White is a
International AVS Center  straightforward but useful AVS filter
Number :  1703  module. You can use this to change all
Language :  C  black pixels in an image to white, or
Submitted :  09/30/93  vica versa, depending on how you set a
Version :  1.000  boolean parameter.
Last Updated :  09/30/93
Ported To :  HP Sun

Topo2irreg

Author :  Wright, E.L.  This module converts a 2D scalar float
U.S Geological Survey  2-coordinate field to an irregular 3-D
Number :  1565  field by adding Z coordinates. The Z
Language :  C  coordinates are the data values of the
Submitted :  01/28/93  input field (typically topography)
Version :  1.000  multiplied by the value of the vertical
Last Updated :  01/28/93  scale factor parameter. The input field
Ported To :  DEC IBM Kubota HP  may be uniform, rectilinear, or

irregular.

FILTER MODULES

UcdColorizeGeom

Author :  Dominique Le Corre  This module takes as input a geometry,
Tethys S.A.  and colors all vertices according to a
Number :  1627  scalar Ucd component and a colormap. It
Language :  C  is similar to the colorize geom module
Submitted :  07/02/93  for fields.
Version :  1.000
Last Updated :  07/02/93
Ported To :  HP DEC Sun

UcdXYZtoFld

Author :  Dominique Le Corre  This module converts the nodes of a UCD
Tethys S.A.  structure to a 1D 3-space (scatter)
Number :  1619  field. Node data are kept into the
Language :  C  output field. Two integer dials (From
Submitted :  05/01/93  and To) can be used to limit the output
Version :  1.000  field to a subset of the input UCD.
Last Updated :  05/01/93
Ported To :  Kubota HP

Unjaggy

Author :  Steve Thorpe  This filter module slightly blurs a
International AVS Center  color image in order to reduce aliasing
Number :  1591  effects, which are also affectionately
Language :  C  known as “the jaggies”. It uses a 3x3
Submitted :  02/23/93  Bartlett window to filter an image. For
Version :  1.000  each (non-edge) red, green, blue and
Last Updated :  02/23/93  alpha component of a pixel in the input
Ported To :  DEC IBM Kubota HP  image, the resulting component in the

corresponding pixel in the output image is based on a weighted average.

WINDOW MATH

Author :  Kathy Pearson  WINDOW MATH – create a field that is a
MCNC Environmental  function of two input fields. The
Programs  WINDOW MATH module takes two sets of 2D
Number :  1820  input data previously sliced on the
Language :  C  specified “X”, “Y”, or “Z” input slice
Submitted :  04/21/94  axis 1/2. The output field contains
Version :  1.000  values that represent the chosen
Last Updated :  04/21/94  “window function” of the two input
Ported To :  DEC HP IBM Kubota Sun  fields. The MCNC Environmental Programs

Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

WIND MINMAX

Author :  Kathy Pearson  WIND MINMAX – output either user
MCNC Environmental  selected min max values for use with a
Programs  color legend or the min max values
Number :  1817  found within the input field of wind
Language :  C  data. The MCNC Environmental Programs
Submitted :  04/21/94  Group has created this module as part
Version :  1.000  of a suite of AVS demo modules,
Last Updated :  04/21/94  networks, and sample data sets to
Ported To :  DEC HP IBM Kubota Sun  demonstrate the functionality of its

visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

WIND SPEED

Author :  Kathy Pearson  WIND SPEED – create a vector for wind
MCNC Environmental  speed from u-v components. The WIND
Programs  SPEED module accepts as input a 4D
Number :  1819  field of 2-vector u-v wind components
Language :  C  and outputs the same information in a
Submitted :  04/21/94  field with an additional vector
Version :  1.000  specifying wind speed, r using the
Last Updated :  04/21/94  formula r = sqrt(u * u + v * v). The
Ported To :  DEC HP IBM Kubota Sun  output field labels are set to UWIND,

VWIND, WIND SPEED, and the units are all set to “m/s”. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

X Crop

Author :  Johan Wiklund  X Crop - This module is a ImageViewer
Number :  1901  Interface to crop it draws a rectange
Language :  C  to crop and the rectangle’s size can be
Submitted :  11/08/95  changed and the rect. can be translated
Version :  1.000
Last Updated :  11/08/95
Ported To :  DEC ultrix DEC alpha

X ray

Author :  Larry Gelberg  Does simple, orthogonal volume
Advanced Visual Systems  visualization on 3D uniform byte
Number :  1116  fields. Outputs a 2D integer field
Language :  C  which can be colorized and displayed as
Submitted :  01/07/92  an image. Essentially, for a given
Version :  1.000  direction, the module looks at the row
Last Updated :  01/07/92  of voxels “behind” each screen pixel
Ported To :  Kubota DEC IBM Sun HP  and, depending of the selected

operation, creates a new pixel based on those voxels. The processing options are SUM (each screen pixel is the sum of the stack of voxels), MEAN (each screen pixel is the sum of the stack of voxels divided by the number of voxels in each stack), MEDIAN (the stack of voxels is sorted by value and the screen pixel gets the center value in the sorted stack), MIN (the screen pixel gets the smallest value in the stack), and MAX (the screen pixel gets the largest value in the stack) This is a pretty fast volume visualization technique and you should be able to easily extend it to cover other cases.

Z slicer

Author :  Wright, E.L.  Z slicer interpolates a 2D irregular
U.S Geological Survey  field at a specified Z value from a 3D
Number :  1566  irregular field with stretched
Language :  C  (“sigma”) vertical coordinates. This is
Submitted :  01/28/93  a common need in 3D atmospheric and
Version :  2.000  oceanic models, where irregular fields
Last Updated :  09/19/94  have fixed X,Y locations for each
Ported To :  DEC IBM Kubota HP  vertical index, but Z values vary with

the thickness of the fluid layer.

aCNN

Author :  Stefano Brozzi, Andrea  The input image is first convolved with
Menozzi, Giulio E.  the control mask then added to the bias
Destri  and to the convolution of the internal
Universita’ degli Studi  state and the feedback mask, then a
di Parma, 43100 Parma,  function f is applied to this to get
Italy  the output image. The output image is
Number :  1872  also the future feedback in case that
Language :  C  more than a iteration is done.
Submitted :  09/06/95
Version :  1.000
Last Updated :  09/06/95
Ported To :  DEC ultrix DEC alpha

add 2 ints

Author :  Steve Thorpe  This short AVS module was written to
International AVS Center  add 2 integers together. When creating
Number :  1769  an animation using animate integer to
Language :  C  1) drive orthogonal slicer’s slice
Submitted :  02/03/94  plane, and 2) same integer to drive an
Version :  1.000  output frame name (as created using
Last Updated :  02/03/94  Terry Myerson’s animate file), I
Ported To :  HP Kubota Sun  wanted an offset to be added to the

file name’s frame number. This offset is provided by a parameter to this module.

add coords 2

Author :  Wes Bethel  Takes as input a 2d uniform field, then
Lawrence Berkeley  adds coordinate information to the
Laboratory  uniform field, “changing” it into an
Number :  1004  irregular field. The notion is that 3-d
Language :  C  coordinate information is assigned to
Submitted :  10/10/91  each input grid point of a 2d uniform
Version :  1.000  field. The user controls the range of
Last Updated :  10/10/91  the coordinates generated by changing
Ported To :  Kubota DEC IBM Sun HP  the values input parameters. There are

no restrictions on the size or range of the xmin, xmax values. A trilinear interpolating scheme is used.

add coords 3

Author :  Wes Bethel  Takes as input a 3d uniform field, then
Lawrence Berkeley  adds coordinate information to the
Laboratory  uniform field, “changing” it into
Number :  1005  either an irregular or rectilinear
Language :  C  field. The notion is that 3-d
Submitted :  10/10/91  coordinate information is assigned to
Version :  1.000  each input grid point. The user
Last Updated :  10/10/91  controls the range of the coordinates
Ported To :  Kubota DEC IBM Sun HP  generated by changing the values of the

dials, xmin, xmax, etc. There are no restrictions on the size or range of the xmin, xmax values.

add cps

Author :  Wes Bethel  Takes as input an irregular 3d
Lawrence Berkeley  3-coordinate scalar field, processing
Laboratory  it using the catmull-rom spline
Number :  1006  formulation, and produces a field
Language :  C  suitable for consumption by the bezier
Submitted :  10/10/91  volume curvilinear volume renderer. The
Version :  1.000  input field consists of a 3d irregular,
Last Updated :  10/10/91  3-coordinate floating point scalar
Ported To :  Kubota DEC IBM Sun HP  field. From the input field, a “voxel”

is defined by the eight vertices which surround it. For each “voxel”, interior control points are computed which define a tricubic bezier solid. Thus, if N points (on a given data axis) are input, this module will produce 3N-2 output points.

add extents

Author :  Steve Larkin  This module copies the input field to
Computer Graphics Unit,  the output field whilst recalculating
University of  the extent information for the data and
Manchester, UK  coordinates supplied in the input
Number :  1602  field. All extra field information is
Language :  C  copied to the output field e.g.,
Submitted :  03/23/93  labels, units, data, coordinates, type
Version :  1.000  The module will accept any type of
Last Updated :  03/26/93  field, with N dimensions and M data
Ported To :  Sun DEC IBM HP  components per data element. This

module is useful when modules further up a network produce fields which do not have the extent information added. EXAMPLE NETWORKS

adjust rgb

Author :  David Haith  Filters image to “tint” or modify color
Power-Media  balance. Each component in RGB space
Number :  1007  can be altered, added or subtracted to
Language :  C  adjust for overall effect. This takes
Submitted :  10/10/91  the place of many other modules,
Version :  1.000  combining “tint” into one operation.
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

alpha mix

Author :  David Haith  Places one image over another and
Power-Media  graduates according to the alpha
Number :  1008  channel contents. Where the alpha
Language :  C  channel permits, blends the images with
Submitted :  10/10/91  the “mix” control.
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

and or merge

Author :  A. Jacobson  And/or merge uses a variety of boolean
UCLA Dept. Pharmacology operators to merge 2 or 3 images, with
/ Crump Institute  2 or 3D ROI (with 2 source images).
Number :  1859  This module can be used to replace the
Language :  C  use of several instances of field math
Submitted :  10/17/94  ip threshold, ip rescale and similar
Version :  1.000  modules. It allows facile merging of
Last Updated :  10/17/94  datasets with output values segmented
Ported To :  DEC alpha HP IBM Sun  to single values for ease of

differentiation when displayed. This is actually 4 modules rolled into 1 to save disk and palette space. The routines are all similar and are only a subset of useful merge possibilities. All data inputs must have the same extents and dimensions. However the routines that take an ROI will take either a 2D ROI, applying the ROI to each “Z” plane if 3D data fields are used, or a 3D ROI, applying each point of the ROI field to the data fields. The ROI field must have the same dimensions as the data inputs, either in x,y (if 2D ROI) or x,y,z (if 3D). In either case, non-zero values in the ROI field are considered “within” the ROI. These 4 routines are described in much greater detail in the documentation accompanying the source code.

arith binary

Author :  International AVS Center  arith binary is an avs module
North Carolina  executable which contains 6 modules
Supercomputing Center  hacked from khoros. Binary Arithmetic
Number :  1500  Operations. ** DEC vabsdiff ** computes
Language :  C  the absolute difference between two
Submitted :  10/01/92  images ** vadd ** adds two images pixel
Version :  0.100  by pixel ** vblend ** forms composite
Last Updated :  10/01/92  image by adding X*(img1 data) to
Ported To :  Sun DEC IBM  (1-X)*(img2data) ** vdiv ** divides two

images pixel by pixel ** vmul ** multiplies two images pixel by pixel ** vsub ** subtracts two images (img1 = img1-img2) pixel by pixel

arith logical

Author :  International AVS Center  arith logical is an avs module
North Carolina  executable which contains 6 modules
Supercomputing Center  hacked from khoros. Logical Operations.
Number :  1501  ** vand ** logically AND two images
Language :  C  pixel by pixel ** vlshift ** shifts the
Submitted :  10/01/92  pixel data in an image left s bit
Version :  0.100  positions ** vor ** performs a logical
Last Updated :  10/01/92  OR between two images ** vreplace **
Ported To :  Sun DEC IBM  replaces pixels of first image with

pixels of second image ** vrshift ** shifts the pixel data in an image right s bit positions ** vxor ** exclusively OR two images pixel by pixel

arith unary

Author :  International AVS Center  arith unary is an avs module
North Carolina  executable which contains 13 modules
Supercomputing Center  hacked from khoros. Unary Arithmetic
Number :  1502  Operations. ** DEC vabs ** computes
Language :  C  absolute value of image ** vclip **
Submitted :  10/01/92  limits range of values that grey levels
Version :  0.100  may occupy ** vconj ** computes the
Last Updated :  10/01/92  complex conjugate of the input image **
Ported To :  Sun DEC IBM  vexp ** computes the base 10

antilogarithm of an image pixel by pixel ** vfloor ** clips the lower levels of the input image to the specified value ** vinvert ** inverts the gray-levels of a byte type image ** vlog ** DEC compliments image by subtracting pixel values from const. level ** vnormal ** normalizes an image ** vnot ** comliments an image by subtracting pixel values from constant level **voffset ** adds a specified offset to each pixel in the input image ** vscale ** multiplies an image by a constant ** vsqrt ** computes the square root of an image pixel by pixel ** vsubstit ** computes the EROSION of a binary image by a structuring element

arithmetic 2

Author :  Anonymous  Combines two fields of integer data,
Stardent Computer  and outputs a new composite field. The
Number :  1014  two input fields are arithmetically
Language :  C  combined, using one of several
Submitted :  10/10/91  operators, selected by setting one of
Version :  1.000  several choice buttons. Both input
Last Updated :  10/10/91  fields must be the same size, but need
Ported To :  Kubota DEC IBM Sun HP  not be the same dimension. The output

field is the same size and type as the input. This module is used to compare similar scalar fields, for instance the difference between the red and green channels of an image, subtract two volumes and visualize the difference, or normalize field data by removal of static background field. This module is open ended, many more buttons and types of operators could easily be added.

backgroundRGB

Author :  John Stephen  backgroundRGB changes the background
Tessella Support  colour of the current camera as set by
Services  3 red - green - blue dials.
Number :  1639
Language :  C
Submitted :  09/24/93
Version :  1.000
Last Updated :  09/24/93
Ported To :  HP IBM Sun

byte image t

Author :  Wes Bethel  Converts an image contained in a byte
Lawrence Berkeley  field to one contained in a floating
Laboratory  point field. This is useful to in
Number :  1126  processing images in which the
Language :  C  algorithms operate on floating point
Submitted :  01/26/92  representations of images. Use of the
Version :  1.000  AVS-supplied converters (field to byte,
Last Updated :  01/26/92  field to float) will produce incorrect
Ported To :  Kubota DEC IBM Sun HP  results. The range of input pixels is 0

¡= value ¡= 255.

classify

Author :  International AVS Center  classify is an avs module executable
North Carolina  which contains 9 modules hacked from
Supercomputing Center  khoros. Classification/Clustering
Number :  1503  Algorithms. ** DEC lrfclass **
Language :  C  classifies an image using the Localized
Submitted :  10/01/92  Receptive Field (LRF) ** lrftrain **
Version :  0.100  trains on an image for the weights used
Last Updated :  10/01/92  with the LRF ** viso2 ** converts image
Ported To :  Sun DEC IBM  into vectors and performs

iso2clustering ** vkmeans ** converts image into vectors and performs kmeans clustering ** vlabel ** performs a labeling in a multiband image or a cluster image ** vmindis ** a simple minimum distance classifier ** vqerr ** computes RMS quantization error with two images ** vquant ** performs N-dimensional vector quantization ** vwmdd ** an Weighted minimum distance detector

closing

Author :  David E. Harris  This module performs the mathematic
Georgia Institute of  morphological closing operation on
Technology  grayscale (or binary) images. See
Number :  1172  Haralick et al. “Image Analysis using
Language :  C  Mathematical Morphology“, IEEE Trans.
Submitted :  04/16/92  on Pattern Analysis and Machine
Version :  1.000  Intelligence, PAMI-9, No. 4, pp.
Last Updated :  04/16/92  532-550, 1987 for a more complete
Ported To :  Kubota  discription. The closing of image A by

structuring element B (a binary mask) is defined as the erosion (by B) of the dilation of A by B. This is done by feeding the output of the DILATE module into the input of the ERODE module and using the same mask for both. The CLOSING module just combines these two modules for convenience. See also the OPENING module.

collage

Author :  Wes Bethel  Copies two images into an output image.
Lawrence Berkeley  The user can control the position of
Laboratory  each of the input images within the
Number :  1020  output image. The portion of the output
Language :  C  image which is not covered by either of
Submitted :  10/10/91  the input images is set to the value of
Version :  2.000  the Gray Fill dial. This was originally
Last Updated :  10/17/94  part of the HIPS image processing
Ported To :  Kubota DEC IBM Sun HP  distribution. This particular version

has been written from scratch to allow for images of varying format. That is, the input images may be grayscale format; single byte, ARGB format; four bytes per pixel, complex format; two bytes per pixel. Version 2 (also by Wes Bethel) provides bug avoidance whereby a percent character in a parm name string caused AVS to dump core.

collapse

Author :  James Pipe  Reduces an N-dimensional data set (1 ¡
University of Michigan  N ¡ 5) to an N-1 dimensional data set
Medical School  by taking the (min—max—mean—std) of the
Number :  1607  data across dimension E (the dimension
Language :  FORTRAN  to be eliminated). This reduction can
Submitted :  04/10/93  take place across the entire span of
Version :  1.000  dimension E, or across some subset of
Last Updated :  04/10/93  it (as specified by the user).
Ported To :  DEC IBM Kubota

collapse 2 num

Author :  James Pipe  Reduces entire data set to a single
University of Michigan  floating point number. This number is
Medical School  either the minimum, maximum, mean, sum,
Number :  1608  standard deviation (std), or count (n)
Language :  FORTRAN  of the input data set. This module will
Submitted :  04/10/93  only work for 4-byte real and 4-byte
Version :  1.000  integer input data types.
Last Updated :  04/10/93
Ported To :  IBM

INTERNATIONAL AVS CENTER MODULE CATALOG

colorizer 2d

Author :  Wes Bethel  Similar to the “colorizer” module in
Lawrence Berkeley  function, this module will take scalar
Laboratory  data along with a colormap, and produce
Number :  1587  color data. Whereas the “colorizer”
Language :  C  module takes a single scalar field and
Submitted :  02/11/93  a single colormap, “colorizer 2d” takes
Version :  1.000  TWO scalar fields (of the same
Last Updated :  02/11/93  dimensions) and TWO colormaps. The
Ported To :  Sun  resulting color vector produced is a

function of the two scalar values and the two colormaps.

colour cones

Author :  Mark Pinches,  The ccone module transforms an AVS
Southampton Parallel  geometry, replacing a set of disjoint
Applications Centre, UK  lines with “cones” constructed out of
and Ian Curington, AVS  eight polygons. The cone tristrip is
Inc.  now upto 2x more efficient, and has a
Number :  1640  subdivision slider control. If the
Language :  C  input has values the module will make
Submitted :  09/24/93  colour coded cones based on the input
Version :  1.000  lines and colormap. Min and Max
Last Updated :  09/24/93  controls allow lines of only certain
Ported To :  HP IBM Kubota Sun  bounded length to have cones, shorter

or longer lines do not get cones.

combine

Author :  Terry Myerson  This module inputs three scalar arrays,
International AVS Center  and uses them to determine the
(NCSC)  coordinates of the output field. No
Number :  1207  data is copied over into the output
Language :  C  field. This module in coordination with
Submitted :  05/13/92  delay can produce fractal systems to be
Version :  1.000  visualized with scatter dots.
Last Updated :  05/13/92
Ported To :  DEC IBM Kubota HP

combine vecto 2

Author :  James Pipe  This module is an extension of the
University of Michigan  supported module ’combine scalars’.
Medical School  This module combines a number of input
Number :  1609  data sets into a single output data set
Language :  FORTRAN  with one vector component of the output
Submitted :  04/10/93  data set corresponding to each vector
Version :  1.000  component of each input data set. Note
Last Updated :  04/10/93  that all of the input data sets MUST
Ported To :  IBM  have same dimensions and data types.

This is different from ’combine scalars’ in the following ways - 1) The input data need not be scalar. For example, a 1-vector, 3-vector, and 2-vector data sets may be combined into a 6-vector output data set. 2) The total vector length of the output data set may be have up to 12 elements. Note there are 12 input ports, so that up to 12 scalar input data sets may be combined. If one or more of the input data sets are an N-vector (N ¿ 1), then not all of the input ports may be used. The limiting factor here is an output data set which is at MOST a 12-vector.

3) You may add labels to each vector component to ’keep track’ of the output data’s vector elements downstream. 4) The vector length of the output data set is automatically set to be the sum of the vector lengths of all of the input data sets.

combine vectors

Author :  Earl W. Hood  The Combine Vectors modules combines up
Convex Computer  to 4 vector and/or scalar fields into
Corporation  one vector field. The input ports of
Number :  1554  the module are processed from
Language :  C  right-to-left. The resulting output
Submitted :  11/16/92  field will have a vector length equal
Version :  1.000  to the sum of all the vector lengths of
Last Updated :  11/16/92  the input field. Any labels and units
Ported To :  IBM  will be carried over to the output

field. Written because I got tired of using a bunch of extract scalar modules when I wanted to process a vector component separately, and then use the combine scalars module. (eg. process the R,G,B components of an image together, and then combine the R,G,B components back with the Alpha channel).

complex math

Author :  Earl W. Hood  The complex math module performs
Convex Computer  binary math operations on 2-vector
Corporation  fields that represent complex numbers.
Number :  1558  The first vector component represents
Language :  C  the real part of the complex number.
Submitted :  11/16/92  The second vector component represents
Version :  1.000  the imaginary part of the complex
Last Updated :  11/16/92  number.
Ported To :  IBM

compute shade

Author :  Robert Mazaika and Larry This module combines the functions of
Gelberg  the colorizer, compute gradient, and
Advanced Visual Systems  gradient shade modules into a single,
Number :  1282  memory efficient module. These modules
Language :  C  are used primarily to make shaded,
Submitted :  06/03/92  ray-traced images. The problem is that
Version :  1.000  they are highly inefficient in terms of
Last Updated :  06/03/92  memory allocation colorizer takes in a
Ported To :  Sun Kubota HP  1 byte per voxel and outputs 4 bytes

per voxel compute gradient takes in 1 byte per voxel and outputs 12 (3 floats) gradient shade outputs 4 bytes per voxel These three modules produce 20 bytes for every input data set byte. It is for this reason that some people have experienced problems trying to render ray castings of large data sets

- the tracing code is pretty computationally efficient - it’s just that most of the system resources go to swapping data, rather than computing the image. The COMPUTE SHADE module does gradient computation, colorizing, and shading on a per slice basis. All in all, it takes less time than running the original three modules in sequence. However, it does take longer than running only gradient shade alone.

concat str

Author :  Urs Meyer  The “concat strings” module will
Swiss Scientific  concatenate two or more strings
Computing Center  together according to the conversion
Number :  1622  specification string given. The format
Language :  C  string parameter is a character string
Submitted :  07/02/93  that controls the concatenation of the
Version :  1.000  string inputs. The conversion
Last Updated :  07/02/93  specification is the same as for
Ported To :  Kubota HP IBM Sun  printf.

cone

Author :  SH Lamson  This module replaces a set of disjoint
General Electric  lines in an AVS geometry with “cones”
Corporate Research and  constructed out of eight polygons. The
Development  user can control the radius of the
Number :  1021  cone’s base. Only values in the range
Language :  C  0.0 - 0.4 produce an acceptable result.
Submitted :  10/10/91  The user can also use the “threshold”
Version :  1.000  parameter to suppress cones for line
Last Updated :  10/10/91  segments that are shorter than the
Ported To :  Kubota DEC IBM Sun  threshold. This is useful when ignoring

the output of the direction of a vector field in regions where the magnitude is small enough as to make the field “negligible” in some sense.

cone 2

Author :  Georg Thallinger,  The cone module transforms an AVS
Joanneum Research Graz, geometry, replacing a set of lines with
Austria  “cones” constructed out of eight
and Scott Lamson,  polygons. The cones are colored based
General Electri  on the colors of the input geomtry.
Number :  1697  Everytime the input geometry changes,
Language :  C  cone tries to find a suitable radius by
Submitted :  09/30/93  calculating the average length of a
Version :  1.000  line segment. The min and max values of
Last Updated :  09/30/93  radius are set to 1/10*r and 10*r
Ported To :  IBM Sun  (where r is the calculated radius).

contour perc

Author :  Terry Myerson  This module inputs a floating point
International AVS Center  number and spits out 10%, 20%, 30%,
(NCSC)  40%, 50%, 60%,7 0%, 80%, and 90%
Number :  1110  of that number. It was originally used
Language :  C  for contour plots.
Submitted :  12/05/91
Version :  1.000
Last Updated :  12/05/91
Ported To :  Kubota DEC IBM Sun HP

convert color

Author :  International AVS Center  convert color is an avs module
North Carolina  executable which contains 6 modules
Supercomputing Center  hacked from khoros. Color
Number :  1504  Transformations. ** vcltrans
Language :  C  **specifies the linear transformation
Submitted :  10/01/92  matricies ** vcluv ** converts to
Version :  0.100  cieucs Luv color spaces ** vcmtrans **
Last Updated :  10/01/92  performs linear color space conversions
Ported To :  Sun DEC IBM  with a 3x3 matrix ** vcrgbhls **

conver6ts from RGB to HLS and vice versa ** vcrgbhsv ** converts from RGB to HSV and vice versa ** vgamut ** takes a 3 band image and generates a pseudo color image

convert data

Author :  International AVS Center  convert data is an avs module
North Carolina  executable which contains 10 modules
Supercomputing Center  hacked from khoros. Data Format
Number :  1505  Conversion. ** DEC vband2vect **
Language :  C  reformats data stored in a band type
Submitted :  10/01/92  image format ** vbandcomb ** combines
Version :  0.100  all bands from m images into one image
Last Updated :  10/01/92  ** vbandspt1 ** takes band from
Ported To :  Sun DEC IBM  multiband image and creates image

w/only that band ** vbandspt3 ** extracts 3 data bands from a multiband image ** DEC vctor ** performs simple mappings of COMPLEX data to REAL form ** vect2band ** rearranges the data stored in vector format to band format ** vect2row ** reformats data so that each vector is represented as a row ** DEC vmos2band ** converts mosaic image into image where subimage makes up a band ** vrect2pol ** converts data file from rectangular to polar coord. and vice versa ** vrtoc ** converts REAL data to COMPLEX data

crop2

Author :  John Tee  CROP2 crops 2-D fields by selecting a
AVS Inc. Consultant  window of the input field. It differs
Number :  1646  from CROP in two ways. The more obvious
Language :  C  difference is that the output field is
Submitted :  09/24/93  defined by the position of a centre
Version :  1.000  pixel, and a field width. The limits of
Last Updated :  09/24/93  the parameters are interdependent, and
Ported To :  HP IBM Kubota Sun  so are updated according to the current

selections. The other reason for the creation of this module is that CROP limits the minimum dimension to 2. This is unacceptable when a 2-D field is required having one dimension set to 1. (ORTHOGONAL SLICER will produce the correct data, but reduces the dimensionality of the output field, which makes it unacceptable as input to many modules.)

cylinder 2

Author :  SH Lamson, General  cylinder converts lines to cylindrical
Electric C RandD  tubes. It transforms an AVS geometry
and Ian Curington,  data type, replacing a set of disjoint
Advanced Visual Systems,  lines with “cylinders” constructed out
Inc  of eight polygons. Unlike “tube”, if
Number :  1714  the input lines have colors, these are
Language :  C  passed through as polygon colors. The
Submitted :  09/30/93  number of polygons representing the
Version :  1.000  cylinder is controlled by the
Last Updated :  09/30/93  resolution slider, from a 2D polygon
Ported To :  HP IBM Kubota Sun  strip to a prism tube, to square, up to

a 20 sided cylinder.

delay

Author :  Terry Myerson  Pushes a 1-D array around its bounds,
International AVS Center  delaying the occurrence of a data value
(NCSC)  until a later point in the array. This
Number :  1138  was initially used with a business
Language :  C  visualization prototype, trying to find
Submitted :  02/26/92  patterns within a linear data set. A
Version :  1.000  sister module named combine was used to
Last Updated :  02/26/92  combine the linear data sets into a 0-D
Ported To :  Kubota DEC IBM Sun HP  3-space field to visualize.uses a 1-D

array around its bounds, delaying the occurrence of a data value until a later point in the array.

demultiplex

Author :  John Tee  DEMULTIPLEX converts a line-interleaved
AVS Inc. Consultant  scalar field to a vector field. This
Number :  1649  module is required when importing data
Language :  C  which is interleaved on a row-by-row
Submitted :  09/29/93  basis from a foreign format file using
Version :  1.000  FILE DESCRIPTOR or DATA DICTIONARY. The
Last Updated :  09/29/93  constant stride variable provided in
Ported To :  HP IBM Kubota Sun  the preceding modules does not permit

line-interleaved data to be decoded.

dilate

Author :  David E. Harris  “Expands” a grayscale (or binary) image
Georgia Institute of  by a structural mask. For binary
Technology  images, this is an OR operation, for
Number :  1171  grayscale images it is a neighborhood
Language :  C  maximum operator. Dilate will tend to
Submitted :  04/16/92  connect islands and fill-in holes.
Version :  1.000
Last Updated :  04/16/92
Ported To :  Kubota

downsize points

Author :  Richard Albrecht  The module divides the max points array
Max-Planck-Institute for  by downsizeval. It is assumed that the
Biochemistry  min points are zero. Now it is possible
Number :  1724  to use ’downsize’ in a correct way with
Language :  C  the ’tracer’. ’Tracer’ uses the
Submitted :  12/02/93  dims-array for adjusting the view-port.
Version :  1.000  In connection with ’downsize’, ’tracer’
Last Updated :  12/02/93  enlarges the output by the
Ported To :  Kubota Sun  downsizevalue (due to the unchanged

points-array). We use this mainly for adjusting the viewing angle and the colors of the ’tracer’ output at high ’downsize’ values ( 4 ). The final calculation is done with 1. Our datasource is a confocal microscope and our normal data size is 256*256*100.

draw pixel

Author :  Ian Curington  Draws pixels into an image. Uses
Stardent Computer  upstream transform information to draw
Number :  1028  pixels over an input image (paint). It
Language :  C  is designed to take an input image, and
Submitted :  10/10/91  copy it to the output port. This is a
Version :  1.000  straight pass-through. If, however,
Last Updated :  10/10/91  once and image is passed, it can
Ported To :  Kubota DEC IBM Sun HP  scribble on top of the image with a

square-block brush in light grey. The mouse (brush) position is fed externally from this module, getting the current mouse position and button state from the “struct upstream transform“ input port. This is designed to be used with “track cursor”, which displays the image and catches the X events.

dual control

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology four-module suite - 1) dual control -
/ Crump Institute  to control flow of field and integers
Number :  1774  to downstream modules, 2) field control
Language :  C  - to control flow of an AVS image to
Submitted :  02/09/94  downstream modules, 3) integer control
Version :  1.000  - to control flow of an integer to
Last Updated :  02/09/94  downstream modules, and 4) real control
Ported To :  HP Kubota Sun  - to control flow of an real to

downstream modules. Dual control will be useful in keeping the flow of data controlable to modules that take integers and fields and fire when either input port (or parameter) is changed. This allows the user the flexibility to update either of the module inputs without the module firing. The only other way to do this with continuing changes would be to connect and disconnect either or both input ports or stop the network flow executive. This module is a combination of the field control and integer control modules. See the documentation for these modules for further details. An example of where this module will be useful is with modules that overwrite files everytime their input is updated. Dual control can be used to only triger such a module as needed.

erode

Author :  David E. Harris  The dual of dilation. It can be thought
Georgia Institute of  of as doing to the background what
Technology  dilate does to the foreground. For
Number :  1170  binary images, this is an AND
Language :  C  operation, for a grayscale image is it
Submitted :  04/16/92  a neighborhood minimum operatior. Erode
Version :  1.000  will remove islands that are smaller
Last Updated :  04/16/92  than the structuring mask.
Ported To :  Kubota

INTERNATIONAL AVS CENTER MODULE CATALOG

excavate

Author :  Anonymous  Copies a volume field from input to
Stardent Computer  output, while zeroing out all data
Number :  1031  within a “corner”, as defined by the
Language :  C  x,y,z parameters. This allows other
Submitted :  10/10/91  mappers to “see” into the volume, to
Version :  1.000  see data values along the interface of
Last Updated :  10/10/91  the “cut”. Typical uses would be
Ported To :  Kubota DEC IBM Sun HP  “tracer” or threshold slice.

exponential

Author :  Earl W. Hood  The exponential filter module computes
Convex Computer  ’e’ raised to each field value, i.e.
Corporation  newvalue = ev̂alue
Number :  1559
Language :  C
Submitted :  11/16/92
Version :  1.000
Last Updated :  11/16/92
Ported To :  IBM

extend

Author :  James Pipe  Turns an N dimensional data set into an
University of Michigan  N+1 dimensional data set. Does this by
Medical School  copying the N-dimensional input data
Number :  1611  set m times across the N+1th dimension,
Language :  FORTRAN  where m (the size of the added
Submitted :  04/10/93  dimension) is specified by the user.
Version :  1.000
Last Updated :  04/10/93
Ported To :  IBM

extrude field

Author :  Earl W. Hood  Extrude field generates a 3D volume by
Convex Computer  applying a transformation matrix
Corporation  defined by the parameters to a 2D
Number :  1555  field. i.e. it “sweeps” a 2D field
Language :  C  through 3-space to generate a 3D
Submitted :  11/16/92  (irreguler) field. There were some
Version :  1.000  researchers who had data sets that
Last Updated :  11/16/92  represented a slice of something
Ported To :  IBM HP  axisymetric in 3-space. With

extrude field, one can recreate the 3D representation. This is also good for people who have 2D data sets, but the modules that they would like to use only accept 3D data sets.

feature

Author :  International AVS Center  feature is an avs module executable
North Carolina  which contains 8 modules hacked from
Supercomputing Center  khoros. Feature Extraction. ** vfractal
Number :  1506  ** performs a fractal feature
Language :  C  extraction of an image ** vpml **
Submitted :  10/01/92  estimates the fractal dimension of an
Version :  0.100  image ** vpolygon ** creates a vector
Last Updated :  10/01/92  image resulting of a linear
Ported To :  Sun DEC IBM  approximation ** vrmatch ** analyzes

two input images that are supposed to be similar ** vshape ** performs a shape analysis on two different types of images ** vspatial ** performs a spatial feature extraction on an image ** vstereo ** analyzes image produced by vrmatch and produces image ** vtexture ** extracts texture features from an image

fft2d

Author :  David E. Harris  This module takes the forward two
Georgia Tech Research  dimensional fast Fourier transform of a
Institute  square array whose dimensions are
Number :  1202  integer powers of two. As a practical
Language :  C  matter, realize that the result is a
Submitted :  05/11/92  field with twice the number of values
Version :  1.000  (real and imaginary components) and
Last Updated :  05/11/92  each is a floating point number. The
Ported To :  Kubota  second output is a grayscale byte field

which represents the log(1+magnitude) of the FFT. The log(1+magnitude) plot shows more detail than a plot of just the magnitude. The companion module to this one is the inverse FFT module ifft2D.

field 2D to 3D

Author :  Chris Allen  Field 2D to 3D converts its input field
Unisys Corporation  from 2 to 3 dimensions, and sets the
Number :  1292  physical-space dimensions to 3-space.
Language :  C  The data remains unchanged. This module
Submitted :  07/11/92  is useful, for example, if you have a
Version :  1.000  3-vector 2D field that you would like
Last Updated :  07/11/92  to view with Hedgehog, which requires
Ported To :  IBM Sun HP  field 3D uniform data.

field averaging

Author :  Earl W. Hood  The Field Averaging module computes the
Convex Computer  average field of a series of multiple
Corporation  fields. The module is useful if there
Number :  1556  exists multiple fields (eg. images)
Language :  C  that represent a series of samplings of
Submitted :  11/16/92  a given object, and an average of the
Version :  1.000  samples will help cancel out the noisy
Last Updated :  11/16/92  data contained in each sample.
Ported To :  IBM

field control

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology four-module suite - 1) dual control -
/ Crump Institute  to control flow of field and integers
Number :  1775  to downstream modules, 2) field control
Language :  C  - to control flow of an AVS image to
Submitted :  02/09/94  downstream modules, 3) integer control
Version :  1.000  - to control flow of an integer to
Last Updated :  02/09/94  downstream modules, and 4) real control
Ported To :  HP Kubota Sun  - to control flow of an real to

downstream modules. Field control will be useful in keeping the flow of data controllable to bottleneck points without requiring the user to repeatedly connect and disconnect modules to control their execution. This module allows the user to control the flow of an AVS field to downstream modules. Often there will be modules, (such as tracer), macros or whole branches of a network that take a long time to execute and may be undesireable in any given run of data through the network. Without this module, the only way to activate or deactivate such a module or branch would be to connect or disconnect it’s input, suspend the flow executive of the whole network, or to deactivate it through its editor dialog. Another example of where this module will be useful is with modules that overwrite files everytime their input is updated. Field control can be used to only triger such a module as needed.

field conver

Author :  Anonymous  Converts an input uniform field into a
Convex Computer  rectilinear or irregular field, based
Number :  1149  on the Field Type parameter.
Language :  C  Non-uniform fields are passed through
Submitted :  03/10/92  unchanged. The output field has the
Version :  1.000  same dimensionality as the input field,
Last Updated :  03/10/92  but is non-uniform. This module has no
Ported To :  Kubota DEC IBM Sun HP  effect on non-uniform input fields.

field label

Author :  anonymous  This module allows the user to modify
Number :  1890  the labels of a field’s components.
Language :  C  When the field is initially connected,
Submitted :  09/07/95  the label parameters will be set to the
Version :  1.000  original values in the data field. The
Last Updated :  09/07/95  the user should set the ’Apply typed in
Ported To :  DEC ultrix DEC alpha  labels’ choice and modify the labels to

their new values. If the ’Apply typed in labels’ parameter is not set, then each time the field is updated upstream of the field label module, any typed in labels will be erased and replaced with the original field values.

field mux

Author :  Ian Curington  Given several (different size or types
Advanced Visual Systems  of) input fields, one is passed to the
Number :  1747  output port, given the choice
Language :  C  selection. All the data from the input
Submitted :  12/15/93  is passed to the output. All 10 input
Version :  1.000  ports are optional, it will work with
Last Updated :  12/15/93  any subset, or just one input. Inputs
Ported To :  Kubota Sun  can be different dimension, different

data types, etc. By default, the names of the input ports on the choice list are “One, Two...” but a string input port can be used to set the choice names from an external procedure. This is an efficient module, as it only passes a pointer, and does not copy data. ADVANCED TECHNIQUE - This routine contains an advanced module writer trick. It does not make a copy of the data. It simply passes a pointer across to form the correct multiplex connection.

field scatte

Author :  Michael Rangitsch  Converts a 3 dimensional field into a 1
Number :  1147  dimensional field and uses the mapping
Language :  C  information of the original field to
Submitted :  03/10/92  compute the 3d mapping space. This
Version :  1.000  works for any type of mapping in the
Last Updated :  03/10/92  original field (uniform, rectilinear or
Ported To :  Kubota DEC IBM Sun HP  irregular). This module is useful in

converting 3d data into a form useable by a ’bubble’ type of visualization module.

field to float parm

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1784  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to int parm,
Last Updated :  02/15/94  field to string parm,
Ported To :  IBM Sun  float parm to field,

int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

field to glo 2

Author :  Raymond E. Flanery Jr.,  This module is used to wrap field data
Director  around a globe. It does a pretty good
Advanced Visualization  job of allowing the user to specify
Research Center, ORNL  where on the globe to put the corners
Number :  1902  of the input field or to use the points
Language :  C  data of an input irregular field as the
Submitted :  11/09/95  longitude and latitude of each point.
Version :  1.000  If the data is 3-D, then concentric
Last Updated :  11/09/95  globes are generated with the radius
Ported To :  DEC ultrix Sun  determined by the “radius min” and
DEC alpha  “radius max” parameters, unless the

data is 3-space and the “use height data“ toggle is on. Scalar fields can also “use height data” as an elevation map, and 2-vector or 3-vector fields can have their data converted from polar-coordinate direction vectors to cartesian vectors.

field to globe

Author :  Ed Bender  This module is used to wrap field data
Convex Computer  around a globe. It does a pretty good
Corporation  job of allowing the user to specify
Number :  1595  where on the globe to put the corners
Language :  C  of the input field or to use the points
Submitted :  03/04/93  data of an input irregular field as the
Version :  3.000  longitude and latitude of each point.
Last Updated :  04/21/94  If the data is 3-D, then concentric
Ported To :  Sun IBM HP  globes are generated with the radius

determined by the “radius min” and “radius max” parameters, unless the data is 3-space and the “use height data“ toggle is on. Version 2 enhancements were made to this module by Scott Ferguson of Cray Computer Corporation, Colorado Springs, CO - Inner loops were restructured to allow vectorization. Also, for scalar input fields, the scalar value can be used as elevation data. And, for 2-vector or 3-vector input fields, the vector data can be converted from polar to cartesian coordinates as well. For example, a model of wind vectors in a global climate model needed to be overlayed on a globe. By selecting “Convert Data”, the data can later be passed into hedgehog to show the arrows wrapped around a sphere. A parameter called “Elev Scale” was added to scale the elevation data when using a scalar input field as an elevation map. Version 3 enhancements were made by Dr. Joachim Biercamp of the German Climate Computer Center. The module now updates the extent field, so downstream modules like hedgehog or sampler don’t get confused

field to int parm

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1785  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

field to irr

Author :  Kathleen Dyer  Converts fields to irregular fields by
Lawrence Livermore  adding grid or coordinate information.
National Laboratory  This is useful for modules which will
Number :  1118  not process uniform or rectilinear
Language :  C  fields. The input must be a 1D, 2D or
Submitted :  01/22/92  3D uniform field, but it may have any
Version :  2.000  data type at each field point. The
Last Updated :  03/04/93  output is a field with the same data,
Ported To :  DEC IBM Sun HP  shape and dimensionality as the

original input field. It is irregular. Version 2.0 update (by Karin Schaefer) allows this module to convert uniform or rectilinear fields to irregular fields by adding a grid.

field to rec

Author :  Kathleen Dyer  Converts fields to rectilinear fields
Lawrence Livermore  by adding grid or coordinate
National Laboratory  information. This is useful for modules
Number :  1119  which will not process uniform or
Language :  C  irregular fields. The input must be a
Submitted :  01/22/92  1D, 2D or 3D uniform field, but it may
Version :  1.000  have any data type at each field point.
Last Updated :  01/22/92  The output is a field with the same
Ported To :  DEC IBM Sun HP  data, shape and dimensionality as the

original input field. It is rectilinear.

field to string parm

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1783  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

field to uni

Author :  Kathleen Dyer  Converts fields to uniform fields by
Lawrence Livermore  the simple expedient of stripping off
National Laboratory  their grid information. The input field
Number :  1120  can be any size, any data, any type.
Language :  C  The output is a uniform field of the
Submitted :  01/22/92  same type.
Version :  1.000
Last Updated :  01/22/92
Ported To :  DEC IBM Sun HP

fill

Author :  Anonymous  Does an area fill on a 2D scalar byte
Stardent Computer  field. The X and Y parameters specify
Number :  1033  the starting position of the fill, the
Language :  C  MIN and MAX parameters specify the
Submitted :  10/10/91  range of data which will get filled in
Version :  1.000  and the NEW parameter specifies the
Last Updated :  10/10/91  data value which will get substituted.
Ported To :  Kubota DEC IBM Sun HP  For instance, to set all connected

pixels in an image which lie in the range 200 to 255 to black, MIN = 200, MAX = 255, NEW = 0, and the the seed position must be set to somewhere in the interior of the region which you wish to fill.

filter 1Dfreq

Author :  International AVS Center  filter 1Dfreq is an avs module
North Carolina  executable which contains 5 modules
Supercomputing Center  hacked from khoros. 1D Frequency Domain
Number :  1507  Filters. ** dfiltbp ** a bandpass
Language :  C  filter design function ** dfiltbs ** a
Submitted :  10/01/92  bandstop filter design program **
Version :  0.100  dfilthp ** a high pass filter design
Last Updated :  10/01/92  function ** dfiltlp ** a low pass IIR
Ported To :  Sun DEC IBM  filter design function ** dzresp **

calculates magnitude, power or phase response of a discrete sys.

filter 1Dtime

Author :  International AVS Center  filter 1Dtime is an avs module
North Carolina  executable which contains 7 modules
Supercomputing Center  hacked from khoros. 1D Time Domain
Number :  1508  Filters. ** dfilconv ** converts 1D
Language :  C  filter coeff from transversal to
Submitted :  10/01/92  lattice realization ** dfilter **
Version :  0.100  filters time domain signals ** dllatt
Last Updated :  10/01/92  ** generates an error, a signal
Ported To :  Sun DEC IBM  estimate, and a PARCOR sequence **

dlsynth ** synthesizes an output sequence ** dltapp ** generates an error, a signal estimate, and a weight sequence ** dmlatt ** generates an error, a signal estimate, and a PARCOR sequence ** dmtapp ** generates an error, a signal estimate, and a weight sequence

filter 2Dfreq

Author :  International AVS Center  filter 2Dfreq is an avs module
North Carolina  executable which contains 8 modules
Supercomputing Center  hacked from khoros. 2D Frequency Domain
Number :  1509  Filters. ** vblse ** performs
Language :  C  band-limited surface extrapolation or
Submitted :  10/01/92  apodization ** vbpf ** performs a
Version :  0.100  band-pass filtering operation ** vbrf
Last Updated :  10/01/92  ** performs a band-reject filtering
Ported To :  Sun DEC IBM  operation ** vhpf ** performs a

high-pass filtering operation on an image ** vinverse ** restores images in the spatial domain by inverse filtering ** vlpf ** performs a low-pass filtering operation on an image ** vsvpsf ** performs restoration of space-varying blurs ** vwiener ** restores images in the spatial domain by Wiener filtering

filter 2Dspatl

Author :  International AVS Center  filter 2Dspatl is an avs module
North Carolina  executable which contains9 modules
Supercomputing Center  hacked from khoros. 2D Spatial Domain
Number :  1510  Filters. ** DEC vconvolve ** computes
Language :  C  two-dimensional spatial convolution of
Submitted :  10/01/92  two images ** vdiff ** performs image
Version :  0.100  differentiation by running 2 NxN
Last Updated :  10/01/92  weighted operations ** vdrf ** optimal
Ported To :  Sun DEC IBM  filter for edge detection ** vgef **

first derivative operator for symmetric exponential filter ** vgrad ** differentiates an image to perform image sharpening ** vhmed ** computes a two dimensional median filter ** vqmed ** computes a two dimensional median filter ** vsdef ** second derivative operator for symmetric exponential filter ** vspeckle ** reduces the speckle index of an image

filter morph

Author :  International AVS Center  filter morph is an avs module
North Carolina  executable which contains 10 modules
Supercomputing Center  hacked from khoros. Morphological
Number :  1511  Operations. ** vm180 ** computes a 180
Language :  C  degree rotation of a morphological
Submitted :  10/01/92  kernel ** vmcdilate ** computes the
Version :  0.100  dilation of a binary image by a
Last Updated :  10/01/92  structuring element ** vmclose **
Ported To :  Sun DEC IBM  computes the closing of an image by a

structuring element B ** vmcustom ** creates a new morphological kernel via keyboard entry ** vmdilate ** computes the dilation of a binary image by a structuring element ** vmerode ** computes the erosion of a binary image by a structuring element ** vmopen ** computes the opening of an image by a structuring element B ** vmoredge ** computes an edge extraction in an xvimage structure ** vmskel ** computes the skeleton of a binary byte image ** vmspan **

FILTER MODULES

fix field edge

Author :  Anonymous  For each cell that in a field (2D
Please forward  2-space 1-vector rectilinear integer)
avs@ncsc.org with author  that has a zero value, this module
info if you have it  searches for the closest neighbor that
Number :  1613  has a positive value and replaces the
Language :  C  zero value with that one.
Submitted :  04/10/93
Version :  1.000
Last Updated :  04/10/93
Ported To :  DEC IBM

float image

Author :  Wes Bethel  Converts an image contained in a
Lawrence Berkeley  floating point field to one contained
Laboratory  in a byte field. This is useful to in
Number :  1161  processing images in which the
Language :  C  algorithms operate on floating point
Submitted :  01/26/92  representations of images. Use of the
Version :  1.000  AVS-supplied converters (field to byte,
Last Updated :  01/26/92  field to float) will produce incorrect
Ported To :  Kubota DEC IBM Sun HP  results. The range of input pixels is 0

¡= value ¡= 255.

float parm to field

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1786  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

float to str

Author :  Urs Meyer  The “float to string” module will
Swiss Scientific  convert a floating point number to a
Computing Center  string according to the conversion
Number :  1623  specification given. The conversion
Language :  C  specification string is given via a
Submitted :  07/02/93  parameter, and it is the same as for
Version :  1.000  printf.
Last Updated :  07/02/93
Ported To :  HP IBM Kubota DEC Sun

fourier filter

Author :  Earl W. Hood and Allen  The FOURIER FILTER module takes a 1-3D
Schiano  input field and filters the field in
Convex Computer Corp.  the frequency-domain and outputs the
and UC Irvine  processed field. The module is
Number :  1548  specifically designed for image and
Language :  C  volume enhancement.
Submitted :  11/04/92
Version :  1.000
Last Updated :  11/04/92

fourier transfor

Author :  Earl W. Hood and Allen  FOURIER TRANSFORM computes the Fourier
Schiano  transform of a field. If the input
Convex Computer Corp.  field is scalar, the modules assumes
and UC Irvine  that the imaginary component of the
Number :  1549  Fourier transform is zero. If the field
Language :  C  is 2-vector, the first component is the
Submitted :  11/04/92  real part, and the second component is
Version :  1.000  the imaginary part. In either case, the
Last Updated :  11/04/92  output field is 2-vector with the real

part in the first component, and the imaginary part in the second component of the vector. The module uses Fast Fourier Transform algorithms to compute the desired Fourier transform. Therefore, the output field might have certain dimensions expanded as a result of the algorithms.

gamma correct

Author :  Wes Bethel  The gamma correct module performs gamma
Lawrence Berkeley  correction on an AVS image. The user
Laboratory  may control the amount of correction to
Number :  1550  apply. The user may also choose how
Language :  C  correct the image, either in a manner
Submitted :  11/04/92  which will result in a hue shift in the
Version :  1.000  image, or so as to preserve hues.
Last Updated :  11/04/92
Ported To :  DEC IBM Kubota HP

geo control

Author :  Massimiliano Bavelloni  This module is useful to create simple
Radioastronomy  animation of any geometry object,
department of C.N.R.  provides simple rotation around
Bologna Italy.  selected axis, invert rotation sense,
Number :  1884  scale and translate geometry objects
Language :  C  inside geometry viewer scene.
Submitted :  09/07/95
Version :  1.000
Last Updated :  09/07/95
Ported To :  DEC ultrix DEC alpha

geomanip

Author :  International AVS Center  geomanip is an avs module executable
North Carolina  which contains 7 modules hacked from
Supercomputing Center  khoros. Geometric Manipulations. **
Number :  1512  vexpand ** expands the size of the
Language :  C  image pointed ** vflip ** performs
Submitted :  10/01/92  reflection or “flip” operations on an
Version :  0.100  image ** vresize ** magnifies or
Last Updated :  10/01/92  reduces an image by using billinear
Ported To :  Sun DEC IBM  interpolation ** vrotate ** rotates an

image by a specfied angle, about a specified point ** vshrink ** decreases the size of an image ** vtranslat ** computes the translation of an xvimage structure by vector (i,j) ** vtranspos ** computes the transpose of the rows and columns in an image

geometrySearch

Author :  John Stephen  This module produces a text string list
Tessella Support  of items currently in a particular
Services  geometry viewer scene. The viewer and
Number :  1658  name search string are matched, and a
Language :  C  full multi-line text string is sent to
Submitted :  09/29/93  the output port.
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Sun

glow white

Author :  David Haith  Blurs white to look like glowing light
Power-Media  from input image to output. Controls
Number :  1039  specify threshold level for glow, and
Language :  C  the mix level for replacement. Both are
Submitted :  10/10/91  integer parameters.
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

gradual mix

Author :  David Haith  Places one image over another with a
Power-Media  graduated fade from left to right. The
Number :  1040  degree of mix on each side of the image
Language :  C  is given individual control. Both
Submitted :  10/10/91  images must be the same size. Four
Version :  1.000  parameters come with the module, all
Last Updated :  10/10/91  integer type - “left min”, “right min”,
Ported To :  Kubota DEC IBM Sun HP  “left mix”, and “right mix”.

grey scales

Author :  Rob Kuyper  The module takes an image as input an
Univeristy of Technology  produces a black and white version of
Delft, The Netherlands  the same image as output. The number of
Number :  1315  grey scales is adjustable by a dial.
Language :  C  The grey level is produced in the same
Submitted :  08/30/92  way as with the command xpr to print
Version :  1.000  windows on a laserjet. The advantage of
Last Updated :  08/30/92  adjusting the number of grey levels is
Ported To :  DEC Sun HP  the fact that the produced image shows

the edges between different colors better. The output can be printed on a laserjet which results in a much better picture then without this module. In order to have white as one of the grey levels, the roundoff works like a ’floor’.

grid scale

Author :  Kathleen Dyer  Allows the grids of 2D and 3D fields to
Lawrence Livermore  be scaled. The I, J, and K grids may be
National Laboratory  scaled independently. There is no
Number :  1122  effect on a uniform field. This module
Language :  C  simply multiplies the various grid
Submitted :  01/22/92  values by the appropriate scale
Version :  1.000  factor(s). Fields which are not entered
Last Updated :  01/22/92  at the origin may appear to shift.
Ported To :  Kubota DEC IBM Sun HP

histogram

Author :  International AVS Center  histogram is an avs modules executable
North Carolina  which contains 7 modules hacked from
Supercomputing Center  khoros. Histogram. ** venhance **
Number :  1513  creates a new image g(x,y) from input
Language :  C  image f(x,y) ** venhpatch ** ** vheq **
Submitted :  10/01/92  performs histogram equalization **
Version :  0.100  vhisto ** computes the histogram of a
Last Updated :  10/01/92  gray-levelimage ** vhsee ** outputs a
Ported To :  Sun DEC IBM  histogram in a VIFF format in

displayable form ** vhstr ** performs a histogram stretch of the grey levels ** vhxray ** enhances an image by using histogram stretch or equalization

hit lumin mi

Author :  David Haith  Mixes according to pixel luminance. The
Power-Media  left input is foreground , the right
Number :  1043  input background. Both images must be
Language :  C  the same size. The mix parameter
Submitted :  10/10/91  controls effect, blend is on pixel
Version :  1.000  luminance of foreground image. This
Last Updated :  10/10/91  module is a variation taken from
Ported To :  Kubota DEC IBM Sun HP  lumin mix, where luminance is (green +

blue *2 ) /3.0, instead of red, green, blue.

hsv to rgb

Author :  Wes Bethel  Converts an image, represented as HSVA
Lawrence Berkeley  (A = alpha) to ARGB (A = alpha). You
Laboratory  might use this module in a network with
Number :  1127  read image, field to float,
Language :  C  rgb to hsv, extract scalar,
Submitted :  01/26/92  field math, combine scalars,
Version :  1.000  field to byte, and colorizer to
Last Updated :  01/26/92  attenuate the V component (luminosity),
Ported To :  Kubota HP  then recombine back into HSV space, and

display the resulting image. The result is a “darker” image. See the .txt file accompanying the module for this example network.

image combine

Author :  Hideo Miyachi  image combine generates a image from
Kubota Corporation  two image datas. This module has 3
Number :  1615  input ports - two images and a
Language :  C  colormap. Colormap is the optional port
Submitted :  05/01/93  which is for background color if the
Version :  1.000  two images have diffrent size. This is
Last Updated :  05/01/93  similar to “collage” but this is very
Ported To :  HP  simple. I used this module when I

generate VTR tape to compare two visualized images showing time dependent data.

image flip

Author :  Bill Sherman  The “image flip” module take an image
National Center for  data field and duplicates the
Supercomputing  information by flipping it over the
Applications  selected edge. This differs from the
Number :  1192  “mirror” module that is provided in the
Language :  C  standard AVS library. The “mirror”
Submitted :  05/05/92  module flips an image, but does not
Version :  1.000  duplicate the data, so the original
Last Updated :  05/05/92  image is not part of the output. “Image
Ported To :  DEC IBM Sun Kubota HP  flip“ keeps the original image, placing

the new mirrored data adjacent to the edge flipped over.

image rotate

Author :  William Hunter  This module is similar to mirror and
GEC Ferranti DSL,  transpose, but combines operations to
Edinborough, Scotland  provide image rotation in 90 degree
Number :  1184  steps, controlled by choice buttons.
Language :  C  The source makes extensive use of
Submitted :  05/04/92  macros to allow for any storage type,
Version :  2.000  such as byte, float, integer, or
Last Updated :  09/30/93  doubles. This is contributed “AS-IS”,
Ported To :  IBM Sun Kubota HP  and implies no liability of the author.

image stack

Author :  Scott Goodyear and Terry The “image stack” module pastes input
Myerson  image 1 and input image 2 together, top
Number :  1835  to bottom, to produce the output image.
Language :  C  Image 1 and Image2 must be the same
Submitted :  07/16/94  width. The dimension of the output
Version :  1.000  image is (height of image 1 + height of
Last Updated :  07/16/94  image 2)*(width of image 1). This was
Ported To :  DEC HP IBM Kubota Sun  used to combine two images for

comparison, or to create a composite mosaic for an aerial photograph GIS application. Derived from Terry Myerson’s side by side module.

image stats

Author :  John Tee  Computes field statistics, but in a
AVS Inc. Consultant  multiband image context. Statistical
Number :  1662  measures are applied to each channel in
Language :  C  the image, and displayed in an
Submitted :  09/29/93  information panel. It computes min,
Version :  1.000  max, mean, and standard deviation. The
Last Updated :  09/29/93  operator may apply to the whole image,
Ported To :  HP IBM Sun  or the ROI region, using the optional

mask port, connected to the AVS5 image viewer.

image tile

Author :  Ian Curington  Creates a replicated (tiled) version of
Advanced Visual Systems  the input image. It uses a simple
Number :  1330  horizontal pixel duplication, to make a
Language :  C  small postage stamp image fill a box.
Submitted :  08/31/92  The input image height and width is
Version :  1.000  reported. This module was used in a
Last Updated :  08/31/92  demo to a design studio that
Ported To :  DEC Sun HP  manufactures ceramic tiles for kitchens

and bathrooms. Maybe you know someone in wallpaper or fabric? It is also used to synthesize a Y dimension, to create an image from 1D data.

info

Author :  International AVS Center  info is an avs module executable which
North Carolina  contains 6 modules hacked from khoros.
Supercomputing Center  File Information. ** dpeakpick ** finds
Number :  1514  peaks or troughs in 1D data vectors **
Language :  C  dstats ** computes 1-D statistics of an
Submitted :  10/01/92  input field ** varviff ** takes as
Version :  0.100  input a field and extracts a single
Last Updated :  10/01/92  data value ** vfileinfo ** prints out
Ported To :  Sun DEC IBM  the header information that describes

input field ** vprdata ** prints field to the screen or to a file in ASCII format ** vstats ** performs statistical calculations

input create1D

Author :  International AVS Center  input create1D is an avs module
North Carolina  executable which contains9 modules
Supercomputing Center  hacked from khoros. Create Signals
Number :  1515  (1D). ** dgexpon ** creates one
Language :  C  dimensional exponential distributed
Submitted :  10/01/92  noise data set ** dggauss ** creates a
Version :  0.100  one dimensional gaussian noise data set
Last Updated :  10/01/92  ** dgpoiss ** creates a one dimensional
Ported To :  Sun DEC IBM  Poisson distributed noise data set

**dgpwl ** creates a one dimensional set of signals **dgrleigh ** creates a one dimensional Rayleigh distributed noise data set **dgsin ** creates a one dimensional ensemble of signals **dgsinc ** creates a one dimensional ensemble of signals **dgunif ** creates a one dimensional uniform noise data set **dgwindow ** creates 1D data sequence based on user selected window function

input create2D

Author :  International AVS Center  input create2D is an avs module
North Carolina  executable which contains12 modules
Supercomputing Center  hacked from khoros. Create Images (2D).
Number :  1516  ** vcustom ** creates a custom kernel
Language :  C  file in the VIFF format ** vgauss **
Submitted :  10/01/92  generates image that contains one or
Version :  0.100  more Gaussian distributions ** vgbox **
Last Updated :  10/01/92  fills buffer with greylevel rep. of
Ported To :  Sun DEC IBM  parallelogram on CCD sensor ** vgcirc

** creates an image of a filled circle ** vgconst ** creates an image of dimension rows bands ** vgfractal ** creates a fractal image with a desired fractal dimension ** vggauss ** creates a two dimensional multi-band Gaussian (noise) image ** vgpwl ** creates a two dimensional piecewise linear image ** vgshot ** corrupts an image uniformly with spikes ** vgsin ** creates a two dimensional image of dimension rows*columns ** vimpulse ** creates image containing 1 or more unit impulses ** vmarr ** creates a convolution filter kernel

int 2 st

Author :  Urs Meyer  The int to string module will convert
Swiss Scientific  an integer to a string according to the
Computing Center  conversion specification given. It is
Number :  1624  very useful and flexible for generating
Language :  C  series of filenames. The format
Submitted :  07/02/93  parameter widget is a character string
Version :  1.000  that controls the conversion of the
Last Updated :  07/02/93  integer value. The conversion
Ported To :  HP IBM Kubota DEC Sun  specification string is the same as for

printf.

int parm to field

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1787  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

integer control

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology four-module suite - 1) dual control -
/ Crump Institute  to control flow of field and integers
Number :  1776  to downstream modules, 2) field control
Language :  C  - to control flow of an AVS image to
Submitted :  02/09/94  downstream modules, 3) integer control
Version :  1.000  - to control flow of an integer to
Last Updated :  02/09/94  downstream modules, and 4) real control
Ported To :  HP Sun  - to control flow of an real to

downstream modules. This module serves the same purpose as Field Control except for integers instead of fields. Its operation is identical. See Field Control for further info. An example of its use is to allow animate integer to remain connected to a number of integer ports on one or several modules while allowing the user to select which of these downstream modules will be effected. Provision is also made for manually sending single integer updates downstream.

integer to real

Author :  Lars M. Bishop  Integer to Real simply takes in an
National Center for  integer-valued input port, and passes
Comp. Elect., Univ. of  its truncation to the output port as a
Illinois  real. This module is part of the the
Number :  1572  Real Math module group, which is a
Language :  C  single file, process-sharing set of
Submitted :  02/11/93  modules that allow basic operations on
Version :  1.000  and between integer and real AVS data
Last Updated :  02/11/93  types. They have been most useful in
Ported To :  DEC IBM  dealing with linkage of several AVS

parameters. AVS has no simple way of performing math operations on reals, even though it does allow math operations on fields.

interp cmap

Author :  Larry Gelberg  interp cmap interpolates between two
Advanced Visual Systems  colormaps. This is particularly useful
Number :  1288  when using the AVS Animator module
Language :  C  because it does not interpolate between
Submitted :  07/03/92  colormaps for you. The dial scale
Version :  1.000  controls the blending between the
Last Updated :  07/03/92  colormaps. When scale = 0.0 the output
Ported To :  Sun HP  is entirely cmap1, when scale = 1.0,

the output is entirely cmap2 and in the middle, the output is interpolated between the two. Every value of every band (hue, saturation, value, and opacity) is evaluated. This module assumes colormaps which are 256 entries long.

irreg 2 reg

Author :  Wes Bethel  Converts an irregular field to one
Lawrence Berkeley  which is uniform by removing the
Laboratory  coordinate information from the
Number :  1047  irregular field. Converting an
Language :  C  irregular field to one which is uniform
Submitted :  10/10/91  is useful when one wants to create a
Version :  1.000  “scalar mesh” (using field2mesh) where
Last Updated :  10/10/91  the elevation is derived from an
Ported To :  Kubota DEC IBM Sun HP  associated scalar value in the field.

Field 2 mesh creates a “scalar mesh” only when the input field is uniform (not when rectilinear or irregular).

last cone

Author :  Mark Pinches,  The lastcone module transforms an AVS
Southampton Parallel  geometry, replacing a set of disjoint
Applications Centre, UK  lines with “cones” constructed out of
and Ian Curington, AVS  polygons. The cone tristrip is now upto
Inc.  2x more efficient, and has a
Number :  1667  subdivision slider control. If the
Language :  C  input has values the module will make
Submitted :  09/29/93  colour coded cones based on the input
Version :  1.000  lines and colormap. The cone is only
Last Updated :  09/29/93  drawn at the end of a polyline
Ported To :  HP IBM Kubota Sun  structure, so it can be used with

cylinder or tube, to create solid

arrows. The cone is sent on one geom output port, while the original polyline is sent on the second port, such that it may be sent to cylinder. Min and Max controls allow lines of only certain bounded length to have cones, shorter or longer lines do not get cones.

linearop1D

Author :  International AVS Center  linearop1D is an avs module executable
North Carolina  which contains 3 modules hacked from
Supercomputing Center  khoros. Linear Operations. ** dacorr **
Number :  1517  generates biased, unbiased or FFT based
Language :  C  autocorrelation estimate ** dconvo **
Submitted :  10/01/92  generates time domain or FFT based
Version :  0.100  circular convolution estimate ** dxcorr
Last Updated :  10/01/92  ** generates biased, unbiased or FFT
Ported To :  Sun DEC IBM  based cross-correlation est.

local area stats

Author :  John Tee  LOCAL AREA STATS is an image processing
AVS Inc. Consultant  module, similar to LOCAL AREA OPS, but
Number :  1666  including other measurements. For all
Language :  C  choices, the pixel value is replaced by
Submitted :  09/29/93  the appropriate statistical measurement
Version :  1.000  of the local area around it, as
Last Updated :  09/29/93  determined by the kernel width
Ported To :  IBM Kubota  parameter. The most useful measurement

provided by this module is the standard deviation, which can be used to create adaptive thresholding fields. Skewness and Kurtosis are defined in the information for STATISTICS which calculates the global values of these parameters over an entire field. The kernel width is constrained to be an odd number, so that it can be centred on a pixel. The data type of the output field is the same as the input field in the cases of min, max, mean and median, but is always of float type for std dev, skewness and kurtosis.

log field

Author :  Wes Bethel  log field computes a logarithm of the
Lawrence Berkeley  input field, and places the result into
Laboratory  the data portion of the output field –
Number :  1304  the output field is the logarithm of
Language :  C  the input field.
Submitted :  08/26/92
Version :  1.000
Last Updated :  08/26/92
Ported To :  DEC Sun HP

lumin mix

Author :  David Haith  Mixes two images according to pixel
Power-Media  luminance. The left input is foreground
Number :  1050  , the right input background. Both
Language :  C  images must be the same size. The mix
Submitted :  10/10/91  parameter controls the effect. The
Version :  1.000  blend is on pixel luminance of
Last Updated :  10/10/91  foreground image.
Ported To :  Kubota DEC IBM Sun HP

luminence

Author :  Earl W. Hood  Luminence computes the luminence of an
Convex Computer  image. However, it differs from the
Corporation  standard AVS luminence module in that
Number :  1557  one can control how much red, green,
Language :  C  and blue contribute to the brightness
Submitted :  11/16/92  of an image. I found this functionality
Version :  1.000  useful when using the module with the
Last Updated :  11/16/92  replace alpha and composite modules
Ported To :  IBM  (eg. I can make any color with red be

fully opaque while other colors are

transparent).

matrix algebra

Author :  International AVS Center  matrix algebra is an avs module
North Carolina  executable which contains10 modules
Supercomputing Center  hacked from khoros. Matrix
Number :  1518  Manipulation. ** DEC mcovar ** computes
Language :  C  maximum likelihood estimate of the
Submitted :  10/01/92  covariance matrix ** meigen ** computes
Version :  0.100  eigenvectors and eigenvalues of real or
Last Updated :  10/01/92  complex sq matrix ** minfo ** obtains
Ported To :  Sun DEC IBM  information about a matrix ** minvert

** computes the inverse of a real or complex matrix ** mlde ** solves system of const. coeff. linear ordinay differential eqs. ** mlse * solves a system of linear silseaneous equations ** mlud ** computes the LU decomposition of a matrix ** mmult ** multiplies two matricies ** msvd ** computes the singular value decomposition of a matrix ** mtrans ** transposes the matrices contained in the input image

matte 3x3

Author :  David Haith  Creates an alpha plane matte which is
Power-Media  antialiased along with the rest of the
Number :  1052  image. Sample ratio is 9 into 1. The
Language :  C  output image is 1/3 size in each
Submitted :  10/10/91  direction. The input image height and
Version :  1.000  width must be a factor of 3, or the
Last Updated :  10/10/91  indexing in this module falls apart.
Ported To :  Kubota DEC IBM Sun HP

matte sample

Author :  David Haith  Creates an alpha plane matte which is
Power-Media  antialiased along with the rest of the
Number :  1051  image. The sample ratio is 4 into 1.
Language :  C  The output image is 1/2 size in each
Submitted :  10/10/91  direction. The input image height and
Version :  1.000  width must be a factor of 2, or the
Last Updated :  10/10/91  indexing in this module falls apart.
Ported To :  Kubota DEC IBM Sun HP

min max fiel

Author :  Terry Myerson  Calculates the smallest and largest
International AVS Center  value in the data segment of a field.
(NCSC)  the field can be of any type, any
Number :  1140  vector. The outputs will always be
Language :  C  floating point values. When used with
Submitted :  03/03/92  continuous stats (another module in the
Version :  1.000  ftp site), this was used to compile a
Last Updated :  03/03/92  running minimum/maximum tally of the
Ported To :  DEC IBM Sun HP  data in a time series. This

minimum/maximum was then used to fix the limits of a colormap.

mix2

Author :  John Tee  MIX2 mixes two fields together by
AVS Inc. Consultant  combining corresponding elements, such
Number :  1669  that output = input1 * weight1 + input2
Language :  C  * ( 1 - weight1) If the two fields have
Submitted :  09/29/93  different dimensions, the operation is
Version :  1.000  performed on the overlapping portion of
Last Updated :  09/29/93  the data - ie using the smaller
Ported To :  HP IBM Kubota Sun  dimension. The module is used, for

example, to produced a “sharpened” image by mixing high-pass filtered data with a proportion of the unprocessed data. It can also be used in conjunction with ANIMATED FLOAT to provide a gradual fade between two images.

modify seq1D

Author :  International AVS Center  modify seq1D is an avs module
North Carolina  executable which contains 9 modules
Supercomputing Center  hacked from khoros. Signal/Sequence
Number :  1519  Modification. ** dcpad ** performs a
Language :  C  constant value padding of a 1D data
Submitted :  10/01/92  sequence ** dextract ** extracts
Version :  0.100  1-dimensional subsignals from the input
Last Updated :  10/01/92  data file ** dinsert ** inserts 1-D
Ported To :  Sun DEC IBM  vectors from input file (img2) into

vectors in img1 ** dmirror ** generates a mirror data file ** dnormal ** normalizes the magnitude of the input data ** dscale ** rescales data for processing thus minimizing roundoff errors ** dshift ** shifts or rotates the 1-D signals in the field ** dsubsamp ** subsamples 1-D vector(s) from the field ** dwindow ** multiplies user selected 1-D data sequence by window function

my mirror

Author :  Michael Rangitsch  Mirrors a 3D data set about one of the
Number :  1131  axes, selected by the user. This is a
Language :  FORTRAN  convenient way to fill out a 3D data
Submitted :  01/27/92  set which was computed using a plane of
Version :  1.000  symmetry. The AVS supplied mirror
Last Updated :  01/27/92  module mirrors the geometry correctly,
Ported To :  DEC IBM Kubota HP  but mirrors the data incorrectly. This

version of my mirror simply negates one of the coordinates, depending on the plane choice. For regular data, it inverts the mapping in the chosen direction.

natural log

Author :  Earl W. Hood  The natural log filter module computes
Convex Computer  log base e of a real field, i.e.
Corporation  newvalue = LOG e (value)
Number :  1560
Language :  C
Submitted :  11/16/92
Version :  1.000
Last Updated :  11/16/92
Ported To :  IBM

no format

Author :  International AVS Center  no format is an avs module executable
North Carolina  which contains 3 modules hacked from
Supercomputing Center  khoros. No File Formats. ** asc2AVS **
Number :  1520  transforms ASCII data into a field **
Language :  C  raw2AVS ** transforms RAW data into a
Submitted :  10/01/92  field vconvert performs data storage
Version :  0.100  type conversions ** AVS2raw ** writes
Last Updated :  10/01/92  ONLY the raw image data present onto
Ported To :  Sun DEC IBM  descriptor output fd

opening

Author :  David E. Harris  This module performs the mathematic
Georgia Institute of  morphological opening operation on
Technology  grayscale (or binary) images. See
Number :  1173  Haralick et al. “Image Analysis using
Language :  C  Mathematical Morphology“, IEEE Trans.
Submitted :  04/16/92  on Pattern Analysis and Machine
Version :  1.000  Intelligence, PAMI-9, No. 4, pp.
Last Updated :  04/16/92  532-550, 1987 for a more complete
Ported To :  Kubota  discription. The opening of image A by

structuring element B (a binary mask) is defined as the dilation (by B) of the erosion of A by B. This is done by feeding the output of the ERODE module into the input of the DILATE module and using the same mask for both. The OPENING module just combines these two modules for convenience. See also the CLOSING module.

output dither

Author :  International AVS Center  output dither is an avs module
North Carolina  executable which contains 3 modules
Supercomputing Center  hacked fomr khoros. Dither. ** vdith45
Number :  1521  ** performs 8x8 ordered block dither
Language :  C  with major axes on 45’ angles **
Submitted :  10/01/92  vdither ** performs 5x5 ordered dithers
Version :  0.100  ** verrdif ** performs dithering
Last Updated :  10/01/92  operation using 4-direction error
Ported To :  Sun DEC IBM  diffusion

output print

Author :  International AVS Center  output print is an avs module
North Carolina  executable which contains 2 modules
Supercomputing Center  hacked from khoros. Print Image
Number :  1522  Hardcopy. ** vlpr ** formats a byte
Language :  C  image for output on a standard line
Submitted :  10/01/92  printer ** vpostscr ** formats a byte
Version :  0.100  image for output on a Postscript laser
Last Updated :  10/01/92  printer
Ported To :  Sun DEC IBM

panel pan

Author :  Anonymous  Uses upstream transform information to
Stardent Computer  crop an input image and display one
Number :  1060  panel at a time. It is designed to take
Language :  C  an input image, and copy it to the
Submitted :  10/10/91  output port. This is a straight
Version :  1.000  pass-through. The size of the output
Last Updated :  10/10/91  image is 1/panel res of the input
Ported To :  Kubota DEC IBM Sun HP  image. The module sensed mouse activity

in borders near the edge of the window, and will flip to the next panel in that direction. The mouse position is fed externally from this module, getting the current mouse position and button state from the “struct upstream transform“ input port. This module is designed to be used with “track cursor”, which displays the image and catches the X events.

pencil sketc

Author :  Larry Gelberg  Simulates a pencil sketched image by
Stardent Computer  randomly sampling an image and
Number :  1061  generating blobs based on the color of
Language :  C  the sample. The number of samples is
Submitted :  10/10/91  controlled by the N-STROKE dial
Version :  1.000  (5000-10000 makes nice pictures). The
Last Updated :  10/10/91  length and widths of the lines are
Ported To :  Kubota DEC IBM Sun HP  controlled by the LENGTH and WIDTH

dials. The lines are all placed down at the same angle (controlled with the ANGLE dial) unless the USE GRADIENT button is turned on in which case, the gradient of the luminance of the sampled pixel is used to determine the line angle. Since all the samples are added into each other, you will quickly max out the data values. In order to avoid this, there is a dial called FILTER SCALE which tones this down. Finally, you can control the color of the background with the GREY PAPER knob. 0.0 represents black and 1.0 represents white.

persp field

Author :  Wes Bethel  Applies a perspective transformation,
Lawrence Berkeley  as specified by its parameters, to each
Laboratory  and every grid point in a 3 space
Number :  1062  irregular field. This module was
Language :  C  designed to be used with the
Submitted :  10/10/91  curvilinear volume renderer, “bezier
Version :  2.000  volume.“ The input field consists of an
Last Updated :  10/10/91  irregular field with 3-space
Ported To :  Kubota IBM  coordinates. The data type and vector

length may be any configuration. The output field is the same size and type as the input field, but with the coordinates transformed according to the parameters.

real control

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology four-module suite - 1) dual control -
/ Crump Institute  to control flow of field and integers
Number :  1777  to downstream modules, 2) field control
Language :  C  - to control flow of an AVS image to
Submitted :  02/09/94  downstream modules, 3) integer control
Version :  1.000  - to control flow of an integer to
Last Updated :  02/09/94  downstream modules, and 4) real control
Ported To :  HP Kubota Sun  - to control flow of an real to

downstream modules. This module serves the same purpose as Integer Control except for floats instead of integers. Its operation is identical. See Field Control for further info. An example of its use is to allow animate float to remain connected to a number of float ports on one or several modules while allowing the user to select which of these downstream modules will be effected. Provision is also made for manually sending single real updates downstream.

real to integer

Author :  Lars M. Bishop  The Real to Integer module simply takes
National Center for  in a real- valued input port, and
Comp. Elect., Univ. of  passes its truncation to the output
Illinois  port as an integer. This module is part
Number :  1573  of the the Real Math module group,
Language :  C  which is a single file, process-sharing
Submitted :  02/11/93  set of modules that allow basic
Version :  1.000  operations on and between integer and
Last Updated :  02/11/93  real AVS data types. They have been
Ported To :  DEC IBM  most useful in dealing with linkage of

several AVS parameters. AVS has no simple way of performing math operations on reals, even though it does allow math operations on fields.

reconstruct slic

Author :  Joseph A. Oswald, III  The reconstruct slice module takes the
Radiation Physics, Duke  result of a “slave slicer” operation,
University Medical  which is a 2D uniform scalar field
Center  representing a slice through a 3D
Number :  1708  scalar field, and constructs a 3-space
Language :  C  field with identical data to this 2D
Submitted :  09/30/93  uniform field, but with coordinates
Version :  1.000  corresponding to the original 3D
Last Updated :  09/30/93  coordinate space. This output is
Ported To :  Kubota  suitable for use with “field to mesh”

to render the 2D slice with a 3D geometry. The three modules in this package, “slave slicer,” “specify slice,“ and ”reconstruct slice,“ are a flexible and efficient alternative to “arbitrary slicer” and Wes Bethel’s “new arbitrary slicer.” Slice specifications are generated by a module separate from the actual slicing module. That permits a single specifying module to specify identical slices through multiple data sets, and also for users to create their own modules to specify slices. The slicing algorithm takes 2D slices of 3D scalar fields (uniform or rectilinear) and outputs them in the most space efficient format, 2D uniform. 2D 3-space irregular output, similar to that output by “new arbitrary slicer” can be generated, if needed, using “reconstruct slice.” The slicing algorithm is generally faster than “new arbitrary slicer,“ at least on the Kubota architecture.

rect2pol

Author :  anonymous  This module will convert any 2D or 3D,
Number :  1891  2 space or 3 space field from
Language :  C  rectangular coordinates to polar, and
Submitted :  09/07/95  vice versa. The output field is always
Version :  1.000  an irregular field, but the input can
Last Updated :  09/07/95  have any uniform or type setting. When
Ported To :  DEC ultrix DEC alpha  converting into polar coordinates, the

output angles will be in the range of 0 to 2*pi. When converting from polar to rectangular, any angles can be used, but it is possible to overlap data if the range of angles is greater than 2*pi. For cylindrical coordinates, points along the z-axis will have a radius equal to zero, and therefore an undefined angle. the atan2 function will usually set this angle to 0 or PI. For that reason, polar to rectangular to polar conversions will not always produce the original input field, though no significant data is lost. For spherical coordinates, this effect only concerns points located at the origin. The Polar Type parameter is removed for 2 space fields since cylindrical coordinates are assumed.

remap field

Author :  Phil McDonald  The remap field module provides the
NOAA/ERL/Forecast  capability of changing a field’s
Systems Laboratory  computational space to physical
Number :  1268  coordinate mapping scheme. For example,
Language :  C  a uniform field may be remapped to an
Submitted :  05/27/92  irregular field, or vice versa. Another
Version :  2.000  example is to change the coordinates of
Last Updated :  07/05/93  an irregular field from one grid system
Ported To :  IBM  to another. The data of one field (the

data field) can be mapped in the same manner as an optional second field (the grid field). The updated version includes several changes to the module and the libraries it uses as well.

remote gis

Author :  International AVS Center  remote gis is an avs modules
North Carolina  executable which contains 7 modules
Supercomputing Center  hacked from khoros. Standard File
Number :  1523  Formats. ** big2AVS ** converts a BIG
Language :  C  image file into a field ** dem2AVS **
Submitted :  10/01/92  converts digital elevation map data
Version :  0.100  (USGS format) to a field ** dlg2AVS **
Last Updated :  10/01/92  converts an optional format digital
Ported To :  Sun DEC IBM  line graph file to a field ** elas2AVS

** converts an ELAS image file format into a field ** elasinfo ** reads and prints an ELAS image fiel format header ** AVS2big ** converts a BIG image file into a field ** vll2utm ** converts geographical to universal transverse mercator coordinates

rgb to cmyk

Author :  Anonymous  Converts from red-green-blue to
Stardent Computer  cyan-magenta-yellow-black color space
Number :  1068  representation. Both AVS filters and
Language :  C  offline standalone program versions are
Submitted :  10/10/91  included. prepend.c and denude.c are
Version :  1.000  used for conversion to and from SCITEX
Last Updated :  10/10/91  CT2T cmyk tape format files. The input
Ported To :  Kubota IBM  and output are both field 2d 4-vector

byte (image data).

rgb to hsv

Author :  Wes Bethel  Converts an image, represented as ARGB
Lawrence Berkeley  (alpha, red, green, blue) to HSVA (hue,
Laboratory  saturation, luminosity, alpha). The H
Number :  1130  component ranges from 0.0 to 1.0. You
Language :  C  might use this module to display only
Submitted :  01/26/92  the V (or luminosity) component of an
Version :  1.000  image, i.e. grayscale. A sample network
Last Updated :  01/26/92  to do this is provided within the .txt
Ported To :  Kubota HP  file accompanying this module.

river fill

Author :  Ken Flurchick  This module is a coroutine which
North Carolina  continuously pads a two-dimensional
Supercomputing Center  array with data until it is turned off.
Number :  1580  This can be used with the “river bed”
Language :  C  module which generates the coordinates
Submitted :  02/11/93  for the array, to show a simulation of
Version :  1.000  concentration over time.
Last Updated :  02/11/93
Ported To :  IBM Kubota HP

scat to tri

Author :  Ian Curington  This is a 2D grid generation routine,
Advanced Visual Systems  that takes random scatter data in a
Number :  1681  plane, and creates a delaunay
Language :  C  triangulation from them. It sends node
Submitted :  09/29/93  position data and connectivity
Version :  1.000  information out via a UCD structure,
Last Updated :  09/29/93  containing only the 2D elements. It
Ported To :  HP Kubota Sun  only grids based on the X and Y

coordinates of the input data, but preserves the Z values on the UCD, so irregular 3D sheets may be generated. The core algorithm is from a public domain library for voronoi and delaunay triangulation at “netlib”, called “sweep2”. The only modifications to the package from netlib is the re-write of “main.c” and “output.c” to create AVS coroutine structures, and to pass data from shared memory rather than file I/O. All data stored at each point in the input field is transfered unchanged to the UCD, for further interpretation or processing.

scat znmo

Author :  Ian Curington  This routine takes two 1D irregular
Advanced Visual Systems  fields, looks up the Z (depth) in the
Number :  1755  first one, finds the corresponding
Language :  C  depth of the second one, and replaces
Submitted :  12/15/93  the X-Y of the first by the second at
Version :  1.000  that point. Linear or Quadratic
Last Updated :  12/15/93  interpolation is available by user
Ported To :  Kubota Sun  choice. Both fields can be different

lengths, and have no depths in common. The first must be contained within the range of the second, or the lookup will fail. All the data from the first is passed to the output. The underlying utility routines could easily be applied in other contexts. NOTE - Advanced AVS Module Techique used. This routine contains an advanced module writer trick, of modifying the input port data inplace, without explicitly or implicitly copying the data out of shared memory. This REQUIRES that a special flag be turned on in your .avsrc file - ReadOnlySharedMemory 0

scatter downsize

Author :  Michael J. Rangitsch  This module samples a scatter data set
Number :  1886  by the input rate. It is useful for
Language :  C  downsizing samplers for use as vector
Submitted :  09/07/95  locations or streamline/particle
Version :  1.000  advection starting points.
Last Updated :  09/07/95
Ported To :  DEC ultrix DEC alpha

scatter to spars

Author :  Penny Rheingans  Convert a 1D field containing scattered
US EPA Scientific  data points into a (potentially)
Visualization Center  sparsely filled uniform field. The
Number :  1309  output field has as many dimensions as
Language :  C  the input field did coordinate
Submitted :  08/30/92  dimensions. For example, if the input
Version :  1.000  is a 1D 3-space field, the output will
Last Updated :  08/30/92  be a 3D 3-space field. Each dimension
Ported To :  DEC Sun HP  of the output field will extend from

the minimum value to the maximum value for that coordinate in the input field. Coordinates in the input field are truncated to integers. Accordingly, if a dimension of the input field extends from 15.3 to 19.0, the output field will have 5 elements in that dimension.

segment

Author :  International AVS Center  segment is an avs module executable
North Carolina  which contains 8 modules hacked from
Supercomputing Center  khoros. Segmentation. ** v1bgamut **
Number :  1524  generates a 1 band image with map of
Language :  C  displayable color levels ** vclose **
Submitted :  10/01/92  tries to close the boundaries **
Version :  0.100  vdistance ** computes the distance from
Last Updated :  10/01/92  the nearest boundary point ** vdyth **
Ported To :  Sun DEC IBM  performs dynamic thresholding on an

image ** vgamth ** generates a binary image (value or 0) ** vgrow ** computes the median axis inverse transform ** vmediaxis ** computes the medial axis transform ** vthresh ** seperates binary image (value or 0) by thresholding input image

set alpha

Author :  David Haith  Sets data in the alpha plane of the
Power Media  output image according to incoming
Number :  1095  luminance of an input image. Copies the
Language :  C  R G B data across unchanged, but
Submitted :  10/15/91  replaces the alpha channel with either
Version :  1.000  the luminance or 0, based on a
Last Updated :  10/15/91  threshold key from the three (R G B)
Ported To :  Kubota DEC IBM Sun HP  parameter dials.

set extents

Author :  Wes Bethel  Maps uniform fields, by default, into
Lawrence Berkeley  world coordinates. Uses the formula
Laboratory  (for each nspace dimension called
Number :  1145  i )- Min[i] = 0, Max[i] =
Language :  C  dimensions[i] - 1. This arrangement
Submitted :  03/10/92  will map a volume of data, where each
Version :  3.000  of the logical i,j,k dimensions of data
Last Updated :  06/20/92  are equal (a 64**3 dataset, for
Ported To :  IBM Sun Kubota HP  example) and map it into a “cube.” For

most purposes, this is fine. However, consider an MR dataset in which the slices have a fine resolution, but there are only a few slices. The slices have thickness which is lost using the above mapping function.

shutter

Author :  Wes Bethel  The shutter module allows the user to
Lawrence Berkeley  control the flow of an AVS image to
Laboratory  downstream modules. This module is
Number :  1571  designed for use in networks in which
Language :  C  animation sequences are created using
Submitted :  01/29/93  the geometry renderer. This module
Version :  1.000  allows for the idiosyncracies of the
Last Updated :  01/29/93  AVS render geometry module which will
Ported To :  DEC IBM Kubota HP  perform several pixmap updates prior to

finishing the rendering of the entire scene. Shutter will be useful in making sure the entire scene has been rendered prior to passing the resultant image downstream to a data output module (which may put the image onto a videodisk, or write the image to a file).

size fit

Author :  John Tee  SIZE FIT is used to make 2 fields the
AVS Inc. Consultant  same size, typically before passing
Number :  1683  them to a module which combines them in
Language :  C  some way. The input field is rescaled,
Submitted :  09/29/93  using linear interpolation, to be the
Version :  1.000  same size as the template field. The
Last Updated :  09/29/93  template field is unaffected.
Ported To :  HP IBM Kubota Sun

slave slicer

Author :  Joseph A. Oswald, III  The slave slicer module extracts a 2D
Radiation Physics, Duke  slice of a 3D scalar field. The slice
University Medical  to be extracted is specified by the
Center  slice spec structure input. The
Number :  1709  slice spec structure will usually be
Language :  C  produced by a separate module that
Submitted :  09/30/93  contains control widgets for specifying
Version :  1.000  the slice in a way suitable for the
Last Updated :  09/30/93  given application. The “specify slice”
Ported To :  Kubota  module, for instance, creates a

slice spec scaled to the input data extents, and specified by Euler angles. Other modules could be written if another method of specification is desired. A module has been written, for instance, to take slices perpendicular to a beam of radiation used to treat cancer. The output is a 2D uniform field, with coordinates specified in the 2D coordinate system. The minimum and maximum X and Y coordinates correspond to the extents given in the slice spec. This data type is the most space-efficient representation for image data. If the coordinates of the data in the original 3D system are desired, the “reconstruct slice” module will reverse the transformation and produce a 2D 3-space irregular scalar field, suitable for visualization with “field to mesh.” The geometry output is a simple rectangular polygon corresponding to the size and position of the 2D slice in 3D space. It is useful for quickly depicting the slice in relation to other objects in 3D. The three modules in this package, “slave slicer,“ ”specify slice,“ and “reconstruct slice,” are a flexible and efficient alternative to “arbitrary slicer“ and Wes Bethel’s ”new arbitrary slicer.“ Slice specifications are generated by a module separate from the actual slicing module. That permits a single specifying module to specify identical slices through multiple data sets, and also for users to create their own modules to specify slices. The slicing algorithm takes 2D slices of 3D scalar fields (uniform or rectilinear) and outputs them in the most space efficient format, 2D uniform. 2D 3-space irregular output, similar to that output by “new arbitrary slicer“ can be generated, if needed, using “reconstruct slice.”

177

FILTER MODULES

smooth

Author :  anonymous  This module will take a 2 dimensional
Number :  1892  field, and perform smoothing operations
Language :  C  on it. The input field will be smoothed
Submitted :  09/07/95  in the Y direction, so the data is
Version :  1.000  expected to be in several independent
Last Updated :  09/07/95  columns (graph viewer format). There
Ported To :  DEC ultrix DEC alpha  are three smoothing methods; Average,

Median, and Weighted Average.

specify slice

Author :  Joseph A. Oswald, III  The specify slice module creates a
Radiation Physics, Duke  slice spec structure, suitable for use
University Medical  with the “slave slicer” module to
Center  extract 2D fields from 3D scalar
Number :  1707  fields. The 3D field can be of any data
Language :  C  type - the field is used only to scale
Submitted :  09/30/93  the extents so that the 2D slice will
Version :  1.000  cover the extents of the field. The
Last Updated :  09/30/93  slice orientation and resolution are
Ported To :  Kubota  controlled by the parameters. The X and

Y resolutions are identical, and the scaling of the slice is selected to cover the “shadow” of the 3D field in this orientation and maintain the “aspect ratio” of the input. That is, parallel slices made by varying the “Z Translate“ parameter will all be the same size, and the geometrical proportions of the input are preserved. The three modules in this package, “slave slicer,” “specify slice,” and “reconstruct slice,” are a flexible and efficient alternative to “arbitrary slicer“ and Wes Bethel’s ”new arbitrary slicer.“ Slice specifications are generated by a module separate from the actual slicing module. That permits a single specifying module to specify identical slices through multiple data sets, and also for users to create their own modules to specify slices. The slicing algorithm takes 2D slices of 3D scalar fields (uniform or rectilinear) and outputs them in the most space efficient format, 2D uniform. 2D 3-space irregular output, similar to that output by “new arbitrary slicer“ can be generated, if needed, using “reconstruct slice.” The slicing algorithm is generally faster than “new arbitrary slicer,” at least on the Kubota architecture. RELATED MODULES - slave slicer, reconstruct slice

spectest1D

Author :  International AVS Center  spectest1D is an avs module executable
North Carolina  which contains 3 modules hacked from
Supercomputing Center  khoros. Spectral Estimation. ** darpsd
Number :  1525  ** determines the 1D autoregressive PSD
Language :  C  and Phase spectral estimate ** dminorm
Submitted :  10/01/92  ** implements the minimum norm for
Version :  0.100  spectral estimation ** dmusic **
Last Updated :  10/01/92  implements the MUSIC for spectral
Ported To :  Sun DEC IBM  estimation

standard

Author :  International AVS Center  standard is an avs module executable
North Carolina  which contains 13 modules hacked from
Supercomputing Center  khoros. Standard File Formats. **
Number :  1526  fits2AVS ** converts a FITS image into
Language :  C  a field ** mat2AVS ** reads single
Submitted :  10/01/92  matrix from file specified and places
Version :  0.100  it in image ** pbm2AVS ** converts a
Last Updated :  10/01/92  PBM image into a field ** rast2AVS **
Ported To :  Sun DEC IBM  converts a SUN raster image into a

field ** tga2AVS ** converts TGA image into a field ** tiff2AVS ** converts TIFF image into a field ** AVS2fits ** converts a field into a FITS fiel format ** AVS2mat ** outputs a matrix in desired format ** AVS2pbm ** converts a field into a PBM image ** AVS2rast ** converts a field into a SUN raster image ** AVS2tiff ** converts a field into a TIFF image ** AVS2xbm ** converts a field into a x11 bitmap format (xbm) ** xbm2AVS ** converts an x11 bitmap file (xbm) into a field

stereogram

Author :  Larry Gelberg  STEREOGRAM produces a version of a
Advanced Visual Systems  “random dot stereogram” based on an
Number :  1700  image which represents distance to the
Language :  C  eye. An easy way to get this kind of
Submitted :  09/30/93  image is to use the Z-buffer output
Version :  1.000  from the AVS geometry viewer’s software
Last Updated :  09/30/93  renderer. A random dot stereogram is an
Ported To :  HP IBM Kubota Sun  image where if you diverge your eyes

while looking at it, the left and right eyes will see a different image and a three-dimensional surface will appear. Two squares are drawn in the top of the display. One method of viewing the stereogram would have you stare “past” the image so that the squares appear to overlap. Hold that for a few seconds and (for some people.) a shape will appear to either protrude from or recede into the image plane.

stick

Author :  David Haith  Sticks one image over another with a
Power Media  variable mix control, and with offset
Number :  1083  translation for image alignment. The
Language :  C  mix has further level key control. The
Submitted :  10/10/91  five parameters, all integer, are “x
Version :  1.000  shift“, ”y shift“, ”mix“, ”level“, and
Last Updated :  10/10/91  “key level”.
Ported To :  Kubota DEC IBM Sun HP

stream to scatte

Author :  Ian Curington  The stream to scatter module
Advanced Visual Systems  generates an output field with the
Number :  1684  coordinates of the stream path. This is
Language :  C  used to create other visual or
Submitted :  09/29/93  analytical processes at the positions
Version :  1.000  along the streamline path. Scatter
Last Updated :  09/29/93  markers could be drawn, or the scatter
Ported To :  HP IBM Sun  field could control the sampling

positions for other interpretation

modules such as hedgehog.

string parm to field

Author :  Joshua Goldstein  This is one of a suite of seven AVS/IDL
Research Systems  modules. The modules are all in the
Incorporated  FILTERS subdirectory at the IAC. They
Number :  1788  are IDL - a UNIX utility that gets IDL
Language :  C  and AVS communicating; and six AVS
Submitted :  02/15/94  modules that convert data between IDL
Version :  1.000  and AVS - field to float parm,
Last Updated :  02/15/94  field to int parm,
Ported To :  IBM Sun  field to string parm,

float parm to field, int parm to field, and string parm to field. These modules have only been tested on the RS 6000 and Sun Sparcstation. Binaries have been compiled on the HP 700 but its usability is unknown. Binary executables are stored at the IAC for all three of these platforms. There are also some demo files at the IAC, stored in the file demo.tar. You must be running IDL 3.1 or later in order to use these. No guarantee is made that this version of avsidl will work with future versions of AVS and/or IDL. In fact, I know that avsidl will be completely reworked for for IDL 4.0.

strip cfd

Author :  Rob Kuyper  strip cfd is a filter module which
Univeristy of Technology  changes a cfd dataset in order to
Delft, The Netherlands  remove the staggering of velocity
Number :  1319  components (in finite volume methods)
Language :  C  or to remove virtual points from the
Submitted :  08/30/92  dataset.
Version :  1.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

subregion

Author :  International AVS Center  subregion is an avs module executable
North Carolina  which contains 5 modules hacked from
Supercomputing Center  khoros. Subimage/Subsignal Operations.
Number :  1527  ** vextcent ** extracts rectangular
Language :  C  image from center of larger image **
Submitted :  10/01/92  vextract ** removes rectangular piece
Version :  0.100  from an image to produce a subimage **
Last Updated :  10/01/92  vinsert ** inserts subimage into the
Ported To :  Sun DEC IBM  input image to form output image **

vpad ** pads a data sequence with a contstant ** vsamp ** samples image along line, given image and a line through it

super sample

Author :  David Haith  Creates an anti-aliased (filtered and
Power Media  downsized) version of an input image.
Number :  1084  It used a simple box filter, averaging
Language :  C  pixels in the sub-region. The sample
Submitted :  10/10/91  ratio is 4 into 1. The output image is
Version :  1.000  1/2 size in each direction. The input
Last Updated :  10/10/91  image height and width must be a factor
Ported To :  Kubota DEC IBM Sun HP  of 2, or the indexing in this module

falls apart.

surface

Author :  International AVS Center  surface is an avs module executable
North Carolina  which contains 6 modules hacked from
Supercomputing Center  khoros. Surface Calculation and
Number :  1528  Modification. ** vcost ** computes
Language :  C  surface arc length from one pixel to
Submitted :  10/01/92  every other pixel ** vdetilt **
Version :  0.100  computes best-fit plane for image, then
Last Updated :  10/01/92  subtracts it from image ** vgettilt **
Ported To :  Sun DEC IBM  computes the least-squares best fit

plane for an image ** vslope **

computes the slope and aspect images from input elevation data ** vsurf ** takes a surface image and produces two output images ** vtilt ** adds specified plane to image to correct for a tilt in luminance

switch 2

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology two-module suite - 1) switchfire, which
/ Crump Institute  switches between two fields, and
Number :  1779  controls flow to downstream modules,
Language :  C  and 2) switch, which switches between
Submitted :  02/09/94  two fields to downstream modules.
Version :  1.000  Switch will be useful in keeping the
Last Updated :  02/09/94  flow of data controllable to bottleneck
Ported To :  HP Kubota Sun  points without requiring the user to

repeatedly connect and disconnect modules to control their source field. The module will send data downstream whenever the chosen input is updated, or the A/B switch is switched. Both inputs are not required for the module to execute. This module allows the user to control the flow of AVS fields to downstream modules. Often it may be desireable to switch the input path between modules, macros or whole branches of a network frequently without having to connect and disconnect inputs repeatedly. This is especially true in larger networks wherein it would be impractical to duplicate the downstream module or tree, and switching inputs may require stopping the flow executive to allow for time to make the change.

switchfire

Author :  A. Jacobson  This module was submitted as one of a
UCLA Dept. Pharmacology two-module suite - 1) switchfire, which
/ Crump Institute  switches between two fields, and
Number :  1778  controls flow to downstream modules,
Language :  C  and 2) switch, which switches between
Submitted :  02/09/94  two fields to downstream modules.
Version :  1.000  Switchfire will be useful in keeping
Last Updated :  02/09/94  the flow of data controllable to
Ported To :  HP Kubota Sun  bottleneck points without requiring the

user to repeatedly connect and disconnect modules to control their source field and execution. The module will switch between two input fields and send data downstream. In control mode, data is sent when triggered by the “send data” button, or, in continuous mode, whenever the chosen input is updated, the A/B choice is switched, or the send data button is triggered. Both inputs are not required for the module to execute. This module allows the user to control the flow of AVS fields to downstream modules. Often it may be desireable to switch the input path between modules, macros or whole branches of a network frequently without having to connect and disconnect inputs repeatedly. Similarly, it is often desireable to stop, start, or trigger data flow and execution to individual modules or branches of a network. This is especially a problem with modules that overwrite existing files whenever they are updated, or are time consuming to run when frequently not necessary.

sysModuleName

Author :  John Stephen  sysModuleName - given the alias name of
Tessella Support  any module in the network this utility
Services  outputs the system’s name for the
Number :  1686  module. This will return the system
Language :  C  name for the module when its alias is
Submitted :  09/29/93  typed into the parameter. For example
Version :  1.000  if you use AVS CLI - “avs¿ module
Last Updated :  09/29/93  “geometry viewer” -alias geomV“, the
Ported To :  HP IBM Sun  system might name the instance of the

geometry viewer as “geometry viewer“.user.17. The module can be referenced by this name or by its alias “geomV” for most purposes, but some commands will only accept the system name. This module will get the system name from the alias which makes an application more robust. Thus in the example, if “geomV” is entered into the parameter, “geometry viewer”.user.17 is output from the module.

table editor

Author :  Ian Curington  Like a mini-spreadsheet application,
Advanced Visual Systems  this module presents a segment of a 2D
Number :  1538  array and allows interactive editng of
Language :  C  the field values. Takes an optional
Submitted :  10/03/92  input field, displays the contents as a
Version :  1.000  2D array table of button cells. Each
Last Updated :  10/03/92  cell or group of cells may be selected
Ported To :  Sun DEC HP  (high-lighted) and new values entered.

The modified field is then sent to the output port. Values can be replaced, added, multiplied, or divided into existing field values. The code shows how to create a page layout from within the module code. With no input port, this module can be used to synthesize new fields. If the input field is multi-vector, each vector can be selected from a choice list, and values can be edited in each channel, while synthetic data can only be created in scalar mode.

thresholdclamp

Author :  A. Jacobson  This module will threshold or clamp the
UCLA Dept. Pharmacology values of a field with or without a 2
/ Crump Institute  or 3D ROI. This module does a few
Number :  1858  things that all preexisting AVS filters
Language :  C  frustratingly don’t do. 1. Thresholding
Submitted :  10/17/94  (segmentation) can be done on a field,
Version :  1.000  setting all values below Min to 0, and
Last Updated :  10/17/94  all values above Max to 255 , with
Ported To :  DEC alpha HP IBM Sun  all values inbetween unchanged. This

can be used to create 3 data ranges, 0, min¡range¡max, and 255. 2. Clamping is performed such that all values below Min are set to Min, except zero, and all values above Max, set to Max.

3. ROI’s can be applied to this data. Any point that is not 0 in the ROI field is filtered in the data field. Areas that are 0 in the ROI can either be left unchanged or set to 0. This module will take either a 2D ROI, applying the ROI to each “Z” plane if a 3D data field is used, or a 3D ROI, applying each point of the ROI field to the data field. The ROI field must have the same dimensions as the data field, either in x,y (if 2D ROI) or x,y,z (if 3D).

transform1D

Author :  International AVS Center  transform1D is an avs module executable
North Carolina  which contains 6 modules hacked from
Supercomputing Center  khoros. Transforms (1D). ** dfft1d **
Number :  1529  performs the 1D Fourier transform on a
Language :  C  set of data sequences ** dfht **
Submitted :  10/01/92  performs a 1-D Fast Hadamard Transform
Version :  0.100  on a signal ** dhart1d ** performs the
Last Updated :  10/01/92  1-D forward Hartley transform of a
Ported To :  Sun DEC IBM  sequence ** difft1d ** performs 1-D

inverse Fourier transform on set of data sequences ** dihart1d ** performs 1-D inverse Hartley transform of a data sequence ** dmpp ** determines the magnitude, power, or phase of each 1-D data point

FILTER MODULES

transform2D

Author :  International AVS Center  transform2D is an avs module executable
North Carolina  which contains 3 modules hacked from
Supercomputing Center  khoros. Tranforms (2D). ** vfft **
Number :  1530  computes the 2 dimensional Fast Fourier
Language :  C  Transform of an image ** vfht **
Submitted :  10/01/92  performs a fast hadamard transform on
Version :  0.100  the image ** vmpp ** computes various
Last Updated :  10/01/92  representations of complex image data
Ported To :  Sun DEC IBM

transform po

Author :  Anonymous  Performs a scale and offset of each
Convex Computer  dimension in coordinate space,
Number :  1148  according to the input parameters
Language :  C  “scale x”, “offset x”, “scale y”, and
Submitted :  03/10/92  “offset y”. It handles 1D, 2D, or 3D
Version :  1.000  input fields, with rectilinear or
Last Updated :  03/10/92  irregular mapping. The input data may
Ported To :  Kubota DEC IBM Sun HP  have any vector length, but must be

non-uniform. The output field has the same dimensionality as the input field.

transpose coords

Author :  Phil McDonald  The transpose coords module provides
NOAA/ERL/Forecast  the capability of transposing the
Systems Laboratory  coordinates of non-uniform fields. This
Number :  1269  capability is convenient for making a
Language :  C  vertical slice horizontal, without
Submitted :  05/27/92  having to resort to the transformation
Version :  1.000  capabilities of the render geometry
Last Updated :  05/27/92  module.
Ported To :  Kubota

tubeness

Author :  Wes Bethel  The tube module transforms an AVS
Lawrence Berkeley  geometry, replacing a set of disjoint
Laboratory  lines with “tubes.” The tubes are
Number :  1791  polygonal approximations to cylinders.
Language :  C  The number of sides used to represent
Submitted :  04/21/94  the cylinder is under user control.
Version :  2.000  Version 2 update - those line segments
Last Updated :  07/15/94  which are “degenerate” will be cause a
Ported To :  DEC HP IBM Sun  sphere of radius “radius scale” to be

generated rather than a flat cylinder.

ucdTri2prism

Author :  Ian Curington  ucd tri to prism - extrudes
Advanced Visual Systems  triangulated surface into 3D prism
Number :  1757  layers, given triangular 2D UCD
Language :  C  elements and a 2D field, loft the
Submitted :  12/15/93  elements into 3D prism elements, with
Version :  1.000  information for the 3rd dimension
Last Updated :  12/15/93  coming from a 2D field, where the first
Ported To :  Kubota Sun  dimension matches the input UCD node

list, the second dimension replicates the structure, adding a Z offset and new data.

ucd extent range

Author :  Ian Curington  The ucd extent range module converts
Advanced Visual Systems  model coordinate ucd extent positions
Number :  1688  into normalized dial range for
Language :  C  ucd slice 2d, etc. It has 3 dials,
Submitted :  09/29/93  which have model coordinate ranges,
Version :  1.000  these are mapped to the the extent in
Last Updated :  09/29/93  X, Y, Z, and normalized -1 to 1 range
Ported To :  HP IBM Sun  floats are output, with appropriate

correction factors to drive

ucd slice2d.

ucd load node da

Author :  Advanced Visual Systems, ucd load node data adds a field of
Inc.  data to a UCD structure. It takes in a
Number :  1689  field of floats as UCD data, and a
Language :  C  field of ints as node ids. It also
Submitted :  09/29/93  takes in an existing UCD structure. For
Version :  1.000  each node id in both the field and the
Last Updated :  09/29/93  UCD it adds the extra data. Hence, not
Ported To :  HP IBM Sun  all of the field float data will

necessarily be added; nor will all of the UCD nodes be guaranteed to receive data (if they do not get data, zeros are inserted).

ucd minmax

Author :  Ian Curington  For Transient UCD data, where data
Advanced Visual Systems  ranges are changing, and/or the mesh is
Number :  1285  static, this module copies UCD input to
Language :  C  UCD output, with almost everything
Submitted :  06/20/92  unchanged, except for the data minimum
Version :  1.000  and maximum header values, and the
Last Updated :  06/20/92  mesh id. This is picked up by
Ported To :  IBM Sun Kubota HP  ucd contour or ucd legend, to colour

the data, using user-specified values for min and max, as set on the parameter dials. This is important for animation, as the colour range must remain constant over time varying data. This module will work with either Node or Cell based data. Make sure the range is outside the actual data values.

ucd particle

Author :  Ian Curington  This module animates spheres moving
Advanced Visual Systems  along the path created by either UCD
Number :  1189  streamlines or the field “streamlines”
Language :  C  module. It post-processes the path
Submitted :  05/04/92  trajectory data, and moves particles
Version :  3.000  along each path. This module will work
Last Updated :  12/15/93  with either UCD or field data, because
Ported To :  Sun Kubota  it inputs neither, it simply uses the

other integration modules to create the paths. The animation will be faster than “particle advector”, because it is not performing the path integration at the same time as the animation. The module has full up and downstream reverse transformation pass-through, so interactive positioning of the streamline control icon is still posible. This module is intended as a source example for other UCD module developers. Version 3.000 update (provided by Mike Rangitsch of Dow Chemical) adds a second (optional) input and moving particle tails.

ucd polar

Author :  anonymous  This module reads in a UCD structure
Number :  1894  and locates all components that have a
Language :  C  vector length of 3. Those data vectors
Submitted :  09/07/95  are then converted from x,y,z format to
Version :  1.000  r,theta,z format. Only those data
Last Updated :  09/07/95  values are changed. Oytherwise the ucd
Ported To :  DEC ultrix DEC alpha  is exactly like the original. This

module always calculates theta around the Z axis. If the user changes the Polar Origin parameters, then the polar coordinates will be computed around this new Origin.

ucd select

Author :  Jonathan Cox  The ucd select adjusts the minimum and
Department of Computer  maximum values of a colormap to those
Science, University of  of a selected scalar component of UCD
Manchester  nodal data. The module is functionally
Number :  1713  similar to the color range module which
Language :  C  performs the same operation for fields.
Submitted :  09/30/93  However, as ucd contour already
Version :  1.000  normalises the colormap this module is
Last Updated :  09/30/93  not required to correctly color ucd
Ported To :  HP  structures. Instead, it is intended

that this module be used to provide a alternative mechanism for using the color legend and label modules with UCD data to that offered by the ucd minmax module. In this respect the ucd color range module offers a superset of the functionality of the ucd minmax module - filtering the ucd input. The ucd minmax module relies on network connections to the generate colormap module in order to set the bounds of the colormap. In contrast ucd select filters the colormap output. This colormap can then be passed to the color legend module. There are two principal advantages to using this module over ucd minmax. Firstly, the module offers more control over what type of changes in the ucd input cause the min/max values to be reset. The other major advantage is that the chosen component is made available on an output port. This can then be used to control the ’Node Data’ parameters of other ucd modules. This makes it possible to select a component throughout a network at a single stroke. This makes it much easier to design ucd networks with a clean user interface. The other output of the module is a title string which can be used with the label module.

ucd split

Author :  John Stephen  This module outputs a portion of the
Tessella Support  input UCD, by selecting lists of nodes,
Services  cells, from an integer field, or by
Number :  1691  typing in cell ranges.
Language :  C
Submitted :  09/29/93
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Kubota Sun

ucd to field

Author :  Ian Curington  This is a UCD filter to convert node
Advanced Visual Systems  data, component lists, node id’s to AVS
Number :  1692  field and string structures for
Language :  C  external processing. The data on the
Submitted :  09/29/93  nodes of the UCD model are brought out
Version :  1.000  to a floating field port, where the
Last Updated :  09/29/93  data can be further processed such as
Ported To :  HP IBM Kubota Sun  with extract scalar, field math, or

statistics, or sent to a graph viewer. The difference between the nodal values on two different UCD structures could be computed, for instance. In “irregular” mode, the output value field is in “scatter” form, so spheres can be placed at node positions with bubbleviz and scatter dots. In addition, the node id’s are converted to an integer field, and the component list is converted to and integer field, with component names and units converted to string output ports. This is very handy for debugging UCD applications, for processing UCD data using field modules, or other user written modules. Note that this is not the compliment of field to ucd, as this is topologicall y impossible.

ucd tri to prism

Author :  Ian Curington  This is a UCD filter to extrude 2D
Advanced Visual Systems  triangle UCD’s to 3D prisms. Given
Number :  1693  triangular 2D UCD elements and a 2D
Language :  C  field, loft the elements into 3D prism
Submitted :  09/29/93  elements, with information for the 3rd
Version :  1.000  dimension coming from a 2D field, where
Last Updated :  09/29/93  the first dimension matches the input
Ported To :  HP  UCD node list, the second dimension

replicates the structure, adding a Z offset and new data. The control field contains all information to create the structure, both coordinates and data, except for cell topology, which comes from the input 2D ucd. This is intended to be generated by a module like scat to tri (delaunay triangulation). Each “row” of the input field represents one 2D surface, where the 4-vector information is - 1) X-Loc, 2) Y-Loc, 3) Time (depth), and 4) Velocity. The X-Y pairs are un-ordered, and represent random surface positions.

un jag

Author :  Tim O. Robinson  This module searches for jagged regions
UC Berkeley  and performs local averaging on pixels
Number :  1275  located at or near a “jag.” A pixel is
Language :  C  determined to be at a jag if either
Submitted :  05/27/92  red, green, or blue values differ from
Version :  1.000  one (of two) surrounding pixel r,g,b
Last Updated :  05/27/92  values by a threshold amount in both
Ported To :  DEC IBM Sun Kubota HP  the x and y directions. The threshold

value is set at 25 in the module. There is no widget for this parameter because adjustment is not required during normal operation. This module is intended for final processing of image data which is limited by screen resolution before printing, its research value is dubious.

unif2irreg

Author :  John Tee  This module takes an input field, such
AVS Inc. Consultant  as a 1D n-vector uniform field, and
Number :  1694  interactively decide which “column” or
Language :  C  channel should be used for each of X,
Submitted :  09/29/93  Y, and Z coordinates of a scatter
Version :  1.000  field. Any remaining quantities, if the
Last Updated :  09/29/93  field had more than 3-vectors, are
Ported To :  HP IBM Kubota Sun  placed on the output field as an n-3

vector field. This can be used to create multidimensional scatter plots, showing channel correlation by spatial clustering. It is much easier to import uniform fields using read field or the ADIA, then use this module to create a scatter, than to define a scatter based on the external file format. This module will work with byte, float, int, double, or short data types.

unif2zscat

Author :  Ian Curington  unif2zscat - uniform 1D to Z-line
Advanced Visual Systems  scatter. This module takes an input
Number :  1848  field, such as a 1D n-vector uniform
Language :  C  field, and creates a new field with all
Submitted :  07/16/94  the same data values, vector length,
Version :  1.000  etc., but in addition adds irregular
Last Updated :  07/16/94  coordinates, creating a scatter field.
Ported To :  DEC HP  The coordinates added have X=0, Y=0,

and Z changing as a function of the data index with uniform spacing. This module is a simple filter used to test the “scatter ribbon” module, for a well log display. This module will work with byte, float, int, double, or short data types.

uniform to i

Author :  Larry Gelberg  Converts a uniform field data set to an
Stardent Computer  identical irregular one. It does this
Number :  1089  by creating explicit 3D coordinates for
Language :  C  all the nodes of the field. This is
Submitted :  10/10/91  sometimes desirable for doing things
Version :  1.000  which the uniform field doesn’t allow.
Last Updated :  10/10/91  In the documentation accompanying the
Ported To :  Kubota DEC IBM Sun HP  module, there is a sample network shown

which helps get around the fact that orthogonal slice creates a slice in 2d instead of 3d.

vec2 to vec3

Author :  Chris Allen  Vec2 to vec3 converts the input field
Unisys Corporation  2-vector into a similar field, with the
Number :  1291  only difference being the data has been
Language :  C  moved into a type 3-vector structure,
Submitted :  07/11/92  with the extra vector component
Version :  1.000  containing a 0. This module is useful,
Last Updated :  07/11/92  for example, if you have a uniform
Ported To :  IBM Sun HP  field of 2-d vectors and would like to

view with Hedgehog, which requires 3-vector uniform data.

vec mag 2d v

Author :  Wes Bethel  Computes a two dimensional gradient
Lawrence Berkeley  from a two dimensional scalar, floating
Laboratory  point field (image). This operation is
Number :  1090  useful in many image processing
Language :  C  applications. There are no parameters
Submitted :  10/10/91  with this module. The documentation
Version :  1.000  accompanying the module shows a network
Last Updated :  10/10/91  that reads in an image (a good test
Ported To :  Kubota DEC IBM Sun HP  case is the “stardent.x” image), then

extracts one of the rgba fields from the image, then computes a 2d gradient from the resulting gray image. Next, they extract either the x- or y-component from the gradient field and display the result as a grayscale image. Switching from channel 0 to channel 1 in extract scalar shows the y-component of the gradient.

INTERNATIONAL AVS CENTER MODULE CATALOG

vect div scal 2

Author :  Adam Usadi  Divides every element of a vector field
Rice University  by a scalar field
Number :  1870
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix DEC alpha

vect div scal 5

Author :  Adam Usadi  Divides every element of a vector field
Rice University  by a scalar field
Number :  1866
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix

vect div scalar

Author :  Adam Usadi  Divides every element of a vector field
Dept. of Space Physics  by a scalar field

and Astronomy Rice

Number :  1862
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix

vector

Author :  International AVS Center  vector is an avs module executable
North Carolina  which contains 1 module hacked from
Supercomputing Center  khoros. Multiband Conversion. **
Number :  1531  vmapdlg ** converts the maps of a field
Language :  C  created using dlg2AVS from actual
Submitted :  10/01/92  geographical attributes to rgb of
Version :  0.100  displayable color
Last Updated :  10/01/92
Ported To :  Sun DEC IBM

FILTER MODULES

vector cross

Author :  Adam Usadi  This module computes either a 3D or 2D
Dept. of Space Physics  vector cross product. You should be
and Astronomy, Rice  able to modify this to allow the
University  calculation of a dot product also.
Number :  1728
Language :  C
Submitted :  12/14/93
Version :  1.000
Last Updated :  12/14/93
Ported To :  IBM Kubota Sun

vector cross 2

Author :  Adam Usadi  INPUTS First Vector - float field of
Dept. of Space Physics  space, any size , any veclen REQUIRED
and Astronomy Rice  Second Vector - float field of any
Number :  1861  space, any size , any veclen REQUIRED
Language :  C  PARAMETERS OUTPUTS Cross Product -
Submitted :  06/27/95  input 3-vectors yield 3-vector output.
Version :  1.000  input 2-vectors yield 1-vector output
Last Updated :  06/27/95  in the third direction
Ported To :  DEC ultrix

vector cross 3

Author :  Adam Usadi  Computes the vector cross product of
Rice University  two float fields
Number :  1864
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix

vector cross 4

Author :  Adam Usadi  Computes the vector cross product of
Rice University  two float fields
Number :  1865
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix

INTERNATIONAL AVS CENTER MODULE CATALOG

vector cross 5

Author :  Adam Usadi  Computes the vector cross product of
Rice University  two float fields
Number :  1869
Language :  C
Submitted :  06/27/95
Version :  2.000
Last Updated :  06/27/95
Ported To :  DEC ultrix DEC alpha

vector dot

Author :  Adam Usadi  Computes the vector dot product of any
Dept. of Space Physics  two float fields of any size and any
and Astronomy, Rice  field coordinates.
Number :  1863
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95

vector dot 3

Author :  Adam Usadi  Computes the vector dot product of any
Rice University  two float fields of any size and any
Number :  1871  field coordinates.
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix DEC alpha

vector dot 6

Author :  Adam Usadi  Computes the vector dot product of any
Rice University  two float fields of any size and any
Number :  1867  field coordinates.
Language :  C
Submitted :  06/27/95
Version :  1.000
Last Updated :  06/27/95
Ported To :  DEC ultrix

vector field

Author :  Terry Myerson  Computes a vector field between two
International AVS Center  planes in physical space. The two
(NCSC)  inputs ports define a initial and
Number :  1150  terminating plane, perpendicular to the
Language :  C  y axis. The vector field that is output
Submitted :  03/10/92  is constructed such that the
Version :  1.000  corresponding corners of each plane
Last Updated :  03/10/92  have vectors pointing to the other one.
Ported To :  Kubota DEC IBM Sun HP  The input data sets consist of an x, y,

z triplet in positions 0, 1, and 2 in the input array. The triplet defines the xz planes that the particles advect from/to.

warp

Author :  International AVS Center  warp is an avs module executable which
North Carolina  contains 4 modules hacked from khoros.
Supercomputing Center  Warp Images. ** vbilinco ** computes
Number :  1532  the coefficients for the two bilinear
Language :  C  equations ** vexpwarp ** alters the
Submitted :  10/01/92  spatial properties of an image **
Version :  0.100  vpolyco ** computes a set of
Last Updated :  10/01/92  bi-polynomial coefficients ** vwarp **
Ported To :  Sun DEC IBM  alters the spatial properties of an

image

xform field

Author :  Wes Bethel  Applies a transformation, as specified
Lawrence Berkeley  from the parameters “X Rotate”, “Y
Laboratory  Rotate“, ”Z Rotate“, ”Ortho Scale“, ”X
Number :  1038  Xlate“, ”Y Xlate“, ”Z Xlate“, and ”flip
Language :  C  Z“. The transformation is applied to
Submitted :  10/10/91  each and every grid point in a 3D, 3
Version :  2.000  space irregular field. This module was
Last Updated :  10/10/91  designed to be used with the
Ported To :  Kubota DEC IBM Sun  curvilinear volume renderer, “bezier

volume.“

INTERNATIONAL AVS CENTER MODULE CATALOG

y join

Author :  David Haith  Joins up two images, with a split line
Power Media  running horizontally. The position of
Number :  1094  the split is controlled by the
Language :  C  parameter “y value”. Both input images
Submitted :  10/14/91  (field 2D 4-vector byte) must be the
Version :  1.000  same size.
Last Updated :  10/14/91
Ported To :  Kubota DEC IBM Sun HP

FILTER MODULES 202

MAPPER MODULES

International AVS Center

MAPPER MODULES

2D bounds

Author :  KGT (Kubota Graphics  The “2D bounds” module generates lines
Technologies) Inc.  that indicate the bounding box of 2D
Number :  1876  field. It is used to get the 2D mesh of
Language :  CandFORTRAN  3D field. For example, you can use this
Submitted :  09/06/95  module to check CFD grid.
Version :  1.000
Last Updated :  09/06/95
Ported To :  DEC alpha

3D contour map

Author :  Krzysztof (Chris) S.  The 3D contour map routine creates
Nowinski  contour map of a given 2D scalar float
Institute of Applied  field together (optionally) with a
Mathematics and  surface equivalent to the surface
Mechanics, Warsaw  generated by the field to mesh routine.
University  The algorithm used is a subset of the
Number :  1781  more familiar 3d marching cubes
Language :  C  algorithm. This is a modification of
Submitted :  02/15/94  the field to con module by Wes
Version :  1.000  Bethel.
Last Updated :  02/15/94
Ported To :  HP Sun

Analog Clock

Author :  Lars M. Bishop  The Analog Clock module is a
National Center for  3-dimensional, graphical readout that
Comp. Elect., Univ. of  presents a “stopwatch” or “piechart”
Illinois  dial display. It is designed to be
Number :  1626  connected to an integer parameter whose
Language :  C  value is changing over time. The Analog
Submitted :  07/02/93  Clock provides a qualitative way of
Version :  1.000  seeing the value change. By setting the
Last Updated :  07/02/93  cyle length of the Analog Clock
Ported To :  DEC IBM Kubota  intelligently, the input parameter’s

change over time can be elucidated. For example, if the cycle length is set to 100, and the Max Value is set to 1000, the fast indicator will cycle for every 100, and the slow indicator will cycle at one tenth this speed. The Analog Clock produces two types of “clock” Pie Chart - This counter uses a “cheese wheel“-shaped clock as its fast indicator, and a change in color of the cheese wheel as its slow indicator. In this version, the colors wrap around past each multiple of Max Value. Analog Clock - This counter uses a standard two-handed analog clock face (without markings) as its fast and slow indicators.

CFD object

Author :  KGT (Kubota Graphics  The “CFD object” module inputs a scalar
Technologies) Inc.  field and creates a geometry(Non-fluid
Number :  1877  Area) for CFD data. This module tells
Language :  FORTRAN  which digit (0 or 1) is in the node as
Submitted :  09/07/95  flag data. The flag data of the nodes
Version :  1.000  is as follows. 1 means the fluid node
Last Updated :  09/07/95  and 0 is the non-fluid node.
Ported To :  DEC alpha

MAPPER MODULES

CFD object conto

Author :  KGT (Kubota Graphics  The “CFD object contour” module inputs
Technologies) Inc.  2-vector field and creates a contour
Number :  1878  surface(on obstacle) for CFD data. This
Language :  FORTRAN  module tells which digit (0 or 1) is in
Submitted :  09/07/95  the node as flag data. The first
Version :  1.000  component data (of 2-vector) is the
Last Updated :  09/07/95  scalar value for color, and the second
Ported To :  DEC alpha  component is the flag data. The flag

data is as follows. 1 means the fluid node and 0 is the non-fluid node.

CFD objectline

Author :  KGT (Kubota Graphics  The “CFD objectline” module inputs a
Technologies) Inc.  scalar field and creates a line
Number :  1879  geometry (obstacle) for CFD data. This
Language :  FORTRAN  module tells which digit (0 or 1) in
Submitted :  09/07/95  the node as flag data. The flag data is
Version :  1.000  as follows. 1 means the fluid node and
Last Updated :  09/07/95  0 is the non-fluid node.
Ported To :  DEC alpha

ChairPlot 2D Txt

Author :  Scott Goodyear (AVS  The ChairPlot module is another
Inc.), Ian Curington  alternative way of visualizing 3D
(AVS Inc.), and Mike  volume data. The volume is displayed
French (E and S)  with 3 orthogonal slice planes, with
Number :  1641  the exterior boundary filled in behind,
Language :  C  so it looks like a section has been
Submitted :  09/24/93  removed from a corner of the volume. It
Version :  1.000  is similar to excavate brick, except
Last Updated :  09/24/93  that it makes 2D polygons, with UV
Ported To :  HP IBM Sun  texture coordinates and resamples the

input data into 2D images that glue onto the slice sections. This way, 2D texture map facilities are used, rather than 3D texture. Several hardware vendors, notably EandS, SGI, Kubota, Vistra, etc., have hardware support for this type of texture mapping. The software renderer also has support for this mode, where no accelerator is available. The controls are virtually identical to excavate brick.

INTERNATIONAL AVS CENTER MODULE CATALOG

Colored bounds

Author :  D. LE CORRE  This mapper module generates a geometry
CONVEX S.A  consisting of 6 orthogonal slices (2 in
Number :  1200  each dimension) from a 3D scalar field.
Language :  C  By default, the 6 ortho slices are
Submitted :  05/08/92  placed at the hull of the input field.
Version :  2.000  The faces of this geometry are colored
Last Updated :  08/30/92  from the scalar field value.
Ported To :  DEC IBM Kubota HP

Contour Lines

Author :  KGT (Kubota Graphics  The “Contour Lines” module inputs a 2D
Technologies) Inc.  field and colormap. This module looks
Number :  1880  like the “contour to geom” module. It
Language :  CandFORTRAN  finds and creates 2D contour lines of
Submitted :  09/07/95  similar value.
Version :  1.000
Last Updated :  09/07/95
Ported To :  DEC alpha

Contour Mesh

Author :  KGT (Kubota Graphics  The “Contour Mesh” module inputs a 2D
Technologies) Inc.  field and colormap. This module looks
Number :  1881  like the “field to mesh” module. This
Language :  FORTRAN  module creates the contour only Fluid
Submitted :  09/07/95  AREA (of CFD).
Version :  1.000
Last Updated :  09/07/95

Draw Text

Author :  Ronald E. Stogner,  Creates a text geometry object from an
Computer Sciences  Label String input. A user can also
Corporation  supply an optional Hour as an integer,
Number :  1736  which will be added into the label.
Language :  C  This is useful if you are animating a
Submitted :  12/15/93  time series of data. This is one of a
Version :  1.000  suite of modules we used to plot the
Last Updated :  12/15/93  emission data for the state of North
Ported To :  Kubota Sun  Carolina. This group of modules allows

displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Field To Lines

Author :  Jeff Wang  This module allows users to draw lines
MCNC, Environmetal  in the geometry viewer. The value of
Programs Group  the line is mapped into color by the
Number :  1737  colormap of AVS. The input field is a
Language :  C  1D 6 space field with (x0, y0, z0, x1,
Submitted :  12/15/93  y1, z1, value). The denomitor “0” means
Version :  1.000  the starting points of the links, while
Last Updated :  12/15/93  “1” means the ending points of the
Ported To :  Kubota Sun  links. This is one of a suite of

modules we used to plot the emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Field To Points

Author :  Jeff Wang  This module allows users to draw points
MCNC, Environmetal  in the geometry viewer. The value of
Programs Group  the point is mapped into color by the
Number :  1739  colormap of AVS. The input field is a
Language :  C  1D 3 space field with (x, y, z, value).
Submitted :  12/15/93  This is one of a suite of modules we
Version :  1.000  used to plot the emission data for the
Last Updated :  12/15/93  state of North Carolina. This group of
Ported To :  Kubota Sun  modules allows displaying points, lines

and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Field To Polygon

Author :  Jeff Wang  This module allows users to draw shaded
MCNC, Environmetal  polygons in the geometry viewer. The
Programs Group  value of the polygo is mapped into
Number :  1738  color by the colormap of AVS. The input
Language :  C  field is a 1D 4 space field with (x, y,
Submitted :  12/15/93  z, n, value), where (x, y, z) is the
Version :  1.000  coordinate of a vertex in the polygon,
Last Updated :  12/15/93  n is the number of vertices in the
Ported To :  Kubota Sun  polygon, value is the value assign to

the polygon. This is one of a suite of modules we used to plot the emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Field grid

Author :  Penny Rheingans  Generates a labelled grid and/or axes
US EPA Scientific  for field data. The extents of the grid
Visualization Center  or axes are the extents of the data (or
Number :  1307  slightly larger).
Language :  C
Submitted :  08/30/92
Version :  1.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

Field to Isosurf

Author :  Chin-I Huang  This module reads a uniform 3D volume
National Chiao Tung  data and passes a threshold level to it
University, Taiwan,  to generate isosurfaces. The algorithm
Republic of China  is based on the Marching Cubes Method
Number :  1630  from Lorensen. Any interesting
Language :  C  threshold level ranged from 0 to 255
Submitted :  07/02/93  can be input and the module will
Version :  1.000  generate isosurfaces corresponding to
Last Updated :  07/02/93  that level. After the generation of
Ported To :  HP IBM DEC Sun  isosurfaces, you can see the number

counts of each table case on the console. Colormap is used for rendering the volume with one color level to a threshold. The module is limited to the data type of char, but the reader can change it into the other data type easily by modifying the source code a little.

GRAPH2D

Author :  Kathy Pearson  GRAPH2D – generate an X-Y plot,
MCNC Environmental  control graph viewer parameters, and
Programs  optionally output PostScript. The
Number :  1800  GRAPH2D module is an entire graphics
Language :  C  management system for 2d plots. It
Submitted :  04/21/94  creates a network including the AVS
Version :  1.000  “graph viewer” and “image to
Last Updated :  04/21/94  postscript“ modules, and it generates
Ported To :  DEC HP IBM Kubota Sun  script files that are played to make

ASCII files required as input to the graph viewer. Up to four 2d windows may be controlled by the module, and the settings will apply to the current window of focus, selected by the user and indicated by the red color band around the window. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

MAPPER MODULES

Get Local IP Num

Author :  Jeff Wang  This module allows users to get the
MCNC, Environmetal  internet IP number for their local
Programs Group  machine. This module is useful when it
Number :  1722  is connected with an editor in the
Language :  C  remote machine. See the related module
Submitted :  09/30/93  Edit File.
Version :  1.000
Last Updated :  09/30/93
Ported To :  DEC IBM Kubota

GraphSound

Author :  Brian Kaplan  The GraphSound module generates a graph
Center for Innovative  of the waveform of a sound produced by
Computer Applications  the Sonify module and is useful as a
Number :  1636  diagnostic tool.
Language :  C
Submitted :  07/05/93
Version :  1.000
Last Updated :  07/05/93

Grid 2 TxtMesh

Author :  Advanced Visual Systems, The texture mesh module serves two
Inc.  needs. First, the need for a robust
Number :  1834  module for mapping 2D data to
Language :  C  quadrilateral meshes is served. The
Submitted :  07/16/94  input 2D data may be a 2D uniform
Version :  1.000  field, in which case the x,y
Last Updated :  07/16/94  coordinates for the output mesh is
Ported To :  HP IBM Sun  derived from the extents information

(if present) or using a mapping from indeces to space if no extents information is present. The z values are obtained from the data in the field. The second need fulfilled by this module is support for texture mapping. Two types are supported. The first, called “dynamic” in AVS parlance, is the traditional type of texture mapping where the coloring of the surface is postponed until rendering time. The second type, called “explicit” here will color the vertices of the field based upon the image on the second input port, along with the uv coordinate ranges specified by the user.

Isoband Slice

Author :  Ian Curington (AVS  A set of contour bands are drawn on a
Inc.), and Mike French  2D slice through a 3D field. The field
(E and S)  input must be a slice through a uniform
Number :  1749  or irregular 3D grid. If the field is
Language :  C  uniform, the plane of the output
Submitted :  12/15/93  contour geometry will have its correct
Version :  1.000  position in 3D space.
Last Updated :  12/15/93
Ported To :  Sun

MAPPER MODULES

Isoband UCD

Author :  Ian Curington (AVS  A set of contour bands are drawn on the
Inc.), and Mike French  exterior faces of a UCD. The UCD must
(E and S)  have a single scalar component of node
Number :  1760  data. To extract a single scalar
Language :  C  component of node data from a composite
Submitted :  12/15/93  ucd use the UCD EXTRACT module. There
Version :  1.000  is no module available to extract a
Last Updated :  12/15/93  scalar from a vector component of node
Ported To :  Kubota Sun  data.

Isoline Slice

Author :  Mike French  A set of contour levels are drawn on a
Evans and Sutherland UK  slice through a field. The field input
Number :  1750  must be a slice through a uniform or
Language :  FORTRAN  irregular 3D grid. If the field is
Submitted :  12/15/93  uniform, the plane of the output
Version :  1.000  contour geometry will have its correct
Last Updated :  12/15/93  position in 3D space.
Ported To :  Kubota

Isoline UCD

Author :  Ian Curington (AVS  A set of contour lines are drawn on the
Inc.), and Mike French  exterior faces of a UCD. The UCD must
(E and S)  have a single scalar component of node
Number :  1765  data. To extract a single scalar
Language :  C  component of node data from a composite
Submitted :  12/15/93  ucd use the UCD EXTRACT module. There
Version :  1.000  is no module available to extract a
Last Updated :  12/15/93  scalar component from a vector
Ported To :  Kubota  component.

Keyframe Ani

Author :  Brian Kaplan  The Keyframe Animator module is used to
Center for Innovative  animate objects which are displayed
Computer Applications  through the Geometry Viewer subsytem.
Number :  1133  It works by supplying a tranformation
Language :  C  matrix to the Geometry Viewer module
Submitted :  02/23/92  for any object which has been loaded
Version :  3.000  into the Geometry Viewer or supplied to
Last Updated :  07/05/93  the Geometry Viewer’s geom input.
Ported To :  Kubota  Generally the Keyframe Animator‘s

geometry output (the transformation matrix) is supplied to the geometry input of the Geometry Viewer module. This is an update from version 2.4 to 3.0 Beta.

LABELS UNITS

Author :  Kathy Pearson  LABELS UNITS – create a label based
MCNC Environmental  upon the input field and input slice
Programs  axis and output a label geometry. The
Number :  1802  MCNC Environmental Programs Group has
Language :  C  created this module as part of a suite
Submitted :  04/21/94  of AVS demo modules, networks, and
Version :  1.000  sample data sets to demonstrate the
Last Updated :  04/21/94  functionality of its visualization
Ported To :  DEC HP IBM Kubota Sun  prototype for an Environmental Decision

Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

LBL bezier v

Author :  Wes Bethel  The bez vol module is an object order
Lawrence Berkeley  volume renderer designed to work with
Laboratory  volumes embedded in warped grids. This
Number :  1016  module assumes that voxels are
Language :  C  topologically equivalent to hexahedra,
Submitted :  10/10/91  the faces of which are not necessarily
Version :  3.000  planar. This solid is manipulated and
Last Updated :  05/19/92  rendered using tricubic Bezier basis
Ported To :  IBM Sun  and rendered using a forward difference

evaluation of the cubic triparametric Bezier representation of a solid.

MASK TILE

Author :  Kathy Pearson  MASK TILE is a mapper module that
MCNC Environmental  produces a geometry from an input field
Programs  and color map. If any data point falls
Number :  1803  below the color scale, this sets its
Language :  C  color to the mask color. For
Submitted :  04/21/94  optimization of Z slices, adjacent
Version :  1.000  squares on the same row of identical
Last Updated :  04/21/94  color will be drawn as a single polygon
Ported To :  DEC HP IBM Kubota Sun  of that color. The MCNC Environmental

Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

MINMAX REPORT

Author :  Kathy Pearson  NAME MINMAX REPORT – generate
MCNC Environmental  statistics report with C-language. All
Programs  floating-point values are written with
Number :  1805  the C-language “%g” format so that
Language :  C  more decimals may be reported than
Submitted :  04/21/94  those using the standard AVS statistics
Version :  1.000  module. The MCNC Environmental Programs
Last Updated :  04/21/94  Group has created this module as part
Ported To :  DEC HP IBM Kubota Sun  of a suite of AVS demo modules,

networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Marker Draw

Author :  Ian Curington  marker draw - generate marker bitmaps
Advanced Visual Systems  from marker input list. This module to
Number :  1836  takes a marker coordinate specification
Language :  C  vector XY-pair list on an input port
Submitted :  07/16/94  from another module, combines this with
Version :  1.000  an an input image, and create a
Last Updated :  07/16/94  composite display image. The input
Ported To :  HP IBM Kubota Sun  image is copied to the output, then

marker polygonal shapes are scan converted and filled using an edge fill algoriythm, with the fill color computed based on hue parameters. It is designed and tested with several other marker modules, such as marker pick or marker select as source The vector marker list is defined by structures in the “marker.h” header file, defining the structure and usage of the field data type for markers. Each marker has a compile-time defined shape, but runtime attributes of position, size, color, etc.

Marker Gen

Author :  Ian Curington  marker gen - generate marker bitmaps
Advanced Visual Systems  from marker input list. This module
Number :  1837  converts vector edge defined marker
Language :  C  shapes into ROI bit masks coverage for
Submitted :  07/16/94  AVS images. The input is compatable
Version :  1.000  with any AVS image, and any fully
Last Updated :  07/16/94  specified marker list 1D integer field.
Ported To :  DEC HP IBM Kubota Sun  The marker list can be specified from a

stored field, or more often, marker select, or marker pick modules. This module scan converts the marker edge list into a binary bit pattern, where the bit pattern at each pixel is either zero for no marker present, or the marker shape index. The output is a full size scalar byte field, the same size as the input image.

Marker Merge

Author :  Ian Curington  marker merge - merge roi marker
Advanced Visual Systems  pattern with full colour image. This
Number :  1838  module modifies a full color input
Language :  C  image by masking it with an ROI byte
Submitted :  07/16/94  field. The marker type code in the ROI
Version :  1.000  is used to control what color is
Last Updated :  07/16/94  replaced inside each marker area, and
Ported To :  DEC HP IBM Kubota Sun  optionally a secondary texture image

can be used as the fill color inside of the marker areas. The ROI field is compatable with other ROI field modules (IP Lib). This module is designed to work with other marker modules, marker gen, marker select, marker draw, marker pick, etc.

Marker Pick

Author :  Ian Curington  Marker Pick - Pick, Edit, Create, and
Advanced Visual Systems  Move Markers on Images. Allow marker
Number :  1839  pick identification and display from an
Language :  C  input marker list and the image viewer
Submitted :  07/16/94  system. Takes as input an AVS field
Version :  1.000  with a list of coordinates for markers,
Last Updated :  07/16/94  displays them over an image (map) in
Ported To :  HP IBM Sun  the image viewer, and reacts to mouse

picks on these markers, sending the pick selection to the output port. Markers are defined by geometric polyline vector lists, not ROI pixel masks. You can externally control position, shape, size, etc. Any marker field sent out on the output port can be read back in to the input field port. This way state save and restore are possible. New marker lists can be made using the “marker pick” module.

Marker Select

Author :  Ian Curington  Marker Select - Define and allow Pick
Advanced Visual Systems  Selection of Markers Allow marker pick
Number :  1840  identification and display from an
Language :  C  input marker list and the image viewer
Submitted :  07/16/94  system. Takes as input an AVS field
Version :  1.000  with a list of coordinates for markers,
Last Updated :  07/16/94  displays them over an image (map) in
Ported To :  HP IBM Sun  the image viewer, and reacts to mouse

picks on these markers, sending the pick selection to the output port. Markers are defined by geometric polyline vector lists, not ROI pixel masks. New marker lists can be made using the “marker pick” module. This module does not allow edition of position or shape, just allows query picking of pre-defined markers.

NEW CROP

Author :  Kathy Pearson  NAME NEW CROP – crop 3d or 4d field to
MCNC Environmental  reduce the range of any of its
Programs  dimensions. Unlike the standard AVS
Number :  1807  crop module, the NEW CROP module allows
Language :  C  the user to crop a field down to a
Submitted :  04/21/94  single grid cell in any dimension. Also
Version :  1.000  unlike the standard crop module, the
Last Updated :  04/21/94  NEW CROP module values will “hold”
Ported To :  DEC HP IBM Kubota Sun  between invocations if the previous

settings are within the latest range for X, Y, Z, or Time. The standard crop module allows cropping on 3 dimensions only, but the NEW CROP module allows the user to crop either 3D or 4D fields.

NEW LEGEND

Author :  Kathy Pearson  NEW LEGEND – create a geometry of a
MCNC Environmental  linear scale using color polygons. The
Programs  NEW LEGEND module creates a color scale
Number :  1811  bar of linear bins in either the
Language :  C  horizontal or vertical with labeling
Submitted :  04/21/94  controlled by C-language formats to
Version :  1.000  give the user more control than that
Last Updated :  04/21/94  found in the standard AVS “color
Ported To :  DEC HP IBM Kubota Sun  legend“ module. Color bins are selected

from the input color map based upon linear increments corresponding to the number of color bins selected. For a typical Newton AVS colormap, the lowest bin will be blue and the highest bin will be red. The bins are labeled according to the Data Minimum and Data Maximum values and the linear increments between. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

NEW PRINTFIELD

Author :  Kathy Pearson  PRINTFIELD – display data in a matrix
MCNC Environmental  format. The NEW PRINTFIELD module
Programs  outputs to a text browser (and
Number :  1808  optionally to an output file) the
Language :  C  values of the input cropped slice of
Submitted :  04/21/94  data. The module also requires that an
Version :  1.000  input string specifying that the input
Last Updated :  04/21/94  data has been sliced along the “X”,
Ported To :  DEC HP IBM Sun  “Y”, or “Z” axis. Each output value is

written in the C-language format set by the user with optional labeling. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

NEW SLICE

Author :  Kathy Pearson  NEW SLICE – slice a field of up to 4
MCNC Environmental  dimensions on X, Y, Z or time. Unlike
Programs  the standard AVS slice module, the NEW
Number :  1809  SLICE module will accept 4D data input
Language :  C  in addition to a 3D data field. It
Submitted :  04/21/94  outputs a field of one dimension less
Version :  1.000  than that input with the same
Last Updated :  04/21/94  coordinate space and the same number of
Ported To :  DEC HP IBM Kubota Sun  vectors as the input. The NEW SLICE

module also outputs the slice axis to indicate to downstream modules whether the slice has occurred in the “X”, “Y”, “Z”, to “Time” dimension. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

NEW TILE

Author :  Kathy Pearson  NEW TILE – create a geometry of a tile
MCNC Environmental  plot displaying 2D 3-space data. The
Programs  NEW TILE module accepts as input a
Number :  1810  field sliced on “X”, “Y”, or “Z” and
Language :  C  displays a tile map in the other two
Submitted :  04/21/94  dimensions. The NEW TILE module was
Version :  1.000  needed because versions available from
Last Updated :  04/21/94  the International AVS Center do not
Ported To :  DEC HP IBM Kubota Sun  work properly for cropped data and do

not function well with any slice axis other than “Z”. Still, much credit is due to David Bock, formerly of MCNC, who is the author of the original, for his pioneering efforts in creating an early AVS version of this module. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Ortho Texture

Author :  Scott Goodyear and Ian  This is a special version of field to
Curington  mesh, for uniform fields, that outputs
Advanced Visual Systems,  just one quadrilateral, with texture
Inc.  from the field, rather than a pair of
Number :  1675  vertex colored triangles for each
Language :  C  sample point. This allows exploitation
Submitted :  09/29/93  of hardware texture mapping. This is
Version :  1.000  prototype unsupported code. The source
Last Updated :  09/29/93  is NOT intended as a working example.
Ported To :  HP IBM Sun

Polygon Volume

Author :  Jeff Wang  This module allows users to draw
MCNC, Environmetal  columns of the irregular shaded
Programs Group  polygon. For example, users can make a
Number :  1731  2 dimensional geographical county map
Language :  C  become 3 dimensional volume using the
Submitted :  12/15/93  population of the county as height.
Version :  1.000  Another application includes building a
Last Updated :  12/15/93  house, showing downtown area of a city
Ported To :  Sun  and plotting economic data. The value

of the polygo is mapped into color by the colormap of AVS. The input field is a 1D 4 space field with (x,y,z,n,value), where (x, y, z) is the coordinate of a vertex in the polygon, n is the number of vertices in the polygon, value is the value assign to the polygon. Portions of the code based on Wes Bethel’s hsv to rgb module, also on the IAC’s anonymous ftp site.

Polyline

Author :  Ian Curington  The polyline module interacts with the
Advanced Visual Systems  AVS5 image viewer, to allow interactive
Number :  1678  line drawing on top of an image. Points
Language :  C  inside the image can be picked, and
Submitted :  09/29/93  line segments are drawn with “rubber
Version :  1.000  banding“ along the pick sequence.
Last Updated :  09/29/93  “Undo” will back up the line segments,
Ported To :  HP IBM Sun  and clear will erase all lines. If

multiple image viewer “views” are open, the rubber banding can be seen in each window at the same time. The resulting picked polyline list is output on a scatter field port, for external processing by other modules.

QUERY TILE

Author :  Kathy Pearson  QUERY TILE – create a geometry of a
MCNC Environmental  cross hair cursor over a tile plot and
Programs  report the value of the point at that
Number :  1812  tile position.
Language :  C
Submitted :  04/21/94
Version :  1.000
Last Updated :  04/21/94
Ported To :  DEC HP IBM Kubota Sun

RESET 2

Author :  Kathy Pearson  RESET – resets geometry viewer to
MCNC Environmental  normalize and scale window to the
Programs  current coordinate system, then
Number :  1798  rescales to reduce its size. The MCNC
Language :  C  Environmental Programs Group has
Submitted :  04/21/94  created this module as part of a suite
Version :  1.000  of AVS demo modules, networks, and
Last Updated :  04/21/94  sample data sets to demonstrate the
Ported To :  DEC HP IBM Kubota Sun  functionality of its visualization

prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Reset Geom

Author :  Jeff Wang  This Module allows users to reset,
MCNC Environmental  rescale and normalize the geometry
Programs Group  viewer. This is one of a suite of
Number :  1740  modules we used to plot the emission
Language :  C  data for the state of North Carolina.
Submitted :  12/15/93  This group of modules allows displaying
Version :  1.000  points, lines and polygons over a map
Last Updated :  12/15/93  of the United States. The module suite
Ported To :  Kubota Sun  is stored as a group in a compressed

tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

Scatter Rect

Author :  Ian Curington  scatter rect - Draw Nested Rectangles
Advanced Visual Systems  from irregular 1D (scatter) fields.
Number :  1846  This module draws 4 variable height
Language :  C  nested coloured rectangles from scatter
Submitted :  07/16/94  input data showing 4-vector data at the
Version :  1.000  specified positions. Each sample point
Last Updated :  07/16/94  gets 4 stacked rectangles. The first
Ported To :  DEC HP Kubota Sun  three are various colors, and fourth is

a wire frame shell around the others. It can be used in place of “scatter dots“. The Scale dial is simply a size multiplier. The height of each nested rectangle is controlled with one of the vector quantities.

TEXT TO GEOM

Author :  Kathy Pearson  NAME TEXT TO GEOM – create a label
MCNC Environmental  geometry from the input string. The
Programs  MCNC Environmental Programs Group has
Number :  1815  created this module as part of a suite
Language :  C  of AVS demo modules, networks, and
Submitted :  04/21/94  sample data sets to demonstrate the
Version :  1.000  functionality of its visualization
Last Updated :  04/21/94  prototype for an Environmental Decision
Ported To :  DEC HP IBM Kubota Sun  Support System (EDSS). You can download

these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

TIME SPAN TEXT

Author :  Kathy Pearson  TIME SPAN TEXT – create a string
MCNC Environmental  containing the starting and ending
Programs  timestamp labels; if they are the same,
Number :  1816  output only the starting timestamp
Language :  C  label The MCNC Environmental Programs
Submitted :  04/21/94  Group has created this module as part
Version :  1.000  of a suite of AVS demo modules,
Last Updated :  04/21/94  networks, and sample data sets to
Ported To :  DEC HP IBM Sun  demonstrate the functionality of its

visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Title

Author :  Roy Wirthlin  This module creates a “label” style
Advanced Visual Systems  text string in GEOM format which can be
Inc  used as a title. There are two
Number :  1287  advantages to using this module over
Language :  C  the labelling facilities in the Geom
Submitted :  06/23/92  Viewer - (1) the information will get
Version :  1.000  saved with a network and (2) the
Last Updated :  06/23/92  optional floating point parameter
Ported To :  DEC IBM Sun Kubota HP  (“Value”) can come from another module

and can represent some important variable such as time, animation step, some parameter, etc. The Title String can contain a ’%f’ ala C programs to include this parameter - for example, the Title String can be “Time Step %f” and the value of “Value” will get transferred to the geom label. Other things you can control are the font number, drop shadows, justification, position, height, and color of the text string.

UCD Node Contour

Author :  Woody Muller  UCD Node Contour creates an array of
Becton Dickinson  floating point RGB values, one color
Research Center  for each node in an unstructured cell
Number :  1830  data set. This module is a replacement
Language :  C  for ucd-contour as supplied by AVS Inc.
Submitted :  07/16/94  AVS Inc’s ucd-contour module turns
Version :  1.000  zones of out-of-range data black,
Last Updated :  07/16/94  rather than clamping it to min and max
Ported To :  DEC HP Sun  values as the user often prefers. This

module is intended as a functional equivalent of ucd-contour, but which clamps out of range values to min and max colors. This allows more detailed coloring of data within the absolute range of the entire UCD without turning any of it black. UCD Node Contour is used to create a color contour of a UCD structure. Its output is passed to ucd-to-geom to produce a colored representation of a UCD structure. UCD Node Contour associates a color with each node of a UCD structure.

UCD grid

Author :  Penny Rheingans  Generates a labelled grid and/or axes
US EPA Scientific  for ucd data. The extents of the grid
Visualization Center  or axes are the extents of the data (or
Number :  1308  slightly larger).
Language :  C
Submitted :  08/30/92
Version :  1.000
Last Updated :  08/30/92
Ported To :  DEC Sun HP

VIEW SHARE

Author :  Terry Myerson  This module is part of the
International AVS Center  collaboratory development being done at
(NCSC)  the North Carolina Supercomputing
Number :  1206  Center. This module allows AVS users,
Language :  C  on up to 9 heterogeneous workstations,
Submitted :  05/12/92  to perform shared visualization. In
Version :  1.000  other words, the data can be located on
Last Updated :  05/12/92  one workstation’s file system, the
Ported To :  Sun DEC IBM  computation could be done on that

user’s workstation, and then the resulting geometry can be broadcast to all of the other users and then rendered on each user’s individual workstation. Each user can then take control of the visualization, rotating and translating not only his view, but that of the other collaborators. In addition, integers, floats, and strings can be broadcast around the collaborative loop to allow for interactive distributed data slicing, contouring, and file switching.

Vector Arrow

Author :  KGT (Kubota Graphics  The “Vector Arrow” module inputs a
Technologies) Inc.  3-vector field (1D,2D,3D) and creates
Number :  1883  arrows. This module looks like the
Language :  C  “hedgehogs” module.
Submitted :  09/07/95
Version :  1.000
Last Updated :  09/07/95
Ported To :  DEC ultrix DEC alpha

WIND PLOT

Author :  Kathy Pearson  The WIND PLOT module accepts as input a
MCNC Environmental  field 3-vector wind field and outputs a
Programs  geometry representation of it. The MCNC
Number :  1818  Environmental Programs Group has
Language :  C  created this module as part of a suite
Submitted :  04/21/94  of AVS demo modules, networks, and
Version :  1.000  sample data sets to demonstrate the
Last Updated :  04/21/94  functionality of its visualization
Ported To :  DEC HP IBM Kubota Sun  prototype for an Environmental Decision

Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Wide Arrow

Author :  Wright, E.L.  This module reads a 1D 3-coordinate
U.S Geological Survey  irregular field containing a single
Number :  1567  floating-point 2-element vector
Language :  C  specifying X and Y velocity components
Submitted :  01/28/93  at the specified (X,Y,Z) coordinate. It
Version :  1.000  then creates the geometry to display a
Last Updated :  01/28/93  wide arrow indicating vector direction

and magnitude at a specified (X,Y,Z) coordinate. The length of the wide arrow is proportional to the vector magnitude. The Scale Factor parameter allows the user to scale the data to produce an arrow of the desired length. This module is useful for displaying wind direction and velocity.

add polar 2D

Author :  Jeff Knerr and Shawn  This module takes in a 2D field and
Mehan  adds coordinate data in (r,theta)
North Carolina  spherical polar coordinates mapped to
Supercomputing Center  (x,y) cartesian as required by the
Number :  1305  geom viewer. The added coordinates are
Language :  C  either uniformly spaced or calculated
Submitted :  08/26/92  from input coordinates, depending on
Version :  1.000  whether the input field is uniform or
Last Updated :  08/26/92  rectilinear. Then, depending on what
Ported To :  DEC Sun HP  the original computational space grid

was, the data can be mapped so that it covers the entire 2*pi grid.

area

Author :  Anonymous  This module takes an input 2d scalar
Stardent Computer  field, and assumes it has a “doughnut”
Number :  1013  or “0” shape in it. The simple
Language :  C  evaluation algorithm searches for the
Submitted :  10/10/91  interior space of the “0”, and computes
Version :  1.000  the cross sectional area of the region.
Last Updated :  10/10/91  The result (a scalar) is passed in a
Ported To :  Kubota IBM Sun HP  1-D output array. This module will

“accumulate”, and pass the history along with the current value. The parameter input controls the current slice, ZERO will reset things. The design was to work with “tubes” with variable cross section. Typical input would come from 3D scalar fields -¿ ortho slice, while typical output would go to the “graph viewer.” This module was written to show how to compute the transverse cross sectional area of a blood vessel in a medical imaging application.

bin field

Author :  Wes Bethel  A grid is constructed in three-space
Lawrence Berkeley  which represents the extents of some
Laboratory  input 3-space field. The grid has a
Number :  1178  user specified number of coordinates,
Language :  C  or bins, along each axis. For every
Submitted :  04/20/92  input point, the count in the
Version :  2.000  three-space bin in which it lies is
Last Updated :  05/19/92  incremented.
Ported To :  Kubota DEC IBM Sun

bivar

Author :  Wes Bethel  The bivar module will resample 1d
Lawrence Berkeley  scatter data in 2 variables. A typical
Laboratory  use for this module is to take a field
Number :  1017  of 2-d coordinate, scalar values, and
Language :  C  resample over a “regular” grid. The
Submitted :  10/10/91  algorithm used is one in which, for
Version :  1.000  each output grid location, all input
Last Updated :  10/10/91  points are scanned to determine if they
Ported To :  Kubota DEC IBM Sun HP  lie within the specified search radius.

If so, that input point’s contribution is weighted as a function of distance from the output grid point and scaled by 1./distance**weightfunction. As such, this module is best suited for use on input data which desn’t necessarily fit a differentiable function. For such data, refer to the sact2d module.

carpet

Author :  Ian Curington  A Carpet Plot surface is made showing
Advanced Visual Systems  surface height variation from one
Number :  1569  field, with color variations from data
Language :  C  in a second field. Each cell is drawn
Submitted :  01/28/93  as per field to mesh, (linear
Version :  3.000  interpolation), or in block mode with
Last Updated :  12/15/93  descrete non-interpolated colors. In
Ported To :  IBM Kubota HP  block mode, it will optionally leave

holes in the carpet where the data goes out of range for the colormap, allowing mask values to control cell visibility, or region boundaries. Version 3.000 adds upgrade to irregular 3space fields for a better hole algorithm.

city scape

Author :  Ian Curington  The city scape module converts a
Advanced Visual Systems  two-dimensional floating point field
Number :  1185  into group of 3D blocks, represented as
Language :  C  a GEOM-format mesh. Each element of the
Submitted :  05/04/92  field is mapped to a “building”. The
Version :  3.000  height of each block above each point
Last Updated :  07/15/94  is proportional to the scalar value of
Ported To :  IBM Sun Kubota HP  the field. This is my attempt at

“Business Graphics”, for a multi-dimensional telecommunications statistics data application, except in interactive 3D. Version 2 is a substantial upgrage from version 1 at the IAC. This module provides 3D block diagram, with averaged statistics on side shadow panels, and annotation. A transparent theshold sheet is superimposed on city. Blocks can be picked for interogation of cell number. parts may be turned on or off for clarity. Two scalar fields can be displayed, one contolling height, the other color, with two independent side stacked side bar statistics. Version 3 update draws the top face of picked block in white hi-lite, adds extra edit list for pick, no normals, new text input for edge labels, added string typeins for edge labels, and other fixes.

color legend

Author :  Wes Bethel  This module takes an input colormap and
Lawrence Berkeley  generates a “color legend,” consisting
Laboratory  of a number of geometry objects which
Number :  1019  may be subsequently rendered using the
Language :  C  geometry viewer module. The range of
Submitted :  10/10/91  colors in the color scale is controlled
Version :  2.000  by the input colormap. All colors in
Last Updated :  08/31/92  the colormap are represented in the
Ported To :  Kubota DEC IBM Sun HP  output color scale. In addition to the

smoothly shaded “color bar”, the user may control various parameters relating to labeling characterisitics, and legend orientation. Version 2 fixes (by Dr. John Tee, AVS Inc consultant) make the code more portable by using standard AVS routine calls.

color tile

Author :  David Bock  The color tile module accepts a 2D
North Carolina  scalar field of either uniform or
Supercomputing Center  rectilinear coordinates of any data
Number :  1151  type and an input colormap and
Language :  C  generates a geometric grid of colored
Submitted :  03/11/92  cells. The resulting grid is generated
Version :  2.000  by creating cells of 4 vertices
Last Updated :  05/19/92  SURROUNDING each data value rather than
Ported To :  DEC IBM Sun Kubota HP  cells with the data values representing

the ENDPOINTS of each cell as is done in other color mapping modules such as field to mesh. The data values within each cell are mapped to colors determined from the input colormap and the input number of steps. This color is used to color the entire cell of the respective data value. By selecting the min/max parameter, the color tile module will calculate the minimum and maximum data values from the input field and use these values to map the data to the color values. Without this selection, the module accepts type-in values for the low and high data values. By selecting the mesh parameter, the module generates and displays an output mesh overlaid upon the output color tile object.

cone bounding

Author :  Massimiliano Bavelloni  Draw geometric cone around 3D scattered
Radioastronomy  points in order to display astronomical
department C.N.R.  points (galaxies). Galaxy coordinates
Bologna Italy.  have been already converted from
Number :  1896  Spherical Coordinates into Cartesian
Language :  C  Coordinates. In the computational
Submitted :  09/07/95  section coordinates are converted into
Version :  1.000  Spherical ones and then converted again
Last Updated :  09/07/95  in Cartesian coordinates. The cone is
Ported To :  DEC ultrix DEC alpha  always centered in the origin (x = 0, y

= 0, z = 0) point.

confocal slicer

Author :  Nick Salmon  Slices through a three dimensional
European Molecular  field, with the slice plane at an
Biology Laboratory  arbitrary orientation. The confocal
Number :  1851  slicer module extracts a 2d slice from
Language :  C  a 3d field. Output consists of a 2d
Submitted :  10/17/94  field or a geometry. The position, size
Version :  1.000  and resolution of the slice plane are
Last Updated :  10/17/94  under user control via the parameters.
Ported To :  DEC alpha HP IBM Sun  For all input fields, the module will

determine “sane” parameter settings for extracting an initial slice. This is basically a hack from the module new arbitrary slicer from IAC (apologies to Wes Bethel). The intention is to use it for slicing data collected from a confocal microscope. This is uniform byte data with one or more values per point in space.

contour field

Author :  Phil McDonald  The contour field module generates
NOAA/ERL/Forecast  contour lines from a 2D field. These
Systems Laboratory  contour lines are output as a geometry
Number :  1193  object with one or more polylines in 3D
Language :  C  space and will nest neatly with a mesh
Submitted :  05/06/92  surface made from the same field. The
Version :  3.000  contours may be at uniform intervals or
Last Updated :  07/05/93  at arbitrary levels. If an optional
Ported To :  IBM  colormap is attached, the contour lines

will be colored according to how the values of the contour lines are mapped in the colormap, otherwise, the contour lines are not colored. The contour lines may be optionally labelled. The interval between labelled contour lines may be uniform or, if arbitrary contour levels have been specified, only specified levels will be labelled. A separate geometry object for these labels is output. This allows the properties of the labels to be entirely different from those of the contour lines. However, the labels can be colored by value only if the contour lines are colored by value. Version 2 of this module includes bug fixes courtesy of Lyle Wiedeman of UC Irvine. Version 3 includes several changes to the module and the libraries it uses as well.

contour to g

Author :  Anonymous  CONTOUR TO GEOM finds and creates
Stardent Computer  contour lines of similar value in a
Number :  1022  scalar byte field. The parameter
Language :  C  “threshold” controls the level of
Submitted :  10/10/91  contour being found. It differs from
Version :  1.000  teh CONTOUR module found in the FILTERS
Last Updated :  10/10/91  directory in that it produces disjoint
Ported To :  Kubota DEC IBM Sun HP  lines which can be rendered with the

geometry viewer. It can also handle 3D data sets and irregular grids.

contour to posts

Author :  Rob Kuyper  “contour to postscript” is the ultimate
University of Technology  contour module. All the things you
Delft, The Netherlands  expect of a contour module are built in
Number :  1586  this one. Automatic scaling, min/max
Language :  C  scaling, typein levels. Colormap port
Submitted :  02/11/93  to color the contours. Postscript
Version :  1.000  output for contours - straight lines on
Last Updated :  02/11/93  the plot . If the dataset is positioned
Ported To :  Kubota HP Sun  in e.g. the y-z plane, only the y-z

coordinates are used in the Postscript output. After compilation the world of interactive contour and Postscript generation opens up to you.

crop and slice

Author :  James Pipe  This module combines the functionality
University of Michigan  of series of the supported ’orthogonal
Medical School  slicer’ and ’crop’ modules and with
Number :  1610  some additional capabilities. For each
Language :  FORTRAN  dimension in the input data set (up to
Submitted :  04/10/93  4 dimensions), one may crop (preserve
Version :  1.000  that dimension but limit its range), or
Last Updated :  04/10/93  slice (eliminate that dimension by
Ported To :  DEC IBM  choosing data from one index into that

dimension) the data along that dimension. If the data is cropped along some dimension, one can specify the range either with ’begin’ and ’end’ indices, or by using ’center’ and ’width’. Changing the value in one pair (begin/end or center/width) causes the module to the values of the other pair. If the data is sliced along any dimension, only the ’dimension center’ widget remains visible; this is similar to the ’slice plane’ parameter in the supported ’orthogonal slicer’ module. For every dimension that is sliced, the dimensionality of the output data is reduced by 1. For instance, a 4D input data set that is sliced along dimensions 2 and 4 results in a two-dimensional output data set. The cropped/sliced data set is present at the LEFT output port. The RIGHT output port contains a ’mask’ data set which is identical to the input data set EXCEPT for the data which was cropped/sliced into the left output port data set; this part of the mask (right port) data is set to zero, so that one can visualize the location in the original data set from which the cropped/sliced data came. For example, if the input is a 2D image, and dimension 2 is sliced, then the left output port will be a 1D 4-vector byte, and the right output port will be the same as the original image except for a vertical black line running through it where dimension 2 was sliced (plus perhaps some cropping along dimension 1).

draw tracks

Author :  John Tee  DRAW TRACKS is a module provided to
AVS Inc. Consultant  demonstrate the use of the geometry
Number :  1650  viewer to display two-dimensional track
Language :  C  information provided as a floating
Submitted :  09/29/93  point 2-vector (x,y) input field. If
Version :  1.000  the input field has more than one
Last Updated :  09/29/93  dimension, multiple tracks are
Ported To :  HP IBM Kubota Sun  displayed. If the field is 3-vector, it

uses the third vector component to color the lines, with the optional colormap input field. The hi-low dials of the colormap control the line color range. DRAW TRACKS also allows the user to select a track of interest by clicking the mouse on it. The selected track is highlighted by changing its colour, and the selected vertex is indicated by positioning a sphere over the vertex, and a line connecting the vertex to the sphere. The coordinates of the selected vertex are displayed in a text window.

field lines

Author :  Rob Kuyper  field lines is a mapper module which
Univeristy of Technology  converts a 2D field in an object of
Delft, The Netherlands  lines and planes. This module will
Number :  1533  create black lines (in optionally i- or
Language :  C  j-direction) with a ’no-lighting’ plane
Submitted :  10/01/92  below it, so the field will look like a
Version :  1.000  serie of non-transparent graphs.
Last Updated :  10/01/92
Ported To :  HP Sun

field o matic

Author :  Phil McDonald  Field o matic.. It slices.. It dices..
NOAA/ERL/Forecast  Actually, it slices and pares. The
Systems Laboratory  field o matic module provides the same
Number :  1265  capabilities as the orthogonal slicer,
Language :  C  crop, and field to mesh modules. A 3D
Submitted :  05/21/92  field may be sliced orthogonally in
Version :  2.000  computational space and this slice may
Last Updated :  07/05/93  be cropped. The resulting slice “zone”
Ported To :  IBM  (window) is output as a 2D field and as

a mesh surface. If an optional colormap is connected, the mesh surface will be colored by this colormap, unless the field is a vector field, in which case the mesh will not be colored. The updated version includes several changes to the module and the libraries it uses as well.

field params

Author :  John Tee  FIELD PARAMS is a module provided for
AVS Inc. Consultant  the Oxford University Osney Labs custom
Number :  1651  software for visualisation of
Language :  C  experimental data. FIELD PARAMS returns
Submitted :  09/29/93  the X and Y dimensions of a field as
Version :  1.000  integer parameters, and the minimum and
Last Updated :  09/29/93  maximum values of the field as float
Ported To :  HP IBM Kubota Sun  parameters. It allows each of these

parameters to be scaled by its own scale factor. FIELD PARAMS is used to provide the GENERATE COLORMAP module with the range of the data values (cf COLOR RANGE), while also allowing the user to modify the colour range. It also provides the X dimension ofthe field as an upper limit to the ANIMATE INT module which controls the position of the viewable window.

field size

Author :  John Tee  FIELD SIZE returns the dimensions of a
AVS Inc. Consultant  field as integer parameters. It allows
Number :  1652  each of these parameters to be scaled
Language :  C  by its own scale factor. FIELD SIZE is
Submitted :  09/29/93  used to limit the range of parameters
Version :  1.000  such as cropping limits to a proportion
Last Updated :  09/29/93  of the size of a field.
Ported To :  HP IBM Kubota Sun

field stats

Author :  John Tee  FIELD STATS returns the minimum,
AVS Inc. Consultant  maximum and mean values of a field as
Number :  1653  floating point parameters. It allows
Language :  C  each of these parameters to be scaled
Submitted :  09/29/93  by its own scale factor. One
Version :  1.000  application of FIELD STATS is to create
Last Updated :  09/29/93  a colormap which is adapted to the
Ported To :  HP IBM Kubota Sun  range of values in the field (as COLOR

RANGE does), but allowing the user to adjust the range manually.

field to con

Author :  Wes Bethel  The field to contour module creates
Lawrence Berkeley  colored contour lines around regions in
Laboratory  which the threshold specified by the
Number :  1024  “level” parameter is crossed. The
Language :  C  module can process only 2d datasets,
Submitted :  10/10/91  but these may may exist in 3-space ( if
Version :  2.000  the field is irregular) or in 2-space
Last Updated :  02/15/94  (rectilinear or uniform). The algorithm
Ported To :  IBM Sun HP  used is a subset of teh more familiar

3d marching cubes algorithm. Version 2 update provided by Krzysztof (Chris) S. Nowinski of Warsaw University’s Institute of Applied Mathematics and Mechanics. Enhancements include Generating contours as polylines and not disjoint lines for faster rendering; Generating whole range of contours by a single module; Generating surfaces (these produced by field to mesh do not fit into dimensions, sometime); Optional plane and/or 3D maps; and Second field and colormap input to color surfaces.

field to map

Author :  Andy Bingham  The field to map module converts 2D
Oxford University, Dept.  scalar field into a surface in 3D
of Earth Sciences  space, represented as a GEOM-format
Number :  1705  mesh. The geometry is annotated with
Language :  C  boundaries, tick marks and a colour
Submitted :  09/30/93  scale, for geographic map
Version :  1.000  visualisation. The input data must be a
Last Updated :  09/30/93  2D field with a scalar data value at
Ported To :  HP IBM Kubota Sun  each element. The extents of the field

are used to annotate the map.

glue

Author :  James Pipe  Continually pastes on incoming data
University of Michigan  onto the end of a growing output data
Medical School  set - always checking to make sure that
Number :  1612  the dimension lengths of the input and
Language :  FORTRAN  output data sets are identical EXCEPT
Submitted :  04/10/93  for the dimension in which the glue
Version :  1.000  operation is proceeding (also checks
Last Updated :  04/10/93  vector lengths). The parameter ’total
Ported To :  DEC IBM Kubota  sets’ is used ONLY for reporting the

total number of data sets glued

together in the current output data set - changing its value manually will do nothing.

grid mesh

Author :  Michael Rangitsch  This module generates an outline or
Number :  1124  grid mesh at a constant topological
Language :  C  plane. The coordinate information of
Submitted :  01/26/92  the input field is used to generate the
Version :  1.000  geometry by connecting adjacent nodal
Last Updated :  01/26/92  positions within a specified plane.
Ported To :  Kubota HP

hog 2D

Author :  Advanced Visual Systems  This module draws vector arrows on the
and Evans and Sutherland lattice of a 2D field, at the node
Number :  1660  positions. It draws arrows in the 2D
Language :  C  plane, given X-Y vector components. The
Submitted :  09/29/93  downsize dial subsamples the mesh, so
Version :  1.000  not all nodes get arrows, for clarity.
Last Updated :  09/29/93  There is never a Z component to the
Ported To :  HP IBM Kubota Sun  vectors, except that the arrow heads

are 3D pyramids. Even if the input field is 3-vector or more, it only draws arrows assuming 2D vectors in first two vector components. The arrows are constructed from disjoint line segments.

interpolating sl

Author :  Wes Bethel  Interpolating slicer is similar in
Lawrence Berkeley  function and form to “orthogonal
Laboratory  slicer“ and ”new ortho slicer“.
Number :  1198  Orthogonal slices are extracted from a
Language :  C  source dataset. Refer to the help for
Submitted :  05/07/92  those modules on the base
Version :  1.000  functionality. The difference is that
Last Updated :  05/07/92  non-integral slicer numbers are
Ported To :  IBM Kubota DEC HP  permitted. In the case of a

non-integral slicer number, the output data is computed to be a linear average of adjacent slices from the input data. For example, if a value of 25.4 is used as the slice number, then the resulting data would be 60% of slice 25 and 40% of slice 26.

irreg2map

Author :  John Tee  IRREG2MAP is intended for applications
AVS Inc. Consultant  in which data from scattered sensors is
Number :  1664  stored in irregular fields, and needs
Language :  C  to be displayed, or processed , as
Submitted :  09/29/93  uniform data. In particular, a time
Version :  1.000  series of measurements may be stored in
Last Updated :  09/29/93  a 2D irregular field, with the one
Ported To :  HP IBM Kubota Sun  compute dimension representing time,

and the other dimension the data values recorded at that time. For the data to be displayed in the Image Viewer, or as a texture map in the Geometry Viewer, this module is needed to create the uniform field required. IRREG2MAP produces a uniform byte field in which the pixels are set to zero, except where data from a selected row or column of the input field exists. The resolution of the output field is determined by the X start, Y start, X stop, Y stop, X step and Y step parameters. The module ignores any Z data in a 3-space field. The X and Y limits determine the range of coordinates, present in the input field, that are mapped into the output field. Values which fall outside the range of coordinates selected for the output field are ignored.

labeler

Author :  Phil McDonald  The labeler module may be used to
NOAA/ERL/Forecast  create label objects from field data or
Systems Laboratory  from arbitrary text. Fields from which
Number :  1621  data labels are derived may contain
Language :  C  data of any vector length and of any
Submitted :  05/01/93  dimensionality. Scalar values may be
Version :  1.000  translated into arbitrary character
Last Updated :  05/01/93  strings by using a value translation
Ported To :  Kubota  file. Title labels may be created from

a single string entered through the module control panel or may be created from the text and attributes contained in an ASCII title file. Title labels may be fixed in the display window or attached to points in the object space. Title labels of this latter type are referred to as “object” labels. Object labels may be specified only through a title file. Object labels and data labels may be offset from their reference points. “Pins” may be used to connect these labels to their reference points. Labels may be colored individually, colored uniformly, or uncolored. Each instance of a labeler module will create either data labels or title labels (including object labels), but not both.

line 1D

Author :  Kathleen Dyer  This module reads in a “field 1D
Lawrence Livermore  2-space or 3-space irregular“ and
National Laboratory  produces a “line”. The line is a GEOM
Number :  1121  object composed of disjoint lines, the
Language :  C  vertices of which are determined from
Submitted :  01/22/92  the grid points of the field. This
Version :  1.000  module was written so that the location
Last Updated :  01/22/92  of an orthogonal slice of a 2D field
Ported To :  Kubota DEC IBM Sun HP  could be shown via the “geometry

viewer“ module.

line probe

Author :  Phil McDonald  The line probe module can be used to
NOAA/ERL/Forecast  sample a 3D field along one or more
Systems Laboratory  lines or paths. The location of the
Number :  1620  probe may be entered manually,
Language :  C  indicated by pointing and clicking the
Submitted :  05/01/93  mouse, selected from a file of sampling
Version :  1.000  locations, or determined by the
Last Updated :  05/01/93  sampling locations in a second field.
Ported To :  Kubota  If either of the sampling options is

used, one sampling location may be selected, creating one probe, or all sampling locations may be selected, creating multiple probes, one for each sampling location. Except for the sample field mode, the probe will be a vertical line running through the height of the data volume at a constant X and Y position. By using the sample field mode, the paths of the probes may be non-linear, like trajectories. If the line probe module has created a single probe, the values found at its intersections with each of the fields layers will be listed in a text browser and output as a 1D field. If multiple probes are created, the output field is 2D (one “row” for each probe) but the values are not shown in the text browser. In either case, a polyline geometry object is created for the line probe(s) and spheres may be used to show the locations of the nodes (sampling points along the probe). If an optional colormap is provided and the field consists of scalars, the line probe and its nodes will be colored based on how the values along the probe map into this colormap. If a colormap is not provided or the contains vectors, the line probe and its nodes will not be colored.

local coord

Author :  John Stephen  This module gives you dial-based
Tessella Support  control over a probe/cropper, etc. It
Services  sets up a coordinate system relative to
Number :  1668  the transformation matrix of a named
Language :  C  object (usually “top”). Dials are used
Submitted :  09/29/93  to translate the derived origin from
Version :  1.000  the parental coordinate system and also
Last Updated :  09/29/93  to rotate its axes. Other dials then
Ported To :  HP IBM Kubota Sun  allow the object to be translated and

rescaled within the new coordinate system.

loop objects

Author :  Phil McDonald  Loop objects is a coroutine module that
NOAA/ERL/Forecast  provides a mild form of animation by
Systems Laboratory  providing the capability of looping a
Number :  1278  series of geometry objects. The result
Language :  C  is that all of the controls (object,
Submitted :  05/29/92  camera, and light transforms, etc.)
Version :  1.000  that are available through the render
Last Updated :  05/29/92  geometry module remain available
Ported To :  Kubota  before, during, and after looping. The

objects to be looped may be single objects or several objects grouped under a parent object. As each object (or family) is received, loop objects makes a note of it and creates a parent for it named “loop objx”, where x is the index of the object in the series. The entire series is given the parent “loop objects”. Animation is accomplished by altering the visibilities of the objects so that only one is visible at a time.

mesh axis

Author :  Tim O. Robinson  This module generates 3D coordinate
UC Berkeley  axes for a surface mesh. It will also
Number :  1276  optionally generate tick marks along
Language :  C  each axis, numeric labels for the tick
Submitted :  05/27/92  marks, and movable mesh planes
Version :  1.000  perpendicular to each of the axes.
Last Updated :  05/27/92
Ported To :  DEC IBM Sun Kubota HP

mesh slicer

Author :  Anonymous  Displays a lattice phase of field data.

Kubota Pacific Computer

Inc

Number :  1156
Language :  C
Submitted :  03/12/92
Version :  1.000
Last Updated :  03/12/92
Ported To :  Kubota DEC IBM Sun HP

mesh to ucd

Author :  Ian Curington  A mesh to ucd makes a 2D UCD surface
Advanced Visual Systems  showing surface height variation from
Number :  1751  one field. Each cell is drawn as per
Language :  C  field to mesh, (linear interpolation),
Submitted :  12/15/93  or in block mode with descrete
Version :  1.000  non-interpolated colors. It will
Last Updated :  12/15/93  optionally leave holes in the UCD
Ported To :  Kubota Sun  output where the data goes out of range

for the float dials, allowing mask values to control cell visibility, or region boundaries. It is basically a variation on field to mesh, allowing more advanced UCD type visualizations and data processing.

mesh with base

Author :  Phil McDonald  The mesh with base module generates a
NOAA/ERL/Forecast  mesh surface from a field, much like
Systems Laboratory  field to mesh does. However, an
Number :  1266  optional base, or “skirt”, can be added
Language :  C  to the surface to make it appear as a
Submitted :  05/21/92  solid, rather than as a simple wavy
Version :  1.000  papery surface. For 1D fields, the
Last Updated :  05/21/92  surface degenerates to a polyline and
Ported To :  Kubota  the base degenerates to a “curtain”. If

an optional colormap is attached, the surface will be colored according to how the field data values map to the colormap, otherwise, the surface will not be colored. Also, the base color may be selected from the colormap.

new abuffer

Author :  Wes Bethel  The new abuffer module is a ray traced
Lawrence Berkeley  volume renderer. Volumes of three
Laboratory  dimensional color information are
Number :  1113  composited on a pixel by pixel basis,
Language :  CandFORTRAN  resulting in an image of a translucent
Submitted :  12/12/91  solid, the signature of volume
Version :  2.000  rendering. The module will compute a
Last Updated :  05/19/92  set of faces which bound the
Ported To :  IBM Sun Kubota  rectangular shaped volume hull, and

perform ray intersections with this volume at each pixel location. For those pixels in which an intersection occurs, the “volume function” is integrated throught the volume, the resulting color information is stored at that pixel location. This module passes the vertices defining the volume hull (boundary) through a viewing transformation, taking the vertices from their native data space into one defined by a canonical view volume. Further, the center of the volume is placed into the center of the resulting image. The user can control the orientation of the volume, as well as specify isometric volume scaling.

new arbitrar

Author :  Wes Bethel  The new arbitrary slicer module
Lawrence Berkeley  extracts a 2d slice from a 3d field.
Laboratory  Output consists of a 2d field (as
Number :  1003  opposed to geometry, as in the “normal”
Language :  C  arbitrary slice module) which may be
Submitted :  10/10/91  piped to other modules which process 2d
Version :  2.000  fields. The position, size and
Last Updated :  02/09/94  resolution of the slice plane are under
Ported To :  Kubota IBM Sun HP  user control via the parameters. For

all input fields, the module will determine “sane” parameter settings for extracting initial slice. The latest update provides PRELIMINARY support for irregular fields. This HAS ITS LIMITATIONS - it uses a brute-force algorithm WHICH IS VERY SLOW; as space subdivision on warped grids is still a research topic.

new bubbleviz

Author :  Wes Bethel  The bubbleviz module generates spheres
Lawrence Berkeley  of various radii and colors at the
Laboratory  element locations of a 1D, 2D or 3D
Number :  1318  field. This is a “cuberille” style of
Language :  C  volume visualization, except that it
Submitted :  08/30/92  uses spheres rather than cubes. This
Version :  2.000  module can be used for non-uniform
Last Updated :  12/15/93  input fields. Version 2 upgrade
Ported To :  Sun HP  provided by Martin Serrano

(mas@epsilon.com) allows this module to work with 1-coord fields and 2-coord fields.

new ortho sl

Author :  Wes Bethel  The orthogonal slicer module takes an
Lawrence Berkeley  n-1 dimensional slice from an n-D
Laboratory  array. It does so by holding the array
Number :  1057  index in one dimension constant, and
Language :  C  does so by letting the other indices
Submitted :  10/10/91  vary. For instances, a data set might
Version :  3.000  include a volume of 5000 points,
Last Updated :  05/19/92  arranged as follows (using FORTRAN
Ported To :  DEC IBM Sun Kubota HP  notation) DATA(I,J,K) I=1,10 J=1,20

K=1,25 You can take a 2D “I-slice” from this data set by setting I=4 and letting the otehr indices vary DATA(4,J,K) J=1,20 K=1,25.This module is capable of slicing through 4D fields.

non ortho

Author :  Tim O. Robinson  This module applies a non-orthogonal
UC Berkeley  transform to a uniform input field. The
Number :  1605  typein real values determine the axis
Language :  C  length and directions. For a typical 3D
Submitted :  04/10/93  field, the a parameter corresponds to
Version :  1.000  the x-axis, the b parameter the y-axis
Last Updated :  04/10/93  and the c parameter the z-axis. The
Ported To :  DEC IBM Kubota  angle alpha is the angle between

vectors b and c, the angle beta is the angle between c and a, and the angle gamma is the angle between a and b. The default values correspond to no transformation whatsoever. This module has been tested briefly with 2D fields. The a, b, and gamma parameters are the only parameters that will have an effect on such a field.

ov slicer

Author :  Anonymous  The orthogonal vector slicer module
Stardent Computer  takes a 2D slice from a 3D array of
Number :  1059  data values and a 3D array of vectors,
Language :  C  overlaying the vector slice on the data
Submitted :  10/10/91  slice as a hedgehog display. It does so
Version :  1.000  by holding teh array index in one
Last Updated :  10/10/91  dimension constant, and letting the
Ported To :  Kubota DEC IBM Sun HP  other indices vary. See new ortho.

plumb line

Author :  Ian Curington  The plumb line module annotates the
Advanced Visual Systems  top surface of an irregular 3D field,
Number :  1677  with red cross hairs over each node.
Language :  C  The cross hairs are pickable, and drive
Submitted :  09/29/93  a line along the remaining axis,
Version :  1.000  through the 3D space. The (I,J) index
Last Updated :  09/29/93  for the picked point is shown and can
Ported To :  HP IBM Kubota Sun  be controlled with the X-Y integer

dials, and are output on two integer output ports. This can be used to control modules such as “orthogonal slicer“ to create a cutting sheet or a graph plot of the selected data section. The picked cross hair changes color to white, and a polyline following the mesh is drawn below.

probe path

Author :  Anonymous  The probe path module displays the
Stardent Computer  numeric data values in a field at a
Number :  1071  location in space. It works for fields
Language :  C  that have been rendered as a n AVS
Submitted :  10/10/91  geometry. It works for uniform,
Version :  1.000  rectilinear, and irregular coordinates,
Last Updated :  10/10/91  and for any data type. It works for
Ported To :  Kubota DEC IBM Sun HP  both scalar and vector fields.

probe path can accumulate the 3D path where the cursor has been, showing a colored line in 3D space of the probe history. Three other colored lines are shown, where the path is “projected” in 2D onto the axis planes. The sampled data is also stored along the path, and passed out in the 1D field structure. This is typically sent to other filter modules or the graph viewer.

real math

Author :  Lars M. Bishop  This module works with the same
National Center for  interface as the Field Math module,
Comp. Elect., Univ. of  with two input ports (one required, one
Illinois  optional), and one output. If the
Number :  1574  operation selected from the control
Language :  C  panel is unary, then the required input
Submitted :  02/11/93  port is used as the operand. If the
Version :  1.000  operation is binary, then both ports
Last Updated :  02/11/93  are used if they are both connected. If
Ported To :  DEC IBM  only the required port is connected,

then the required port, and the scalar parameter are used as the two operands. It is the user’s responsibility to assure that the values sent to this module are in the domain of the selected function. This module is part of the the Real Math module group, which is a single file, process-sharing set of modules that allow basic operations on and between integer and real AVS data types. They have been most useful in dealing with linkage of several AVS parameters. AVS has no simple way of performing math operations on reals, even though it does allow math operations on fields.

MAPPER MODULES

rgbtogray

Author :  Wes Bethel  The rgbtogray module will receive an
Lawrence Berkeley  AVS image as input, then compute an
Laboratory  output field which is a gray scale
Number :  1536  version of the rgb components of the
Language :  C  input image.
Submitted :  10/03/92
Version :  1.000
Last Updated :  10/03/92
Ported To :  DEC HP

rubberbonds

Author :  Nihar Gokhale  Rubberbonds forms a stretchable bond
Argonne National  between two atoms, given a molecule,
Laboratories  and the atoms within the molecule. The
Number :  1730  cylinders start with a default radius
Language :  C  for the equilibrium condition. As the
Submitted :  12/14/93  bond gets stronger (shorter distance),
Version :  1.000  the radius is enlarged proportionately.
Last Updated :  12/14/93  As bonds get weaker (longer distance),
Ported To :  Kubota Sun  the cylinders become long and thin,

and, at some point “pop” (i.e. are no longer visible).

scat 2d

Author :  Wes Bethel  The scat2d module will resample 1d
Lawrence Berkeley  scatter data in 2 variables. A typical
Laboratory  use for this module is to take a field
Number :  1077  of 2 corrdinate, scalar values, and
Language :  C  resample over a “regular” grid.
Submitted :  10/10/91
Version :  2.000
Last Updated :  10/10/91
Ported To :  DEC IBM Sun

scat 3d

Author :  Wes Bethel  The sact3d module will resample
Lawrence Berkeley  sacttered data in three variables. A
Laboratory  typical use for this module is to take
Number :  1078  a field of 3-coordinate, scalar values,
Language :  C  and resample over a “regular” grid.
Submitted :  10/10/91
Version :  1.000
Last Updated :  10/10/91
Ported To :  Sun

INTERNATIONAL AVS CENTER MODULE CATALOG

scat base la

Author :  Ian Curington  If your vector data is 1D n-vector
Advanced Visual Systems  float irregular, this module will draw
Number :  1754  a text label just above the first
Language :  C  point, showing the data label in the
Submitted :  12/15/93  AVSfield header. This was used in one
Version :  1.000  application to show the names of wells,
Last Updated :  12/15/93  floating above the well heads.
Ported To :  Kubota Sun

scat bub

Author :  Michael Rangitsch  Visualize a scatter data set (field 1D
Number :  1284  real 3-space) by drawing spheres at the
Language :  C  coordinate locations of the field and
Submitted :  06/20/92  optionally coloring the spheres based
Version :  1.000  on the field values and optionally
Last Updated :  06/20/92  dropping vertical lines in each
Ported To :  DEC IBM Sun Kubota HP  coordinate plane.

scatter arrows

Author :  Ian Curington  This module draws vectors with arrow
Advanced Visual Systems  heads from scatter input showing vector
Number :  1326  data at the specified positions. The
Language :  C  module will draw both 2D and 3D
Submitted :  08/31/92  vectors, as required. The Scale dial is
Version :  2.000  simply a length multiplier. Version 2.0
Last Updated :  02/21/93  mods submitted by Steve Larkin of the
Ported To :  Sun HP  University of Manchester.

scatter bounds

Author :  Ian Curington  This module draws a white bounding box
Advanced Visual Systems  from scatter input (1D irregular
Number :  1334  fields, or random collections of
Language :  C  points) showing data min/max positions.
Submitted :  08/31/92  It can be used with “scatter dots”.
Version :  1.000
Last Updated :  08/31/92
Ported To :  DEC Sun HP

scatter cubes

Author :  Ian Curington  This module draws variable size
Advanced Visual Systems  coloured cubes from scatter input
Number :  1333  showing scalar data at the specified
Language :  C  positions. It can be used in place of
Submitted :  08/31/92  “scatter dots”. The Scale dial is
Version :  1.000  simply a size multiplier. The size of
Last Updated :  08/31/92  each cube is taken by the first vector
Ported To :  DEC Sun HP  if “Variable”, otherwise it is set to

“Scale” size in “Uniform” mode. The colour of each cube is independently controled from the remaining three vector quantities. If the input vector length is 1 (scalar), the colour will default to white.

scatter rib 2

Author :  Ian Curington  If your vector data is 1D n-vector
Advanced Visual Systems  float irregular, this module will draw
Number :  1756  a ribbon going along the side of the
Language :  C  polyline path, from joining up the
Submitted :  12/15/93  specified positions in the scatter
Version :  1.000  field coord array. This module is used
Last Updated :  12/15/93  for Well Log data display, showing
Ported To :  Kubota Sun  various properties down a deviated well

shaft. The width, and color of the ribbon can be controlled to create a ribbon data representation. Tick marks and labels will annotate the graph. Interpolation can be turned on and off.

scatter ribbon

Author :  Ian Curington  This module may be used for 3D Well Log
Advanced Visual Systems  data display, showing various
Number :  1682  properties down a deviated well shaft.
Language :  C  If your vector data is 1D n-vector
Submitted :  09/29/93  float irregular, this module will draw
Version :  1.000  a ribbon going along the side of the
Last Updated :  09/29/93  polyline path, from joining up the
Ported To :  HP IBM Sun  specified positions in the scatter

field coord array. The width, and color of the ribbon can be controlled to create a ribbon data representation. Tick marks and labels will annotate the graph. Interpolation can be turned on and off.

show flow

Author :  Phil McDonald  The show flow module generates a
NOAA/ERL/Forecast  geometry object named “flow symbols”
Systems Laboratory  that depicts the flow in a field of 1-,
Number :  1195  2-, or 3-vectors of any data type and
Language :  C  in any grid configuration. Show flow
Submitted :  05/06/92  provides the same basic function as
Version :  1.100  hedgehog but it has been enhanced to
Last Updated :  04/05/93  serve the meteorolog- ical community
Ported To :  IBM Kubota Sun  where wind barbs, rather than arrows,

may be the desired flow symbol. Flow speed values are the resultant values of the field n-vector nalues. The symbols may be rotated about their reference points and their aspect with respect to the viewer can be frozen. While symbols of either type can be arbitrarily sized, arrows can be further scaled by a divisor which is the maximum of the field, the average of the field, or entered manually. The symbols will be colored by their speeds or the values in a second, optional field if an optional colormap is supplied.

str to geom

Author :  Urs Meyer  This module creates a “label” style
Swiss Scientific  text string in the GEOM format which
Computing Center  can be used as a title. It takes a
Number :  1625  string value as input. It can be used
Language :  C  to represent some important variable
Submitted :  07/02/93  such as time, animation step, some
Version :  1.000  parameter, etc. It has the advantage
Last Updated :  07/02/93  over the labeling facilities in the
Ported To :  HP IBM Kubota DEC Sun  Geom Viewer in that the information

will get saved with a network. Other things you can control are the font number, drop shadows, justification, position, height, and color of the text string.

strip charts

Author :  Alan Barnum-Scrivener  I got sick of people saying that AVS
Advanced Visual Systems  won’t do mutiple strip charts, even
Inc.  though we all know “it can’t not do
Number :  1847  it,“ so I whipped up this demo. (It
Language :  C  only works with float data, and assumes
Submitted :  07/16/94  a data range of 0.0 to 256.0, but it
Version :  1.000  does make a convincing demo.) The
Last Updated :  07/16/94  module is a subroutine module that
Ported To :  DEC HP Kubota Sun  takes multiple input signal lines from

a float 2D field, and creates a line map showing the progression of the signal amplitude. The representation is created using AVS geometry, and may be sent to the geometry viewer, “tube” module, etc. An extent box is drawn around the border of the plot area. By using an animated integer controlling a moving crop region, the signal can be animated to simulate a traditional chart recorder.

surface mesh

Author :  Phil McDonald  The surface mesh module creates a 3D
NOAA/ERL/Forecast  wire mesh from a polytriangle object of
Systems Laboratory  a 3D field. Typically, polytriangle
Number :  1270  objects are used to represent surfaces
Language :  C  such as those created by the isosurface
Submitted :  05/27/92  and field to mesh modules. The surface
Version :  1.000  mesh module creates a wire mesh of such
Last Updated :  05/27/92  a surface by passing planes that are
Ported To :  Kubota  perpendicular to the coordinate axes

and finding the line segments that are the intersections between the surface and these planes. The mesh resolution (or distance between the planes) along each axis is independent from that along the other axes. The edges of the polygons defined by the polytriangles may also be shown.

surface web

Author :  Phil McDonald  The surface web module creates a 3D
NOAA/ERL/Forecast  wire web from a surface object created
Systems Laboratory  from a 3D field. Modules that create
Number :  1271  surfaces, like the isosurface and field
Language :  C  to mesh modules, represent the surfaces
Submitted :  05/27/92  with polytriangle strips. The web is
Version :  1.000  created by running “strings”, or lines,
Last Updated :  05/27/92  from each vertex of each polygon
Ported To :  Kubota  defined by these strips to the line

segment opposite the vertex. The web density is the number of strings run from each vertex. The edges of the polygons are always shown, even if the web density is zero.

symmetrize

Author :  Terry Myerson  The symmetrize module inputs a list of
International AVS Center  r and z values ( cylindrical
(NCSC)  coordinates ) which define a
Number :  1137  axisymmetric object, and then
Language :  C  constructs a mesh object around the Z
Submitted :  02/24/92  axis using this list of coordinates to
Version :  1.000  defines the vertices. The number of
Last Updated :  02/24/92  divisions cylindrically around the
Ported To :  Kubota DEC IBM Sun HP  Z-axis, in the theta direction, is

determined by the divisions parameter.

texture mesh

Author :  Wes Bethel  The texture mesh module serves two
Lawrence Berkeley  needs. First, the need for a robust
Laboratory  module for mapping 2D data to
Number :  1295  quadrilateral meshes is served. The
Language :  C  input 2D data may be a 2D uniform
Submitted :  08/04/92  field, in which case the x,y
Version :  2.000  coordinates for the output mesh is
Last Updated :  12/14/93  derived from the extents information
Ported To :  IBM Sun HP  (if present) or using a mapping from

indeces to space if no extents

information is present. The z values are obtained from the data in the field.

tracer bounds

Author :  Lars M. Bishop  Tracer Volume Bounds - This module
National Center for  takes in a 3D Uniform 4 vector Byte
Comp. Elect., Univ. of  field, and adds a white frame to the
Illinois  edges. It is designed to be applied to
Number :  1570  a volume of data right before it is
Language :  C  sent to the tracer module, and allows
Submitted :  01/28/93  the addition of lines that look like
Version :  1.000  the geometry viewer’s volume bounds,
Last Updated :  01/28/93  which are normally inaccessible to
Ported To :  IBM  volume-traced data. It also includes

the ability to fill in the min X,Y, or Z walls, as the original volume bounds module does.

transform tool

Author :  Tim O. Robinson  The transform tool module is used to
UC Berkeley  obtain a copy of the current
Number :  1606  transformation matrix from the geometry
Language :  C  viewer and apply it to an oject in
Submitted :  04/10/93  another geometry viewer window. For
Version :  1.000  example, to capture the transformation
Last Updated :  04/10/93  for the top object in one geometry
Ported To :  DEC IBM Kubota  viewer window and apply it to the top

object in another geometry viewer window, set both typein values to %top. Currently, the transformation is only used after motion of the control geometry stops. Experienced users may experiment with the flags setting in the GEOMedit transform mode call in the module code to alter this behavior.

trivar

Author :  Wes Bethel  The bivar module will resample 1d
Lawrence Berkeley  scatter data in 3 variables. A typical
Laboratory  use for this module is to take a field
Number :  1088  of 3-coordinate, scalar values, and
Language :  C  resample over a “regular” grid. The
Submitted :  10/10/91  algorithm used is one in which, for
Version :  2.000  each output grid location, all input
Last Updated :  05/19/92  points are scanned to determine if they
Ported To :  DEC IBM Sun Kubota  lie within the specified search radius.

If so, that input point’s contribution is weighted as a function of distance from the output grid point and scaled by 1./distance**weightfunction. As such, this module is best suited for use on input data which doesn’t necessarily fit a diffentiable function. For such data, refer to the scat3d module.

ucd cell arrows

Author :  Ian Curington  This module draws vectors with arrow
Advanced Visual Systems  heads from UCD input showing CELL
Number :  1186  vector data. This is a special ucd
Language :  C  mapper module designed to make
Submitted :  05/04/92  representations based on CELL based
Version :  1.000  values rather than the more typical
Last Updated :  05/04/92  NODE based values. If you have CELL
Ported To :  DEC IBM Sun HP  based values, you should also consider

the ucd cell to node module. The module will draw both 2D and 3D vectors, as required. The Scale dial is simply a length multiplier. This module is intended as a source example for other UCD module developers.

ucd cell cones

Author :  Ian Curington  This module draws cones from UCD input
Advanced Visual Systems  showing CELL vector data. The length
Number :  1187  and direction of the cones are
Language :  C  representative of the vector quantity
Submitted :  05/04/92  passed in through the cell data ucd
Version :  1.000  structure. Scale and width dials
Last Updated :  05/04/92  control magnification factors on the
Ported To :  DEC IBM Sun HP  raw data, to get the cones sized for

visibility. The subdivision control defines the course to fine tessalation. This is a special ucd mapper module designed to make representations based on CELL based values rather than the more typical NODE based values. If you have CELL based values, you should also consider the ucd cell to node module. The module will draw both 2D and 3D vectors, as required. The Scale dial is simply a length multiplier. This module is intended as a source example for other UCD module developers.

ucd cell geometr

Author :  Ian Curington  This module draws geometric
Advanced Visual Systems,  representation of UCD structures
Inc.  showing CELL data, CELL Material ID, or
Number :  1188  Cell Number Sequence, Node Number
Language :  C  Sequence, scalar Node data, Cell
Submitted :  05/04/92  topology type, Face order, node
Version :  2.000  positions, or face dip, by color. It
Last Updated :  09/30/93  shows all cells in either solid or wire
Ported To :  HP IBM Sun  frame, unlike ucd to geom and

colors the model directly, without using and external module such as ucd contour or ucd cell colors. It allows picking and moving nodes within the model, and has transformation icons, drag lines, and position markers to aid position editing. The edit information is available in a small field, containing new x-y-z node position, node number, and cell number. Exterior faces only may be displayed with the remove interior option. Downsize will subsample the structure and only display a fraction of cells, for faster interaction. The color range is defined by the colormap hi-low, and DOES NOT autonormalize to the input UCD range, so allows user refinement of the colormap. The “Dip” option allows pseudo-color of surface angle or curvature away from each axis. This is a major update from version 1.0. This module colours each cell ONE color, without color interpolation, showing a single scalar quantity. If your cell data is n-vector, this module will only color the first one, so use ucd extract to get the one you want.

ucd line

Author :  Jonathan Cox  This module produces an arbitrary line
Department of Computer  of sampled node data from a 3D UCD mesh
Science, University of  and outputs it as a field. This can be
Manchester  a conventional AVS 1D field with
Number :  1535  irregular 1D coordinates or a 2D field
Language :  C  suitable for connection to the graph
Submitted :  10/01/92  viewer module. It is also possible to
Version :  3.000  output a uniformly spaced 1D field.
Last Updated :  09/30/93  Version 2 fixes bug concerning
Ported To :  Sun DEC HP IBM  disappearing input ports. Ported to IBM

12/18/92 SRT. Version 3 changes - Code is ANSI C, therefore it requires AVS 5. Modules have been restructured to make them cooperative and reentrant. They are now designed to be compiled in a single executable. If you want to split them up on the FTP site refer to the README file to find which files are required for each module.

ucd line 2

Author :  Jonathan Cox  This module produces an arbitrary line
Department of Computer  of sampled node data from a 3D UCD mesh
Science, University of  and outputs it as a field. This can be
Manchester  a conventional AVS 1D field with
Number :  1767  irregular 1D coordinates or a 2D field
Language :  C  suitable for connection to the graph
Submitted :  12/15/93  viewer module. It is also possible to
Version :  1.000  output a uniformly spaced 1D field. The
Last Updated :  12/15/93  position and orientation of the line is
Ported To :  Sun  arbitrary and is controllable either by

the direct manipulation of a probe object or by manipulating a collection of dial widgets. The original version of this module (also on the IAC’s anonymous ftp site) only works with AVS 4.0, for this enhanced version AVS 5.0 is required.

ucd planes

Author :  Jonathan Cox  The UCD PLANES module supports the
Department of Computer  interactive sub-sampling of node data
Science, University of  held within a 3D UCD structure. The
Manchester  region sampled is selected
Number :  1534  interactively in the region of a
Language :  C  transparent ’probe’ object. Several
Submitted :  10/01/92  different types of probes are
Version :  3.000  supported, all based upon a collection
Last Updated :  09/29/93  of planes. Version 2 fixes bug
Ported To :  Sun DEC HP IBM  concerning disappearing input ports.

Ported to IBM 12/18/92 SRT. Version 3 - Support for the generation of 2D contours has been added. Contour lines can be superimposed on the output of any of the probes inc. the 3D graph. The range of values used can be set using ucd legend in a similar manner to ucd isolines. Also, Sampled region can be output as a UCD structure consisting of triangular elements. A default layout is included in the description function. There are now too many widgets to fit comfortably onto a single control panel. The default layout splits the widgets into three subpanels. Parameter modifications are interpreted more intelligently to avoid needlessly regenerating the sample when it is not needed. Code is ANSI C, therefore it requires AVS 5. Modules have been restructured to make them cooperative and reentrant. They are now designed to be compiled in a single executable. If you want to split them up on the FTP site refer to the README file to find which files are required for each module.

ucd planes 2

Author :  Jonathan Cox  The UCD PLANES module supports the
Department of Computer  interactive sub-sampling of node data
Science, University of  held within a 3D UCD structure. The
Manchester  region sampled is selected
Number :  1768  interactively in the region of a
Language :  C  transparent ’probe’ object. Several
Submitted :  12/15/93  different types of probes are
Version :  1.000  supported, all based upon a collection
Last Updated :  12/15/93  of planes. For each probe type, the
Ported To :  Sun  sampled region can be output either as

various forms of geometry, as a UCD structure or as both. When a geometry is output, 2D contours of a scalar data component can be superimposed on the sample object. The selection of probe type is made from the ’Probe Type’ choice widget on the module’s control panel. Possible probes include a single plane (SIMPLE), two and three crossed orthogonal planes (SHEAF and CROSSED) and multiple parallel planes (MULTIPLE). These probe types produce a ’sample’ that is the same shape as the probe object but is coloured according to a scalar component of the structure’s nodal data. The original version of this module (also on the IAC’s anonymous ftp site) only works with AVS 4.0, for this enhanced version AVS 5.0 is required.

ucd plumbline

Author :  Ian Curington  The plumb line module annotates the
Advanced Visual Systems  top surface of a UCD structure with red
Number :  1690  cross hairs over each node. The cross
Language :  C  hairs are pickable, and change colour.
Submitted :  09/29/93  The node number and x-y coordinate
Version :  1.000  value of the picked node is shown and
Last Updated :  09/29/93  can be controlled with the an integer
Ported To :  HP IBM Kubota Sun  typein, and are output on one integer

and two float output ports. This can be used to control modules that use position dials, perhaps to create a cutting sheet or a graph plot of the selected data section. Downsize control allows a subset of node positions to be visible, while size and offset controls specify the marker shape and position.

ucd profile

Author :  Ian Curington  This module draws Vector Arrows from
Advanced Visual Systems  UCD Cell Surface Boundary data as
Number :  1761  intersected by a surface cross-section
Language :  C  profile. Airflow Lift-Drag analysis
Submitted :  12/15/93  plots from surface pressure can be
Version :  1.000  manipulated or animated dynamically, by
Last Updated :  12/15/93  dragging the profile plane through a
Ported To :  Kubota Sun  large model. The profile sampled data

can be sent to the graph viewer, or other field modules. This mapper module computes the exterior boundary surface of the UCD from node connectivity and material regions, then cuts these faces with an interactive slicing profile plane. It works with 2D shell elements or 3D elements.

ucd streak

Author :  Ian Curington  The ucd streak module generates surface
Advanced Visual Systems  constrained stream lines. It is an
Number :  1762  advanced variation of the
Language :  C  ucd streamlines module in AVS5. It
Submitted :  12/15/93  contains an optional mode where the
Version :  1.000  streamlines follow the boundary of the
Last Updated :  12/15/93  flow domain, tracking along the
Ported To :  Sun  surface, as long at the velocity field

is into the boundary surface, and are not displayed when they leave the surface. This simulates oil drop streak lines from exterior flow wind tunnel tests. The flow is integrated through the 3D boundary elements, and optional integration step markers can be shown. Although the streamlines are only displayed at the surface, the whole interior 3D streamline calculation is taking place, so this method is quite different from a 2D surface only particle track algorithm.

ucd to prim

Author :  Penny Rheingans  This module takes an input ucd
US EPA Scientific  structure and generates a geometric
Visualization Center  primitive (sphere or prism) for each
Number :  1312  node of the structure. Primitives are
Language :  C  colored according to a selected value
Submitted :  08/30/92  at the node and an optional input
Version :  1.000  colormap. If no colormap is provided,
Last Updated :  08/30/92  primitives will be mapped with a grey
Ported To :  DEC Sun HP  scale.

volume mesh

Author :  Ed Bender  The volume mesh module is similar to
Convex Computer  the volume bounds module. The
Corporation  difference is that it generates a mesh
Number :  1593  using a colormap for the desired faces
Language :  C  of the irregular field instead of a
Submitted :  03/04/93  wireframe. This replaces up to six
Version :  2.000  ortho slice/field to mesh
Last Updated :  03/23/93  combinations. By placing the crop
Ported To :  IBM Kubota HP  module upstream, it is easy to control

the location of the slice planes.

Version 2.000 update generates normals.

waffle

Author :  Ian Curington  The waffle module converts a
Number :  1286  two-dimensional floating point field
Language :  C  into group of 3D cross section planes,
Submitted :  06/20/92  represented as a GEOM-format mesh, with
Version :  1.000  the base at zero. Each X or Y line of
Last Updated :  06/20/92  the field is mapped to a “fence”. The
Ported To :  IBM Sun Kubota HP  height of each fence post above each

point is proportional to the scalar value of the field. This is my attempt at “Business Graphics”, for a multi-dimensional telecommunications statistics data application, while allowing 3D interaction.

wire bender

Author :  Phil McDonald  The wire bender module bends line
NOAA/ERL/Forecast  vectors to conform with a surface
Systems Laboratory  defined by a 2D field. The source of
Number :  1194  the input vectors may be either a
Language :  C  geometry object or an ASCII vector
Submitted :  05/06/92  file. If no field is specified, the
Version :  1.000  input vectors are passed through
Last Updated :  05/06/92  unmodified, retaining their original 2D
Ported To :  Kubota  or 3D shapes.

zone fence

Author :  Ian Curington  This module draws a fence panel
Advanced Visual Systems  connecting two wells, joining up common
Number :  1763  zones, and texture map image data onto
Language :  C  the fence panels. This routine takes
Submitted :  12/15/93  three geometric field inputs, one the
Version :  1.000  zonation table, containing a 2D array
Last Updated :  12/15/93  of wells and zones, and the deviation
Ported To :  Sun  scatter path data for two wells. It

creates a geometric representation of the zonation region between the wells. The depths are resampled using linear interpolation from the deviation data to create a 3D sheet. Two optional inputs, texture image A and B, if present will texture map onto alternate panels of the fence.

zone picker

Author :  Ian Curington  This routine takes two 1D inputs, one
Advanced Visual Systems  the zonation table, and second, the
Number :  1764  deviation file for one 3D well track.
Language :  C  It creates a geometric representation
Submitted :  12/15/93  of the zonation picks along the well
Version :  1.000  bore, as golden rings, and allows
Last Updated :  12/15/93  picking and dragging of the positions
Ported To :  Kubota Sun  for editing. It looks at the labels on

the well dev field to figure out which well it is, then indexes the zonations by this, so there is no menu selection needed here. The zone icons can be picked and dragged along the well, the new positions recorded in the memory file structure.

DATA OUTPUT MODULES

International AVS Center

DATA OUTPUT MODULES

Bit Movie

Author :  Ian Curington  The bit movie module is used to create
Advanced Visual Systems  and playback single bit black and white
Number :  1832  movie loops. It takes source images
Language :  C  from any image source in AVS. The
Submitted :  07/16/94  bit movie module saves up movie loop
Version :  1.000  sequences in “The X server” memory, and
Last Updated :  07/16/94  then allows for playback after storage.
Ported To :  DEC HP IBM Sun  The motivation is to create the longest

animation sequences possible, at the highest possible display rate, without going to disk.

CREATE MINI DATA

Author :  Kathy Pearson  The CREATE MINI DATA module is used in
MCNC Environmental  conjunction with the AVS module “write
Programs  field“, and other specially designed
Number :  1794  modules include “FIND MAP”,
Language :  C  “MULTI extract”, and “WRITE
Submitted :  04/21/94  TIMESTAMP“. The MCNC Environmental
Version :  1.000  Programs Group has created this module
Last Updated :  04/21/94  as part of a suite of AVS demo modules,
Ported To :  DEC HP IBM Kubota Sun  networks, and sample data sets to

demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

CREATE USER MINI

Author :  Kathy Pearson  CREATE USER MINI DATA – creates a
MCNC Environmental  local data set from a remote data file
Programs  and place it in the users own directory
Number :  1795  structure. The MCNC Environmental
Language :  C  Programs Group has created this module
Submitted :  04/21/94  as part of a suite of AVS demo modules,
Version :  1.000  networks, and sample data sets to
Last Updated :  04/21/94  demonstrate the functionality of its
Ported To :  DEC HP IBM Sun  visualization prototype for an

Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

Create MPEG

Author :  Alex Knowles  Create MPEG is a subroutine module
The University of  which will create MPEG movies from a
Manchester  series of AVS images. It uses Andy
Number :  1789  Hung’s ’MPEG’ which is available by
Language :  C  anonymous ftp from Stanford. Because of
Submitted :  04/21/94  the way MPEG movies are encoded by
Version :  2.000  ’MPEG’, many temporary files need to be
Last Updated :  07/15/94  stored, bear this in mind when choosing
Ported To :  DEC HP IBM Sun  which directory to make your movie in.

Version 2 fixes bugs.

Draw Map

Author :  Jeff Wang  This module allows users to draw the
MCNC, Environmetal  United States map with state boundary
Programs Group  and county boundary. With the
Number :  1741  distribution of the module, two map
Language :  C  data files are also included.
Submitted :  12/15/93  “allcounty.txt” is ASCII data for
Version :  1.000  boundaries of ALL counties in the
Last Updated :  12/15/93  United States, including Alaska, Hawaii
Ported To :  Sun  and Perto Rico. “countryd.txt” is ASCII

data for boundaries of the states. Users may construct their own boundary map according to the format. This is one of a suite of modules we used to plot the emission data for the state of North Carolina. This group of modules allows displaying points, lines and polygons over a map of the United States. The module suite is stored as a group in a compressed tar file at the International AVS Center, with some sample networks and datasets included. You may find it easier to access the modules this way. The compressed tar file is approximately 1.2MB in size, and can be found via anonymous ftp to avs.ncsc.org, in /sample data/avs data/Wang GIS mods s.tar.Z.

GEOM2IV

Author :  YOSHIKAWA, Hiroyuki  This module interprets AVS geometry
FUJITSU NAGANO SYSTEMS data into Inventor format(V2.0 ascii)
ENGINEERING LTD  and writes it out to a new file.
Number :  1898
Language :  C
Submitted :  09/11/95
Version :  1.000
Last Updated :  09/11/95
Ported To :  DEC ultrix DEC alpha

GenMask

Author :  Stefano Brozzi, Andrea  The module GenMask is an input module
Menozzi, Giulio E.  created to generate the templates, both
Destri  the control and the feedback template,
Universita’ degli Studi  to apply to the image when the image
di Parma, 43100 Parma,  proc essing is executing. As a input
Italy  module GenMask does not have any input
Number :  1874  and it gives as output a field 2-D
Language :  C  scalar float that is a 2-dimensional
Submitted :  09/06/95  array of float. Since the module aCNN
Version :  1.000  works with templates of dimension 3x3
Last Updated :  09/06/95  even the output array of GenMask will
Ported To :  DEC ultrix DEC alpha  be 3x3.

Geom to Wavefron

Author :  Steve Chall  “Geom to Wavefront” accepts as input an
MCNC Visualization Group AVS GEOMedit list struct which
Number :  1790  contains a pointer to a linked list of
Language :  C  GEOMedit structs. The list is traversed
Submitted :  04/21/94  and each element which is of type
Version :  1.000  GEOM EDIT GEOMETRY is examined to
Last Updated :  04/21/94  determine its GEOMobj type–as opposed
Ported To :  DEC HP IBM Sun  to GEOMedit type–where the possible

types are mesh, polyhedron,

polytriangle strip array, sphere, or label. An equivalent Wavefront .obj file is written to disk for each polyhedron and polytri in the edit list. Note that a single AVS geometry may map to multiple .obj files.

HdB1

Author :  Han de Bruijn  This product was developed for C3820,
TUD Computing Centre,The ConvexOS/10, Fortran Compiler 6.0. A
Netherlands  so-called GEOM driver for the
Number :  1297  Application Visualization System (AVS)
Language :  CandFORTRAN  has been built into the classical
Submitted :  08/04/92  RC-plotroutines (Calcomp), as an AVS
Version :  1.000  course material spin-off. This means
Last Updated :  08/04/92  that AVS *.geom files can be produced

by the plotroutines, which can be read in and visualized by the Geometry Viewer.

Image to Postscr

Author :  Lars M. Bishop  The Image to Postrscript module is a
National Center for  module that prints an Adobe Postscript
Comp. Elect., Univ. of  file containing the image that is
Illinois  connected to the input port. It is
Number :  1578  designed to give the user a wide
Language :  C  variety of image sizes, orientations,
Submitted :  02/11/93  positions, and qualities. Note that the
Version :  1.000  Dump to File button must be pressed to
Last Updated :  02/11/93  dump the image to the file, as the dump
Ported To :  DEC IBM Kubota  may take several seconds.

Record Anima

Author :  Brian Kaplan  Used to record AVS animations to
Center for Innovative  videotape or laserdisc, one frame at a
Computer Applications  time. It works by talking with either a
Number :  1096  Lyon-Lamb Minivas videotape controller,
Language :  C  or a Sony Laser Videodisc
Submitted :  10/16/91  recorder/player. Record Animation sends
Version :  2.300  the appropriate commands to initialize
Last Updated :  10/16/91  the recording device, to tell the
Ported To :  Kubota  recording device when to record a

frame, and to wait for the recording device to complete the frame record before continuing with the next frame. Usually, the Record Animation module is used in conjunction with an animation module of some sort which produces geometry for each frame of the animation. This geometry is then fed into the “render geom” module to be rendered as a pixmap output. The pixmap would then be connected to “display pixmap“ and also to the pixmap input of the Record Animation module. The Record Animation module uses the pixmap to tell when the current scene has changed so that it will know to send the appropriate record command to the recording device.

Spline Animator

Author :  Mitchell Roth  Version 2.0 is basically just the text
University of Alaska  version (i.e. 1.0) with a Motif
Number :  1860  interface dropped on top. The goal is
Language :  C  to provide easier script creation and
Submitted :  06/01/95  adding, as well as a few new features.
Version :  2.000  Interested users should consult the
Last Updated :  06/01/95  Instructions.ps file for a general

guide on creating scripts with v2.0. Not covered in the instructions is the Convert script which is intended to allow v1.0 users to port their scripts to v2.0 format. The v2.0 format will be used in the final AVS/Motif version. The Convert script is simple to use - Convert source file dest file

Trace Render

Author :  Ian Curington, Advanced  This is a display module, it takes a 2D
Visual Systems, Inc.  floating point field, treats each row
Danny Harvey, University  as a 1D signal, and draws the signal
of Colorado  traces next to each other. The offset,
Number :  1687  amplitude, signal spacing, etc. can be
Language :  C  controlled for any range of data. The
Submitted :  09/29/93  signals can be plotted vertically,
Version :  1.000  horizontally, or in 180 degree polar
Last Updated :  09/29/93  configurations. Additional controls
Ported To :  HP Sun  allow for Seismic moveout

interpretation, given an Time Offset (T0) and a Velocity, linear or hyperbolic moveout functions may be overlayed and optionally applied to the trace data. The display is like wiggle trace, but the areas under the curve are not filled in. The module catches mouse events in the window, but these are currently commented out. One possibility is to provide dynamic pan and zoom under mouse control, or allow signal picking. If the input is 1D only, it will draw one signal line, and may be used like a real-time oscilliscope display.

WRITE ANY IMAGE

Author :  Terry Myerson  This module writes an image from an AVS
International AVS Center  Network in a variety of formats. All
(NCSC)  formats which the San Diego
Number :  1212  Supercomputing Center’s image tools
Language :  C  support, this module supports. With
Submitted :  05/21/92  this module you can write out an AVS
Version :  2.000  image in MacPaint format for inclusion
Last Updated :  05/21/92  in a to-be published document .. Any of
Ported To :  Sun DEC IBM HP  the following image file formats can be

read in by this module eps Encapsulated PostScript file, gif Compuserve Graphics image file, hdf Hierarchical Data File, icon Sun Icon and Cursor file, iff Sun TAAC Image File Format, mpnt Apple Macintosh MacPaint file, pbm Portable Bit Map file, pcx ZSoft IBM PC Paintbrush file, pgm Portable Gray Map file, pic PIXAR picture file, pict Apple Macintosh QuickDraw/PICT file, pix Alias image file, pnm Portable aNy Map file, ppm Portable Pixel Map file, ps PostScript file, ras Sun Rasterfile, rgb SGI RGB image file, rla Wavefront raster image file, rle Utah Run length encoded image file, rpbm Raw Portable Bit Map file, rpgm Raw Portable Gray Map file, rpnm Raw Portable aNy Map file, rppm Raw Portable Pixel Map file, synu Synu image file, tiff Tagged image file, x Stardent AVS X image file, xbm X11 bitmap file, xwd X Window System window dump image file. Version 2.0 frees allocated memory after writing each image.

WRITE TIMESTAMP

Author :  Kathy Pearson  WRITE TIMESTAMP – write into the
MCNC Environmental  specified file name the input timestamp
Programs  string. The WRITE TIMESTAMP module
Number :  1821  creates an ASCII file and writes into
Language :  C  it the set of timestamp labels input as
Submitted :  04/21/94  one string. This file is needed to
Version :  1.000  indicate the date and time represented
Last Updated :  04/21/94  by AVS data fields that have been
Ported To :  DEC HP IBM Sun  created by extracting pieces from

original model files of type ROM, RADM, UAM and ALPHA. The MCNC Environmental Programs Group has created this module as part of a suite of AVS demo modules, networks, and sample data sets to demonstrate the functionality of its visualization prototype for an Environmental Decision Support System (EDSS). You can download these modules individually from the IAC using anonymous ftp to the appropriate directories /avs modules at avs.ncsc.org. If you would like to download the entire suite of modules, networks, and sample datasets at the same time, the IAC recommends that you use anonymous ftp to grab the following file from avs.ncsc.org - /sample data/avs data/MCNC Env Pgms s/MCNC Env Pgms.tar.Z This contains all source code to all local and remote modules, as well as all of the sample datasets, networks, documentation, installation, and usage notes.

WrUcdAnim

Author :  D. LE CORRE  This DATA output module generates a
CONVEX S.A  sequence of UCD data time steps onto a
Number :  1317  (set of) file(s), which may then be
Language :  C  read by the twin module, RdUcdAnim. For
Submitted :  08/30/92  each time step, it can store node
Version :  1.200  positions, node data, cell data, model
Last Updated :  12/30/92  data or a subset of these. It is fully
Ported To :  Sun HP  compatible with the standard read ucd

format, so the first time step may be read by read ucd.

DATA OUTPUT MODULES

avs2vrml

Author :  John Evans  The avs2vrml module saves an AVS
U.S. Geological Survey  geometry data structure as an
Number :  1899  equivalent VRML Version 1.0 compliant
Language :  C  *.wrl file. Meshes, polyhedrons,
Submitted :  09/21/95  spheres, and polytriangles are
Version :  1.000  fully-supported. Labels don’t quite
Last Updated :  09/21/95  work yet.
Ported To :  DEC alpha

avs2vrml

Author :  John Evans  The avs2vrml module saves an AVS
U.S. Geological Survey  geometry data structure as an
Number :  1900  equivalent VRML Version 1.0 compliant
Language :  C  *.wrl file. Meshes, polyhedrons,
Submitted :  09/21/95  spheres, and polytriangles are
Version :  1.000  fully-supported. Labels don’t quite
Last Updated :  09/21/95  work yet.
Ported To :  DEC ultrix DEC alpha

comput vol

Author :  Andy R. Haas  This module accepts a geometry object
National Center for  (mesh, polyhedron, polytriangle, or
Toxicological Research  sphere type) and outputs the object’s
Number :  1868  volume and surface area. The name of
Language :  C  the object and its and area values are
Submitted :  06/27/95  also printed to the terminal window in
Version :  1.000  which AVS was started.
Last Updated :  06/27/95
Ported To :  DEC ultrix DEC alpha

compute volume

Author :  Andy R. Haas  This module accepts a geometry object
Number :  1897  (mesh, polyhedron, polytriangle, or
Language :  C  sphere type) and outputs the object’s
Submitted :  09/07/95  volume and surface area. The name of
Version :  1.000  the object and its volume and area
Last Updated :  09/07/95  values are also printed to the terminal
Ported To :  DEC ultrix DEC alpha  window in which AVS was started.

continuous s

Author :  Terry Myerson  Keeps a running tally of the biggest
International AVS Center  and smallest numbers to come through
(NCSC)  its input ports. The text string
Number :  1136  parameters will update each time a
Language :  C  smaller of larger value comes through
Submitted :  02/24/92  one of its ports. If the reset button
Version :  1.000  is turned on, then the running tally is
Last Updated :  02/24/92  reset to have to current minimum or
Ported To :  Kubota DEC IBM Sun HP  maximum. For example, if on one

execution the module received 0.020 and -0.020, and at the next execution the module receives 0.039 and -0.05, then the string parameters will display 0.039 and -0.020. If you have a time series of data and you want to find out the absolute minimum and maximum of the data values over time, this module allows that capability.

create unix pipe

Author :  John Stephen  This module creates and sets up a unix
Tessella Support  pipe file, for communication between an
Services  AVS module and an external application.
Number :  1645  It makes unix system calls to create
Language :  C  the pipe file, based on a typein name.
Submitted :  09/24/93  It helps manage the use of the pipe,
Version :  1.000  with options to flush, etc. It outputs
Last Updated :  09/24/93  the full pathname of the pipe, once
Ported To :  HP IBM Sun  established, for downstream modules to

use for the communication.

deleteGeom

Author :  John Stephen  This module deletes objects from the
Tessella Support  geometry viewer on command, based on
Services  the name (alias) of the module that
Number :  1647  created the geometry.
Language :  C
Submitted :  09/29/93
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Kubota Sun

delete object

Author :  John Stephen  Deletes a geometry object having a user
Tessella Support  supplied name.

Services

Number :  1648
Language :  C
Submitted :  09/29/93
Version :  1.000
Last Updated :  09/29/93
Ported To :  HP IBM Sun

dira2avs

Author :  Massimiliano Bavelloni  Read data from ascii file, data may be
Radioastronomy  integer or floating point numbers
department C.N.R.  separated by spaces.

Bologna,Italy

Number :  1875
Language :  C
Submitted :  09/06/95
Version :  1.000
Last Updated :  09/06/95
Ported To :  DEC ultrix DEC alpha

display ARGB vol

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1823  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  There are actually two modules here,
Last Updated :  07/15/94  which have identical functionality.
Ported To :  HP  They both takes in a word/voxel dataset

(with packed ARGB) and a transform matrix and renders a volume in a window. The only difference is that the first module uses the PRE SCALED ALPHA data form, while the second does not. For an explanation of this CUBIC POLYPOINT option see the man page for the module.

display lit volu

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1822  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  This module takes in a byte/voxel
Last Updated :  07/15/94  dataset, a colormap, and a transform
Ported To :  HP  matrix and renders a volume in a

window. The difference with this module is that it implements a form of gradient shading. For coding convenience and speed the light source direction has been limited to one of the 26 vectors which pass directly through an adjacent voxel. The gradient magnitude is approximated as the magnituded difference of voxel under computation and the adjacent voxel along the light source vector. This magnitude is store in a second byte (per voxel) along with the original voxel scalar magnitude byte. This 2 byte per voxel field is fed to the CUBIC POLYPOINT gescape along with a table which contains a lighting model applied to the original colormap supplied to the model. This lighting model is driven by ambient, diffuse, and specular weighting and theta/phi light source angles supplied as input parameters.

display tracker

Author :  Ian Curington  Track 2D cursor motion over displayed
Advanced Visual Systems  image. This module takes an input image
Number :  1327  and displays it in an X-window. Inputs
Language :  C  an image and allows the user create
Submitted :  08/31/92  mouse clicks and moves in the image
Version :  1.000  window, with event information passed
Last Updated :  08/31/92  to other AVS modules using the
Ported To :  “upstream data” concept. This module

acts similarly to “display image” or “display tracker” in that it accepts an input image and displays this in an X-window. It also then accepts asynchronous mouse events in the window, and passes these mouse events ( in an upstream transform structure ) to an output port. This allows applications to be built using this module to do all the X dependent code, leaving the actual operation (such as pan, draw, etc.) to an upstream module, with no concern of X events.

display volume

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1824  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  This module takes in a byte/voxel
Last Updated :  07/15/94  dataset, a colormap, and a transform
Ported To :  HP  matrix and renders a volume in a

window.

dump icc

Author :  Wes Bethel  This module converts an input pixmap to
Lawrence Berkeley  the so-called ICC format. This image
Laboratory  format is used by local device drivers
Number :  1029  which communicate with the Kodak XL7700
Language :  C  printer. The single parameter is a file
Submitted :  10/10/91  browser that allows you to specify the
Version :  1.000  name of the ICC file to be created.
Last Updated :  10/10/91  After the file is written, the filename
Ported To :  Sun HP  is reset to null. This prevents

subsequent changes upstream in the network from automatically triggering the rewriting of the file. A new file is written only after the user has entered a filename.

dump ps

Author :  Wes Bethel  This module converts a pixmap to the
Lawrence Berkeley  PostScript page description language
Laboratory  and stores it in a file. After the file
Number :  1030  is written, the filename is reset to
Language :  C  NULL. This prevents subsequent changes
Submitted :  03/20/92  upstream in the network from
Version :  2.000  automatically triggering the rewriting
Last Updated :  10/10/91  of the file. A new file is written only
Ported To :  Sun Kubota HP  after the user has entered a filename.

Two types of PostScript output are supported, PostScript-compatible monochrome laser printer and PostScript-compatible color laser printer.

dump sunras

Author :  Wes Bethel  This module produces a sun rasterfile
Lawrence Berkeley  from an input AVS format pixmap. This
Laboratory  module is familiar with 8,24 and 32-bit
Number :  1015  pixmaps. It will write an 8-bit
Language :  C  rasterfile when presented an 8-bit
Submitted :  10/10/91  pixmap as input, otherwise a 24-bit BGR
Version :  1.000  (“RT STANDARD”) rasterfile is
Last Updated :  10/10/91  produced.
Ported To :  Sun HP

field2 to Math

Author :  Tom Wickham-Jones  The “field2 to Math” module allows a
Wolfram Research  two-dimensional scalar field to be
Number :  1539  imported into Mathematica from AVS.
Language :  C  When the “field2 to Math” module is
Submitted :  10/14/92  invoked an xterm is started running
Version :  1.000  Mathematica with the input and output
Last Updated :  10/14/92  of the xterm directed to and from the
Ported To :  Sun DEC Kubota  Mathematica session main loop. Entering

Mathematica commands into the window will result in the appropriate Mathematica results. The session is initialised with a package of Mathematica commands which have started a MathLink communication channel to AVS. Other initialisations are made so that the Mathematica command AVSReadField[ ] will read a two-dimensional scalar field from AVS. Please note that at the present time the IAC does not have Mathematica installed on a , Sun, or Kubota platforms. As a result this module was placed on the ftp site using the Makefiles provided by the author, without testing.

field info

Author :  Kathleen Dyer  Calculates the maxima and minima of
Lawrence Livermore  input field data and of the grids. If
National Laboratory  the field has more than three
Number :  1117  dimensions, grid information is
Language :  C  obtained for only the first three. The
Submitted :  01/22/92  information is used to produce GEOM
Version :  1.000  text objects. The min/max parameters
Last Updated :  01/22/92  allow individual objects to be toggled
Ported To :  Kubota DEC IBM Sun HP  on or off. The “Text Height” parameter

sets the height of the GEOM text as a fraction of the window height.

field to EXCEL

Author :  Ian Curington  ASCII Excel spreadsheet format files
Advanced Visual Systems  are produces containing field data in a
Number :  1654  table format. These files can be
Language :  C  imported into a suitable excel format
Submitted :  09/29/93  for analysis and plotting. Separator
Version :  1.000  characters may be specified for
Last Updated :  09/29/93  differing output versions, three
Ported To :  HP IBM Sun  possible floating point output formats

are selectable, and the maximum number of values per line is adjustable. The module only outputs the file if the file browser changes. This keeps the output file from beeing written too often.

fix points

Author :  Larry Gelberg  “orthoslice” and “field to mesh”
Stardent Computer  communicate through a secret code
Number :  1107  stored in the “points” array of the
Language :  C  field structure. This will be an array
Submitted :  11/27/91  of 6 floating point numbers
Version :  1.000  representing [minx, maxx, miny, maxy,
Last Updated :  11/27/91  minz, maxz] for the 2D (3-space) field
Ported To :  Kubota DEC IBM Sun HP  which orthoslice produces. FIELD TO

MESH looks at these searching for when the min == max to determine which direction (I, J, or K) the slice was taken from. This is the means that FIELD TO MESH uses to orient and position the slices properly in the volume. The way to defeat this is to impose your own point information on the existing ones. For instance, you have a 64x64x64 volume and you’re taking the 12th K slice. The points array will look like - [0, 63, 0, 63, 12, 12]. If you overwrite the Z mins/maxes with 0.0’s, the slice will appear positioned on the XY plane (not 12 units above it). Similarly, the 23rd I slice points array looks like [23, 23, 0, 63, 0, 63] and will produce a slice parallel to the YZ plane at X=23. By overwriting the 23’s with 0’s, you will get it to lie right on the YZ plane. This should put between the orthoslice and the field to mesh modules in your network.

geom parent

Author :  Ed Bender  The geom parent module allows an
Convex Computer  interactive user to reparent any and
Corporation  all objects in the geometry viewer
Number :  1592  system. This is useful if you want some
Language :  C  objects to move together, but not all.
Submitted :  03/04/93  The “Group Name” parameter is used to
Version :  1.000  define the group to add objects to.
Last Updated :  03/04/93  When the “Include Current Object”
Ported To :  DEC IBM Kubota HP  oneshot is pushed, the current object

selected within the geometry viewer is reparented to the “Group Name”.

gui page

Author :  John Stephen  To help place, size, and maintain user
Tessella Support  interface panels and widgets, this
Services  module gives dial and text parameter
Number :  1659  control over the “layout” of any item,
Language :  C  without using the layout editor. Once
Submitted :  09/29/93  an application is built, this module
Version :  1.000  can will enable the layout to be
Last Updated :  09/29/93  dynamically changed by the end user, or
Ported To :  HP IBM Kubota Sun  used to prototype a layout, which can

then be saved. Panel visibility can be controlled, as well as parent, place, and size. It used the CLI as the control path.

image 2 icc

Author :  Wes Bethel  Converts its input image to the
Lawrence Berkeley  so-called ICC format. This image format
Laboratory  is used by local device drivers which
Number :  1045  communicate with the Kodak XL7700
Language :  C  printer. The single parameter is a file
Submitted :  10/10/91  browser that allows you to specify the
Version :  1.000  name of the ICC file to be created.
Last Updated :  10/10/91  After the file is written, the filename
Ported To :  DEC IBM Sun HP  is reset to NULL. This prevents

subsequent changes upstream in the network from automatically triggering the rewriting of the file. A new file is written only when you enter a filename.

image to PPM

Author :  P. Fulconis  This is one of two modules written to
CICG computer center of  translate AVS animations into mpeg
Grenoble University,  format - image to YUV and
France  image to PPM. Both formats can be
Number :  1772  read by mpeg encoder. A ppm image will
Language :  C  have a memory size twice bigger than
Submitted :  02/09/94  the yuv one, but it can be visualised
Version :  1.000  by xv. You will need to make use of
Last Updated :  02/09/94  mpeg encode and mpeg play, both
Ported To :  IBM  public domain utilities available via

anonymous ftp from the University of California at Berkeley. These modules also make use of the San Diego Supercomputer Center’s Image Tools, also available via anonymous ftp.

image to YUV

Author :  P. Fulconis  This is one of two modules written to
CICG computer center of  translate AVS animations into mpeg
Grenoble University,  format - image to YUV and
France  image to PPM. Both formats can be
Number :  1773  read by mpeg encoder. A ppm image will
Language :  C  have a memory size twice bigger than
Submitted :  02/09/94  the yuv one, but it can be visualised
Version :  1.000  by xv. You will need to make use of
Last Updated :  02/09/94  mpeg encode and mpeg play, both
Ported To :  IBM  public domain utilities available via

anonymous ftp from the University of California at Berkeley. These modules also make use of the San Diego Supercomputer Center’s Image Tools, also available via anonymous ftp.

image to can

Author :  Hideo Miyachi  This module prints out an AVS image to
Kubota Corporation  a laser shot printer. You can specify
Number :  1766  the actual command to print in the
Language :  C  parameter “Command”. If you specify
Submitted :  12/15/93  “cat ¿ /tmp/filename” in “Command”, you
Version :  1.000  can save the image in the file in lbp
Last Updated :  12/15/93  format.
Ported To :  Kubota Sun

INTERNATIONAL AVS CENTER MODULE CATALOG

multi geom view

Author :  John Stephen  This coroutine takes upstream transform
Tessella Support  events from any geometry viewer scene,
Services  and outputs two transformation control
Number :  1671  ports, such that a small iconic scene
Language :  C  can have remote control over two or
Submitted :  09/29/93  more main geometry viewer windows. Once
Version :  1.000  the position is defined from the icon
Last Updated :  09/29/93  window, pushing the “snap” button will
Ported To :  HP Sun  send the transformation over to the

other window.

multiformat writ

Author :  A. Jacobson  This module writes an image to disk
UCLA Dept. Pharmacology from AVS in any of a variety of
/ Crump Institute  formats. It also includes the
Number :  1850  multiformat read code, and both will
Language :  C  compile into a single binary. Some of
Submitted :  10/12/94  the formats which this module supports
Version :  1.000  are from San Diego Supercomputing
Last Updated :  10/12/94  Center’s image tools package, others
Ported To :  DEC HP IBM Sun  are supported from John Cristy’s

ImageMagick package, still others require tools from GNU’s Ghostscript, and the NetPBM package. Depending on how the module is configured, it can write about 30 to 40 image formats, including , AVS X, Alias, Apple Macintosh MacPaint, Apple Macintosh QuickDraw/PICT, CCIR 601, Compuserve Graphics, Flexible Image Transport System (monochrome only), Hierarchical Data File, JFIF Rev1.02, Khoros VIFF, Magick, Mark VanDeWettering’s MTV ray tracer image format, Microsoft Windows bitmap, Microsoft Windows cursor, Microsoft Windows icon, PIXAR picture, Photo CD, PostScript, Radiance, Radiance, SDSC Synu, SGI RGB, Sun Icon and Cursor, Sun Rasterfile, Sun TAAC Image, Tagged, Truevision Targa, Utah Run length encoded, Wavefront raster, X Window System window dump, X11 bitmap, X11 pixmap, ZSoft IBM PC Paintbrush, binary or ASCII PBM Portable Bit Map, binary or ASCII PBM Portable Gray Map, binary or ASCII PBM Portable Pixel Map, binary or ASCII PBM Portable aNy Map. This module contains a superset of the formats in the READ ANY IMAGE module, with several new formats added and improved support for several of the original file types. Please note that this module calls several executable files, including imconv, convert, and gzip. Platform specific binaries for several of these programs are available from the IAC ftp site with this module. See the .txt file of this module or the file multiio.h for further details on the sources for this module’s tools. The tools are stored at the IAC in one large compressed tar file, which is 11MB and is stored at avs.ncsc.org in /misc tools/multiio tools/multiio to ools.tar.Z.

296

object opera

Author :  Kathleen Dyer  Recreates the functionalities of some
Lawrence Livermore  of the transformation options available
National Laboratory  in the geometry viewer. While this may
Number :  1123  seem redundant at first glance, it has
Language :  C  proven to be a very useful module. The
Submitted :  01/22/92  module is mainly used in networks where
Version :  1.000  the Layout Editor is used to customize
Last Updated :  01/22/92  the user interface. The locations and
Ported To :  DEC IBM Sun HP  appearances of this module’s control

widgets can be altered, while their corresponding widgets in the Geometry Viewer cannot. The module has no input or output ports. To include it in a network, simply drag it into the workspace.

object propertie

Author :  John Stephen  This module creates a menu structure
Tessella Support  for all or part of the geometry viewer
Services  scene, and allows alteration of the
Number :  1674  object properties by simple color or
Language :  C  transparency selection without going to
Submitted :  09/29/93  the geometry viewer controls, for a
Version :  1.000  simpler user interface.
Last Updated :  09/29/93
Ported To :  HP IBM Sun

object visibilit

Author :  John Stephen  This module takes input strings for the
Tessella Support  name of the object, the name of the
Services  view, and has a boolean toggle control
Number :  1673  to control the visible/invisible, or
Language :  C  hide/show on that object in the view.
Submitted :  09/29/93  It uses the CLI to perform the control,
Version :  1.000  hence it has no output port.
Last Updated :  09/29/93
Ported To :  HP IBM Sun

output 8bit

Author :  Wes Bethel  Takes as input an AVS image field and
Lawrence Berkeley  creates a postscript file suitable for
Laboratory  printing on a monochrome postscript
Number :  1073  printer. The postscript file contains a
Language :  C  representation of the input color image
Submitted :  10/10/91  using the “image” postscript command,
Version :  2.000  using 8bits of gray. The color image is
Last Updated :  10/10/91  squashed down to 8 bits of gray by
Ported To :  Kubota DEC IBM Sun HP  weighting each of the color channels in

an attempt to approximate luminance. The postscript image is scaled so as to be centered on the page and to occupy as much area on the page as possible. The nice thing about the file produced is that it will work and produce similar results on, for example, a 200dpi device (like an Apple Laserwriter) or on a 600dpi device (like a Verityper). After executing, the module resets the filename parameter to NULL. This has the effect of requiring the user to enter a new filename for each subsequent invocation of this module before executing.

output a60

Author :  Wade Smith  The output a60 module takes an AVS
Convex Computer  image data structure as input, converts
Number :  1190  it into Abekas YUV format and sends it
Language :  C  to an Abekas a60 digital disk recorder.
Submitted :  05/05/92  The user can specify the hostname of
Version :  1.000  the a60 digital disk recorder on the
Last Updated :  05/05/92  network, the insert point and the
Ported To :  DEC HP IBM  number of times to repeat each frame.

The insert point may be entered in any of four time formats (hh mm ss ff, mm ss ff, ss ff, or ff).

output color

Author :  Wes Bethel  Takes as input an AVS image field and
Lawrence Berkeley  creates a postscript file suitable for
Laboratory  printing on a color postscript printer.
Number :  1072  After executing, the module resets the
Language :  C  filename parameter to NULL. This has
Submitted :  10/10/91  the effect of requiring the user to
Version :  1.000  enter a new filename for each
Last Updated :  10/10/91  subsequent invocation of this module
Ported To :  Kubota DEC IBM Sun  before executing.

point sample

Author :  Anonymous  Extracts a single scalar value out of
Stardent Computer  the input 2D scalar floating point
Number :  1066  field. The X and Y position parameters
Language :  FORTRAN  set the index point for the sample to
Submitted :  10/10/91  be taken from the field. The numerical
Version :  1.000  field value is then sent to stdout,
Last Updated :  10/10/91  usually into the window where avs was
Ported To :  Kubota IBM HP  invoked. The one line message consist

of the current X and Y sample index position, and the floating point value of the field at that point. The position values are clamped to the field dimensions internally, without regard to the dial min-max. This module does virtually no work, so “immediate” mode is recommended on the dials. The author has used this for debugging other networks and filters, by linking it in with a “t” junction to other more complex networks.

print iv pick

Author :  Ian Curington  Print the contents of the “iv pick”
Advanced Visual Systems  output port of the AVS4 image viewer
Number :  1324  module. This shows the results from
Language :  C  picking on images with the left mouse
Submitted :  08/31/92  button. It is a basic module showing
Version :  1.000  the possibilities, and used for
Last Updated :  08/31/92  debugging of other image pick
Ported To :  DEC HP  application modules. This module relies

on features in AVS4, and will not function with earlier releases. This module will be useful for those building interactive image processing systems.

print real

Author :  Lars M. Bishop  Print real creates a widget on the
National Center for  control panel that will contain the
Comp. Elect., Univ. of  value of the module’s input port. This
Illinois  module is part of the the Real Math
Number :  1575  module group, which is a single file,
Language :  C  process-sharing set of modules that
Submitted :  02/11/93  allow basic operations on and between
Version :  1.000  integer and real AVS data types. They
Last Updated :  02/11/93  have been most useful in dealing with
Ported To :  DEC IBM  linkage of several AVS parameters. AVS

has no simple way of performing math operations on reals, even though it does allow math operations on fields.

print verts

Author :  Terry Myerson  This module inputs any geometry, and
International AVS Center  outputs to the avs console window all
(NCSC)  of the vertices in the geometry objects
Number :  1196  present in the geometry edit list. This
Language :  C  module initially can serve as a
Submitted :  05/06/92  debugging tool for modules that output
Version :  1.000  geometries. It also serves as an
Last Updated :  05/06/92  example module on how to process
Ported To :  DEC IBM Sun Kubota HP  geometries on a module’s input port. It

most useful purpose however, is probably to extract quantitative information regarding surfaces, from visualization displays.

print xfrm

Author :  Ian Currington  Is a diagnostic tool, used to show the
Advanced Visual Systems  contents of “upstream transform” data
Number :  1069  types, sent upstream for interactive
Language :  C  control in AVS3. It uses “user defined
Submitted :  10/10/91  data types“.
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

print xgeom

Author :  Ian Currington  Is a diagnostic tool, used to show the
Advanced Visual Systems  contents of “upstream geom” data types,
Number :  1070  sent upstream for interactive control
Language :  C  in AVS3. It uses “user defined data
Submitted :  10/10/91  types“.
Version :  1.000
Last Updated :  10/10/91
Ported To :  Kubota DEC IBM Sun HP

psfilter

Author :  Wes Bethel  The “output 8bit ps” module takes as
Lawrence Berkeley  input a scalar (byte, single channel)
Laboratory  field and creates a postscript file
Number :  1177  suitable for printing on a monochrom
Language :  C  postscript printer. The postscript file
Submitted :  10/10/91  contains a representation of the input
Version :  2.000  color image using the “image”
Last Updated :  05/19/92  postscript command, using 8bits of
Ported To :  DEC IBM Sun Kubota HP  gray. The postscript image is scaled so

as to be centered on the page and to occupy as much area on the page as possible. The nice thing about the file produced is that it will work and produce similar results on, for example, a 200dpi device (like an Apple Laserwriter) or on a 600dpi device (like a Verityper).

remote lit volum

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1825  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  This module prototypes an ability to
Last Updated :  07/15/94  make use an HP graphics devices volume
Ported To :  HP  rendering capability remotely. The idea

is to run AVS on the HP with the X display routed to a networked server, but make use of the HP workstation and one of the above listed display devices to render volume dataset on the HP then send the results (as AVS images) to the remote X display. To use this module you’ll need to setup the environment variable COMPUTE SERVER to point at the X server on the HP server that will be performing the rendering. This includes the same gradient lighting model described for the disp lit vol module.

remote volume

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1826  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  This module prototypes an ability to
Last Updated :  07/15/94  make use an HP graphics devices volume
Ported To :  HP  rendering capability remotely. The idea

is to run AVS on the HP with the X display routed to a networked server, but make use of the HP workstation and one of the above listed display devices to render volume dataset on the HP then send the results (as AVS images) to the remote X display. To use this module you’ll need to setup the environment variable COMPUTE SERVER to point at the X server on the HP server that will be performing the rendering.

statistics

Author :  Anonymous  Is a field diagnostic and analysis
Stardent Computer  module, similar to VOLSTAT, used to
Number :  1082  obtain global statistical information
Language :  C  about field data. This module scans the
Submitted :  10/10/91  data in the volume, and produces a
Version :  1.000  small table, sent to stdout, usually
Last Updated :  10/10/91  into the window where avs was invoked,
Ported To :  Kubota HP  unless redirected. The table includes

the following information about a field- Dimensions, Min/Max, Mean, Median, Standard Deviation, Skewness, and Kurtosis.

stereo lit volum

Author :  Howard Stroyan  This AVS module is designed to provide
Hewlett-Packard Advanced  access to the HP specific back-to-front
Systems Division  Volume Rendering graphics functionality
Number :  1827  of the Starbase graphics API. This
Language :  C  module can only be successfully built
Submitted :  07/15/94  and executed on an HP700 workstation.
Version :  1.000  This module assumes that a stereo ready
Last Updated :  07/15/94  monitor and viewing system (i.e.
Ported To :  HP  StereoGraphics’ Crystal Eyes) are

available. It will render a stereoscopic version of the disp lit vol module. This module make use of a similar environment variable as the COMPUTE SERVER variable used by the remote volume module. This module uses the environment variable STEREO SERVER. Set this variable to point at the stereo-ready device. [In order to use this module well it is required that a second X display of some kind is available.] Run AVS with the DISPLAY set to the non-stereo screen or server. This module will open a full screen window on the display pointed to by STEREO SERVER and will place the display in stereo mode (via the Starbase STEREO gescape). Stereo mode can be toggled off and on with the “Stereo” control button. (Turn off the stereo mode before deleting this module or exiting AVS, or your stereo display will be left in stereo mode.)

string to file

Author :  Ian Curington  Sends the contents of the string input
Advanced Visual Systems  port to an output file, opened with a
Number :  1561  file browser. The string can replace
Language :  C  the current contents of the file, or be
Submitted :  11/16/92  appended to the file since opened
Version :  2.000  during the session, as a log of strings
Last Updated :  09/30/93  in the network, controlled by a mode
Ported To :  IBM HP  toggle. This is designed to be used to

send command strings via a named pipe (see mknod pipe.file p) to another application program, where command input has been redirected to the same pipe. This allows an AVS user interface modules to control an external command-line driven application.

ucd statistics

Author :  anonymous  Ucd statistics is a diagnostic and
Number :  1895  analysis module, similar to the field
Language :  C  module statistics, used to obtain
Submitted :  09/07/95  statistical information about field
Version :  1.000  data. This module scans the data in the
Last Updated :  09/07/95  volume, and produces a small table,
Ported To :  DEC ultrix DEC alpha  sent to a scrollable text window in the

module control panel. The table contains the statistics for each data component calculated separately. Then if there was more than one component, the table contains the statistics for all of the data components grouped together. It will create statistics for both node and cell data. It will also generate overall statistics for node data and cell data separately.

ucd to wave

Author :  Ion Barosan  The ucd to wave module writes a ucd
Eindhoven University of  structure to disk, in wavefront format,
Technology, The  which is supported by Data Visualizer.
Netherlands  Ucd to wave works only with node data,
Number :  1597  therefore if you have cell data, you
Language :  C  have to transform the cell data to node
Submitted :  03/04/93  data using the cell to node module.
Version :  1.000  Because Data Visulizer doesn’t use
Last Updated :  03/04/93  mid-edge nodes, at this version, the
Ported To :  DEC IBM  ucd to wave module generally doesn’t

use the mid-edge nodes. One exception has been made - if the ucd cell is triangle with all mid-edge nodes present, ucd to wave module will split the triangle in 4 triangles. The user can change the module to support another ucd type which use mid-edge nodes. The wavefront file that will be written to disk is an ASCII file.

unix command

Author :  A. Jacobson  Often there will be a need to convert a
UCLA Dept. Pharmacology string in AVS to a command or argument
/ Crump Institute  in unix to be performed at the command
Number :  1831  line. This module allows the user to
Language :  C  output strings and integers to the
Submitted :  07/16/94  command line for execution by the shell
Version :  1.000  in the directory that AVS was run from.
Last Updated :  07/16/94
Ported To :  DEC HP IBM Kubota Sun

volume stati

Author :  Ian Curington  Is a diagnostic and analysis module,
Stardent Computer  used to obtain global statistical
Number :  1091  information about scalar volume data.
Language :  C  Often the result of numerical analysis
Submitted :  10/10/91  and visualization leads to an interest
Version :  1.000  in just a few numbers such as the
Last Updated :  10/10/91  average value in the field. This
Ported To :  Kubota IBM HP  “volume statistics” module scans the

data in the volume, and produces a small table, sent to stdout, usually into the window where avs was invoked, unless redirected. The table shows the dimensional size of the field, total number of data points, the minimum and maximum scalar values, and the mean. This is a very simple module, and could be extended to measure other parameters such as standard deviation, median, etc. This module does virtually no computationally intensive work, has no parameters. The author has used this for debugging other networks and filters, by linking it in with a “t” junction to other more complex networks. In particular, when using “interpolate”, it is nice to know the resulting volume dimensions.

write A60 yu

Author :  Terry Myerson  Intended to be used with the AVS
International AVS Center  Animator application. The module hooks
(NCSC)  right up to the Animator module or the
Number :  1106  read frame seq module. This module
Language :  C  accepts images, one at a time. The
Submitted :  11/27/91  image is then converted into the Abekas
Version :  3.000  A60 yuv format - with the same
Last Updated :  02/09/94  dimensions of the Abekas A60 yuv format
Ported To :  DEC IBM Sun HP  - and output to a file in the Frame

directory with a name corresponding to the frame number. ( i.e. Frame Number 20 – Filename - 20 ) If the image is not the same size as the A60 image standard, then the AVS image is centered within the view of the Abekas. This works well when you consider the degradation of the image borders on its path to NCSC. This module was adapted from Ian Currington’s write abekas module and Abekas Inc.’s sample to yuv program in their manual. It writes out a temporary file in between conversions. The module works fine - but it isn’t pretty. Version 3 update provided by the UCLA School of Medicine’s Andy Jacobson. The update includes bug fixes, and features wherein the user can name the file as with the animate filename module, but with a fixed .yuv siffix. The user also has the choice of compressing the (huge.) output files. In some nets it may be desireable to allow or prevent overwriting of files, so a switch is now provided for this. Also you can supply your favorite compression program to this module, if desired.

write Dore i

Author :  Anonymous  Interprets AVS image data into Dore
Kubota Pacific Computer  format and writes it out to a new file.

Inc

Number :  1154
Language :  C
Submitted :  03/12/92
Version :  1.000
Last Updated :  03/12/92
Ported To :  Kubota DEC IBM Sun HP

write KSWAD

Author :  Anonymous  Interprets AVS image data into KSWAD
Kubota Pacific Computer  format and writes it out to a new file.

Inc

Number :  1157
Language :  C
Submitted :  03/12/92
Version :  1.000
Last Updated :  03/12/92
Ported To :  Kubota HP

write MooV

Author :  Charles J Williams III  This module is a first shot at
Naval Research Lab  incorporating QuickTime(tm) into AVS as
Number :  1544  a method for producing convenient,
Language :  C++  compact, and high quality video
Submitted :  10/14/92  segments. QuickTime(tm) is rather
Version :  1.000  convenient since there is likely to be
Last Updated :  10/14/92  more support on other systems, and
Ported To :  Sun  there is a growing base of

QuickTime(tm) cheaply available for the Macintosh. This code currently requires, the QuickTime Movie Exchange Toolkit, which is available from Apple, Inc, and a C++ compiler (you also need the C++ compiler to build the QuickTime(tm) libraries. Please note - since the IAC doesn’t have the Exchange Toolkit, we weren’t able to test this module out here.

write abekas

Author :  Ian Currington  This module provides I/O capability to
Stardent Computer  images in the native machine format
Number :  1076  used by the abekas A60 digital video
Language :  C  disk recording system. The image file
Submitted :  10/10/91  specification is for the PAL (50 Hz,
Version :  1.000  625 line) version, but can easily be
Last Updated :  10/10/91  modified to support the NTSC version,
Ported To :  Kubota DEC IBM Sun HP  or both. The abekas A60 holds 750

frames, or 30 seconds of live broadcast standard video imagery in the digital storage format dictated by the CCIR-601 specification. Although the disk format is Y-uv, the disk has firmware processing to encode/decode to 3-byte Red-Green-Blue images on-the-fly during tcp/ip file transfers.

write compressed

Author :  Wes Bethel  The “write compressed image” module
Lawrence Berkeley  provides the same functionality as the
Laboratory  AVS-supplied “write image” module, with
Number :  1162  the added twist that the image is in a
Language :  C  “compressed” format. An AVS image may
Submitted :  03/15/92  be compressed using the UNIX(tm)
Version :  2.000  compress utility, or by using the
Last Updated :  05/19/92  sibling module “write compressed
Ported To :  Kubota IBM Sun DEC HP  image.“ Compression ratios of 1 to 10

are easily achieved using this

compression scheme.

write eps image

Author :  Terry Myerson  This module writes an image in the eps
International AVS Center  format from an AVS Network.
(NCSC)  Encapsulated PostScript color,
Number :  1239  grayscale, and black-and- white image
Language :  C  files are used by a variety of
Submitted :  05/21/92  PostScript tools in order to include
Version :  1.000  PostScript diagrams and images within
Last Updated :  05/21/92  other documents. Encapsulated
Ported To :  Sun DEC IBM  PostScript files are recognized by the

follow- ing filename suffixes .eps,

.epi, .epsi, and .epsf.

DATA OUTPUT MODULES

write g 2

Author :  John Langner  This is yet another module to write a
Stardent Computer  Compuserve GIF image file.
Number :  1294
Language :  C
Submitted :  07/11/92
Version :  1.000
Last Updated :  07/11/92
Ported To :  IBM Sun HP

write geom

Author :  Georg Thallinger  The write geom module saves the
Joanneum Research Graz,  geometry in a file. Currently there is
Austria  a choice of three formats - avs-text,
Number :  1698  avs-binary, dxf (drawing exchange
Language :  C  format by AUTODESK). The dxf-routines
Submitted :  09/30/93  supply a rudimentary header, every
Version :  1.000  object is placed on a single layer
Last Updated :  09/30/93  (PRISMS converts every layer in an
Ported To :  IBM  object). The dxf-output is at last

readable by PRISMS, ALIAS and of course AutoCAD.

write gif image

Author :  Terry Myerson  This module writes an image in the gif
International AVS Center  format from an AVS Network. GIF
(NCSC)  (Graphics Interchange Format) is
Number :  1240  CompuServe’s standard for generalized
Language :  C  color raster images. This standard is a
Submitted :  05/21/92  mechanism to exchange and display
Version :  1.000  high-quality, high- resolution graphics
Last Updated :  05/21/92  images. CompuServe gif files are
Ported To :  Sun DEC IBM  recognized by the filename suf- fixes

.gif and .giff.

write hdf image

Author :  Terry Myerson  This module writes an image in the hdf
International AVS Center  format from an AVS Network. hdf is a
(NCSC)  generic, tagged Hierarchical Data File
Number :  1241  format developed by the National Center
Language :  C  for Supercomputing Applica- tions
Submitted :  05/21/92  (NCSA). hdf files may contain images,
Version :  1.000  scientific data sets, and miscellaneous
Last Updated :  05/21/92  data items. Such files may be created
Ported To :  Sun DEC IBM  by several NCSA tools. See the NCSA

documentation for details on how to use these tools. hdf format handling within the SDSC image library is limited to images of certain depths and storage methods. hdf files are recognized by these filename suffixes .hdf, .df, and .ncsa.

write icon image

Author :  Terry Myerson  This module writes an image in the icon
International AVS Center  format from an AVS Network. icon image
(NCSC)  files are used by Sun Microsystem’s
Number :  1242  SunView win- dow system, NeWS window
Language :  C  system, OpenWindows NeWS tool set, and
Submitted :  05/21/92  X11 XView tool set for the storage of
Version :  1.000  icons, cursors, fill patterns, and
Last Updated :  05/21/92  pieces of widgets (like button check-
Ported To :  Sun DEC IBM  marks). Sun icon files can be most

easily generated using Sun’s iconedit(1) icon and cursor editor. The Sun operating sys- tem release includes a directory of standard icons, cursors, background patterns, and widget pieces in icon format in the directory /usr/include/images. See the Sun documentation set for details on how to use the tools dealing with Sun icon files. Sun icon files are recognized by these filename suffixes .icon, .cursor, and .pr.

write iff image

Author :  Terry Myerson  This module writes an image in the iff
International AVS Center  format from an AVS Network. iff image
(NCSC)  files are generated by Sun Microsystems
Number :  1264  TAAC software such as voxvu(1) and
Language :  C  cloudvu(1). See the TAAC-1 Application
Submitted :  05/21/92  Accelerator Software Reference Manual
Version :  1.000  for information on how to use these
Last Updated :  05/21/92  programs. Note that image files
Ported To :  Sun DEC IBM  compressed with the Sun-TAAC utility

make movie(1) cannot be read by the SDSC image library. iff files are recognized by the following filename suffix .iff.

write irreg

Author :  Wes Bethel  Writes an ASCII representation of an
Lawrence Berkeley  input irregular field to a file. The
Laboratory  file format is the “LBL” irregular
Number :  1092  field file format, and is readable by
Language :  C  the “read irreg” input module. This
Submitted :  10/10/91  module is useful for looking at
Version :  1.000  irregular fields, for example, after
Last Updated :  10/10/91  some processing step in a data flow
Ported To :  Kubota DEC IBM Sun HP  network. A browser is supplied to let

the user specify the filename to which the ascii representation will be written.

write jpeg

Author :  John Langner  Compresses an image with the JPEG
Stardent Computer  compression standard and writes to a
Number :  1102  file.The best and most readily
Language :  C  available introduction to the JPEG
Submitted :  11/22/91  compression algorithm is Wallace’s
Version :  1.000  article in the April ’91 CACM- Wallace,
Last Updated :  11/22/91  Gregory K. “The JPEG Still Picture
Ported To :  IBM Sun HP  Compression Standard“, Communications

of the ACM, April 1991 (v. 34 no. 4), pp. 30-44. (Adjacent articles in that issue discuss MPEG motion picture compression, applications of JPEG, and related topics.) I highly recommend reading that article before looking at any of the JPEG code. For more detail about the JPEG standard you pretty much have to go to the draft standard, which is not nearly as intelligible as Wallace’s article. The current version is ISO/IEC Committee Draft CD 10918-1 dated 1991-03-15.

write mpnt image

Author :  Terry Myerson  This module writes an image in the mpnt
International AVS Center  format from an AVS Network. The
(NCSC)  MacPaint mpnt file is the standard
Number :  1243  Apple Macintosh mono- chrome bitmap
Language :  C  image file format. It can be read by
Submitted :  05/21/92  many Macintosh graphics applications,
Version :  1.000  and many Macintosh applica- tions that
Last Updated :  05/21/92  can export bitmap graphics do so in the
Ported To :  Sun DEC IBM  MacPaint mpnt file format. mpnt files

are recognized by these filename suffixes .mpnt, .macp, and .pntg.

write p 2

Author :  John Langner  Writes a PC Paintbrush PCX image file.
Stardent Computer  Note that a PCX file can contain a
Number :  1293  maximum of 256 colors. For maximum
Language :  C  flexibility, a separate module is used
Submitted :  07/11/92  to pick those colors.
Version :  1.000
Last Updated :  07/11/92
Ported To :  IBM Sun HP

write pbm image

Author :  Terry Myerson  This module writes an image in the pbm
International AVS Center  format from an AVS Network. pbm image
(NCSC)  files are used by various tools in Jef
Number :  1244  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pbm files are
Version :  1.000  recognized by the filename suffix .pbm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

write pcx image

Author :  Terry Myerson  This module writes an image in the pcx
International AVS Center  format from an AVS Network. The pcx
(NCSC)  image file format was invented by ZSoft
Number :  1245  for use in its IBM PC PC Paintbrush
Language :  C  tool series. It has become a defacto
Submitted :  05/21/92  standard in the IBM PC world and is
Version :  1.000  regularly used for the storage of
Last Updated :  05/21/92  monochrome and color pixel information
Ported To :  Sun DEC IBM  by paint-type tools. See the

documentation for each of the IBM PC tools for details on how to use them. ZSoft pcx files are recognized by these filename suffixes .pcx and .pcc.

write pgm image

Author :  Terry Myerson  This module writes an image in the pgm
International AVS Center  format from an AVS Network. pgm image
(NCSC)  files are used by various tools in Jef
Number :  1246  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pgm files are
Version :  1.000  recognized by the filename suffix .pgm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

write pic image

Author :  Terry Myerson  This module writes an image in the pic
International AVS Center  format from an AVS Network. pic image
(NCSC)  files are generated by PIXAR
Number :  1247  programming software, such as
Language :  C  ChapVolumes and ChapReyes, the PIXAR
Submitted :  05/21/92  Image Runtime Library called Pirl, and
Version :  1.000  the PIXAR rendering tool RenderMan. See
Last Updated :  05/21/92  the PIXAR documentation set for details
Ported To :  Sun DEC IBM  on how to use these applications and

tools. PIXAR’s pic file format is sometimes referred to as PICIO in PIXAR documentation. pic and PICIO mean the same thing. PIXAR’s xpic is not the same as pic and is not supported by the SDSC image library. Note PIXAR’s RenderMan always saves its image files with .pic filename suffixes. However, depending upon output defaults, RenderMan can generate .pic files with PICIO (same as pic) data or TIFF data. .pic files with TIFF data will confuse the SDSC image library. For you to avoid generating this type of file, we recommend that you configure RenderMan defaults to generate .pic files with PICIO data. PIXAR pic files are recognized by the following filename SDSC Last change March 4, 1991 1 IMPIC(3IM) SDSC IMAGE LIBRARY IMPIC(3IM) suffixes .pic, .picio, and .pixar.

write pict image

Author :  Terry Myerson  This module writes an image in the pict
International AVS Center  format from an AVS Network. The PICT
(NCSC)  file is the standard Apple Macintosh
Number :  1248  image file format. It can be read by
Language :  C  almost any Macintosh graphics
Submitted :  05/21/92  application. Most Macintosh
Version :  1.000  applications that can export graphics
Last Updated :  05/21/92  do so in the PICT file format. pict
Ported To :  Sun DEC IBM  files are recognized by these filename

suffixes .pict and .pict2.

write pict seque

Author :  Michael Rangitsch and  This module writes a series of images
Terry Myerson  in the pict format from an AVS Network.
Dow Chemical and  The file names are constructed from the
International AVS Center  base file name and the frame number as
Number :  1577  such ¡base file name¿.¡frame
Language :  C  number¿.pict The frame number is padded
Submitted :  02/11/93  to the left with “0” to allow proper
Version :  1.000  sequencing of frames. The intent of
Last Updated :  02/11/93  this is to allow conversion of a
Ported To :  HP IBM Sun  sequence of images into the pict format

for video production.

write pix image

Author :  Terry Myerson  This module writes an image in the pix
International AVS Center  format from an AVS Network. pix image
(NCSC)  files are generated by the rendering
Number :  1249  and painting tools of Alias Research,
Language :  C  Inc., such as renderer, raytracer, and
Submitted :  05/21/92  paint. See the Alias documentation set
Version :  1.000  for details on how to use these tools.
Last Updated :  05/21/92  ΦN Φo Φt Φe Alias quickpaint,
Ported To :  Sun DEC IBM  available on Silicon Graphics, Inc.,

IRIS workstations, uses Silicon

Graphic’s rgb image file format rather than the Alias pix image file format. See the imrgb(3IM) man page for details on the rgb format. Alias pix files are recognized by these following filename suffixes .alias and .pix.

write pnm image

Author :  Terry Myerson  This module writes an image in the pnm
International AVS Center  format from an AVS Network. pnm image
(NCSC)  files are used by various tools in Jef
Number :  1250  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. pnm files are
Version :  1.000  recognized by the filename suffix .pnm.
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

INTERNATIONAL AVS CENTER MODULE CATALOG

write ps image

Author :  Terry Myerson  This module writes an image in the ps
International AVS Center  format from an AVS Network. PostScript
(NCSC)  color, grayscale, and black-and-white
Number :  1251  image files are used by a variety of
Language :  C  PostScript laser printers and win-
Submitted :  05/21/92  dowing system tools. PostScript files
Version :  1.000  are recognized by the following
Last Updated :  05/21/92  filename suffixes .ps and .postscript.
Ported To :  Sun DEC IBM

write ras image

Author :  Terry Myerson  This module writes an image in the ras
International AVS Center  format from an AVS Network. ras image
(NCSC)  files are used by various Sun
Number :  1252  Microsystems Inc. tools, such as
Language :  C  screendump(1) and screenload(1). See
Submitted :  05/21/92  the Sun documentation set for details
Version :  1.000  on how to use these tools. Sun ras
Last Updated :  05/21/92  files are recognized by any of the
Ported To :  Sun DEC IBM  following filename suffixes .ras, .scr,

.sr, and .sun.

write rectilinea

Author :  Wes Bethel  This module write data in ascii format
Lawrence Berkeley  from a “rectilinear field” data
Laboratory  structure. The disk file is structured
Number :  1296  to contain information about the field
Language :  C  at the beginning of the file, followed
Submitted :  08/04/92  by the data.
Version :  1.000
Last Updated :  08/04/92
Ported To :  DEC IBM Sun HP

write reg

Author :  Wes Bethel  Writes an ASCII representation of an
Lawrence Berkeley  input uniform field to a file. The file
Laboratory  format is the “LBL” uniform field file
Number :  1093  format, and is readable by the “read
Language :  C  uniform“ input module. This module is
Submitted :  10/10/91  useful for looking at uniform fields,
Version :  1.000  for example, after some processing step
Last Updated :  10/10/91  in a data flow network. The uniform
Ported To :  Kubota DEC IBM Sun HP  input field can be any type, size, and

dimensionality. A browser is supplied to let the user specify the filename to which the ascii representation will be written.

write rgb image

Author :  Terry Myerson  This module writes an image in the rgb
International AVS Center  format from an AVS Network. rgb image
(NCSC)  files are generated by Silicon
Number :  1253  Graphics, Inc., software such as
Language :  C  icut(1) and snapshot(1). See the
Submitted :  05/21/92  Silicon Graphics documentation for
Version :  1.000  information on how to use these and
Last Updated :  05/21/92  other Silicon Graphics programs.
Ported To :  Sun DEC IBM  Silicon Graphics rgb files are

recognized by the following filename suffixes .rgb, .iris, and .sgi.

write rgb sequen

Author :  Michael Rangitsch and  This module writes a series of images
Terry Myerson  in the rgb format from an AVS Network.
Dow Chemical and  The file names are constructed from the
International AVS Center  base file name and the frame number as
Number :  1576  such ¡base file name¿.¡frame
Language :  C  number¿.rgb The frame number is padded
Submitted :  02/11/93  to the left with “0” to allow proper
Version :  1.000  sequencing of frames. The intent of
Last Updated :  02/11/93  this is to allow conversion of a
Ported To :  HP IBM Sun  sequence of images into the rgb format

for video production.

write rla image

Author :  Terry Myerson  This module writes an image in the rla
International AVS Center  format from an AVS Network. rla image
(NCSC)  files are generated by Wavefront
Number :  1254  Technologies, Inc., software such as
Language :  C  image and tdv. See the Wavefront
Submitted :  05/21/92  documentation for information on how to
Version :  1.000  use these and other Wavefront programs.
Last Updated :  05/21/92  rla is a subset of the newer Wavefront
Ported To :  Sun DEC IBM  rlb specification. Programs that can

read and write rlb files can also read and write rla files. The SDSC image library reader/writer is written to accommodate both types of files. Wavefront rla files are recognized by the following filename suffixes .rla and .rlb.

write rle image

Author :  Terry Myerson  This module writes an image in the rle
International AVS Center  format from an AVS Network. rle image
(NCSC)  files are generated by the tools of
Number :  1255  Utah’s Raster Toolkit. See the Utah
Language :  C  documentation set for details on how to
Submitted :  05/21/92  use these tools. Utah rle files are
Version :  1.000  only recognized by the following
Last Updated :  05/21/92  filename suffix .rle
Ported To :  Sun DEC IBM

write rpbm image

Author :  Terry Myerson  This module writes an image in the rpbm
International AVS Center  format from an AVS Network. rpbm image
(NCSC)  files are used by various tools in Jef
Number :  1256  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpbm files are
Version :  1.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpbm.
Ported To :  Sun DEC IBM

DATA OUTPUT MODULES

write rpgm image

Author :  Terry Myerson  This module writes an image in the rpgm
International AVS Center  format from an AVS Network. rpgm image
(NCSC)  files are used by various tools in Jef
Number :  1257  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpgm files are
Version :  1.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpgm.
Ported To :  Sun DEC IBM

write rpnm image

Author :  Terry Myerson  This module writes an image in the rpnm
International AVS Center  format from an AVS Network. rpnm image
(NCSC)  files are used by various tools in Jef
Number :  1258  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rpnm files are
Version :  1.000  recognized by the filename suffix
Last Updated :  05/21/92  .rpnm.
Ported To :  Sun DEC IBM

write rppm image

Author :  Terry Myerson  This module writes an image in the rppm
International AVS Center  format from an AVS Network. rppm image
(NCSC)  files are used by various tools in Jef
Number :  1259  Poskanzer’s PBM+ tool suite. See the
Language :  C  PBM documentation set for details on
Submitted :  05/21/92  how to use these tools. rppm files are
Version :  1.000  recognized by the filename suffix
Last Updated :  05/21/92  .rppm.
Ported To :  Sun DEC IBM

write synu image

Author :  Terry Myerson  This module writes an image in the synu
International AVS Center  format from an AVS Network. synu is the
(NCSC)  image file format output by SDSC’s synu
Number :  1260  (Syn- thetic Universe) portable
Language :  C  renderer. synu files are recognized
Submitted :  05/21/92  only by the following filename suffix
Version :  1.000
Last Updated :  05/21/92
Ported To :  Sun DEC IBM

write tiff image

Author :  Terry Myerson  This module writes an image in the tiff
International AVS Center  format from an AVS Network. tiff is a
(NCSC)  generic Tagged Image File Format
Number :  1261  developed by Aldus and Microsoft in
Language :  C  conjunction with leading scanner and
Submitted :  05/21/92  printer manufacturers. tiff files may
Version :  1.000  contain images and miscellaneous other
Last Updated :  05/21/92  image-related items. Such files may be
Ported To :  Sun DEC IBM  created and manipulated by a variety of

Tagged Image File Format tools. See the Tagged Image File Format documentation for details on how to use these tools. tiff support within the SDSC image library is limited to images of certain depths and storage methods. tiff files are recognized by the following filename suf- fixes .tiff and .tif.

write xbm image

Author :  Terry Myerson  This module writes an image in the xbm
International AVS Center  format from an AVS Network. xbm bitmap
(NCSC)  image files are generated by MIT’s X
Number :  1262  Window Sys- tem, version 11 (hereafter
Language :  C  referred to as X11). xbm files are
Submitted :  05/21/92  generated by the X11 bitmap(1) bitmap
Version :  1.000  editor and used by most X11 tools to
Last Updated :  05/21/92  define cursors, icons, and other mono-
Ported To :  Sun DEC IBM  chrome glyphs. See the X11

documentation set for details on how to use tools and subroutines dealing with X11 bitmaps. X11 xbm files are recognized by the following filename suf- fixes .xbm and .bm.

write xwd image

Author :  Terry Myerson  This module writes an image in the xwd
International AVS Center  format from an AVS Network. xwd window
(NCSC)  dump image files are used by the xwd(1)
Number :  1263  and xwud(1) tools of MIT’s X Window
Language :  C  System, version 11 (hereafter referred
Submitted :  05/21/92  to as X11). See the X11 documentation
Version :  1.000  set for details on how to use these
Last Updated :  05/21/92  tools. X11 xwd files are recognized by
Ported To :  Sun DEC IBM  the following filename suf- fixes .xwd

and .x11.

Catalog

vi

Data Input Modules  0
3D axis  ACE to UCD  AFS extract  AMS extract  ANIMATE DUAL
ARC MAP  Application Menu BGF Protein Disp BP Coroutine  ConstImage
DMOL to field Dials Contro  Dials Control  Dipole  Edit File
Edit String  Extrusion  FIND DATA AND MA FIND MAP  FITS read
GET MODEL MAP Generate label  Get Terrain Data Graph slider  Image Sequen
LBF extract  Label Axis  Life  Life WT  Load Simulation
MOSS LINES MOSS Markers  MOSS POLYGONS Menger sponge  Mouse Trap
Mouse Trap IV QUAL field  RADM Credits  READ ANY IMAGEREAD TIMESTAMP
READ TIN  READ WIND TIMEST RdUcdAnim  ReadLat  ReadRadiance
ReadSeqAnim Read 2  Read CFD  Read DXF  Read EJ65
Read Hologram Read SDRC  Read SEG Y  Read Unversal FI Sin Wav
Sonify  Splash  Tick Marks  Vector Label  Vernier Dial
WriteSeqAnim Zoomer  animate file  animate floa  animate inte
animated boolean animated str  animated track  application mana awais
cabinet maker ccyl pipe  choice 2  choice alias  choice selec
color cube  color palette  component menu compress file  coord2geom
crystal  curdle  cylinders  draw arrow  draw polygon
endif  extruder  fast animate  field2 from Math field3 from Math
file path and na file to field  flight path  fractal 3d f  fractal field
gaussian field  generate contine  generate gri  genloadmd  geom from Math
geometry sphere graduate32  grid generator  harwell colo  if 2
image snap  int positiv  integer list  integer step  jigsaw
lathe  lines to geom  load md  load mdn  macro menu
mandelbrot  mask generator  menu example  microscope  mk vel2
mk vel  multiformat read named toggle  newfilla  noise source
pdb to geom  phoenics int  pick center  place camera  plate
plato super  plot xyz  plot xyz col  r tif ra  r tifvol ra
readFLOW3D read 16 bit  read ACR NEMA read AIFF  read Dore im
read F3D  read FLOW3D  read GOCAD  read HDF SDS  read Leica confo
read abekas  read blokjes  read compressed  read dem  read dyna3d
read eps image read gif2  read gif  read gif image  read grd to fiel
read hdf image read icon image  read iff image  read irregul  read jpeg
read mpnt image read netcdf  read netcdf example read p 2  read pbm image
read pcx image read pgm image  read pic image  read pict image  read pix image
read plot3d  read pnm image  read points  read ras image  read rectilinear
read rgb image read rla image  read rle image  read rpbm image read rpgm image
read rpnm image read rppm image read semper  read shak  read sunras
read synu image read tiff image  read tiff sectio  read tiff vol  read ucd points
read uniform  read xbm image  read xwd image  river bed  shaker
sololoadmd  spaceball master  spaceball slave  sphere to geom  string example
string list  string macros  strip chart  taperer  teapot
track cursor  transform xyz  tree  trigger int  trigger int 2
ucd color ra  ucd comp  uncompress file  uv rect  uv sector

waves demo

4D to 3D Sli  ColorToBW  Color Legend  Cylinder  Delta Int
Downsize 2  E01SAF  Extract label  Field Concat  Field Formula
Field Math  GRID MATH  GeomXYZtoFld  Geom Dup n Mirro Geom Duplicate
Geom Unique Name Geos Image  IDL  IMAGE WINDOW Input Formula
Int to string  Intervals Field  Intervals UCD  Iterate  LatLon UTM
MINMAX 2 3 MOVING AVG  MathAVS  MinMax 2  Nvect to 2d
Quantize Image Reset Legend  SIDE BY SIDE  Shadow  SlowDown
Sobel operat  Sphere  Stepper  Toggle Black Whi Topo2irreg
UcdColorizeGeom UcdXYZtoFld  Unjaggy  WINDOW MATH WIND MINMAX
WIND SPEED X Crop  X ray  Z slicer  aCNN
add 2 ints  add coords 2  add coords 3  add cps  add extents
adjust rgb  alpha mix  and or merge  arith binary  arith logical
arith unary  arithmetic 2  backgroundRGB  byte image t  classify
closing  collage  collapse  collapse 2 num  colorizer 2d
colour cones  combine  combine vecto 2  combine vectors  complex math
compute shade concat str  cone  cone 2  contour perc
convert color  convert data  crop2  cylinder 2  delay
demultiplex  dilate  downsize points  draw pixel  dual control
erode  excavate  exponential  extend  extrude field
feature  fft2d  field 2D to 3D  field averaging  field control
field conver  field label  field mux  field scatte  field to float parm
field to glo 2  field to globe  field to int parm  field to irr  field to rec
field to string parm field to uni  fill  filter 1Dfreq  filter 1Dtime
filter 2Dfreq  filter 2Dspatl  filter morph  fix field edge  float image
float parm to fieldfloat to str  fourier filter  fourier transfor  gamma correct
geo control  geomanip  geometrySearch  glow white  gradual mix
grey scales  grid scale  histogram  hit lumin mi  hsv to rgb
image combine image flip  image rotate  image stack  image stats
image tile  info  input create1D  input create2D  int 2 st
int parm to field integer control  integer to real  interp cmap  irreg 2 reg
last cone  linearop1D  local area stats  log field  lumin mix
luminence  matrix algebra  matte 3x3  matte sample  min max fiel
mix2  modify seq1D  my mirror  natural log  no format
opening  output dither  output print  panel pan  pencil sketc
persp field  real control  real to integer  reconstruct slic  rect2pol
remap field  remote gis  rgb to cmyk  rgb to hsv  river fill
scat to tri  scat znmo  scatter downsize  scatter to spars  segment
set alpha  set extents  shutter  size fit  slave slicer
smooth  specify slice  spectest1D  standard  stereogram
stick  stream to scatte  string parm to field strip cfd  subregion
super sample  surface  switch 2  switchfire  sysModuleName
table editor  thresholdclamp  transform1D  transform2D  transform po
transpose coords tubeness  ucdTri2prism  ucd extent range  ucd load node da
ucd minmax  ucd particle  ucd polar  ucd select  ucd split
ucd to field  ucd tri to prism  un jag  unif2irreg  unif2zscat
uniform to i  vec2 to vec3  vec mag 2d v  vect div scal 2  vect div scal 5
vect div scalar vector  vector cross  vector cross 2  vector cross 3
vector cross 4  vector cross 5  vector dot  vector dot 3  vector dot 6
vector field  warp  xform field  y join

viii

2D bounds  3D contour map  Analog Clock  CFD object  CFD object conto
CFD objectline ChairPlot 2D Txt Colored bounds  Contour Lines  Contour Mesh
Draw Text  Field To Lines  Field To Points  Field To Polygon Field grid
Field to Isosurf GRAPH2D  Get Local IP Num GraphSound  Grid 2 TxtMesh
Isoband Slice  Isoband UCD  Isoline Slice  Isoline UCD  Keyframe Ani
LABELS UNITS LBL bezier v  MASK TILE  MINMAX REPORT Marker Draw
Marker Gen  Marker Merge  Marker Pick  Marker Select  NEW CROP
NEW LEGEND NEW PRINTFIELD NEW SLICE  NEW TILE  Ortho Texture
Polygon Volume Polyline  QUERY TILE  RESET 2  Reset Geom
Scatter Rect  TEXT TO GEOM TIME SPAN TEXT Title  UCD Node Contour
UCD grid  VIEW SHARE  Vector Arrow  WIND PLOT  Wide Arrow
add polar 2D  area  bin field  bivar  carpet
city scape  color legend  color tile  cone bounding  confocal slicer
contour field  contour to g  contour to posts  crop and slice  draw tracks
field lines  field o matic  field params  field size  field stats
field to con  field to map  glue  grid mesh  hog 2D
interpolating sl irreg2map  labeler  line 1D  line probe
local coord  loop objects  mesh axis  mesh slicer  mesh to ucd
mesh with base new abuffer  new arbitrar  new bubbleviz  new ortho sl
non ortho  ov slicer  plumb line  probe path  real math
rgbtogray  rubberbonds  scat 2d  scat 3d  scat base la
scat bub  scatter arrows  scatter bounds  scatter cubes  scatter rib 2
scatter ribbon show flow  str to geom  strip charts  surface mesh
surface web  symmetrize  texture mesh  tracer bounds  transform tool
trivar  ucd cell arrows  ucd cell cones  ucd cell geometr  ucd line
ucd line 2  ucd planes  ucd planes 2  ucd plumbline  ucd profile
ucd streak  ucd to prim  volume mesh  waffle  wire bender
zone fence  zone picker

Bit Movie  CREATE MINI DATA CREATE USER MINI Create MPEG  Draw Map
GEOM2IV  GenMask  Geom to Wavefron HdB1  Image to Postscr
Record Anima Spline Animator  Trace Render  WRITE ANY IMAGE WRITE TIMESTAMP
WrUcdAnim  avs2vrml  avs2vrml  comput vol  compute volume
continuous s  create unix pipe  deleteGeom  delete object  dira2avs
display ARGB voldisplay lit volu  display tracker  display volume  dump icc
dump ps  dump sunras  field2 to Math  field info  field to EXCEL
fix points  geom parent  gui page  image 2 icc  image to PPM
image to YUV image to can  multi geom view  multiformat writ  object opera
object propertie object visibilit  output 8bit  output a60  output color
point sample  print iv pick  print real  print verts  print xfrm
print xgeom  psfilter  remote lit volum  remote volume  statistics
stereo lit volum string to file  ucd statistics  ucd to wave  unix command
volume stati  write A60 yu  write Dore i  write KSWAD  write MooV
write abekas  write compressed write eps image  write g 2  write geom
write gif image write hdf image  write icon image  write iff image  write irreg
write jpeg  write mpnt image write p 2  write pbm image write pcx image
write pgm image write pic image  write pict image  write pict seque  write pix image
write pnm image write ps image  write ras image  write rectilinea  write reg
write rgb image write rgb sequen  write rla image  write rle image  write rpbm image
write rpgm image write rpnm image write rppm image write synu image write tiff image
write xbm image write xwd image